### v2

# On Topology and Geometry of Hessian Manifolds

Abstract
For a differentiable manifold $M$, a pair $(M,\nabla)$ is termed an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. A Riemannian metric $g$ on $M$ is called a Hessian metric on affine manifold $(M,\nabla)$, if $g$ can be expressed locally as the Hessian quadratic form of some smooth potential $f$. An affine manifold equipped with a Hessian metric is termed a Hessian manifold. We study the topology and geometry of Hessian manifolds.
Key topological constraints are derived, including the vanishing of the Euler characteristic for any compact Hessian manifold and the proof that its fundamental group must be infinite and torsion-free. We demonstrate that compact Hessian manifolds with an abelian fundamental group are homeomorphic to the flat torus. The theory of flat line bundles is developed to introduce the canonical and obstruction bundles, which are used to prove that a curved compact orientable Hessian manifold must be a mapping torus.
A Hessian manifold $((M,\nabla),g)$ is said to be of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $g=\nabla\eta$. We place emphasis on Hessian manifolds of Koszul type, establishing an equivalence between their existence and that of positive flat line bundles, space-like Lagrangian immersions, and hyperbolic affine structures. A duality with radiant manifolds, analogous to the classical Legendre transformation, is explored. We also construct a non-trivial family of compact Hessian manifolds of Koszul type with negative scalar curvature, and classify all complete Hessian surfaces.
keywords:
Hessian manifold, Affine differential geometry, Flat line bundle, Legendre transformation, Lagrangian immersion.
1 Basics on Hessian Geometry
1.1 Introduction and Background
The study of Hessian manifolds originates with the foundational work of J.-L. Koszul in 1961 [1]. While their roots lie in affine differential geometry, in recent decades, these geometric structures have garnered significant attention and emerged as indispensable tools across a diverse spectrum of mathematical and scientific disciplines. Their influence spans from the theoretical underpinnings of information geometry and complex analysis to practical applications in optimization and mathematical statistics, and even touches upon the abstract realms of homological mirror symmetry. This resurgence of interest highlights the deep connections these manifolds forge between seemingly disparate areas, offering a unifying geometric language to explore sophisticated phenomena.
In affine differential geometry and information geometry, Hessian metrics underpin the dually flat spaces used in the study of statistical manifolds, where the Amari–Chentsov tensor codifies divergences and dual connections. Hessian metrics also arise naturally in convex optimization, where barrier functions induce complete Hessian metrics on convex domains. For more details, we refer to [2] and [3].
Also, regarding the applications to mirror symmetry, the seminal work of Hitchin [4], Gross–Siebert [5], and Kontsevich–Soibelman [6] has shown that affine manifolds with singularities and Hessian metrics model the base of torus fibrations in the Strominger-Yau-Zaslow conjecture.
On the other hand, Cheng and Yau’s solution [7] to the real Monge–Ampère equation constructs complete affine hyperspheres of which Blaschke metrics are precisely global Hessian metrics with prescribed volume forms. Remarkably, these examples are linked to convex projective manifolds via the work of Loftin [8], Benoist–Hulin [9], and that of Danciger–Guéritaud–Kassel [10].
However, despite this rich tapestry, fundamental questions about Hessian manifolds remain open, particularly regarding the interplay between global topology, curvature properties, and algebraic obstructions. In this manuscript, we undertake a study of Hessian manifolds from both topological and geometric perspectives:
After briefly summarizing the history and background of the theory of Hessian manifolds, we recall the definitions of affine atlases and flat connections before formally defining Hessian metrics. While the literature contains several well-known examples of Hessian metrics from information theory, such as the canonical metrics on the probability simplexes [11] and the family of Gaussian distributions [11], this collection has remained quite limited and tied to specific contexts. To address this, we introduce a powerful and explicit warped product formula (Theorem1.8), a systematic construction that generates a vast new class of Hessian manifolds and demonstrates that structures like the Robertson-Walker solutions from general relativity are part of this geometric framework, significantly expanding the known examples beyond their original domains.
Our investigation then begins by establishing topological constraints that provide a rigid global framework for any Hessian structure. Foundational results by Shima and Yagi [12] in 1997 established that complete Hessian manifolds are aspherical, as their universal coverings are convex domains. Building on this, and on partial results by Kostant–Sullivan [13] in 1975 and Fried [14] in 1982, our work confirms Chern’s longstanding 1955 conjecture in this general setting by proving that the Euler characteristic of any compact Hessian manifold must vanish (Proposition2.17). We also provide the first proof that the fundamental group of a complete Hessian manifold is torsion-free (Proposition2.19), and give two different proofs of the fact that the fundamental group of a compact Hessian manifold is infinite (Corollary2.7 and Remark2.21). These new algebraic results possess immediate classificatory power, allowing us to prove that any compact Hessian manifold with an abelian fundamental group is homeomorphic to a $n$-torus (Theorem2.22).
Having established this broad landscape, we switch from these foundational constraints to a more refined analysis by developing the theory of flat line bundles as our primary analytical tool. We introduce this as a novel framework for studying Hessian geometry and define two crucial new objects within it: the canonical line bundle, an intrinsic invariant of the affine structure, and the obstruction line bundle. The latter is constructed to precisely measure the failure of locally defined Hessian potentials to patch together into a global function; we show its gauge theoretic triviality is equivalent to the existence of a global potential function. As a direct consequence, we also establish that any orientable compact Hessian manifold with vanishing first Betti number must be flat (Theorem2.40).
We then continue to study the natural question that in which case the Hessian metric admits a globally defined potential function. Using sheaf cohomology and spectral sequences technique, we show that a Hessian manifold admits a single global potential function if its fundamental group is finite (Theorem2.4). Moreover, it is readily seen that a compact Hessian manifold cannot support a global potential function. This problem brings a historically significant class of Hessian manifolds into sharp focus: those of Koszul type, which admit a global 1-form potential. For this class, we prove using Hodge theory that the first Betti number of a Hessian manifold of Koszul type is always positive (Proposition2.9).
A central achievement of our work is to reveal that Hessian metrics of Koszul-type are part of a much larger network of connections than previously understood. Whilst the work of Shima [15] established an important equivalence between the existence of a Koszul type Hessian metric and that of a hyperbolic affine structure, we establish a new, powerful list of equivalences, demonstrating that these conditions are also entirely equivalent to two other seemingly disparate properties: the existence of a positive flat line bundle (Corollary2.48) and admitting a space-like Lagrangian immersion (Corollary3.24) into its own cotangent bundle, equipped with the Riemann extension metric first studied by Patterson and Walker in 1952 [16], thereby building a new bridge between affine differential geometry and para-Kähler geometry.
To get a finer understanding of the moduli of hyperbolic affine structures on an oriented Hessian manifold of Koszul type, we define canonical class as the de-Rham cohomology class corresponding to the canonical line bundle. In 1968, J.-L.Koszul proved that the hyperbolicity is stable under small perturbations of the affine structure [17]. Using group cohomology with Lie algebra coefficients, we show that any hyperbolic affine structure can be deformed into another of which canonical class is rational (Theorem3.33).
This cohesive picture is further enriched by developing a duality with radiant affine manifolds, i.e. differentiable manifolds of which transition functions are linear, which builds upon the classical theory of dual connections in information geometry pioneered by Amari and Chentsov [2]. As a new contribution, we provide an explicit expression for the $1$-form potential of the dual of a radiant Hessian manifold of Koszul type, resembling the classical Legendre transformation in theoretical mechanics, and use this to situate the celebrated Cheng–Yau metric [18] within our framework.
Next, in a digression section, we also develop a new link to pseudoconvexity in complex analysis of several variables by showing that the Sasaki lifting of a complete Hessian manifold is a Stein manifold (Proposition3.3), and correspondingly, positive flat line bundles are lifted to positive holomorphic line bundles (Lemma3.2). A further observation then yields a stronger real analytic embedding theorem for this class of manifolds (Theorem3.4) of which dimension bound is more optimal than the classical Whitney embedding.
The synthesis of these tools now culminates in powerful structural theorems. By applying D.Tischler’s classical 1970 theorem on fibrations [19] to the first Koszul form, we prove that any curved compact orientable Hessian manifold must be a mapping torus (Theorem2.43), revealing a remarkably rigid underlying structure. This result is further strengthened by a de Rham-type splitting theorem, which shows that the universal cover of a compact Hessian manifold isometrically splits as a product of another contractible Riemannian manifold and the real line (Proposition3.41).
After obtaining a number of theorems, we recall that constructing explicit examples of non-trivial curved compact Hessian manifolds has been a significant challenge, with most known examples being flat Bieberbach manifolds. We therefore address this by explicitly constructing a non-trivial family of compact radiant Hessian manifolds of Koszul type that possess negative scalar curvature 1-1, suggesting that our theory is not vacuous.
Finally, as one of the highlights of this article, we adopt our general theory to low dimensions. Applying classical algebraic topology techniques and the solvability of Heisenberg group $H_{3}(\mathbb{R})$, we provide the first complete topological classification of all complete Hessian surfaces (Theorem4.5). We also derive severe constraints on the intersection form for Hessian 4-manifolds, and then prove that any closed orientable Hessian-Einstein manifold of dimension four must be flat (Proposition4.20). Last but not least, a novel characterization for potentials of $2$-dimensional flat metrics is established, showing that a pseudo-Riemannian metric
$g=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is flat if and only if the vector $(f_{xx},f_{xy},f_{yy})$ of second derivatives lies on a cone surface in the Euclidean space $\mathbb{R}^{3}$ (Proposition4.10).
Before closing this section, we provide a concise road map of the present article.
This paper is organized as follows:
Section1 reviews fundamental definitions and introduces a warped product construction as first examples. Section2 suggests key topological obstructions, concerning the vanishing of Euler characteristic and the infinitude of fundamental group. We then develop the theory of flat line bundles, proving that curved compact orientable Hessian manifolds must be mapping tori. Section3 establishes the generalised Legendre transform, and a central equivalence between Hessian metrics of Koszul type and several other geometric structures, including positive flat line bundles and space-like Lagrangian immersions. Curvature properties and link to complex analysis are also discussed. Finally, Section4 applies these results to low dimensions, yielding a complete topological classification of Hessian surfaces and deriving strong constraints for Hessian 4-manifolds. An appendix is included to support our classification of compact Hessian manifolds with Abelian fundamental groups.
1.2 Affine and Hessian Manifolds
For each $i\in\{1,\ldots,n\}$, denote by $\pi^{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ the projection onto the $i$-th coordinate.
Let $M$ be an $n$-dimensional differentiable manifold. As usual, for any chart $(U,\varphi)$ of $M$, the smooth functions $x^{1},\ldots,x^{n}\in C^{\infty}(U)$ defined by
$x^{i}:U\longrightarrow\mathbb{R},\quad p\longmapsto\pi^{i}(\varphi(p))$
are said to be the local coordinates associated to the chart $(U,\varphi)$. An atlas
$\mathscr{A}=\left\{\left(U_{\alpha},\varphi_{\alpha}\right)\mid\alpha\in I\right\}$
of $M$ is said to be an affine atlas of $M$, if for any $\alpha,\beta\in I$ there exist $A\in\operatorname{GL}_{n}(\mathbb{R})$ and $\bf{a}\in$ $\mathbb{R}^{n}$ such that
$\varphi_{\beta}\left(\varphi_{\alpha}^{-1}(\bf{x})\right)=A{\bf{x}}+{\bf{a}}$
for all ${\bf{x}}\in\varphi_{\alpha}\left(U_{\alpha}\cap U_{\beta}\right)$. An element of an affine atlas of $M$ is then called an affine chart of $M$, and the local coordinates associated to an affine chart of $M$ are termed affine coordinates.
For any two affine atlases $\mathscr{A}_{1}$ and $\mathscr{A}_{2}$ of $M$, we write $\mathscr{A}_{1}\sim\mathscr{A}_{2}$ if and only if $\mathscr{A}_{1}\cup\mathscr{A}_{2}$ is also an affine atlas of $M$. An affine atlas of $M$ is said to be maximal, if for any other affine atlas $\mathscr{A}^{\prime}$ of $M$ we have that $\mathscr{A}^{\prime}\sim\mathscr{A}$ implies $\mathscr{A}^{\prime}\subseteq\mathscr{A}$. Traditionally, a maximal affine atlas on $M$ is also termed an affine structure of the smooth manifold $M$.
Recall that a connection on a vector bundle over $M$ is said to be flat, if its curvature 2-form vanishes identically.
Consider an arbitrary connection $D$ on the tangent bundle $TM\rightarrow M$. By Cartan-Ambrose-Hicks theorem, if $D$ is flat and torsion-free then there exists a maximal affine atlas $\mathscr{A}_{D}:=\left\{\left(U_{\alpha},\varphi_{\alpha}\right)\mid\alpha\in I\right\}$ of $M$, such that
$D(dx_{\alpha}^{1})=\cdots=D(dx_{\alpha}^{n})=0$
for all $\alpha\in I$, where $x_{\alpha}^{1},\ldots,x_{\alpha}^{n}$ are the local coordinates associated to the chart $\left(U_{\alpha},\varphi_{\alpha}\right)$. Moreover, the map $D\longmapsto\mathscr{A}_{D}$ is a bijection from the set of all flat and torsion-free connections on the tangent bundle $TM\rightarrow M$ to the collection of all maximal affine atlases of $M$, and the atlas $\mathscr{A}_{D}$ is termed the adapted affine atlas of $D$. Therefore, fixing an affine structure on $M$ is equivalent to specifying a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. In view of this fact, by a slightly abuse of notation we also refer to a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$ as an affine structure of $M$, and refer to a pair $(M,\nabla)$ as an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$.
Remark 1.1.
All affine manifolds are in particular analytic, as their transition functions are by definition real analytic mappings.
On an affine manifold, there exist two important sheaves of Abelian groups:
Definition 1.2.
Let $(M,\nabla)$ be an affine manifold. The sheaf $\mathcal{A}:=\mathcal{A}_{M}$ of affine functions on $(M,\nabla)$ is defined by
$\mathcal{A}(U):=\{u\in C^{\infty}(U)\mid\nabla(du)=0\}$
where $U$ is any open subset of $M$. Sections of $\mathcal{A}$ are termed affine functions.
Definition 1.3.
Let $(M,\nabla)$ be an affine manifold. We denote by $\mathcal{P}:=\mathcal{P}_{M}$ the sheaf of parallel 1-forms on $(M,\nabla)$.
Recall from differential geometry that, on an $n$-dimensional affine manifold, the sheaf of parallel 1-forms is locally constant of rank $n$.
We shall also recall the following standard definition of Hessian manifolds:
Definition 1.4.
Let $(M,\nabla)$ be an affine manifold. A Riemannian metric $h$ on $M$ is called a Hessian metric on $(M,\nabla)$, if for any point $p\in M$, there exists an open neighborhood $U$ of $p$ in $M$, such that $h|_{U}=\nabla(df)$ for some $f\in C^{\infty}(U)$. A pair $((M,\nabla),h)$ is termed a Hessian manifold, if $h$ is a Hessian metric on the affine manifold $(M,\nabla)$.
As Hessian manifolds are affine, we first present a theorem on topological invariants of affine manifolds, which of course applicable to all Hessian manifolds.
Proposition 1.5.
All Pontryagin classes and rational Chern classes of an affine manifold must vanish.
Proof.
Note that the tangent bundle of an affine manifold admits a connection of which curvature vanishes identically. The proof of the desired result is then a direct application of the Chern–Weil construction. ∎
It is also worth noting the following famous conjecture of Chern.
Conjecture 1.6 (Chern,1955).
The Euler characteristic of a compact affine manifold is zero.
Chern’s conjecture is known to hold in various cases. For example, the 2-dimensional case is settled by John Milnor [20] in 1957. Moreover, in [21], Bruno Klingler proved that if a compact affine manifold admits a parallel nowhere vanishing top differential form then its Euler characteristic is zero. In latter sections we will also confirm this conjecture for all compact Hessian manifolds.
1.3 Warped Product Construction
In this section, we shall show that any warped product over the real line $\mathbb{R}$ with Hessian fibers is again Hessian, which allow us to construct new Hessian manifolds from known ones. In particular, we obtain that hyperbolic spaces and sufficiently small domains in round $n$-spheres are Hessian.
The following trivial observation serves as a base case in the recursive constructions of Hessian metrics.
Example 1.7.
Let $(M,g)$ be an one dimensional Riemannian manifold, and let $\nabla$ be an arbitrary connection on the tangent bundle $TM\rightarrow M$. Then, since $\Omega^{2}(M)=0$, the connection $\nabla$ is automatically flat and torsion-free. Moreover, by the theory of ordinary differential equations, we have that $g$ is always a Hessian metric on $(M,\nabla)$.
Now we state the fundamental construction theorem of warped products over a Hessian metric.
Theorem 1.8.
Let $\Omega$ be a domain in $\mathbb{R}^{n}$, and let $\varphi=\varphi(x_{1},\dots,x_{n})\in C^{\infty}(\Omega)$. Let $f=f(t)\in C^{\infty}(\mathbb{R})$ be a smooth function with nowhere vanishing first derivative, so that in particular
  1. 1.
    $f$ admits an inverse $F\colon f(\mathbb{R})\rightarrow\mathbb{R}$, i.e. the smooth function satisfying $f\circ F=id$;
  2. 2.
    $y_{1}:=\operatorname{exp}(f)x_{1},\,\dots\,,y_{n}:=\operatorname{exp}(f)x_{n},\,y_{n+1}\equiv y:=\operatorname{exp}(f)$ form a coordinate system of an open domain in the upper half space $\{(y_{1},\dots,y_{n},y)\in\mathbb{R}^{n+1}:y>0\}$.
Then, the smooth function
$\hat{\varphi}(y_{1},\dots,y_{n},y):=\varphi(y_{1}/y,\dots,y_{n}/y)y+\int_{0}^{y}F^{\prime}(\operatorname{log}(t))^{2}(y-t)t^{-2}dt$
is well-defined, and it holds that
$\sum_{i=1}^{n+1}\sum_{j=1}^{n+1}\frac{\partial^{2}\hat{\varphi}}{\partial y_{i}\partial y_{j}}dy_{i}\otimes dy_{j}=dt^{2}+\operatorname{exp}(f)\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}dx_{i}\otimes dx_{j}.$
Proof.
The proof of this theorem is a straightforward computation and hence omitted. ∎
Combining Example1.7 and Theorem1.8, one may conclude that many important metric tensors in geometry and theoretical physics are actually Hessian. Examples include but not limited to sufficiently small geodesic disks in space forms, and the Robertson–Walker solutions to Einstein’s field equation.
The following is just one particular example of many corollaries of the powerful theorem above.
Example 1.9.
On the upper half space $\mathbb{H}^{n}:=\{(x_{1},\dots,x_{n})\in\mathbb{R}^{n}:x_{n}>0\}$, consider the coordinates
$y_{1}:=\frac{2x_{1}}{x_{n}^{2}},\,\dots,\,y_{n-1}:=\frac{2x_{n-1}}{x_{n}^{2}},\,y_{n}:=\frac{1}{x_{n}^{2}}.$
Let $\nabla$ be the affine structure on $\mathbb{H}^{n}$ satisfying $\nabla(dy_{i})=0$ for all $i=1,\dots,n$. Define smooth function
$\varphi(y_{1},\dots,y_{n}):=\frac{y_{1}^{2}+\dots+y_{n-1}^{2}}{8y_{n}}-\frac{1}{4}\operatorname{log}(y_{n})$
on $\mathbb{H}^{n}$. Then the Hessian quadratic form of $f$ is precisely
$h:=\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}dy_{i}\otimes dy_{j}=\frac{dx_{1}^{2}+\dots+dx_{n}^{2}}{x_{n}^{2}}.$
In particular, $((\mathbb{H}^{n},\nabla),h)$ is a Hessian manifold.
2 Global Invariants for Hessian manifolds
2.1 Obstructions from Sheaf Cohomology
In this section, we study the very natural question that in which case there exists a global potential $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. Before we can give an answer to this question, we need to state and prove the following two lemmata from algebraic topology and sheaf theory.
Lemma 2.1.
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. If $M$ is simply connected, then the first sheaf cohomology group $\mathrm{H}^{1}(M;\mathcal{A})$ vanishes.
Proof.
Since the holonomy of $(M,\nabla)$ is trivial and the sheaf $\mathcal{P}$ of parallel 1-forms on $M$ is a locally constant sheaf of rank $n$, by abuse of notation we identify $\mathcal{P}$ with $\underline{\mathbb{R}}^{n}$. It is then readily seen that there exists a short exact sequence
$0\longrightarrow\mathbb{R}\longrightarrow\mathcal{A}\overset{d}{\longrightarrow}\mathbb{R}^{\oplus n}=\mathcal{P}\longrightarrow 0$
of sheaves, which induces a long exact sequence
$\cdots\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})\longrightarrow\mathrm{H}^{1}(M;\mathcal{A})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})^{\oplus n}\longrightarrow\cdots$
of cohomology groups. But since $M$ is simply connected, we have that the first de-Rham cohomology group of $M$ vanishes. Therefore, we obtain $\mathrm{H}^{1}(M;\mathcal{A})=0$. ∎
Lemma 2.2.
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. Suppose that $\pi\colon\tilde{M}\rightarrow M$ is a universal covering. Then $\pi^{-1}\mathcal{A}$ is the sheaf of affine functions on the affine manifold $(\tilde{M},\pi^{*}\nabla)$, and there exists a natural isomorphism
$\mathrm{H}^{1}(M;\mathcal{A})\cong\mathrm{H}^{1}(G;\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))$
of $\mathbb{R}$-vector spaces, where $G:=\pi_{1}(M)$ is the fundamental group of $M$.
Proof.
It is routine to check that $(\tilde{M},\pi^{*}\nabla)$ is an affine manifold and $\pi^{-1}\mathcal{A}$ is its sheaf of affine functions.
To prove the existence of the natural isomorphism, we use Cartan-Leray spectral sequence, of which $E_{2}$-page terms are
$E_{2}^{p,q}=\mathrm{H}^{p}(G,\mathrm{H}^{q}(\tilde{M};\pi^{-1}\mathcal{A}))\implies\mathrm{H}^{p+q}(M;\mathcal{A}).$
Recall from homological algebra that the following sequence
$0\longrightarrow E^{1,0}_{2}\longrightarrow\mathrm{H}^{1}(M;\mathcal{A})\longrightarrow E^{0,1}_{2}\longrightarrow E^{2,0}_{2}\longrightarrow\cdots$
is exact. Since $\tilde{M}$ is simply connected, by Lemma2.1 $\mathrm{H}^{1}(\tilde{M};\pi^{-1}\mathcal{A})=0$. Therefore we have that
$E^{0,1}_{2}=\mathrm{H}^{0}(G;0)=0.$
This proves that
$\mathrm{H}^{1}(M;\mathcal{A})=E^{1,0}_{2}=\mathrm{H}^{1}(G,\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))$
as required. ∎
The following vanishing theorem of sheaf cohomology group turns out to be useful for our purpose:
Proposition 2.3.
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. If the fundamental group of $M$ is finite, then the first sheaf cohomology group $\mathrm{H}^{1}(M;\mathcal{A})$ vanishes.
Proof.
Take any universal covering $\pi\colon\tilde{M}\rightarrow M$. Then by Lemma2.2
$\mathrm{H}^{1}(M;\mathcal{A})\cong\mathrm{H}^{1}(G;\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))$
as $\mathbb{R}$-vector spaces, where $G:=\pi_{1}(M)$ is the fundamental group of $M$. But, recall from the theory of group cohomology that, the first cohomology of a finite group is a torsion group. We therefore conclude that $\mathrm{H}^{1}(G;\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))=0$. ∎
We are now ready to state and prove the main theorem in this section.
Theorem 2.4.
Let $((M,\nabla),h)$ be a Hessian manifold. If the fundamental group of $M$ is finite, then there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$.
Proof.
By the definition of a Hessian metric, there exists an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a smooth function $f_{\alpha}\in C^{\infty}(U_{\alpha})$, such that for each $\alpha\in I$ it holds that $h_{\alpha}=\nabla(df_{\alpha})$, where $h_{\alpha}$ is the restriction of $h$ on $U_{\alpha}$.
For every $\alpha,\beta\in I$, define $g_{\alpha\beta}:=f_{\alpha}-f_{\beta}\in C^{\infty}(U_{\alpha}\cap U_{\beta})$. Then $g=\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ is a Čech 1-cocycle with coefficients in the sheaf $\mathcal{A}$ of affine functions on $M$. It remains to prove that the Čech cohomology class $g]\in\check{\mathrm{H}}^{1}(M;\mathcal{A})$ vanishes. But this immediately follows from Proposition2.3. ∎
Example 2.5.
There does exist Hessian manifold with finite but non-trivial fundamental group. The simplest example is the punctured affine space $\mathbb{R}^{3}-\{0\}$ quotient by the antipodal map $\textup{{x}}\mapsto-\textup{{x}}$, of which fundamental group is precisely $\mathbb{Z}/2\mathbb{Z}$.
Before we proceed to the next section, also notice that we have the following obvious example for the non-existence of global potential for a pseudo-Riemannian metric.
Proposition 2.6.
Let $(M,g)$ be a compact pseudo-Riemannian manifold, and let $\nabla$ be a connection on the cotangent bundle $T^{*}M\rightarrow M$. Then $g\neq\nabla(df)$ for all $f\in C^{\infty}(M)$.
Proof.
Take any $f\in C^{\infty}(M)$ such that the Hessian quadratic form $\nabla(df)$ of $f$ is everywhere non-degenerate. In particular $f$ is not a constant function. Since $M$ is compact, by the extreme value theorem, there exists distinct points $p,q\in M$ such that $f$ attains a global maximum at $p$ and attains a global minimum at $q$. But then by the second derivative test, the signatures of the Hessian quadratic form of $f$ must be different at points $p$ and $q$. ∎
Proposition2.6 leads immediately to an interesting result:
Corollary 2.7.
The fundamental group of a compact Hessian manifold is infinite.
Proof.
Let $h$ be a Hessian metric on affine manifold $(M,\nabla)$. Assume the contrary that $\pi_{1}(M)$ is a finite group. Then by Proposition2.4, there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. But this contradicts Proposition2.6 as $M$ is assumed to be compact. ∎
Remark 2.8.
In [22], it is proved that there exists a unique closed orientable flat Riemannian 3-manifold with vanishing first Betti number, called the Hantzsche–Wendt manifold. Therefore, in particular, although by Corollary2.7 a compact Hessian manifold admits an infinite fundamental group, its first Betti number can still be zero.
However, in contrast, we have the following proposition:
Proposition 2.9.
Let $(M,\nabla)$ be an affine manifold, and let $h$ be a Riemannian metric on $M$. Suppose that there exists $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$. Then $\eta$ is a closed 1-form. Moreover, if it is assumed further that $M$ is compact, then $\eta$ is not exact, and in particular the first Betti number of $M$ is positive.
Proof.
Since $\nabla$ is torsion-free, the exterior differential $d\eta$ is the anti-symmetrization of $\nabla\eta$. But the anti-symmetrization of $\nabla\eta$ is identically zero as $\nabla\eta=h$ is symmetric. Therefore, $\eta$ is a closed 1-form.
Now, assume that $M$ is compact. By Hodge decomposition theorem, there exists a smooth function $f\in C^{\infty}(M)$ and a harmonic 1-form $\theta$ such that $\eta=df+\theta$. Since
$\nabla\theta=\nabla\eta-\nabla(df)=h-\nabla(df)\neq 0$
by Proposition2.6, we conclude that $\theta\neq 0$. Now it suffices to recall from Hodge theory that $\mathrm{H}^{1}(M;\mathbb{R)}$ is isomorphic to the space of the harmonic 1-forms on $M$. ∎
Proposition2.9 motivates the following definition.
Definition 2.10.
Let $(M,\nabla)$ be an affine manifold, and let $h$ be a Riemannian metric on $M$. The pair $((M,\nabla),h)$ is termed a Hessian manifold of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$, and in this case, the 1-form $\eta$ is termed a $1$-form potential for $((M,\nabla),h)$.
By Poincaré’s lemma, Hessian manifolds of Koszul type are indeed Hessian. As explained in the introduction section, this is, actually, the ultimate origin of the notion of Hessian structures.
2.2 Fundamental Group and Euler Number
For each $n\in\mathbb{N}$, we denote by $\operatorname{Aff}_{n}(\mathbb{R})$ the group of affine transformations on $\mathbb{R}^{n}$.
Definition 2.11.
A Hessian manifold $((M,\nabla),h)$ is said to be complete, if $h$ is a complete Riemannian metric on $M$.
First, we recall the definition of affine mappings: For affine manifolds $(M_{1},\nabla_{1})$ and $(M_{2},\nabla_{2})$, a smooth mapping $\Phi:M_{1}\rightarrow M_{2}$ is said to be affine if $\hat{\nabla}(d\Phi)=0$, where
$\hat{\nabla}=\nabla_{1}\otimes\operatorname{Id}+\operatorname{Id}\otimes\Phi^{*}\nabla_{2}$
is the induced connection on the tensor bundle $T^{*}M_{1}\otimes\Phi^{*}TM_{2}$.
Let $(M,\nabla)$ be an affine manifold. It is readily seen that, a vector-valued smooth function
$F\colon M\longrightarrow\mathbb{R}^{m},\quad p\longmapsto(F_{1}(p),\dots,F_{m}(p))$
is an affine mapping between $(M,\nabla)$ and the flat Euclidean space, if and only if $F_{1},\dots,F_{m}$ are global affine functions on $M$.
For an affine manifold $(M,\nabla)$, since $\nabla$ is torsion-free, the exterior differential $d\omega$ of a $1$-form $\omega$ is the anti-symmetrization of the covariant derivative $\nabla\omega$. Therefore, on an affine manifold, all parallel 1-forms are closed.
Moreover, for an affine manifold $(M,\nabla)$ with finite fundamental group, parallel 1-forms admit no periods as the first de-Rham cohomology group of $M$ vanishes. Therefore, the following notion of affine development is well-defined:
Definition 2.12.
Let $(M,\nabla)$ be a simply connected $n$-dimensional affine manifold, and fix a reference point $p_{0}\in M$. The affine development of $(M,\nabla)$ is the smooth mapping $\Phi\colon M\rightarrow\mathcal{P}(M)^{\vee}=\mathbb{R}^{n}$ defined by
$\langle\omega,\Phi(p)\rangle=\int_{\gamma}\omega$
for $\omega\in\mathcal{P}(M)$ and $p\in M$, where the bracket denotes the canonical pairing between $\mathcal{P}(M)$ and its dual space $\mathcal{P}(M)^{\vee}$, and where $\gamma\colon[0,1]\rightarrow M$ is a smooth curve such that $\gamma(0)=p_{0}$ and $\gamma(1)=p$.
See for example [15] for a proof of the fact that an affine development is indeed an affine mapping from an affine manifold to the flat Euclidean space.
In [12], using the concept of affine development, H.Shima and K.Yagi proved the following important theorem on convexity of complete Hessian metrics.
Theorem 2.13 (Shima-Yagi, 1997[12]).
Let $((M,\nabla),h)$ be a simply connected complete Hessian manifold. Then the affine development $\Phi\colon M\rightarrow\mathbb{R}^{n}$ of $(M,\nabla)$ is a diffeomorphism onto its image such that $\Phi(M)$ is a convex domain in $\mathbb{R}^{n}$.
Theorem2.13 provides a strong restriction on the topology of complete Hessian manifolds.
Corollary 2.14.
Let $((M,\nabla),h)$ be a complete Hessian manifold. Then $M$ is aspherical, i.e. the homotopy group $\pi_{k}(M)$ vanishes for all $k\geq 2$.
Proof.
Recall that being an aspherical space is equivalent to possessing a contractible universal cover. The desired result then follows immediately from Theorem2.13. ∎
In [14] and [13], Kostant et al. and Fried verified Conjecture1.6 in some special cases.
Theorem 2.15 (Kostant-Sullivan,1975[13]).
Let $(M,\nabla)$ be a compact affine manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. If $\nabla$ is geodesically complete, or equivalently the affine development $\Phi\colon\tilde{M}\rightarrow\mathbb{R}^{n}$ of $(\tilde{M},\pi^{*}\nabla)$ is a diffeomorphism, then the Euler characteristic $\chi(M)=0$.
Theorem 2.16 (Fried,1982[14]).
[…]
Combine Theorem2.13 and Theorem2.15&2.16, we immediately get the following result, confirming Conjecture1.6 for Hessian manifolds.
Proposition 2.17.
The Euler characteristic of a compact Hessian manifold is zero.
Proof.
[…]
Corollary2.7 shows that compact Hessian manifolds has infinite fundamental group. We shall now improve the result in Corollary2.7.
Lemma 2.18.
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$, and let $G$ be a group acting freely on $\Omega$. Then $G$ is torsion-free, that is, it contains no non-trivial element of finite order.
Proof.
Suppose that $g\in G$ enjoys finite order $k\in\mathbb{N}$. Take any arbitrary point $\textbf{x}\in\Omega$. Since $\Omega$ is convex and $g^{1}(\textbf{x}),\dots,g^{k}(\textbf{x})\in\Omega$, the barycenter
$\textbf{b}:=\frac{1}{k}\sum_{i=0}^{k-1}g^{i}(\textbf{x})$
is also a point in $\Omega$. But by the very construction we have that $g(\textbf{b})=\textbf{b}$. Therefore $g$ is the identity element of $G$, as the $G$-action is by assumption free. ∎
The following results are direct consequences of Lemma2.18:
Proposition 2.19.
The fundamental group a complete Hessian manifold is torsion-free.
Proof.
By Theorem2.13, the fundamental group of a complete Hessian manifold acts freely on a convex domain. The desired result then follows from Lemma2.18. ∎
Corollary 2.20.
The fundamental group a complete Hessian manifold is either trivial or infinite.
Proof.
This follows from Proposition2.19. ∎
Remark 2.21.
Let $((M,\nabla),h)$ be a compact Hessian manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. Then $((\tilde{M},\pi^{*}\nabla),\pi^{*}h)$ is a complete Hessian manifold. Provided that the fundamental group of $M$ is finite, then $\tilde{M}$ is compact, which contradicts Theorem2.13. This gives another proof of Corollary2.7.
Before we close this section, we note that Proposition2.19 also straightforwardly implies the following interesting theorem:
Theorem 2.22.
Let $((M,\nabla),h)$ be a compact Hessian manifold of dimension $n$. If the fundamental group of $M$ is Abelian, then $M$ is homeomorphic to the flat $n$-torus $\mathbb{T}^{n}$.
Proof.
By Proposition2.19, the fundamental group $\pi_{1}(M)$ of $M$ is isomorphic to $\mathbb{Z}^{k}$ for some $k\in\mathbb{N}$. By Corollary2.14 and the uniqueness of Eilenberg-MacLane space, $M$ is homotopy equivalent to the $k$-torus $\mathbb{T}^{k}$. Comparing the top homology groups of $M$ and $\mathbb{T}^{k}$ yields that $k=n$. The desired result then follows from TheoremA.3 in the appendix. ∎
2.3 Flat Line Bundles
2.3.1 The Picard group
In this section, we study Hessian manifolds using the concept of flat line bundles. We also define the Picard group, and use it to prove that a curved compact orientable Hessian manifold has positive first Betti number.
We shall now recall the notion of a flat $\mathbb{R}$-line bundle:
Definition 2.23.
For a $\mathbb{R}$-vector bundle $\mathscr{E}\rightarrow M$ over a differentiable manifold $M$, a pair $(\mathscr{E},D)$ is termed a flat $\mathbb{R}$-vector bundle over $M$ if $D$ is a flat connection on $\mathscr{E}\rightarrow M$. A flat $\mathbb{R}$-vector bundle of rank 1 is then called a flat $\mathbb{R}$-line bundle. In particular, we call $(M\times\mathbb{R},d)$ the trivial flat $\mathbb{R}$-line bundle over $M$.
Let $\mathscr{L}\rightarrow M$ be a $\mathbb{R}$-line bundle over a differentiable manifold $M$. By the local triviality of $\mathscr{L}\rightarrow M$, there exists an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a local section $s_{\alpha}$ of $\mathscr{L}\rightarrow M$ defined on $U_{\alpha}$, such that $s_{\alpha}$ is non-vanishing for all $\alpha\in I$. The collection $\{s_{\alpha}\}_{\alpha\in I}$ is traditionally termed a local frame field of $\mathscr{L}\rightarrow M$.
Lemma 2.24.
Let $M$ be a differentiable manifold and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Then $\mathscr{L}\rightarrow M$ admits a local frame field $\{s_{\alpha}\}_{\alpha\in I}$ such that $Ds_{\alpha}=0$ for all $\alpha\in I$. Moreover, for each $\alpha,\beta\in I$ there exists a constant $g_{\alpha\beta}\in\mathbb{R}^{\times}$ such that
$s_{\beta}(p)=s_{\alpha}(p)\cdot g_{\alpha\beta}$
for all $p\in U_{\alpha}\cap U_{\beta}$, where $U_{\alpha}$ here is the domain of definition of $s_{\alpha}$ and $U_{\beta}$ is the domain of definition of $s_{\beta}$.
Proof.
Since the connection $D$ is flat, its connection 1-forms are locally exact. Without lose of generality, we may assume that there exists $f\in C^{\infty}(U_{\alpha})$ such that the restriction of $D$ on $U_{\alpha}$ is $d+df$.
Now, it suffices to show that, up to a real constant multiple, the first order partial differential equation $ds+sdf=0$ for $s\in C^{\infty}(U_{\alpha})$ has a unique positive solution. But by the method of integration factor, we have
$e^{f}(ds+sdf)=d(e^{f}s)$
and hence the general solution of differential equation $ds+sdf=0$ is precisely $s=Ce^{-f}$, where $C\in\mathbb{R}$ is an integration constant. This proves the desired result. ∎
Definition 2.25.
Let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over a differentiable manifold $M$. A local frame field $\{s_{\alpha}\}_{\alpha\in I}$ of $\mathscr{L}\rightarrow M$ is termed an affine local frame field of $(\mathscr{L},D)$, if $Ds_{\alpha}=0$ for all $\alpha\in I$.
Now we introduce the isomorphism classes in the category of flat $\mathbb{R}$-line bundle over $M$.
Definition 2.26.
Two flat $\mathbb{R}$-line bundles $(\mathscr{L}_{1},D_{1})$ and $(\mathscr{L}_{2},D_{2})$ over $M$ are said to be gauge equivalent, if there exists a $\mathbb{R}$-vector bundle isomorphism $\phi\colon\mathscr{L}_{1}\rightarrow\mathscr{L}_{2}$ such that $D_{1}-\phi^{*}D_{2}\in\Omega^{1}(M)$ is an exact 1-form on $M$.
Theorem 2.27.
Let $M$ be a differentiable manifold. Up to gauge equivalence, flat $\mathbb{R}$-line bundles over $M$ are classified by $\operatorname{Hom}(\pi_{1}(M),\mathbb{R}^{\times})$, or equivalently, by the additive group $\mathrm{H}^{1}(M;\mathbb{R})\oplus\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$ if $M$ is aspherical.
Proof.
Notice that by the universal coefficients and Hurewicz’s theorem,
$\mathrm{H}^{1}(M;\mathbb{R})\oplus\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\cong\operatorname{Hom}(\pi_{1}(M),\mathbb{R})\oplus\operatorname{Hom}(\pi_{1}(M),\mathbb{Z}/2\mathbb{Z})=\operatorname{Hom}(\pi_{1}(M),\mathbb{R}^{\times})$
where $\pi_{1}(M)$ is the fundamental group of $M$.
Let $\rho\colon\pi_{1}(M)\rightarrow\mathbb{R}^{\times}$ be a representation. Consider the associated bundle $\mathscr{L}:=\tilde{M}\times_{\rho}\mathbb{R}$ where $\pi\colon\tilde{M}\rightarrow M$ is a universal cover, and each element $g$ of $\pi_{1}(M)$ acts diagonally by
$g\cdot(p,x):=(g(p),\rho(g)x)$
for every $p\in\tilde{M}$ and $x\in\mathbb{R}$. By construction, the connection $d$ on the trivial bundle is $\pi_{1}(M)$-equivariant, and hence descends to a flat connection on $\mathscr{L}$.
Conversely, for any flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$, since the curvature of $D$ vanishes identically, the holonomy representation sending a homotopy class represented by a smooth loop $\gamma$ in $M$ to the parallel transport along $\gamma$ is a well-defined group representation $\rho\colon\pi_{1}(M)\rightarrow\operatorname{GL}_{1}(\mathbb{R})=\mathbb{R}^{\times}$.
Moreover, if $D$ be a flat connection on $\mathscr{L}$ and $f\in C^{\infty}(M)$, then the curvature 2-form of $D+df$ is
$ddf+df\wedge df=0$
and by the gradient theorem the holonomy of $D$ coincides with that of $D+df$. ∎
Corollary 2.28.
Let $M$ be a differentiable manifold. The collection of gauge equivalence classes of flat $\mathbb{R}$-line bundle over $M$ equipped with the tensor product $\otimes$ as binary operation forms a multiplicative group.
Proof.
This is essentially a reformulation of Theorem2.27. ∎
Definition 2.29.
Let $M$ be a differentiable manifold. The group of gauge equivalence classes of flat $\mathbb{R}$-line bundles over $M$ is termed the Picard group of $M$.
Recall that a line bundle is trivial if and only if it admits a nowhere vanishing global section. The following proposition gives a refinement in the case of flat line bundles.
Proposition 2.30.
Let $M$ be an differentiable manifold, and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Then $(\mathscr{L},D)$ is gauge equivalent to the trivial flat $\mathbb{R}$-line bundle over $M$ if and only if it admits a nowhere vanishing parallel global section.
Proof.
Assume that $(\mathscr{L},D)$ is gauge equivalent to the trivial flat $\mathbb{R}$-line bundle over $M$. Then we identify $\mathscr{L}\equiv M\times\mathbb{R}$, and there exists an exact form $\theta\in\Omega^{1}(M)$ such that $D=d+\theta$. Take $f\in C^{\infty}(M)$ such that $\theta=df$. Then the global section
$s=\operatorname{exp}(-f)\in\mathrm{H}^{0}(M;M\times\mathbb{R})=C^{\infty}(M)$
is nowhere vanishing and satisfies that
$Ds=d\operatorname{exp}(-f)+e^{-f}\theta=e^{-f}(\theta-df)=0.$
To prove the converse statement, we assume that $(\mathscr{L},D)$ admits a nowhere vanishing parallel global section $s$. Then in particular we have $\mathscr{L}\cong M\times\mathbb{R}$. Moreover, since
$D(fs)=df\otimes s+f\cdot Ds=df\otimes s$
for all $f\in C^{\infty}(M)$, we conclude that $D=d$. ∎
2.3.2 The canonical line bundle
Amongst all flat $\mathbb{R}$-line bundle over an affine manifold, the following one, which will be called the canonical line bundle, is of great importance.
Definition 2.31.
Let $(M,\nabla),$ be an affine manifold, and denote also by $\nabla$ the connection on tensor bundles of $M$ induced by $\nabla$. The flat $\mathbb{R}$-line bundle
$K_{\nabla}=(\bigwedge^{n}(T^{*}M),\nabla)$
over $M$ is termed the canonical line bundle of $(M,\nabla)$.
Note that, although the top exterior power of the cotangent bundle of an orientable manifold is always trivial, the canonical line bundle of an orientable affine manifold in general is non-trivial as a flat $\mathbb{R}$-line bundle.
Recall that a Calabi-Yau manifold is a Kähler manifold with trivial canonical line bundle, or equivalently, admitting a holomorphic volume form. The following corollary exhibits an close analogy.
Corollary 2.32.
Let $(M,\nabla)$ be an affine manifold of dimension $n$. Then the canonical line bundle $K_{\nabla}$ of $(M,\nabla)$ is trivial if and only if $M$ admits a nowhere vanishing $n$-form $\Omega$ such that $\nabla\Omega=0$.
Proof.
This directly follows from Proposition2.30. ∎
We now introduce the notion of Koszul forms of a Hessian line bundle.
Definition 2.33.
Let $M$ be a differentiable manifold. A Hessian line bundle over $M$ is a pair $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, where $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$ and $\langle\cdot,\cdot\rangle$ is a bundle metric of $\mathscr{L}\rightarrow M$.
Definition 2.34.
Let $M$ be a differentiable manifold, and let $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ be a Hessian line bundle over $M$. The first Koszul form $\varkappa\in\Omega^{1}(M)$ of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is defined by
$\varkappa|_{U_{\alpha}}:=-\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
where $\{s_{\alpha}\}_{\alpha\in I}$ is an affine local frame of $(\mathscr{L},D)$ and $U_{\alpha}$ is the domain of definition of $s_{\alpha}$ for each $\alpha\in I$.
By the properties of affine local frame, the first Koszul form of a Hessian line bundle is well-defined, and is closed as it is by definition locally exact.
Definition 2.35.
Let $((M,\nabla),h)$ be a Hessian manifold and denote by $\langle\cdot,\cdot\rangle_{h}$ the bundle metric on tensor bundles of $M$ induced by $h$. The first Koszul form of the Hessian line bundle $(K_{\nabla},\langle\cdot,\cdot\rangle_{h})$ is also said to be the first Koszul form of $((M,\nabla),h)$.
For the convenience of the reader, we review the standard notion of first Koszul forms of a Hessian manifold.
For an oerientable Riemannian manifold $(M,h)$, we denote by $\operatorname{vol}_{h}$ its volume form.
Proposition 2.36.
Let $((M,\nabla),h)$ be an oriented Hessian manifold, and let $\varkappa$ be its first Koszul form. Then it holds that $2\varkappa=d\operatorname{log}(|h|)$, and
$\nabla_{X}\operatorname{vol}_{h}=\varkappa(X)\operatorname{vol}_{h}$
for every vector field $X$ on $M$, where $|h|$ is the determinant of the matrix of which $(i,j)$-entry is $h(\partial/\partial x^{i},\partial/\partial x^{j})$ for some oriented affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$.
Proof.
The proof is straightforward and hence omitted. ∎
In [12], H.Shima and K.Yagi also established the following criterion for flatness of Hessian manifolds in terms of the first Koszul form.
Proposition 2.37 (Shima-Yagi, 1997[12]).
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold. Let $\bar{\nabla}$ be the Levi-Civita connection of $(M,h)$, and let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((M,\nabla),h)$. Then it holds that $\bar{\nabla}\varkappa=0$.
Theorem 2.38 (Shima-Yagi, 1997[12]).
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold of dimension $n$. Then the following statements are equivalent:
  1. 1.
    The first Koszul form of $((M,\nabla),h)$ vanishes identically;
  2. 2.
    $\nabla$ is the Levi-Civita connection of $(M,h)$;
  3. 3.
    $M$ admits a nowhere vanishing $n$-form $\Omega$ such that $\nabla\Omega=0$.
Recall that, in Remark2.8, we found an example of an orientable compact Hessian manifold $((M,\nabla),h)$ with vanishing first Betti number, but in that case $(M,h)$ is in fact a flat Riemannian manifold. This is not just a coincidence.
Lemma 2.39.
Let $M$ be a differentiable manifold, and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Suppose that $\mathscr{L}$ is a trivial $\mathbb{R}$-line bundle, and there exists a positive integer $k$ such that $(\mathscr{L}^{\otimes k},D)$ is a trivial flat $\mathbb{R}$-line bundle. Then $(\mathscr{L},D)$ itself is a trivial flat $\mathbb{R}$-line bundle.
Proof.
Let $s\in\mathrm{H}^{0}(M;\mathscr{L})$ be a nowhere-vanishing smooth section. Then by assumption and Proposition2.30, there exists $f\in C^{\infty}(M)$ such that $\operatorname{exp}(kf)s^{\otimes k}$ is parallel. Then by Leibniz rule of differentiation,
$k(\operatorname{exp}(f)s)^{\otimes k-1}\otimes D(\operatorname{exp}(f)s)=D(\operatorname{exp}(kf)s^{\otimes k})=0$
and hence $D(\operatorname{exp}(f)s)=0$, as $k(\operatorname{exp}(f)s)^{\otimes k-1}$ is nowhere vanishing. Therefore $\operatorname{exp}(f)s$ is a nowhere vanishing parallel section of $(\mathscr{L},D)$. Again by Proposition2.30, we conclude that $(\mathscr{L},D)$ is trivial as a flat $\mathbb{R}$-line bundle. ∎
Theorem 2.40.
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold. If the first Betti number of $M$ is zero, then $(M,h)$ is a flat Riemannian manifold and $\nabla$ is its Levi-Civita connection.
Proof.
Since $\mathrm{H}^{1}(M;\mathbb{R})=0$, by Theorem2.27 we have that the Picard group of $M$ is isomorphic to $\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$, and hence is 2-torsion. In particular, the canonical line bundle $K$ of $(M,\nabla)$ satisfies $K^{\otimes 2}=0$. Since $M$ is orientable, the $\mathbb{R}$-line bundle $\bigwedge^{n}(T^{*}M)$ is trivial. Therefore, by virtue of Lemma2.39, the canonical line bundle of $M$ is trivial as a flat $\mathbb{R}$-line bundle. Corollary2.32 then implies that $M$ admits a volume form $\Omega$ such that $\nabla\Omega=0$. The desired result follows immediately from Theorem2.38. ∎
As an application of the first Koszul forms, we suggest several important splitting theorems of Hessian manifolds, which all relies on the following fundamental result of Tischler on fibration over circles.
Theorem 2.41 (Tischler, 1970[19]).
Let $M$ be a compact differentiable manifold. If $M$ admits a nowhere vanishing closed 1-form, then $M$ is a mapping torus.
Theorem 2.42.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then $M$ is a mapping torus.
Proof.
Let $\eta\in\Omega^{1}(M)$ be a $1$-form potential of $((M,\nabla),h)$. View $\eta\colon M\rightarrow T^{*}M$ as an immersion. Then, $\eta$ vanishes at $p\in M$ if and only if $p$ is an intersection point of $\eta\colon M\rightarrow T^{*}M$ and the zero section $\sigma$ of the cotangent bundle of $M$. Since $\nabla\eta=h>0$, each intersection point of $\eta$ and $\sigma$ is a nondegenerate transverse intersection, and the intersection numbers at those intersection points are all positive. But, on the other hand, the sum of the intersection numbers at the intersection points of $\eta$ and $\sigma$ is precisely
$\langle e(T^{*}M),[M]\rangle=\chi(M)=0$
by Proposition2.17. This proves that the sections $\eta$ and $\sigma$ must have no intersection, i.e. $\eta$ is a nowhere vanishing closed 1-form on $M$. The desired result then follows from Theorem2.41. ∎
An stronger result can be obtained if the manifold under consideration is orientable.
Theorem 2.43.
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold. Then either $(M,h)$ is a flat Riemannian manifold, or $M$ is a mapping torus.
Proof.
Let $\varkappa$ be the first Koszul form of $((M,\nabla),h)$. If $\varkappa=0$, then $(M,h)$ is a flat Riemannian manifold by Theorem2.38. We therefore assume $\varkappa\neq 0$. But then since $\bar{\nabla}\varkappa=0$ by Proposition2.37, we have that $\varkappa$ is nowhere vanishing by parallel transport. Therefore by Theorem2.41 we conclude that $M$ is a mapping torus. ∎
2.3.3 The obstruction line bundle
Resembling the notion of positive holomorphic line bundles in complex geometry, we now define and study the positivity of a Hessian line bundle.
Definition 2.44.
Let $(M,\nabla)$ be an affine manifold. A flat line bundle $(\mathscr{L},D)$ over $M$ is said to be a positive flat line bundle over $(M,\nabla)$ if there exists a bundle metric $\langle\cdot,\cdot\rangle$ of $\mathscr{L}\rightarrow M$ such that $\nabla\varkappa>0$, where $\varkappa$ is the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$.
By the very definition, if an affine manifold $(M,\nabla)$ admits a positive flat line bundle $(\mathscr{L},D)$, then $((M,\nabla),\nabla\varkappa)$ is a Hessian manifold of Koszul type, where $\varkappa$ is the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ and $\langle\cdot,\cdot\rangle$ is a bundle metric of $\mathscr{L}\rightarrow M$ such that $\nabla\varkappa>0$.
Besides the canonical line bundle, another important flat $\mathbb{R}$-line bundle over a Hessian manifold is its obstruction line bundle, which we shall now define.
Definition 2.45.
Let $((M,\nabla),h)$ be a Hessian manifold. Consider an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a smooth function $f_{\alpha}\in C^{\infty}(U_{\alpha})$, such that for each $\alpha\in I$ it holds that $h_{\alpha}=\nabla(df_{\alpha})$, where $h_{\alpha}$ is the restriction of $h$ on $U_{\alpha}$. The flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$, where $\mathscr{L}\rightarrow M$ is the $\mathbb{R}$-line bundle over $M$ defined by the 1-cocycle $\{\operatorname{exp}(f_{\alpha}-f_{\beta})\}_{\alpha,\beta\in I}$ and where $D$ is the flat connection on $\mathscr{L}\rightarrow M$ defined by the connection 1-forms $\{df_{\alpha}\}_{\alpha\in I}$, is termed the obstruction line bundle of $((M,\nabla),h)$.
Attentive readers are invited to check the fact that the obstruction line bundle of a Hessian manifold is uniquely defined up to gauge equivalence.
Also note, for an obstruction line bundle $(\mathscr{L},D)$, since the structure group of $\mathscr{L}$ reduces to $\operatorname{SO}(1)$, we have that $\mathscr{L}$ is always a trivial $\mathbb{R}$-line bundle.
Now, we can reformulate the conditions on the existence of global potential for a Hessian manifold in terms of its obstruction line bundle.
Proposition 2.46.
Let $((M,\nabla),h)$ be a Hessian manifold of Koszul type, and $(\mathscr{L},D)$ its obstruction line bundle. Then there exists a bundle metric $\langle\cdot,\cdot\rangle$ on $\mathscr{L}\rightarrow M$ such that the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is a $1$-form potential of $((M,\nabla),h)$. Moreover, if it is assumed further that there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$, then $(\mathscr{L},D)$ is a trivial flat $\mathbb{R}$-line bundle.
Proof.
Let $\eta$ be a $1$-form potential of $((M,\nabla),h)$. Since $\eta$ is closed by Proposition2.9, by Poincaré’s lemma there exists an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a smooth function $f_{\alpha}\in C^{\infty}(U_{\alpha})$, such that for each $\alpha\in I$ it holds that $\eta_{\alpha}=df_{\alpha}$, where $\eta_{\alpha}$ is the restriction of $\eta$ on $U_{\alpha}$. Then $\mathscr{L}\rightarrow M$ is the $\mathbb{R}$-line bundle over $M$ defined by the 1-cocycle $\{\operatorname{exp}(f_{\alpha}-f_{\beta})\}_{\alpha,\beta\in I}$ and $D$ is the flat connection on $\mathscr{L}\rightarrow M$ defined by the connection 1-forms $\{df_{\alpha}\}_{\alpha\in I}$. For any $\alpha,\beta\in I$, since $df_{\alpha}-df_{\beta}=0$, we have that the function $\operatorname{exp}(f_{\alpha}-f_{\beta})$ is constant.
Let $\{s_{\alpha}\}_{\alpha\in I}$ be an affine local frame of $(\mathscr{L},D)$ and for each $\alpha\in I$ denote by $U_{\alpha}$ the domain of definition of $s_{\alpha}$, so that for each $\alpha,\beta\in I$ there exists a constant $g_{\alpha\beta}\in\mathbb{R}^{\times}$ such that $s_{\beta}(p)=s_{\alpha}(p)g_{\alpha\beta}$ for all $p\in U_{\alpha}\cap U_{\beta}$. Define a bundle metric $\langle\cdot,\cdot\rangle$ on $\mathscr{L}$ by
$\langle s_{\alpha},s_{\alpha}\rangle:=\operatorname{exp}(-2f_{\alpha})$
for every $\alpha\in I$. Then, $\langle\cdot,\cdot\rangle$ is well-defined, and by construction the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is precisely $\eta$.
Now, assume further that $\eta$ is exact. Then $D=d+\eta$ differs from $d$ by an exact 1-form. By definition, $(\mathscr{L},D)$ is a trivial flat $\mathbb{R}$-line bundle. ∎
Corollary 2.47.
Let $((M,\nabla),h)$ be a Hessian manifold. Then the obstruction line bundle of $((M,\nabla),h)$ is a trivial flat $\mathbb{R}$-line bundle if and only if there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$.
Proof.
By Proposition2.46, it remains to prove that if obstruction line bundle $(\mathscr{L},D)$ of $((M,\nabla),h)$ is a trivial flat $\mathbb{R}$-line bundle then there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. But, if
$(\mathscr{L},D)=(M\times\mathbb{R},d+df)$
for some $f\in C^{\infty}(M)$, then it is readily seen that $h=\nabla(df)$. ∎
Corollary 2.48.
An affine manifold admits a Hessian metric of Koszul type if and only of it admits a positive flat line bundle.
Proof.
This follows from Proposition2.46 and the very definition of a positive flat line bundle. ∎
3 Interaction with Classical Geometric Structures
3.1 Digression on Complex Geometry
This digression demonstrates the utility of our framework beyond affine geometry. By lifting the Hessian structure to the tangent bundle, we can leverage powerful theorems from complex analysis.
We prove that every positive flat line bundles over an affine manifold $(M,\nabla)$ can be lifted to a positive holomorphic line bundles over the Sasaki lifting of $(M,\nabla)$, and that the Sasaki lifting of a complete Hessian manifold is Stein, which finally results in an analytic embedding theorem for complete Hessian manifolds.
Let $(M,\nabla)$ be an affine manifold, and let $\pi\colon TM\rightarrow M$ be the tangent bundle of $M$. Denote by $\mathcal{A}$ the sheaf of affine functions on $(M,\nabla)$.
Consider any coordinate map
$U\longrightarrow\mathbb{R}^{n},\quad p\longmapsto(x^{1}(p),\dots,x^{n}(p))$
of $M$ such that $x^{1},\dots,x^{n}\in\mathcal{A}(U)$ are affine functions, and that $TM\rightarrow M$ is trivialized over $U$.
Then, it is readily seen that
$\pi^{-1}(U)\longrightarrow\mathbb{R}^{2n},\quad(p,v^{1}\partial/\partial x^{1}+\cdots+v^{n}\partial/\partial x^{n})\longmapsto(x^{1}(p),\dots,x^{n}(p);v^{1},\dots,v^{n})$
is a coordinate map of $TM$.
Now, the canonical almost complex structure $J_{\nabla}$ on the tangent bundle $T(TM)\rightarrow TM$ is defined by
$J_{\nabla}(\partial/\partial x^{i}):=-\partial/\partial v^{i},\quad J_{\nabla}(\partial/\partial v^{i}):=\partial/\partial x^{i}$
for all $i=1,\dots,n$. By the classical construction in [23], the linear operator $J_{\nabla}$ is a well-defined integrable complex structure, i.e. $(TM,J_{\nabla})$ is a complex manifold. The complex manifold $(TM,J_{\nabla})$ is said to be the Sasaki lifting of $(M,\nabla)$.
Moreover, for each flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$, it is readily seen that the anti-holomorphic part $\bar{\partial}_{D}$ of the $\mathbb{C}$-linear extension of the connection $D$ is a Dolbeault operator on the pullback bundle $\mathscr{L}_{\mathbb{C}}:=\pi^{*}(\mathscr{L}\otimes\mathbb{C})$.
We thus arrived at the following definition:
Definition 3.1.
For an affine manifold $(M,\nabla)$, the complex manifold $(TM,J_{\nabla})$ constructed above is termed the Sasaki lifting of the affine manifold $(M,\nabla)$, and moreover, for a flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$, the holomorphic line bundle $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ over $(TM,J_{\nabla})$ constructed above is termed the holomorphic pullback of $(\mathscr{L},D)$ along $\nabla$, where a holomorphic line bundle here is equivalently defined as a pair consists of a smooth $\mathbb{C}$-line bundle and a Dolbeault operator on it.
For the complex geometric aspect of affine manifolds, the following result is of fundamental importance.
Lemma 3.2.
Let $(M,\nabla)$ be an affine manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Suppose that $(\mathscr{L},D)$ is a positive flat line bundle over $(M,\nabla)$. Then the holomorphic pullback $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ of $(\mathscr{L},D)$ along $\nabla$ is a positive holomorphic line bundle over $(TM,J_{\nabla})$.
Proof.
By the positivity of $(\mathscr{L},D)$, there exists a bundle metric $\langle\cdot,\cdot\rangle$ of $\mathscr{L}\rightarrow M$ such that the first Koszul form $\varkappa$ of the Hessian line bundle $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ satisfies $\nabla\varkappa>0$.
Take any arbitrary point $p\in M$. Since the first Koszul form $\varkappa$ of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is closed, by Poincaré’s lemma there exists an open neighborhood $U$ of $p$ in $M$ and a smooth function $f\in C^{\infty}(U)$ such that $\varkappa|_{U}=df$. Then, by the very definition of the first Koszul form of a Hessian line bundle, there exists a non-vanishing parallel local section $s$ of $\mathscr{L}\rightarrow M$ defined on $U$, such that $\langle s,s\rangle=\operatorname{exp}(-2f)$.
Denote by $\pi$ the bundle map of the tangent bundle $TM\rightarrow M$. We equip $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ with the pullback bundle metric $\langle\cdot,\cdot\rangle_{\mathbb{C}}:=\pi^{*}\langle\cdot,\cdot\rangle$. Then the curvature 2-form $\Theta$ of Chern connection of the Hermitian line bundle $((\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D}),\langle\cdot,\cdot\rangle_{\mathbb{C}})$ is given by
$2\pi\sqrt{-1}\Theta=\partial\bar{\partial}\operatorname{log}(\langle\pi^{*}s,\pi^{*}s\rangle_{\mathbb{C}})=2\partial\bar{\partial}(\pi^{*}f).$
Now it remains to prove that $\pi^{*}f$ is a plurisubharmonic function. But this follows immediately from the assumption that $\nabla(df)=\nabla\varkappa|_{U}>0$. ∎
Although a tangent bundle of positive dimension is never compact, we may hope that the Sasaki liftings are Stein. Fortunately, for the the Sasaki liftings of complete Hessian manifolds, this is indeed true.
Proposition 3.3.
Let $((M,\nabla),h)$ be a complete Hessian manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Then $(TM,J_{\nabla})$ is Stein.
Proof.
Let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. By considering the Sasaki lifting of $(\tilde{M},\pi^{*}\nabla)$, it is readily seen that there exists a holomorphic covering map $T\tilde{M}\rightarrow TM$. Therefore it suffices to show that $T\tilde{M}$ is Stein.
But, by Theorem2.13, $T\tilde{M}$ is a convex tube domain, and is therefore pesudoconvex by complex analysis of several variables. ∎
Now, we are ready to state and prove the main theorem of this section.
Theorem 3.4.
Let $((M,\nabla),h)$ be an $n$-dimensional complete Hessian manifold. Then there exists a proper real analytic embedding of $M$ into $\mathbb{R}^{d}$ for any integer $d>(3n+1)/2$.
Proof.
By Proposition3.3, the Sasaki lifting $(TM,J_{\nabla})$ of $(M,\nabla)$ is Stein. By Stein’s embedding theorem [24], there exists a proper holomorphic embedding $\Phi$ of $(TM,J_{\nabla})$ into $\mathbb{C}^{d}$ for any integer $d>(3n+1)/2$. Now, it suffices to restrict the real part $\operatorname{Re}(\Phi)$ of $\Phi$ to the totally real submanifold $M$ of $(TM,J_{\nabla})$. ∎
Remark 3.5.
Recall that Whitney’s embedding theorem states that all real analytic $n$-manifolds can be analytically embedded into $\mathbb{R}^{2n}$. Theorem3.4 suggests a more efficient bound on analytical embedding dimension for complete Hessian manifolds.
3.2 Legendre Transformation
It turns out that there exists a natural duality between radiant Hessian manifolds and Hessian manifolds of Koszul type. Also, for a radiant Hessian manifold $((M,\nabla),h)$ of Koszul type, we prove that its dual $((M,\nabla^{*}),h)$ is also a radiant Hessian manifold of Koszul type, and provide an explicit expression of the $1$-form potential of $((M,\nabla^{*}),h)$ in terms of that of $((M,\nabla),h)$, resembling the classical Legendre transformation.
In this section, we introduce radiant manifolds and then study Hessian manifold in terms of the generalized Legendre duality mentioned above.
Definition 3.6.
Let $(M,\nabla)$ be an affine manifold. A vector field $H$ on $M$ is termed an Euler vector field of $(M,\nabla)$, if $\nabla_{X}H=X$ for every vector field $X$ on $M$. An affine manifold is said to be radiant, if it admits an Euler vector field.
Remark 3.7.
It is readily seen that, if $X$ and $Y$ are vector fields on an affine manifold such that $X$ is Eulerian, then $Y$ is also Eulerian if and only if XYX-Y is parallel.
Definition 3.8.
Let $M$ be a differentiable manifold. An atlas $\mathscr{A}=\left\{\left(U_{\alpha},\varphi_{\alpha}\right)\mid\alpha\in I\right\}$ of $M$ is said to be a radiant affine atlas, if for any $\alpha,\beta\in I$ there exist $A\in\operatorname{GL}_{n}(\mathbb{R})$ such that
$\varphi_{\beta}\left(\varphi_{\alpha}^{-1}(\bf{x})\right)=A{\bf{x}}$
for all ${\bf{x}}\in\varphi_{\alpha}\left(U_{\alpha}\cap U_{\beta}\right)$. An element of a radiant affine atlas of $M$ is called a radiant affine chart of $M$, and the local coordinates associated to a radiant affine chart of $M$ are termed radiant affine coordinates.
The following theorem shows that the above two definitions of radiant affine manifolds are equivalent to each other.
Theorem 3.9.
An affine manifold $(M,\nabla)$ is radiant if and only if the adapted affine atlas of $\nabla$ is a radiant affine atlas of $M$.
Proof.
See for example [15] and [25]. ∎
The following proposition characterizes the universal coverings of compact radiant affine manifolds as open cones in the Euclidean spaces.
Proposition 3.10.
[…]
Proof.
By Theorem2.13, we may identify $\Omega:=\Phi(\tilde{M})=\tilde{M}$.
Since, up to a translation, $\Omega$ is a cone with apex at the origin, all affine automorphisms of $(\Omega,\pi^{*}\nabla)$ are in fact linear. Now, it suffices to descend the standard Euler vector field
$x^{1}\frac{\partial}{\partial x^{1}}+\cdots+x^{n}\frac{\partial}{\partial x^{n}}$
defined on $(\Omega,\pi^{*}\nabla)$ to an Euler vector field of $(M,\nabla)$. ∎
The converse statement also holds if we assume compactness.
Proposition 3.11.
[…]
Proof.
By Theorem2.13, we may identify $\Omega:=\Phi(\tilde{M})=\tilde{M}$. An Euler vector field $H$ of $(M,\nabla)$ lifts to a standard Euler vector field
$\pi^{*}H=\sum_{i=1}^{n}(x^{i}-a^{i})\frac{\partial}{\partial x^{i}}$
on $(\Omega,\pi^{*}\nabla)$, where $a^{1},\dots,a^{n}\in\mathbb{R}$. Moreover, by Poincaré-Hopf index theorem, we have $(a^{1},\dots,a^{n})\notin\Omega$. Up to an affine change of coordinates, we may assume $a^{1},\dots,a^{n}=0$. Since $M$ is compact, it follows that $H$ is a complete vector field. Since $\pi\colon\Omega\rightarrow M$ is a covering space, we obtain that $\pi^{*}H$ is also a complete vector field as $H$ is. Now it remains to recall that the integral curves of the standard Euler vector field
$x^{1}\frac{\partial}{\partial x^{1}}+\cdots+x^{n}\frac{\partial}{\partial x^{n}}$
are rays emitting from the origin of $\mathbb{R}^{n}$. ∎
We shall also recall the core concept of information geometry, that is, the notion of dual connections.
Definition 3.12.
Let $(M,g)$ be a pseudo-Riemannian manifold. For a connection $D$ on the tangent bundle $TM\rightarrow M$, its conjugate connection $D^{*}$ with respect to $g$ is defined to be the unique affine connection on the tangent bundle of $M$ such that the equation
$Zg(X,Y)=g\left(D_{Z}X,Y\right)+g\left(X,D_{Z}^{*}Y\right)$
holds for all smooth vector fields $X,Y,Z$ on $M$.
See [11] for a proof of the fact that the conjugate connection is well-defined for every connection on the tangent bundle of a pseudo-Riemannian manifold.
Theorem 3.13.
Let $(M,\nabla)$ be an affine manifold, and let $g$ be a pseudo-Riemannian metric on $(M,\nabla)$. Then $g$ is a Hessian metric on $(M,\nabla)$ if and only if $\nabla^{*}$ is also flat and torsion-free, or equivalently, the $(0,3)$-tensor $\nabla g$ is totally symmetric. Moreover, in this case
$\bar{\nabla}:=\frac{1}{2}\left(\nabla^{*}+\nabla\right)$
is the Levi-Civita connection of $(M,g)$.
Proof.
See [15] for a proof of this statement. ∎
Lemma 3.14.
Let $(M,g)$ be a pseudo-Riemannian manifold, and let $\nabla$ be a connection on the tangent bundle $TM\rightarrow M$. Then for any 1-form $\theta\in\Omega^{1}(M)$ and any vector field $X$ on $M$, it holds that $\nabla_{X}(\theta^{\sharp})=(\nabla^{*}_{X}\theta)^{\sharp}$, and in particular $\nabla^{*}\theta=0$ if and only if $\nabla(\theta^{\sharp})=0$, where $\sharp$ is the musical isomorphism.
Proof.
Take arbitrary vector fields $X,Y$ on $M$. Then
$(\nabla^{*}_{X}\theta)(Y)=X(\theta(Y))-\theta(\nabla^{*}_{X}Y).$
Now, since $\theta(Y)=g(\theta^{\sharp},Y)$, we have
$(\nabla^{*}_{X}\theta)(Y)=X(g(\theta^{\sharp},Y))-g(\theta^{\sharp},\nabla^{*}_{X}Y).$
Moreover, since
$X(g(\theta^{\sharp},Y))=g(\nabla_{X}\theta^{\sharp},Y)+g(\theta^{\sharp},\nabla^{*}_{X}Y),$
substituting this into the previous expression, we obtain
$(\nabla^{*}_{X}\theta)(Y)=(g(\nabla_{X}\theta^{\sharp},Y)+g(\theta^{\sharp},\nabla^{*}_{X}Y))-g(\theta^{\sharp},\nabla^{*}_{X}Y)=g(\nabla_{X}\theta^{\sharp},Y).$
This proves that $\nabla_{X}\theta^{\sharp}=(\nabla^{*}_{X}\theta)^{\sharp}$ holds for all vector fields $X$ on $M$. ∎
The following two fundamental theorems illustrate the duality between radiant Hessian manifolds and Hessian manifolds of Koszul type.
Proposition 3.15.
Let $((M,\nabla),h)$ be a Hessian manifold of Koszul type. Then $((M,\nabla^{*}),h)$ is a radiant Hessian manifold. In fact, if $\eta$ is a $1$-form potential of $((M,\nabla),h)$, then $\eta^{\sharp}$ is an Euler vector field of $(M,\nabla^{*})$, where $\sharp$ is the musical isomorphism.
Proof.
Consider affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. For each $i=1,\dots,n$, define $p_{i}:=\eta(\partial/\partial x^{i})$, so that
$\eta=\sum_{i=1}^{n}p_{i}dx^{i}.$
Since $\partial/\partial x^{i}$ is parallel, we obtain that
$h=\nabla\eta=\sum_{i=1}^{n}dp_{i}\otimes dx^{i}.$
Then for every $i,j\in\{1,\dots,n\}$ we have that $h(\partial/\partial x^{i},\partial/\partial p_{j})=\delta^{i}_{j}$ and hence
$h(\nabla^{*}_{X}(\partial/\partial p_{j}),\partial/\partial x^{i})=$ $h(\nabla_{X}(\partial/\partial x^{i}),\partial/\partial p_{j})+h(\partial/\partial x^{i},\nabla^{*}_{X}(\partial/\partial p_{j}))$
$=$ $Xh(\partial/\partial x^{i},\partial/\partial p_{j})$
$=$ $X\delta^{i}_{j}$
$=$ $0$
for every vector field $X$ on $M$. Therefore $(dx^{i})^{\sharp}=\partial/\partial p_{i}$ and $\nabla^{*}(\partial/\partial p_{i})=0$ for all $i=1,\dots,n$. This proves that
$\eta^{\sharp}=\sum_{i=1}^{n}p_{i}(dx^{i})^{\sharp}=\sum_{i=1}^{n}p_{i}\frac{\partial}{\partial p_{i}}$
and it readily seen that the latter is an Euler vector field of $(M,\nabla^{*})$. ∎
Proposition 3.16.
Let $((M,\nabla),h)$ be a radiant Hessian manifold. Then $((M,\nabla^{*}),h)$ is a Hessian manifold of Koszul type. In fact, if $H$ is an Euler vector field of $(M,\nabla)$, then $H^{\flat}$ is a $1$-form potential of $((M,\nabla^{*}),h)$, where $\flat$ is the musical isomorphism.
Proof.
Consider radiant affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. For each $i=1,\dots,n$, since $\nabla(dx^{i})=0$, by Lemma3.14 we have $\nabla^{*}(\partial/\partial x^{i})^{\flat}=0$. Recall that parallel 1-forms on an affine manifold is closed. By Poincaré’s lemma, there exist local coordinates $p_{1},\dots,p_{n}$ of $M$, such that $dp_{i}=(\partial/\partial x^{i})^{\flat}$ for all $i=1,\dots,n$. Therefore we arrive at the expression
$h=\sum_{i=1}^{n}dp_{i}\otimes dx^{i}$
for the Hessian metric $h$.
Since $x^{1},\dots,x^{n}$ are radiant affine, it is readily seen that the vector field
$X:=\sum_{i=1}^{n}x^{i}\frac{\partial}{\partial x^{i}}$
is a well-defined Euler vector field of $(M,\nabla)$. By the uniqueness of Euler vector field, there exists a parallel vector field $Y$ on $(M,\nabla)$, such that $H=X+Y$. Therefore, again by Lemma3.14, we obtain that
$\nabla^{*}(H^{\flat})=\nabla^{*}(Y^{\flat})+\nabla^{*}\sum_{i=1}^{n}x^{i}(\partial/\partial x^{i})^{\flat}=0+\sum_{i=1}^{n}dx^{i}\otimes dp_{i}=\sum_{i=1}^{n}dp_{i}\otimes dx^{i}=h$
as required. ∎
Corollary 3.17.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type. Then the Hessian manifold $((M,\nabla^{*}),h)$ is also radiant of Koszul type.
Proof.
This follows immediately from Proposition3.15 and Proposition3.16. ∎
Now, in view of Corollary3.17, for a radiant Hessian manifold $((M,\nabla),h)$ of Koszul type, it is natural to ask for an explicit expression for the $1$-form potential for its dual $((M,\nabla^{*}),h)$. We will now give an satisfactory answer to this question.
For a vector field $X$, as per usual, we denote by $\mathcal{L}_{X}$ the Lie derivative along $X$.
Definition 3.18.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type, with Euler vector field $H$ and $1$-form potential $\eta$. The differential form $\eta^{*}:=-\eta+\mathcal{L}_{H}\eta$ is said to be the Legendre transform of $\eta$ along direction $H$.
Theorem 3.19.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type, with Euler vector field $H$ and $1$-form potential $\eta$. Then the Legendre transform $\eta^{*}$ of $\eta$ along direction $H$ satisfies $h=\nabla^{*}\eta^{*}$.
Proof.
Consider radiant affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$ so that
$H=\sum_{i=1}^{n}x^{i}\frac{\partial}{\partial x^{i}}$
is Eulerian. For each $i\in\{1,\dots,n\}$ define $p_{i}:=\eta(\partial/\partial x^{i})$. Then
$\eta=\sum_{i=1}^{n}p_{i}dx^{i}$
and $\nabla^{*}(dp_{1})=\cdots=\nabla^{*}(dp_{n})=0$. By Cartan’s magic formula, we obtain
$\eta^{*}=-\eta+\mathcal{L}_{H}\eta=d(\eta(H))-\eta=d(x^{1}p_{1}+\cdots+x^{n}p_{n})-\sum_{i=1}^{n}p_{i}dx^{i}=\sum_{i=1}^{n}x^{i}dp_{i}$
and hence
$\nabla^{*}\eta^{*}=\sum_{i=1}^{n}dx^{i}\otimes dp_{i}=h$
as required. ∎
Remark 3.20.
If the Hessian manifold $((M,\nabla),h)$ under consideration satisfies
  1. 1.
    $M=\Omega$ is a convex domain in $\mathbb{R}^{d}$;
  2. 2.
    $\nabla$ is the Levi-Civita connection of the $d$-dimensional flat Euclidean space;
  3. 3.
    $h$ is the Hessian quadratic form of a strictly convex potential,
then the construction in Definition3.18 recovers the Legendre transformation in classical mechanics.
3.3 Space-like Lagrangian Immersions
In this section, we characterize Hessian manifolds of Koszul type as space-like Lagrangians.
Throughout this section we use Einstein’s summation convention.
Definition 3.21 (Patterson-Walker,1952[16]).
For a differentiable manifold $M$ and a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$ with Christoffel symbol $\Gamma^{k}_{ij}$, the pseudo-Riemannian metric
$g:=\frac{1}{2}(dx^{i}\otimes dp_{i}+dp_{i}\otimes dx^{i})-\Gamma^{k}_{ij}p_{k}dx^{i}dx^{j}$
on $T^{*}M$ is called the Riemann extension of $\nabla$, where $x^{1},\dots,x^{n},p_{1},\dots,p_{n}$ here are canonical coordinates on $T^{*}M$.
Note that, the Riemann extension of a flat torsion-free connection is of neutral signature.
Definition 3.22.
For a differentiable manifold $M$ and a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$, a $1$-form $\eta$ on $M$, i.e. a smooth section $\eta:M\rightarrow T^{*}M$ of the cotangent bundle $T^{*}M\rightarrow M$, is said to be a space-like Lagrangian immersion of $(M,\nabla)$, if $\eta:M\rightarrow T^{*}M$ is an immersed Lagrangian submanifold of $(T^{*}M,\omega)$ and the symmetric $(0,2)$-tensor $\eta^{*}g$ is positive definite, where $\omega$ is the standard symplectic form on $T^{*}M$ and $g$ is the Riemann extension of $\nabla$.
The following theorem shows that, in the category of affine manifolds, space-like immersed Lagrangians are precisely those Hessian manifolds of Koszul type.
Theorem 3.23.
Let $M$ be a differentiable manifold, and let $\nabla$ be a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$. Then, for a $1$-form $\eta$ on $M$, the following statements are equivalent:
  1. 1.
    $\eta$ is a space-like Lagrangian immersion of $(M,\nabla)$;
  2. 2.
    $\eta$ is closed, and $\nabla\eta>0$.
Moreover, for a closed $1$-form $\eta$, the $(0,3)$-tensor $\nabla(\nabla\eta)$ is totally symmetric if and only if $R^{m}_{ijk}\eta_{m}=0$, where $R$ here is the curvature tensor of $\nabla$.
Proof.
Recall from differential topology that a smooth cross-section of a fiber bundle is always an immersion.
Denote by $\tau$ the tautological 1-form of $T^{*}M$, and denote by $\omega$ the standard symplectic form on $T^{*}M$. Then we have $\omega=-d\tau$ and $\eta^{*}\tau=\eta$. Since $d\eta=d(\eta^{*}\tau)=\eta^{*}d\tau=-\eta^{*}\omega$, we obtain that $\eta\in\Omega^{1}(M)$ is closed if and only if $\eta:M\rightarrow T^{*}M$ is a Lagrangian immersion.
Consider canonical coordinates $x^{1},\dots,x^{n},p_{1},\dots,p_{n}$ on $T^{*}M$. Denote by $g$ the Riemann extension of $\nabla$. Now, it suffices to notice that, by the very definition, we have
$\nabla\eta=(\partial_{i}\eta_{j}-\Gamma^{k}_{ij}\eta_{k})dx^{i}\otimes dx^{j}=d\eta_{i}\otimes dx^{i}-\Gamma^{k}_{ij}\eta_{k}dx^{i}dx^{j}=\eta^{*}g$
where $\Gamma^{k}_{ij}$ is the Christoffel symbol of $\nabla$. Denote $h:=\nabla\eta$ and denote $e_{i}:=\partial/\partial x^{i}$. Then
$\nabla_{e_{k}}h=(\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik})dx^{i}\otimes dx^{j}.$
Therefore, $\nabla h$ is totally symmetric if and only if $\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik}$ is invariant under permutations of indices $i,j,k$. Define $S^{m}_{ijk}:=\partial_{k}\Gamma^{m}_{ij}-\Gamma^{m}_{il}\Gamma^{l}_{jk}-\Gamma^{m}_{jl}\Gamma^{l}_{ik}$. Then a straightforward computation yields that
$\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik}=\partial_{i}\partial_{j}\eta_{k}-S^{m}_{ijk}\eta_{m}-(\Gamma^{m}_{ij}\partial_{k}\eta_{m}+\Gamma^{m}_{jk}\partial_{i}\eta_{m}+\Gamma^{m}_{ik}\partial_{j}\eta_{m}).$
Notice that $\partial_{i}\partial_{j}\eta_{k}$ and $\Gamma^{m}_{ij}\partial_{k}\eta_{m}+\Gamma^{m}_{jk}\partial_{i}\eta_{m}+\Gamma^{m}_{ik}\partial_{j}\eta_{m}$ are already invariant under permutations of indices $i,j,k$. Therefore, the expression $\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik}$ is invariant under permutations of indices $i,j,k$ if and only if $S^{m}_{ijk}\eta_{m}$ is. But, by construction, we have $S^{m}_{ijk}-S^{m}_{ikj}=-R^{m}_{ijk}$. This concludes the proof. ∎
Corollary 3.24.
Let $(M,\nabla)$ be an affine manifold, and $\eta$ a $1$-form on $M$. Then, the following statements are equivalent:
  1. 1.
    $\eta$ is a space-like Lagrangian immersion of $(M,\nabla)$;
  2. 2.
    $\eta$ is closed, and $\nabla\eta>0$, i.e. $((M,\nabla),\nabla\eta)$ is a Hessian manifold of Koszul type.
Proof.
Recall Theorem3.13 and then apply Theorem3.23. ∎
We remark that, in [26], J.Zhang and G.Khan derived a similar condition for totally symmetry of the $(0,3)$-tensor described in Theorem3.23.
3.4 Hyperbolic Affine Manifolds
For completeness, we recall yet another necessary and sufficient condition for a compact affine manifold to admit a Hessian metric of Koszul type.
Definition 3.25.
A convex domain $\Omega$ in $\mathbb{R}^{d}$ is said to be regular, if it contains no straight line, i.e. a $1$-dimensional affine subspace of $\mathbb{R}^{d}$.
Definition 3.26.
Let $(M,\nabla)$ be an affine manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. The affine manifold $(M,\nabla)$ is said to be hyperbolic, if the affine development $\Phi\colon\tilde{M}\rightarrow\mathbb{R}^{n}$ of $(\tilde{M},\pi^{*}\nabla)$ is a diffeomorphism onto its image, and its image is a regular convex domain in $\mathbb{R}^{n}$.
Theorem 3.27.
A compact affine manifold is hyperbolic if and only if it admits a Hessian metric of Koszul type.
Proof.
This theorem is proved in [15]. ∎
To summarize, we suggest the following theorem:
Theorem 3.28.
For a compact affine manifold, the following statements are equivalent:
  1. 1.
    it is hyperbolic;
  2. 2.
    it admits a Hessian metric of Koszul type;
  3. 3.
    it admits a positive flat line bundle;
  4. 4.
    it admits a space-like Lagrangian immersion.
Proof.
Combine Corollary2.48&3.24 with Theorem3.27. ∎
We shall now introduce an important property of compact hyperbolic affine manifolds.
[…]
Definition 3.29.
Let $(M,\nabla)$ be an oriented affine manifold, and let
[…]
be the holonomy of $\nabla$. The homomorphism
$\pi_{1}(M)\rightarrow\mathbb{R},\quad g\mapsto\operatorname{log}(\operatorname{det}(\rho(g)))$
as a group cohomology class in $\mathrm{H}^{1}(\pi_{1}(M);\mathbb{R})$, is termed the canonical class of $(M,\nabla)$.
The canonical class of an oriented affine manifold is by no means an unfamiliar object: We now show that, when the oriented affine manifold is equipped with a complete Hessian metric, its first Koszul form represents precisely the canonical class.
Lemma 3.30.
Let $((M,\nabla),h)$ be a complete oriented Hessian manifold. Then, under the identification $\operatorname{H}^{1}(M;\mathbb{R})=\operatorname{H}^{1}(\pi_{1}(M);\mathbb{R})$, the cohomology class represented by the first Koszul form of $((M,\nabla),h)$ is the canonical class of $(M,\nabla)$.
Proof.
Denote by $\varkappa\in\Omega^{1}(M)$ the first Koszul form of $((M,\nabla),h)$. Since
$\nabla_{X}\operatorname{vol}_{h}=\varkappa(X)\operatorname{vol}_{h}$
for every vector field $X$ on $M$, it follows that $\varkappa$ is the connection 1-form of the induced connection $\nabla$ on $\bigwedge^{n}T^{*}M$. Therefore
$\operatorname{log}(\operatorname{det}(\rho(g)))=\int_{\gamma}\varkappa$
[…]
The following corollary is natural.
Corollary 3.31.
Let $((M,\nabla),h)$ be a complete oriented Hessian manifold. Suppose that the canonical class of $(M,\nabla)$ is a rational cohomology class. Then the canonical class of $(M,\nabla^{*})$ is also rational.
Proof.
The desired result follows immediately from Lemma3.30 and the fact that the determinant of $h$ in affine coordinates of $(M,\nabla^{*})$ is the reciprocal of the determinant of $h$ in affine coordinates of $(M,\nabla)$. ∎
We shall show that, on an oriented compact manifold, an hyperbolic affine structure can always be deformed into another hyperbolic affine structure of which canonical class is rational.
To proceed, we adapt a theorem of J.-L.Koszul to our setting.
Theorem 3.32 (Koszul, 1968[17]).
[…]
Applying Theorem3.32 together with Lie algebra-valued group cohomology, we can prove the following density theorem:
Theorem 3.33.
Let $(M,\nabla)$ be a compact oriented hyperbolic affine manifold. Then there exists a flat and torsion-free connection $\nabla^{\prime}$ on the tangent bundle $TM\rightarrow M$ such that $(M,\nabla^{\prime})$ is a hyperbolic affine manifold of which canonical class is rational.
Proof.
[…]
$\operatorname{Hom}(G,\operatorname{GL}^{+}_{n}(\mathbb{R}))\rightarrow\operatorname{Hom}(G,\mathbb{R}^{+}),\quad\rho\mapsto(g\mapsto\operatorname{det}(\rho(g)))$
is a submersion, or equivalently, the induced map
$\operatorname{tr}_{*}\colon\mathrm{H}^{1}(G;\mathfrak{gl}_{n}(\mathbb{R}))\rightarrow\mathrm{H}^{1}(G;\mathbb{R})$
of the trace homomorphism
$\operatorname{tr}\colon\mathfrak{gl}_{n}(\mathbb{R})\rightarrow\mathbb{R}$
is surjective, which is a direct consequence of the fact that the short exact sequence
$0\rightarrow\mathfrak{sl}_{n}(\mathbb{R})\rightarrow\mathfrak{gl}_{n}(\mathbb{R})\rightarrow\mathbb{R}\rightarrow 0$
splits. This concludes the proof. ∎
3.5 Riemannian Geometric Aspect
While previous sections focus on the topology of Hessian manifolds, in this section we delve into geometric aspect.
Throughout this section we use Einstein’s summation convention.
The following notion is classical in information geometry.
Definition 3.34.
The Amari–Chentsov tensor $A$ of a Hessian manifold $((M,\nabla),h)$ is defined to be the $(0,3)$-tensor satisfying
$A(X,Y,Z)=(\nabla_{Z}h)(X,Y)=h(\nabla^{*}_{X}Y-\nabla_{X}Y,Z)$
for all vector fields $X,Y,Z$ on $M$.
We first note that the Riemann curvature tensor of a Hessian manifold $((M,\nabla),h)$ enjoys a simplified form, which also provides a trivial bound for Ricci curvature in terms of the Amari–Chentsov tensor.
Lemma 3.35.
Let $((M,\nabla),h)$ be a Hessian manifold, and $A$ its Amari–Chentsov tensor. Then $A$ is totally symmetric and
$4R_{ijkl}=A_{il}^{m}A_{jkm}-A_{ik}^{m}A_{jlm},$
where $A_{ij}^{k}:=A_{ijl}h^{kl}$, and $R$ is the Riemann curvature tensor of $(M,h)$.
Proof.
See for example [27] for a proof. ∎
The followings are immediate consequences of Lemma3.35.
Corollary 3.36.
Let $((M,\nabla),h)$ be an $n$-dimensional Hessian manifold, and let $A$ be its Amari–Chentsov tensor. Then, in an orthonormal frame, the Riemann curvature tensor $R$ and Ricci tensor $\operatorname{Ric}$ of $(M,h)$ satisfies $4R_{ijkl}=[A_{i},A_{j}]_{kl}$ and
$4\operatorname{Ric}_{ij}=\operatorname{tr}(A_{i}A_{j})-\sum_{k=1}^{n}\operatorname{tr}(A_{k})A_{ijk}$
where $A_{k}$ is the $n\times n$ matrix of which $(i,j)$-th entry is $A_{ijk}$.
Corollary 3.37.
Let $f$ be a smooth function defined on a domain in $\mathbb{R}^{2}$. Suppose that
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a pseudo-Riemannian metric. Then Gaußian curvature $K$ of the metric $g$ satisfies
$K=-\frac{\left|\begin{array}[]{ccc}f_{xx}&f_{xxx}&f_{xxy}\\ f_{xy}&f_{xxy}&f_{xyy}\\ f_{yy}&f_{xyy}&f_{yyy}\end{array}\right|}{4\left|\begin{array}[]{cc}f_{xx}&f_{xy}\\ f_{xy}&f_{yy}\end{array}\right|^{2}}=\frac{f_{xx}\{f_{yy},f_{xy}\}+f_{yy}\{f_{xy},f_{xx}\}+f_{xy}\{f_{xx},f_{yy}\}}{4(f_{xx}f_{yy}-f_{xy}f_{xy})^{2}}$
where $\{\cdot,\cdot\}$ is the standard Poisson bracket normalized by $\{x,y\}=1$.
Corollary 3.38.
Let $((M,\nabla),h)$ be an $n$-dimensional Hessian manifold, and let $A$ be its Amari–Chentsov tensor. Then the Ricci tensor $\operatorname{Ric}$ of $M$ satisfies
$\frac{n-1}{2}|A|^{2}h(X,X)\leq\operatorname{Ric}(X,X)\leq\frac{n-1}{4}|A|^{2}h(X,X)$
for every vector field $X$ on $M$.
Proof.
Apply the Cauchy–Schwarz inequality to Lemma3.36. ∎
The following simple observations on sectional curvature shows that there exists restrictions on other Riemannian metrics that a Hessian manifold can possess.
Proposition 3.39.
Let $((M,\nabla),h)$ be a compact Hessian manifold. Then for any Riemannian metric $g$ on $M$, the Ricci tensor of $(M,g)$ is not everywhere positive.
Proof.
If the Ricci tensor of $(M,g)$ is positive, then by Bonnet–Myers theorem, the fundamental group of $M$ is finite, which contradicts Corollary2.7. ∎
Proposition 3.40.
Let $((M,\nabla),h)$ be a compact Hessian 4-manifold. Then for any Riemannian metric $g$ on $M$, the Riemannian 4-manifold $(M,g)$ is not of negative sectional curvature.
Proof.
If $(M,g)$ is of negative sectional curvature, then, by Hopf’s conjecture in dimension four which is proved using Gauß–Bonnet formula, we have that $\chi(M)>0$, a contradiction to Proposition2.17. ∎
Finally, we note an application of de Rham splitting theorem to Hessian manifolds.
Proposition 3.41.
Let $((M,\nabla),h)$ be a $(n+1)$-dimensional compact Hessian manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. Then there exist a complete Riemannian metric $g$ on $\mathbb{R}^{n}$ such that $(\tilde{M},\pi^{*}h)$ is isometric to the product of Riemannian manifolds $(\mathbb{R}^{n},g)$ and $(\mathbb{R},dt^{2})$.
Proof.
By considering the orientation double cover of $M$, without lose of generality we may assume $M$ is orientable.
If $(M,h)$ is a flat Riemannian manifold, then $(\tilde{M},\pi^{*}h)$ is isometric to the $(n+1)$-dimensional flat Euclidean space, so the required property is trivially satisfied.
From now on, we assume that $(M,h)$ is not a flat Riemannian manifold. By Theorem2.13, we may assume further that $\tilde{M}=\mathbb{R}^{n+1}$. Denote by $\bar{\nabla}$ the Levi-Civita connection of $(M,h)$. Then $\pi^{*}\bar{\nabla}$ is the Levi-Civita connection of $(\mathbb{R}^{n+1},\pi^{*}h)$. By Proposition2.37 and Theorem2.38, the first Koszul form $\varkappa$ of $((M,\nabla),h)$ satisfies $\varkappa\neq 0$ and $\bar{\nabla}\varkappa=0$. Now, apply de Rham splitting theorem to the non-trivial parallel vector field $\pi^{*}\varkappa$ on the complete Riemannian manifold $(\mathbb{R}^{n+1},\pi^{*}h)$. ∎
Corollary 3.42.
Let $((M,\nabla),h)$ be a compact Hessian manifold of dimension $n\geq 2$. Then $(M,h)$ is not of negative sectional curvature.
Proof.
Let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. Assume the contrary that $(M,h)$ is of negative sectional curvature. Then $(\tilde{M},\pi^{*}h)$ is also of negative sectional curvature, but this contradicts Proposition3.41. ∎
3.6 The Cheng-Yau Metric
Amongst all compact orientable affine manifolds, those being simultaneously radiant and hyperbolic are special, as they admit canonical Hessian metrics. We shall explain this observation in the present section.
For a smooth function $u$ defined on a domain in $\mathbb{R}^{n}$, we write
$\nabla u:=\left[\begin{array}[]{ccc}\partial_{1}u\\ \vdots\\ \partial_{n}u\end{array}\right],\quad\operatorname{Hess}(u):=\left[\begin{array}[]{ccc}\partial_{1}\partial_{1}u&\cdots&\partial_{1}\partial_{n}u\\ \vdots&\ddots&\vdots\\ \partial_{n}\partial_{1}u&\cdots&\partial_{n}\partial_{n}u\end{array}\right],$
where $\partial_{i}=\partial/\partial x^{i}$ for $i=1,\dots,n$, and $x^{1},\dots,x^{n}$ are the standard coordinates of $\mathbb{R}^{n}$.
Recall the following existence and uniqueness result on real Monge–Ampère equation due to Cheng and Yau.
Theorem 3.43 (Cheng-Yau, 1982[18]).
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$. Then there exists a unique convex function $u\in C^{\infty}(\Omega)$ satisfying the following properties:
  1. 1.
    $\operatorname{det}(\operatorname{Hess}(u))=\operatorname{exp}(2u)^{n}$, and $u(\textup{{x}})\rightarrow\infty$ as $\textup{{x}}\rightarrow\partial\Omega$;
  2. 2.
    $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a complete Riemannian metric on $\Omega$;
  3. 3.
    $L^{*}g=g$ for every linear automorphism $L$ of $\Omega$.
Definition 3.44.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$. The convex function $u\in C^{\infty}(\Omega)$ satisfying the properties listed in Theorem3.43 is termed the Cheng-Yau potential of $\Omega$.
The differential of a Cheng-Yau potential enjoys one more important property, namely, it is always a unit covector field.
Lemma 3.45.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$, and let $u\in C^{\infty}(\Omega)$ be the Cheng-Yau potential of $\Omega$. Then the holds that $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv 1$.
Proof.
For any $t>0$, define function $u_{t}\in C^{\infty}(\Omega)$ by $u_{t}(\textup{{x}}):=u(t\cdot\textup{{x}})+\operatorname{log}(t)$. Then straightforward computation yields that $u_{t}$ is also a Cheng-Yau potential of $\Omega$ for every $k>0$. Therefore, by the uniqueness part of Theorem3.43, we obtain the functional equation
$u(\textup{{x}})-u(t\cdot\textup{{x}})=\operatorname{log}(t)$
where $\textup{{x}}\in\Omega$ and $k>0$. Now, for any $\textup{{x}}\in\Omega$, since
$\frac{u(\textup{{x}}+t\cdot\textup{{x}})-u(\textup{{x}})}{t}=-\operatorname{log}(1+t)/t\rightarrow-1$
as $t\rightarrow 0$, we have that
$\langle\nabla u(\textup{{x}}),\textup{{x}}\rangle\equiv-1.$ (1)
Differentiate both sides of equation(1) again, we obtain that
$\operatorname{Hess}(u)^{-1}\nabla u(\textup{{x}})=-\textup{{x}}$ (2)
holds for all $\textup{{x}}\in\Omega$. Substitute equation(2) back into equation(1), we finally arrive at $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv 1$ as required. ∎
Using Lemma3.45, we can prove that Cheng-Yau potentials are harmonic functions.
Proposition 3.46.
Let $\Omega$ be a domain in $\mathbb{R}^{d}$, and let $u\in C^{\infty}(\Omega)$. Suppose that $\operatorname{Hess}(u)$ is nowhere degenerated, so that $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a pseudo-Riemannian metric on $\Omega$. Then, the Laplace–Beltrami operator $\Delta^{g}$ of $(\Omega,g)$ satisfies
$\Delta^{g}\phi=\operatorname{tr}(\operatorname{Hess}(u)^{-1}\operatorname{Hess}(\phi))-\frac{1}{2}\langle\nabla\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u))),\operatorname{Hess}(u)^{-1}\nabla\phi\rangle$
for all $\phi\in C^{\infty}(\Omega)$.
Proof.
The proof of this theorem is a straightforward computation and hence omitted. ∎
Proposition 3.47.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{d}$, and let $u\in C^{\infty}(\Omega)$ be the Cheng-Yau potential of $\Omega$, so that $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a Riemannian metric on $\Omega$. Then $u$ is a harmonic function on $(\Omega,g)$.
Proof.
Since $u$ is a Cheng-Yau potential, we have $2d\cdot u=\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u)))$. Denote by $\Delta^{g}$ the Laplace–Beltrami operator of $(\Omega,g)$. Then by Lemma3.45 and Proposition3.46,
$\Delta^{g}u=$ $\operatorname{tr}(\operatorname{Hess}(u)^{-1}\operatorname{Hess}(u))-\frac{1}{2}\langle\nabla\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u))),\operatorname{Hess}(u)^{-1}\nabla u\rangle$
$=$ $\operatorname{tr}(\operatorname{Id})-d\cdot\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle$
$=$ $d-d\cdot 1$
$=$ $0$
as required. ∎
Adapt Cheng and Yau’s result to our setting, we immediately obtain the following crucial observation.
Proposition 3.48.
Let $(M,\nabla)$ be an $n$-dimensional compact orientable affine manifold. Suppose that $(M,\nabla)$ is both radiant and hyperbolic. Then, there exists a unique space-like Lagrangian immersion $\eta$ of $(M,\nabla)$ such that $\varkappa=n\eta$, where $\varkappa$ is the first Koszul form of the Hessian manifold $((M,\nabla),\nabla\eta)$.
Proof.
[…]
$n\cdot df=\frac{1}{2}d\operatorname{log}(\operatorname{det}(\operatorname{Hess}(f)))$
descends to $n\cdot\eta=\varkappa$ on $(M,\nabla)$ so that $\eta$ is a space-like Lagrangian immersion of $(M,\nabla)$ and $\varkappa$ is the first Koszul form of $((M,\nabla),\nabla\eta)$. The uniqueness of $\eta$ also follows immediately from the uniqueness of the Cheng-Yau potential. ∎
4 Topics in Low Dimensional Hessian Geometry
4.1 Complete Hessian surfaces
It is a well-known fact that an orientable closed Hessian surface is homeomorphic to the 2-torus $\mathbb{T}^{2}$. The ultimate objective of this section is to classify all complete Hessian surfaces topologically. To achieve this goal, the following preparations are needed:
Definition 4.1.
Let $(M,\nabla)$ be an affine manifold. We denote by $\operatorname{Aff}(M;\nabla)$ the group of affine automorphisms of $(M,\nabla)$. Whenever $M=\Omega$ is a convex domain in $\mathbb{R}^{d}$, and $\nabla$ is the Levi-Civita connection of the $d$-dimensional flat Euclidean space, by convention we omit $\nabla$ to write $\operatorname{Aff}(\Omega):=\operatorname{Aff}(\Omega;\nabla)$.
Lemma 4.2.
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$ and $\operatorname{recc}(\Omega)$ its recession cone. Let $f\in\operatorname{Aff}_{2}(\mathbb{R})$ and let $T$ be the linear part of $f$, that is, $f(\textup{{x}})=T\textup{{x}}+f(0)$ for all $\textup{{x}}\in\mathbb{R}^{d}$. Suppose that $\Omega=f(\Omega)$. Then $T\textup{{x}}\in\operatorname{recc}(\Omega)$ for all $\textup{{x}}\in\operatorname{recc}(\Omega)$.
Proof.
Take any $\textbf{x}\in\operatorname{recc}(\Omega)$ and $\textbf{y}\in\Omega$. By the definition of a recession cone, it suffices to prove that $T\textbf{x}+\textbf{y}\in\Omega$. But indeed, we have
$T\textbf{x}+\textbf{y}=T\textbf{x}+Tf^{-1}(\textbf{y})+f(0)=f(\textbf{x}+f^{-1}(\textbf{y}))\in\Omega$
as required. ∎
The following proposition concerning compactness of the affine automorphism group of a bounded convex domain is itself interesting and useful.
Proposition 4.3.
Let $\Omega$ be a bounded convex domain in $\mathbb{R}^{n}$. Then $\operatorname{Aff}(\Omega)$ is isomorphic to a compact Lie subgroup of $\operatorname{GL}_{n}(\mathbb{R})$.
Proof.
It is readily seen that $G:=\operatorname{Aff}(\Omega)$ is a closed subgroup of $\operatorname{Aff}_{n}(\mathbb{R})$. By Cartan’s theorem on closed subgroups, $G$ is a Lie subgroup of $\operatorname{Aff}_{n}(\mathbb{R})$.
Denote by $\overline{\Omega}$ the closure of $\Omega$. By Heine-Borel theorem and the assumption on $\Omega$, we have that $\overline{\Omega}$ is a convex body. Denote by $E_{\Omega}$ the inner Löwner–John ellipsoid of $\overline{\Omega}$. By the uniqueness of the inner Löwner–John ellipsoid, $E_{\Omega}=g(E_{\Omega})$ for every $g\in G$. By linear algebra, if a non-degenerate affine transform $g$ maps an ellipsoid $E$ onto $E$ itself, then $g$ fix the center of $E$. It follows that all elements of $G$ share a common fixed point, namely the center of the ellipsoid $E_{\Omega}$. Without lose of generality, we may assume that the ellipsoid $E_{\Omega}$ is centered at the origin $0\in\mathbb{R}^{n}$. Therefore $G$ is a subgroup of $\operatorname{GL}_{n}(\mathbb{R})$.
Denote by $\|\cdot\|_{\operatorname{op}}$ the operator norm. Denote by $\ell_{1}$ and $\ell_{2}$ the diameter and the width of $E_{\Omega}$, respectively. Then it is readily seen that $\|g\|_{\operatorname{op}}\leq\ell_{1}/\ell_{2}$ for all $g\in G$. In particular, $G$ is a bounded subset of $\mathbb{R}^{n\times n}$. Now, again by Heine-Borel theorem, $G$ is compact. ∎
We will also use the following classical theorem on crystallographic groups of D. Fried and W. Goldman.
Theorem 4.4 (Fried-Goldman,1983[28]).
Let $G$ be a subgroup of $\operatorname{Aff}(\mathbb{R}^{2})$. Suppose that $G$ acts proper discontinuously on $\mathbb{R}^{2}$ by affine transform. Then $G$ admits a solvable subgroup of finite index.
Proof.
This is a main theorem proved in [28]. ∎
Now we are fully prepared to prove the main theorem of this section.
Theorem 4.5.
Let $((M,\nabla),h)$ be a complete Hessian surface. Then $M$ is diffeomorphic to one of the following smooth manifolds:
  1. 1.
    the plane $\mathbb{R}^{2}$;
  2. 2.
    the cylinder $\mathbb{S}^{1}\times\mathbb{R}$;
  3. 3.
    the Möbius strip;
  4. 4.
    the 2-torus $\mathbb{T}^{2}$;
  5. 5.
    the Klein bottle.
Proof.
Since $((M,\nabla),h)$ is a complete Hessian manifold of dimension 2, by Theorem2.13, there exists a planar convex domain $\Omega$ and a discrete subgroup $G$ of the affine group $\operatorname{Aff}_{2}(\mathbb{R})$, such that $G$ acts freely and proper discontinuously on $\Omega$ by affine transform, and that $M\cong\Omega/G$.
If $G$ is trivial, then $M=\Omega$ is homeomorphic to $\mathbb{R}^{2}$. If $M$ is closed, then by Theorem2.17, we have that $\chi(M)=0$ and hence $M$ is homeomorphic to the 2-torus or the Klein bottle. Recall from low-dimensional topology that the fundamental group of a non-compact surface is always free. From now on, we assume that $M$ is non-compact and $G$ is a non-trivial free group.
Denote by $\operatorname{recc}(\Omega)$ the recession cone of $\Omega$. By elementary Euclidean geometry, any convex cone in $\mathbb{R}^{2}$, and hence in particular $\operatorname{recc}(\Omega)$, is congruent to one of the following sets:
  1. 1.
    a point in $\mathbb{R}^{2}$;
  2. 2.
    a straight line, i.e. an 1-dimensional affine subspace of $\mathbb{R}^{2}$;
  3. 3.
    a wedge, i.e. the set $W_{\theta}:=\{z\in\mathbb{C}^{1}:0\leq\operatorname{Arg}(z)\leq\theta\}$ for $0\leq\theta\leq\pi$;
  4. 4.
    the whole plane $\mathbb{R}^{2}$.
Assume that $\operatorname{recc}(\Omega)$ is a point. Then, by convex geometry, we have that $\Omega$ is bounded. Therefore, by Proposition4.3, $\operatorname{Aff}(\Omega)$ is a compact Lie group. But, by assumption, $G$ is a infinite discrete subgroup of $\operatorname{Aff}(\Omega)$, which leads to a contradiction. Therefore $\operatorname{recc}(\Omega)$ cannot be a point.
Assume that $\operatorname{recc}(\Omega)$ is the whole plane $\mathbb{R}^{2}$. Then we have $\Omega=\operatorname{recc}(\Omega)=\mathbb{R}^{2}$. Therefore $G$ admits a solvable subgroup $G_{0}$ of finite index by Theorem4.4.
Assume that $\operatorname{recc}(\Omega)$ is a straight line in $\mathbb{R}^{2}$ or a wedge $W_{\theta}$ with $0\leq\theta\leq\pi$. Consider the group homomorphism
$\phi\colon\operatorname{Aff}_{2}(\mathbb{R})\rightarrow\operatorname{GL}_{2}(\mathbb{R}),\quad f\mapsto(\textbf{x}\mapsto f(\textbf{x})-f(0))$
that is, the map sending every affine transform to its linear part. Then by Lemma4.2, we have that $\phi(G)$ acts on $\operatorname{recc}(\Omega)$ by linear automorphism. In particular, $\phi(G)$ preserves the boundary of $\operatorname{recc}(\Omega)$. Therefore, $\phi(G)$ admits a subgroup $H$ of index at most 2, such that all elements of $H$ share at least one eigenvector, and hence $G_{0}:=\phi^{-1}(H)\cap G$ is isomorphic to a subgroup of the group $B$ of invertible upper-triangular real $3\times 3$ matrices. Recall from linear algebra that the length of the derived sequence of Heisenberg group $H_{3}(\mathbb{R})$ is exactly $2$, and observe that $B/H_{3}(\mathbb{R})\cong(\mathbb{R}^{\times})^{3}$ is Abelian. By group theory we obtain that $B$ is solvable. Since $B$ is solvable and $G_{0}$ is a subgroup of $G$ of index at most 2, we again have that $G$ admits a solvable subgroup $G_{0}$ of finite index.
This proves that $G$ always admits a solvable subgroup $G_{0}$ of finite index.
Since $G$ is infinite, we have $G_{0}$ ditto. By Nielsen–Schreier theorem, $G_{0}$ is also free. Since any free group of rank $r>1$ is not solvable, the rank of $G_{0}$ is at most 1. Therefore, we conclude that $G_{0}\cong\mathbb{Z}$. Finally, by Schreier’s index formula, we arrive at that in fact the rank of $G$ is also 1, i.e. $G$ is isomorphic to $\mathbb{Z}$.
Therefore, the non-compact surface $M$ has fundamental group $\pi_{1}(M)=\mathbb{Z}$.
By the classification theorem of non-compact surfaces, we have that $M$ is a closed surface $N$ with $k\geq 1$ punctures.
Assume that $N$ is orientable of genus $g$. Then by Seifert-van Kampen theorem, $\pi_{1}(M)$ is a free group of rank r=2g+k1r=2g+k-1. Since $r=1$ and $k>0$, we have that $g=0$ and $k=2$. Therefore $M$ is a sphere with $2$ punctures, i.e. a cylinder.
Assume that $N$ is non-orientable. Then, by the classification theorem of closed surfaces, $N$ is the connected sum of $l$ real projective planes for some integer $l\geq 1$. Again by Seifert-van Kampen theorem, $\pi_{1}(M)$ is a free group of rank r=l+k1r=l+k-1. Since $r=1$ and $k>0$, we obtain $l=k=1$. Therefore $M$ is a punctured $\mathbb{R}P^{2}$, i.e. a Möbius strip.
Quod erat demonstrandum. ∎
Example 4.6.
A complete Hessian surface of constant Gaußian curvature 1-1 can be isometric to a hyperbolic funnel, that is, the quotient of the Lobachevsky plane by an expanding homothety $z\mapsto\lambda\cdot z$ where $\lambda>1$.
In the polar coordinates $(r,\theta)$ of the complex plane, the $1$-form potential of a hyperbolic funnel is precisely
$\eta=\frac{dr}{2r}+\left(\frac{1}{2}\frac{\operatorname{sin}(\theta)}{\operatorname{cos}(\theta)}-\frac{\operatorname{sin}(\theta)}{\operatorname{cos}(\theta)^{3}}\right)d\theta$
defined on an open annulus centered at the origin. A straightforward computation shows that $\eta$ is closed, but not exact.
4.2 Compact Radiant Examples of Koszul Type
The flat tori, or more generally, Bieberbach manifolds are examples of compact Hessian manifolds, but they are trivial in the sense that their affine structures are given by their Levi-Civita connections. In is section, we construct concretely a family of compact radiant Hessian manifolds of Koszul type with negative scalar curvature.
Denote by $\mathbb{H}^{2}:=\{z\in\mathbb{C}:\operatorname{Im}(z)>0\}$ the upper half plane in $\mathbb{C}^{1}$. Consider the product Riemannian metric
$ds^{2}:=d\rho^{2}+\frac{|d\tau|^{2}}{\operatorname{Im}(\tau)^{2}}$
on $\mathbb{H}^{2}\times\mathbb{R}$, where $\tau$ is the coordinate on $\mathbb{H}^{2}$ and $\rho$ is the coordinate on $\mathbb{R}$.
[…]
Consider the quadratic function
$Q\colon\mathbb{R}^{3}\rightarrow\mathbb{R},\quad(x,y,t)\mapsto t^{2}-x^{2}-y^{2}$
and denote $\Omega:=\{(x,y,t)\in\mathbb{R}^{3}:Q(x,y,t)>0\}$. Also, define the smooth function $f\in C^{\infty}(\Omega)$ by
$f(x,y,t):=-\frac{1}{2}\operatorname{log}(t^{2}-x^{2}-y^{2}).$
Then computation yields that the Hessian quadratic form
$2\frac{(tdt-xdx-ydy)^{2}}{(t^{2}-x^{2}-y^{2})^{2}}-\frac{dt^{2}-dx^{2}-dy^{2}}{t^{2}-x^{2}-y^{2}}$
of $f$ is a Riemannian metric on $\Omega$.
Denote by $\nabla$ the Levi-Civita connection of the Euclidean 3-space. Then, by construction, $\nabla(df)$ is a Hessian metric on $(\Omega,\nabla)$.
A tedious but straightforward computation shows that
$\Phi\colon\mathbb{H}^{2}\times\mathbb{R}\rightarrow\Omega,\quad(\tau,\rho)\mapsto\frac{\operatorname{exp}(\rho)}{2\operatorname{Im}(\tau)}(\tau+\bar{\tau},|\tau|^{2}-1,|\tau|^{2}+1)$
is an isometry between $(\mathbb{H}^{2}\times\mathbb{R},ds^{2})$ and $(\Omega,\nabla(df))$.
By the additivity of scalar curvature, we obtain that the scalar curvature of $(\Omega,\nabla(df))$ is constantly 1+0=1-1+0=-1.
Denote by $\tilde{\nabla}:=\Phi^{*}\nabla$ the pullback connection. Then $((\mathbb{H}^{2}\times\mathbb{R},\tilde{\nabla}),ds^{2})$ is a Hessian manifold as $((\Omega,\nabla),\nabla(df))$ is.
From now on, for simplicity, we identify $((\Omega,\nabla),\nabla(df))$ with $((\mathbb{H}^{2}\times\mathbb{R},\tilde{\nabla}),ds^{2})$.
Consider any arbitrary cocompact Fuchsian group action
$(\Gamma,\mathbb{H}^{2})\rightarrow\mathbb{H}^{2},\quad(\gamma,\tau)\mapsto\gamma\cdot\tau$
on $\mathbb{H}^{2}$ so that $\mathbb{H}^{2}/\Gamma=\Sigma_{g}$ is a closed orientable surface of genus $g>1$. Let $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ be a periodic mapping class. By Nielsen’s realization theorem, without lose of generality, we may assume that $\varphi$ lifts to an $\Gamma$-equivariant isometry $\tilde{\varphi}$ of the Lobachevsky plane. Then $G:=\Gamma\rtimes_{\varphi}\mathbb{Z}$ acts on $\mathbb{H}^{2}\times\mathbb{R}$ by isometries via
$G\times(\mathbb{H}^{2}\times\mathbb{R})\rightarrow\mathbb{H}^{2}\times\mathbb{R},\quad((\gamma,n),(\tau,\rho))\mapsto(\gamma\cdot\varphi^{n}(\tau),\rho+n).$
The $G$-action is by construction free and properly discontinuous.
Recall from the theory of Lie groups that the isometry group $\operatorname{PSL}_{2}(\mathbb{R})$ of $\mathbb{H}^{2}$ is isomorphic to the Lorentz group $\operatorname{SO}^{+}(2,1)$, whilst the latter is a Lie subgroup of $\operatorname{GL}_{3}(\mathbb{R})$. It follows that $G$ acts on the convex cone $\Omega$ by linear automorphisms. Therefore, $\tilde{\nabla}$ descends to a flat and torsion-free connection $D$ on the tangent bundle of the mapping torus $M_{\varphi}=(\mathbb{H}^{2}\times\mathbb{R})/G$ of $\varphi$, so that $(M_{\varphi},D)$ is an affine manifold. Moreover, $(M_{\varphi},D)$ is radiant by Proposition3.10.
Since it always holds that
$(|z|^{2}+1)^{2}-(|z|^{2}-1)^{2}-(z+\bar{z})^{2}=4\operatorname{Im}(z)^{2}$
for every $z\in\mathbb{C}$, we obtain that
$Q(\Phi(\gamma\cdot\varphi^{n}(\tau),\rho+n))=Q(\Phi(\tau,\rho))e^{2n}$
and hence
$f(\Phi(\gamma\cdot\varphi^{n}(\tau),\rho+n))=-\frac{1}{2}\operatorname{log}(Q(\Phi(\tau,\rho))\operatorname{exp}(2n))=f(\Phi(\tau,\rho))-n$
for every $(\gamma,n)\in G$ and every $(\tau,\rho)\in\mathbb{H}^{2}\times\mathbb{R}$. Therefore the 1-form potential $df$ of $((\Omega,\nabla),\nabla(df))$ is invariant under the $G$-action, and thus descends to the quotient $M_{\varphi}$. In particular, the product Riemannian metric $ds^{2}$ on $\mathbb{H}^{2}\times\mathbb{R}$ descends to a Riemannian metric $h$ on $M_{\varphi}$.
This proves that $((M_{\varphi},D),h)$ is a compact radiant Hessian manifold of Koszul type with constant negative scalar curvature 1-1, as claimed.
Remark 4.7.
The example above is consistent with the results established earlier. As predicted by Theorem2.41, for a closed orientable surface $\Sigma_{g}$ of genus $g>1$, the mapping torus $M_{\varphi}$ of a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ is indeed a fiber bundle over $\mathbb{S}^{1}$. The fundamental group of $M_{\varphi}$ is a semidirect product of $\mathbb{Z}$ and $\pi_{1}(\Sigma_{g})$, which is indeed infinite and torsion free, as stated in Proposition2.7 and Proposition2.19.
Remark 4.8.
It is obvious that the product of two Hessian manifolds is naturally a Hessian manifold. However, the converse statement is far from being correct, as in the example above, the product of the circle $\mathbb{S}^{1}$ and a closed orientable surface $\Sigma_{g}$ of genus $g>1$ admits an affine structure, but $\Sigma_{g}$ itself does not admit any affine structure if $g\neq 1$ by the proof of Chern’s conjecture in dimension two [20].
Therefore, in particular, the example presented in this section is different to those constructed using Example1.7 and Theorem1.8.
4.3 Exotic Potentials for the Euclidean plane
Hessian potential functions for a metric, when exist, are in general not unique. For instance, besides the half norm square, the flat metric tensor of Euclidean or Minkowski spaces may well be the Hessian quadratic form of other potential functions.
Here, we shall provide a convenient equivalent condition for a smooth function defined on a planar domain to be a Hessian potential of a flat metric.
Lemma 4.9.
Let $\Omega$ be a domain in $\mathbb{R}^{2}$, and let
$\textup{{x}}\colon\Omega\rightarrow\mathbb{R}^{3},\quad(u,v)\mapsto\textup{{x}}(u,v)$
be a parametrization of a regular smooth surface $S$ in $\mathbb{R}^{3}$. Suppose that $S$ does not pass through the origin, and the vectors $\textup{{x}},\textup{{x}}_{u},\textup{{x}}_{v}$ are everywhere $\mathbb{R}$-linearly dependent. Then, there exists a homogeneous function $F$ in three variables such that $F(\textup{{x}})\equiv 0$. Consequently, $S$ lies on a cone of which apex is the origin.
Proof.
This is a well-known fact in classical differential geometry in $\mathbb{R}^{3}$. We nevertheless give a sketch of proof.
Consider the smooth mapping $\phi:=\textbf{x}/\|\textbf{x}\|$. By the assumption on vectors $\textbf{x},\textbf{x}_{u},\textbf{x}_{v}$, a straightforward computation shows that $\phi_{u}\times\phi_{v}\equiv 0$. Therefore, the image $\phi(\Omega)$ is a smooth curve on the $2$-sphere $\mathbb{S}^{2}:=\{(x,y,z)\in\mathbb{R}^{3}|x^{2}+y^{2}+z^{2}=1\}$. By the implicit function theorem, there exists a smooth function $f\colon\mathbb{S}^{2}\rightarrow\mathbb{R}$ such that $\phi(\Omega)$ is locally the zero locus of $f$. Now, define homogeneous function $F$ of degree zero via
$F(x,y,z):=f(x/r,y/r,z/r)$
where $r=\sqrt{x^{2}+y^{2}+z^{2}}$. Then, by the very construction, we have $F(\textbf{x})\equiv 0$. ∎
Proposition 4.10.
Let $f$ be a smooth function defined on a domain in $\mathbb{R}^{2}$ such that $\operatorname{Hess}(f)$ is everywhere non-degenerate, so in particular
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a pseudo-Riemannian metric. Then, $g$ is flat if and only if there exists a homogeneous function $F$ in three variables such that $F(f_{xx},f_{xy},f_{yy})\equiv 0$.
Proof.
Combine Corollary3.37 with Lemma4.9. ∎
Example 4.11.
Let $f$ be a smooth homogeneous function of degree $d$ defined on a domain in $\mathbb{R}^{2}$ such that its Hessian quadratic form $H:=\operatorname{Hess}(f)$ is everywhere non-degenerate. Then, by Euler’s theorem on homogeneous functions, we have
$(2-d)H+x\frac{\partial H}{\partial x}+y\frac{\partial H}{\partial y}=0$
that is, the matrices $H,H_{x},H_{y}$ are everywhere $\mathbb{R}$-linearly dependent. Therefore, by Corollary3.37, we have that
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a flat pseudo-Riemannian metric.
The following interesting example shows that nonlinear harmonic functions are Hessian potentials for the flat $1+1$ Minkowski plane.
Example 4.12.
Let $\Omega$ be a domain in $\mathbb{C}^{1}$, and let $f\in C^{\infty}(\Omega)$ be a nonlinear harmonic function. Then, $\operatorname{tr}(\operatorname{Hess}(f))=f_{xx}+f_{yy}=0$ and
$\operatorname{det}(\operatorname{Hess}(f))=f_{xx}f_{yy}-f_{xy}^{2}=-f_{yy}^{2}-f_{xy}^{2}<0.$
By Cayley–Hamilton theorem, we obtain that $\operatorname{Hess}(f)$ is non-degenerate and indefinite, and hence
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a Lorentzian metric. Moreover, by Proposition4.10, $g$ is flat, i.e. locally the metric of the $1+1$ Minkowski space.
Example 4.13.
Let $\mathbb{H}^{2}:=\{(x,y)\in\mathbb{R}^{2}:y>0\}$ be the upper half plane and define $f\in C^{\infty}(\mathbb{H}^{2})$ by
$f(x,y):=\frac{x^{2}}{2y}+\frac{1}{4}\operatorname{log}(y)y.$
Then the coordinate change $(x(r,\theta),y(r,\theta))=(r^{2}\theta,r^{2})$ brings the Hessian quadratic form
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
of $f$ to the flat metric tensor $dr^{2}+r^{2}d\theta^{2}$ of the Euclidean plane in its polar coordinates. The flatness of $g$ can be proved also by noticing that
$f_{xx}^{2}-4f_{xx}f_{yy}+4f_{xy}^{2}\equiv 0$
and then apply Proposition4.10.
Recall from Example4.11 that, in dimension $2$, all smooth homogeneous functions with non-singular Hessian are potentials for flat metrics. However, a computation in [29] indicates that this is no longer true in higher dimension.
Before we close this section, we cite from [29] an example given by Dubrovin, to show that techniques from algebraic geometry, for example the Clebsch covariant, might be useful for finding homogeneous potentials.
Example 4.14.
Let $f(x,y,z):=x^{6}+y^{6}+z^{6}-10(x^{3}y^{3}+y^{3}z^{3}+z^{3}x^{3})$ be the Maschke sextic. Then the Hessian quadratic form
$g=\sum_{i=1}^{3}\sum_{j=1}^{3}\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}dx_{i}\otimes dx_{j}$
of $f$ is a flat metric on $\mathbb{R}^{3}$.
4.4 Topology of Hessian 4-Manifolds
In this section, we list several properties of Hessian manifolds in dimension four.
For an oriented closed topological 4-manifold $M$, as per usual, we denote by $Q_{M}$ its intersection form.
Lemma 4.15.
Let $((M,\nabla),h)$ be an oriented closed Hessian 4-manifold. Then $\chi(M)=\operatorname{sign}(Q_{M})=0$. In particular, the first Betti number of $M$ is positive and the second Betti number of $M$ is even.
Proof.
The first claim follows from Proposition1.5&2.17 and Hirzebruch’s signature theorem. Moreover, since $\operatorname{dim}M=4$ is even and $\chi(M)=0$, by Poincaré duality the first Betti number of $M$ is positive and the second Betti number of $M$ is even. ∎
Corollary 4.16.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. Then $M$ is null-cobordant, i.e. there exist an orientable smooth 5-manifold $N$ with bundary $\partial N=M$.
Proof.
See [30] for a proof of the fact that an oriented closed smooth 4-manifold is null-cobordant if and only if it has signature zero. ∎
Recall that all closed orientable 4-manifolds have $\operatorname{spin}^{c}$ structures, and therefore their third integral Stiefel-Whitney classes vanish. The following property is specific for Hessian 4-manifolds.
Proposition 4.17.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. If $M$ is spin then it admits an almost complex structure.
Proof.
Recall that $M$ is spin if its second Stiefel–Whitney class $w_{2}(TM)\in\mathrm{H}^{2}(M;\mathbb{Z}/2\mathbb{Z})$ vanishes.
Moreover, by Ehresmann-Wu theorem, $M$ admits an almost complex structure if and only if there exists $c\in\mathrm{H}^{2}(M;\mathbb{Z})$ such that
$3\operatorname{sign}\left(Q_{M}\right)+2\chi(M)=c^{2}$ (3)
and $w_{2}(TM)$ is the $\bmod 2$ reduction of $c$. For our case, by Lemma4.15, $\operatorname{sign}\left(Q_{M}\right)+2\chi(M)=0$. If $w_{2}(TM)=0$, then $c=0$ satisfies equation(3) and its $\bmod 2$ reduction is trivially $w_{2}(TM)$. ∎
Furthermore, the following proposition shows that there are only two possibilities for the intersection form of an oriented closed Hessian 4-manifold, depending on the existence of $\operatorname{spin}$ structure.
Proposition 4.18.
Let $((M,\nabla),h)$ be an oriented closed Hessian 4-manifold with the second Betti number $b_{2}(M)=2m$. If $M$ is spin, then it holds that $Q_{M}=H^{\oplus m}$ where
$H:=Q_{\mathbb{S}^{2}\times\mathbb{S}^{2}}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right];$
if $M$ is not spin, then $Q_{M}=\operatorname{diag}(\underbrace{1,\dots,1}_{m},\underbrace{-1,\dots,-1}_{m})$.
Proof.
Recall that $M$ is spin if its second Stiefel–Whitney class $w_{2}:=w_{2}(TM)$ vanishes. By Wu’s formula, we have $a\smile a=a\smile w_{2}$ for all $a\in\mathrm{H}^{2}(M;\mathbb{Z}/2\mathbb{Z})$. It is then readily seen that $Q_{M}$ is even if and only if $w_{2}=0$. The desired result then follows from Hasse-Minkowski classification theorem of indefinite unimodular symmetric bilinear form over $\mathbb{Z}$, and the fact that $\operatorname{sign}(Q_{M})=0$. ∎
Finally, we characterize closed orientable Einstein-Hessian 4-manifolds.
Definition 4.19.
A Hessian manifold $((M,\nabla),h)$ is said to be a Hessian-Einstein manifold, if $h$ is an Einstein metric on $M$.
Proposition 4.20.
Let $((M,\nabla),h)$ be a closed Hessian-Einstein manifold of dimension four. Then $(M,h)$ is a flat Riemannian manifold, and $\nabla$ is its Levi-Civita connection.
Proof.
By considering the orientation double cover, we may assume without lose of generality that $M$ is orientable.
Now, assume the contrary that the Einstein manifold $(M,h)$ is not flat. Then, since by Lemma4.15
$\chi(M)=0=\frac{3}{2}\left|\operatorname{sign}\left(Q_{M}\right)\right|,$
by Hitchin–Thorpe inequality we have that the universal covering of $M$ is $\pi\colon X\rightarrow M$ where $X$ is a K3 surface. Since K3 surfaces are compact and without lose of generality assume $\pi$ is orientation-preserving, we have that $d:=\operatorname{deg}(\pi)<\infty$ is a positive integer. But as $\chi(X)=d\cdot\chi(M)$ by Riemann-Hurwitz formula, and it is well-known in algebraic geometry that $\chi(X)=24$, we arrive at $\chi(M)>0$, which contradicts Proposition2.17. ∎
Appendix A On Homotopy Tori
Theorem2.22 crucially uses the result in geometric topology that a homotopy torus is in fact homeomorphic to the flat torus, which we will now produce a proof for completeness. We shall first recall the following two propositions from low-dimensional topology.
Proposition A.1.
Let $M$ be a closed $3$-manifold. Suppose that there exists a closed surface $B$ with infinite fundamental group, and a finitely generated normal subgroup $G$ of $\pi_{1}(M)$, such that the following sequence
$1\longrightarrow G\longrightarrow\pi_{1}(M)\longrightarrow\pi_{1}(B)\longrightarrow 1$
is exact. Then $G=\mathbb{Z}$ and $M$ is an $\mathbb{S}^{1}$-bundle over $B$.
Proof.
See [31] for a combinatorial proof of this statement assuming Poincaré’s conjecture, while the latter is confirmed by G.Perelman in [32], [33] and [34] (c.f. [35]). ∎
Proposition A.2.
Let $f\colon M\rightarrow N$ be a homotopy equivalence between two closed aspherical 4-manifolds with polycyclic fundamental group. Then $f$ is homotopic to a homeomorphism.
Proof.
This is proved in [36]. ∎
Now we are ready to prove the main theorem regarding homotopy tori.
Theorem A.3.
Let $M$ be a closed topological manifold. If $M$ is homotopy equivalent to the flat $n$-torus $\mathbb{T}^{n}$, then $M$ is actually homeomorphic to $\mathbb{T}^{n}$.
Proof.
Comparing the top homology groups yields that $M$ is orientable and $\operatorname{dim}M=n$. For $n\leq 2$, it is well-known that closed $n$-manifolds are diffeomorphic if they are of the same homotopy type. For $n\geq 5$, the desired result is proved in [37]. It remains to consider the case when $n=3$ or $4$.
Assume that $n=4$. Since $\mathbb{Z}^{4}$ is polycyclic and the closed 4-manifold $\mathbb{T}^{4}$ is aspherical, by PropositionA.2 we conclude that $M$ is homeomorphic to $\mathbb{T}^{4}$.
Assume that $n=3$. By PropositionA.1, it is readily seen that $M$ is a $\mathbb{S}^{1}$-bundle over $\mathbb{T}^{2}$. But by the theory of 3-manifolds, two $\mathbb{S}^{1}$-bundles over $\mathbb{T}^{2}$ are homeomorphic if and only if their Euler class coincide. Assume that $M$ is the $\mathbb{S}^{1}$-bundles over $\mathbb{T}^{2}$ with Euler class $k\in\mathbb{Z}$. Then $\mathrm{H}_{1}(M;\mathbb{Z})=\mathbb{Z}^{2}\oplus\mathbb{Z}/k\mathbb{Z}$. But by Hurewicz’s theorem, $\mathrm{H}_{1}(M;\mathbb{Z})=\mathbb{Z}^{3}$ as the Abelianization of $\pi_{1}(M)$. Therefore we obtain that $k=0$, and hence $M$ is the trivial $\mathbb{S}^{1}$-bundles over $\mathbb{T}^{2}$, i.e. $M$ is homeomorphic to the 3-torus $\mathbb{T}^{3}$. ∎
Acknowledgement
The author wants to express the author’s most sincere gratitude to Weiyi Zhang for very useful discussions and advices along the preparation of this manuscript, and to Maxwell Stolarski for a careful proofreading. The author also wants to thank Jun Zhang for pointing out a potential connection with Riemann extension, which results in an improvement of section3.3.
The author thanks the reviewers for various suggestions.
This research is completed while the author is studying at the mathematics institute of University of Warwick. The author therefore would like to thank the hospitality of University of Warwick.
Statements and Declarations
No funding was received to assist with the preparation of this manuscript.
The author certify that the author has no affiliations with or involvement in any other organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
The author hereby provide consent for the publication of the manuscript detailed above.