### v4

# On Geometry and Topology of Hessian Manifolds

Abstract
For a differentiable manifold $M$, a pair $(M,\nabla)$ is termed an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. A Riemannian metric $g$ on $M$ is called a Hessian metric on $(M,\nabla)$, if $g$ can be expressed locally as the Hessian quadratic form $\nabla(df)$ of some smooth potential $f$. An affine manifold equipped with a Hessian metric is termed a Hessian manifold. We study the topology and geometry of Hessian manifolds.
Key topological constraints are derived, including the vanishing of the Euler characteristic for any compact Hessian manifold and infinite torsion-freeness of its fundamental group. We demonstrate that compact Hessian manifolds with an abelian fundamental group are homeomorphic to the flat torus. The theory of flat line bundles is developed to introduce the Picard group and the canonical class, which are used to prove that a curved compact orientable Hessian manifold must be a mapping torus. Curvature properties are then discussed. As an application, we prove that all compact Hessian-Einstein manifolds are flat.
A Hessian manifold $((M,\nabla),g)$ is said to be of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $g=\nabla\eta$. We place emphasis on Hessian manifolds of Koszul type, establishing an equivalence between their existence and that of positive flat line bundles, space-like Lagrangian embeddings, and hyperbolic affine structures. A duality with radiant manifolds, analogous to the classical Legendre transformation, is explored. We also develop a geometric uncertainty principle on Hessian manifolds of Koszul type, and classify all complete Hessian surfaces and closed Hessian $3$-folds.
keywords:
Hessian metric, Affine differential geometry, regular convex domain, Flat line bundle, Legendre transform, Lagrangian embedding, para-Kähler manifold.
1 Basics on Hessian Geometry
1.1 Introduction and Background
The study of Hessian manifolds originates with the foundational work of J.-L. Koszul in 1961 [1]. While their roots lie in affine differential geometry, in recent decades, these geometric structures have garnered significant attention and emerged as indispensable tools across a diverse spectrum of mathematical and scientific disciplines. Their influence spans from the theoretical underpinnings of information geometry and complex analysis to practical applications in optimization and mathematical statistics, and even touches upon the abstract realms of homological mirror symmetry. This resurgence of interest highlights the deep connections these manifolds forge between seemingly disparate areas, offering a unifying geometric language to explore sophisticated phenomena.
In affine differential geometry and information geometry, Hessian metrics underpin the dually flat spaces used in the study of statistical manifolds, where the Amari–Chentsov tensor codifies divergences and dual connections. Hessian metrics also arise naturally in convex optimization, where barrier functions induce complete Hessian metrics on convex domains. For more details, we refer to [2] and [3].
Also, regarding the applications to mirror symmetry, the seminal work of Hitchin [4], Gross–Siebert [5], and Kontsevich–Soibelman [6] has shown that affine manifolds with singularities and Hessian metrics model the base of torus fibrations in the Strominger-Yau-Zaslow conjecture.
On the other hand, Cheng and Yau’s solution [7] to the real Monge–Ampère equation constructs complete affine hyperspheres of which Blaschke metrics are precisely global Hessian metrics with prescribed volume forms. Remarkably, these examples are linked to convex projective manifolds via the work of Loftin [8], Benoist–Hulin [9], and that of Danciger–Guéritaud–Kassel [10].
However, despite this rich tapestry, fundamental questions about Hessian manifolds remain open, particularly regarding the interplay between global topology, curvature properties, and algebraic obstructions. In this manuscript, we undertake a study of Hessian manifolds from both topological and geometric perspectives:
After briefly summarizing the history and background of the theory of Hessian manifolds, we recall the definitions of affine atlases and flat connections before formally defining Hessian metrics. While the literature contains several well-known examples of Hessian metrics from information theory, such as the canonical metrics on the probability simplexes [11] and the family of Gaussian distributions [11], this collection has remained quite limited and tied to specific contexts. To address this, we introduce a powerful and explicit warped product formula (Theorem 1.8), a systematic construction that generates a vast new class of Hessian manifolds and demonstrates that structures like the Robertson-Walker solutions from general relativity are part of this geometric framework, significantly expanding the known examples beyond their original domains.
Our investigation then begins by establishing topological constraints that provide a rigid global framework for any Hessian structure. Foundational results by Shima and Yagi [12] in 1997 established that complete Hessian manifolds are aspherical, as their universal coverings are convex domains. Building on this, and on partial results by Kostant–Sullivan [13] in 1975 and Fried [14] in 1982, our work confirms Chern’s longstanding 1955 conjecture in this general setting by proving that the Euler characteristic of any compact Hessian manifold must vanish (Proposition 2.17). We also provide the first proof that the fundamental group of a complete Hessian manifold is torsion-free (Proposition 2.19), and give two different proofs of the fact that the fundamental group of a compact Hessian manifold is infinite (Corollary 2.7 and Remark 2.21). These new algebraic results possess immediate classificatory power, allowing us to prove that any compact Hessian manifold with an abelian fundamental group is homeomorphic to a $n$-torus (Theorem 2.22).
Having established this broad landscape, we switch from these foundational constraints to a more refined analysis by developing the theory of flat line bundles as our primary analytical tool. We introduce this as a novel framework for studying Hessian geometry and define two crucial new objects within it: the canonical class, an intrinsic invariant of the affine structure, and the locally constant Picard group. The latter classifies all flat $\mathbb{R}$-line bundles over a differentiable manifold up to gauge equivalence. As a direct consequence, we also establish that any orientable compact Hessian manifold with vanishing first Betti number must be flat (Theorem 2.45).
We then continue to study the natural question that in which case the Hessian metric admits a globally defined potential function. Using sheaf cohomology and spectral sequences technique, we show that a Hessian manifold admits a single global potential function if its fundamental group is finite (Theorem 2.4). Moreover, it is readily seen that a compact Hessian manifold cannot support a global potential function. This problem brings a historically significant class of Hessian manifolds into sharp focus: those of Koszul type, which admit a global 1-form potential. For this class, we prove using Hodge theory that the first Betti number of a Hessian manifold of Koszul type is always positive (Proposition 2.9).
A central achievement of our work is to reveal that Hessian metrics of Koszul type are part of a much larger network of connections than previously understood. Whilst the work of H.Shima [15] established an important equivalence between the existence of a Koszul-type Hessian metric and that of a hyperbolic affine structure, we establish a new, powerful list of equivalences, demonstrating that these conditions are also entirely equivalent to two other seemingly disparate properties: the existence of a positive flat line bundle (Proposition 2.50) and admitting a space-like Lagrangian embedding (Corollary 3.32) into its own cotangent bundle, equipped with the Riemann extension metric first studied by Patterson and Walker in 1952 [16], thereby building a new bridge between affine differential geometry and para-Kähler geometry.
To get a finer understanding of the moduli of hyperbolic affine structures on an oriented Hessian manifold of Koszul type, we define canonical class as the de-Rham cohomology class corresponding to the canonical line bundle. In 1968, J.-L.Koszul proved that the hyperbolicity is stable under small perturbations of the affine structure [17]. Using group cohomology with Lie algebra coefficients, we show that any hyperbolic affine structure can be deformed into another of which canonical class is rational (Theorem 3.50).
This cohesive picture is further enriched by developing a duality with radiant affine manifolds, i.e. differentiable manifolds of which transition functions are linear, which builds upon the classical theory of dual connections in information geometry pioneered by Amari and Chentsov [2]. As a new contribution, we provide an explicit expression for the $1$-form potential of the dual of a radiant Hessian manifold of Koszul type, resembling the classical Legendre transformation in theoretical mechanics, and use this to situate the celebrated Cheng–Yau metric [18] within our framework.
Next, in a digression section, we also develop a new link to pseudoconvexity in complex analysis of several variables by showing that the Sasaki lifting of a complete Hessian manifold is a Stein manifold (Proposition 3.5), and correspondingly, positive flat line bundles are lifted to positive holomorphic line bundles (Lemma 3.2). A further observation then yields a stronger real analytic embedding theorem for this class of manifolds (Theorem 3.6) of which dimension bound is more optimal than the classical Whitney embedding.
The synthesis of these tools now culminates in powerful structural theorems. By applying D.Tischler’s classical 1970 theorem on fibrations [19] to the first Koszul form, we prove that any curved compact orientable Hessian manifold must be a mapping torus (Theorem 2.48), revealing a remarkably rigid underlying structure, particularly in low dimensions (Corollary 5.15). This result is further strengthened by a de Rham-type splitting theorem, which shows that the universal cover of a compact Hessian manifold isometrically splits as a product of another contractible Riemannian manifold and the real line (Proposition 4.8). Using this splitting structure and tools from geometric analysis, we are able to prove that a compact Hessian manifold cannot support a metric of negative curvature(Theorem 4.10). As yet another important application, we then prove that all compact Hessian-Einstein manifolds are flat (Theorem 4.13). An interesting geometric uncertainty principle is also established on Hessian manifolds, revealing links to quantum mechanics.
After obtaining a number of theorems, we recall that constructing explicit examples of non-trivial curved compact Hessian manifolds has been a significant challenge, with most known examples being flat Bieberbach manifolds. We therefore address this by explicitly constructing a non-trivial family of compact radiant Hessian manifolds of Koszul type that possess negative scalar curvature 1-1, suggesting that our theory is not vacuous. Furthermore, this careful investigation in low dimension finally results in the first complete topological classification of closed orientable Hessian 3-manifolds(Theorem 5.15), namely, they are always mapping tori of periodic mapping classes.
Finally, as one of the highlights of this article, we adopt our general theory to low dimensions. Applying classical algebraic topology techniques and using the solvability of Heisenberg group $H_{3}(\mathbb{R})$, we provide the first complete topological classification of all complete Hessian surfaces (Theorem 5.11). We also derive severe constraints on the intersection form for Hessian $4$-manifolds, and then prove that any Einstein metric on a closed Hessian $4$-manifold must be flat (Proposition 5.22). Last but not least, a novel characterization for potentials of $2$-dimensional flat metrics is established, showing that a pseudo-Riemannian metric
$g=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is flat if and only if the vector $(f_{xx},f_{xy},f_{yy})$ of second derivatives lies on a cone surface in the Euclidean space $\mathbb{R}^{3}$ (Proposition 5.2).
Before closing this section, we provide a concise road map of the present article.
This paper is organized as follows: Section 1 reviews fundamental definitions and introduces a warped product construction as first examples. Section 2 establishes key topological obstructions, such as the vanishing of the Euler characteristic and the infinitude of the fundamental group. We then develop the theory of flat line bundles, proving that curved compact orientable Hessian manifolds must be mapping tori. Section 3 explores the interaction with other geometric fields, establishing a central equivalence between Hessian metrics of Koszul type and several other structures, including positive line bundles and space-like Lagrangian embeddings. A duality via a generalized Legendre transform and a link to complex analysis are also discussed. Section 4 delves into the geometric analysis of Hessian metrics, studying their curvature tensors and proving that a compact Hessian manifold cannot support a metric of negative curvature, and that all compact Hessian-Einstein manifolds are flat. Finally, Section 5 applies these results to low dimensions, yielding a complete topological classification of complete Hessian surfaces and closed orientable Hessian 3-folds, and deriving strong constraints for Hessian 4-manifolds. An appendix on homotopy tori is included to support our classification of compact Hessian manifolds with Abelian fundamental groups, while another appendix is devoted to the geometric analysis techniques used in our study of curvature properties.
1.2 Affine and Hessian Manifolds
Throughout this article, all smooth manifolds are tacitly assumed to be connected and without boundary.
For each $i\in\{1,\ldots,n\}$, denote by $\pi^{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ the projection onto the $i$-th coordinate.
Let $M$ be an $n$-dimensional differentiable manifold. As usual, for any chart $(U,\varphi)$ of $M$, the smooth functions $x^{1},\ldots,x^{n}\in C^{\infty}(U)$ defined by
$x^{i}:U\longrightarrow\mathbb{R},\quad p\longmapsto\pi^{i}(\varphi(p))$
are said to be the local coordinates associated to the chart $(U,\varphi)$. An atlas
$\mathscr{A}=\left\{\left(U_{\alpha},\varphi_{\alpha}\right)\mid\alpha\in I\right\}$
of $M$ is said to be an affine atlas of $M$, if for any $\alpha,\beta\in I$ there exist $A\in\operatorname{GL}_{n}(\mathbb{R})$ and $\bf{a}\in$ $\mathbb{R}^{n}$ such that
$\varphi_{\beta}\left(\varphi_{\alpha}^{-1}(\bf{x})\right)=A{\bf{x}}+{\bf{a}}$
for all ${\bf{x}}\in\varphi_{\alpha}\left(U_{\alpha}\cap U_{\beta}\right)$. An element of an affine atlas of $M$ is then called an affine chart of $M$, and the local coordinates associated to an affine chart of $M$ are termed affine coordinates.
For any two affine atlases $\mathscr{A}_{1}$ and $\mathscr{A}_{2}$ of $M$, we write $\mathscr{A}_{1}\sim\mathscr{A}_{2}$ if and only if $\mathscr{A}_{1}\cup\mathscr{A}_{2}$ is also an affine atlas of $M$. An affine atlas of $M$ is said to be maximal, if for any other affine atlas $\mathscr{A}^{\prime}$ of $M$ we have that $\mathscr{A}^{\prime}\sim\mathscr{A}$ implies $\mathscr{A}^{\prime}\subseteq\mathscr{A}$. Traditionally, a maximal affine atlas on $M$ is also termed an affine structure of the smooth manifold $M$.
Recall that a connection on a vector bundle over $M$ is said to be flat, if its curvature 2-form vanishes identically.
Consider an arbitrary connection $D$ on the tangent bundle $TM\rightarrow M$. By Cartan-Ambrose-Hicks theorem, if $D$ is flat and torsion-free then there exists a maximal affine atlas $\mathscr{A}_{D}:=\left\{\left(U_{\alpha},\varphi_{\alpha}\right)\mid\alpha\in I\right\}$ of $M$, such that
$D(dx_{\alpha}^{1})=\cdots=D(dx_{\alpha}^{n})=0$
for all $\alpha\in I$, where $x_{\alpha}^{1},\ldots,x_{\alpha}^{n}$ are the local coordinates associated to the chart $\left(U_{\alpha},\varphi_{\alpha}\right)$. Moreover, the map $D\longmapsto\mathscr{A}_{D}$ is a bijection from the set of all flat and torsion-free connections on the tangent bundle $TM\rightarrow M$ to the collection of all maximal affine atlases of $M$, and the atlas $\mathscr{A}_{D}$ is termed the adapted affine atlas of $D$. Therefore, fixing an affine structure on $M$ is equivalent to specifying a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. In view of this fact, by a slightly abuse of notation we also refer to a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$ as an affine structure of $M$, and refer to a pair $(M,\nabla)$ as an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$.
Remark 1.1.
All affine manifolds are in particular analytic, as their transition functions are by definition real analytic mappings.
On an affine manifold, there exist two important sheaves of Abelian groups:
Definition 1.2.
Let $(M,\nabla)$ be an affine manifold. The sheaf $\mathcal{A}:=\mathcal{A}_{M}$ of affine functions on $(M,\nabla)$ is defined by
$\mathcal{A}(U):=\{u\in C^{\infty}(U)\mid\nabla(du)=0\}$
where $U$ is any open subset of $M$. Sections of $\mathcal{A}$ are termed affine functions.
Definition 1.3.
Let $(M,\nabla)$ be an affine manifold. We denote by $\mathcal{P}:=\mathcal{P}_{M}$ the sheaf of parallel 1-forms on $(M,\nabla)$.
Recall from differential geometry that, on an $n$-dimensional affine manifold, the sheaf of parallel 1-forms is locally constant of rank $n$.
Moreover, since $\nabla$ is always assumed to be torsion-free, the exterior differential $d\omega$ of a $1$-form $\omega$ is the anti-symmetrization of the covariant derivative $\nabla\omega$. Therefore, in particular, on an affine manifold, all parallel 1-forms are closed.
We shall also recall the following standard definition of Hessian manifolds:
Definition 1.4.
Let $(M,\nabla)$ be an affine manifold. A Riemannian metric $h$ on $M$ is called a Hessian metric on $(M,\nabla)$, if for any point $p\in M$, there exists an open neighborhood $U$ of $p$ in $M$, such that $h|_{U}=\nabla(df)$ for some $f\in C^{\infty}(U)$. A pair $((M,\nabla),h)$ is termed a Hessian manifold, if $h$ is a Hessian metric on the affine manifold $(M,\nabla)$.
As Hessian manifolds are affine, we first present a theorem on topological invariants of affine manifolds, which of course applicable to all Hessian manifolds.
Proposition 1.5.
All Pontryagin classes and rational Chern classes of an affine manifold must vanish.
Proof.
Note that the tangent bundle of an affine manifold admits a connection of which curvature vanishes identically. The proof of the desired result is then a direct application of the Chern–Weil construction. ∎
It is also worth noting the following famous conjecture of Chern.
Conjecture 1.6 (Chern,1955).
The Euler characteristic of a compact affine manifold is zero.
Chern’s conjecture is known to hold in various cases. For example, the 2-dimensional case is settled by John Milnor [20] in 1957. Moreover, in [21], Bruno Klingler proved that if a compact affine manifold admits a parallel nowhere vanishing top differential form then its Euler characteristic is zero. In latter sections we will also confirm this conjecture for all compact Hessian manifolds.
1.3 Warped Product Construction
In this section, we shall show that any warped product over the real line $\mathbb{R}$ with Hessian fibers is again Hessian, which allow us to construct new Hessian manifolds from known ones. In particular, we obtain that hyperbolic spaces and sufficiently small domains in round $n$-spheres are Hessian.
The following trivial observation serves as a base case in the recursive constructions of Hessian metrics.
Example 1.7.
Let $(M,g)$ be an one dimensional Riemannian manifold, and let $\nabla$ be an arbitrary connection on the tangent bundle $TM\rightarrow M$. Then, since $\Omega^{2}(M)=0$, the connection $\nabla$ is automatically flat and torsion-free. Moreover, by the theory of ordinary differential equations, we have that $g$ is always a Hessian metric on $(M,\nabla)$.
Now we shall state the fundamental construction theorem of Hessian warped products.
Theorem 1.8.
Let $\Omega$ be a domain in $\mathbb{R}^{n}$, and let $\varphi=\varphi(x_{1},\dots,x_{n})\in C^{\infty}(\Omega)$. Let $f=f(t)\in C^{\infty}(\mathbb{R})$ be a smooth function with nowhere vanishing first derivative, so that in particular
  1. 1.
    $f$ admits an inverse $F\colon f(\mathbb{R})\rightarrow\mathbb{R}$, i.e. the smooth function satisfying $f\circ F=id$;
  2. 2.
    $y_{1}:=\operatorname{exp}(f)x_{1},\,\dots\,,y_{n}:=\operatorname{exp}(f)x_{n},\,y_{n+1}\equiv y:=\operatorname{exp}(f)$ form a coordinate system of an open domain in the upper half space $\{(y_{1},\dots,y_{n},y)\in\mathbb{R}^{n+1}:y>0\}$.
Then, the smooth function
$\hat{\varphi}(y_{1},\dots,y_{n},y):=\varphi(y_{1}/y,\dots,y_{n}/y)y+\int_{0}^{y}F^{\prime}(\operatorname{log}(t))^{2}(y-t)t^{-2}dt$
is well-defined, and it holds that
$\sum_{i=1}^{n+1}\sum_{j=1}^{n+1}\frac{\partial^{2}\hat{\varphi}}{\partial y_{i}\partial y_{j}}dy_{i}\otimes dy_{j}=dt^{2}+\operatorname{exp}(f)\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}dx_{i}\otimes dx_{j}.$
Proof.
The proof of this theorem is a straightforward computation and hence omitted. ∎
Combining Example 1.7 and Theorem 1.8, one may conclude that many important metric tensors in geometry and theoretical physics are actually Hessian. Examples include but not limited to sufficiently small geodesic disks in space forms, and the Robertson–Walker solutions to Einstein’s field equation.
The following is just one particular example of many corollaries of the powerful theorem above.
Example 1.9.
On the upper half space $\mathbb{H}^{n}:=\{(x_{1},\dots,x_{n})\in\mathbb{R}^{n}:x_{n}>0\}$, consider the coordinates
$y_{1}:=\frac{2x_{1}}{x_{n}^{2}},\,\dots,\,y_{n-1}:=\frac{2x_{n-1}}{x_{n}^{2}},\,y_{n}:=\frac{1}{x_{n}^{2}}.$
Let $\nabla$ be the affine structure on $\mathbb{H}^{n}$ satisfying $\nabla(dy_{i})=0$ for all $i=1,\dots,n$. Define smooth function
$\varphi(y_{1},\dots,y_{n}):=\frac{y_{1}^{2}+\dots+y_{n-1}^{2}}{8y_{n}}-\frac{1}{4}\operatorname{log}(y_{n})$
on $\mathbb{H}^{n}$. Then the Hessian quadratic form of $f$ is precisely
$h:=\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}dy_{i}\otimes dy_{j}=\frac{dx_{1}^{2}+\dots+dx_{n}^{2}}{x_{n}^{2}}.$
In particular, $((\mathbb{H}^{n},\nabla),h)$ is a Hessian manifold.
2 Global Invariants for Hessian manifolds
2.1 Obstructions from Sheaf Cohomology
In this section, we study the very natural question that in which case there exists a global potential $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. Before we can give an answer to this question, we need to state and prove the following two lemmata from algebraic topology and sheaf theory.
Lemma 2.1.
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. If $M$ is simply connected, then the first sheaf cohomology group $\mathrm{H}^{1}(M;\mathcal{A})$ vanishes.
Proof.
Since the holonomy of $(M,\nabla)$ is trivial and the sheaf $\mathcal{P}$ of parallel 1-forms on $M$ is a locally constant sheaf of rank $n$, by abuse of notation we identify $\mathcal{P}$ with $\underline{\mathbb{R}}^{n}$. It is then readily seen that there exists a short exact sequence
$0\longrightarrow\mathbb{R}\longrightarrow\mathcal{A}\overset{d}{\longrightarrow}\mathbb{R}^{\oplus n}=\mathcal{P}\longrightarrow 0$
of sheaves, which induces a long exact sequence
$\cdots\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})\longrightarrow\mathrm{H}^{1}(M;\mathcal{A})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})^{\oplus n}\longrightarrow\cdots$
of cohomology groups. But since $M$ is simply connected, we have that the first de-Rham cohomology group of $M$ vanishes. Therefore, we obtain $\mathrm{H}^{1}(M;\mathcal{A})=0$. ∎
Lemma 2.2.
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. Suppose that $\pi\colon\tilde{M}\rightarrow M$ is a universal covering. Then $\pi^{-1}\mathcal{A}$ is the sheaf of affine functions on the affine manifold $(\tilde{M},\pi^{*}\nabla)$, and there exists a natural isomorphism
$\mathrm{H}^{1}(M;\mathcal{A})\cong\mathrm{H}^{1}(G;\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))$
of $\mathbb{R}$-vector spaces, where $G:=\pi_{1}(M)$ is the fundamental group of $M$.
Proof.
It is routine to check that $(\tilde{M},\pi^{*}\nabla)$ is an affine manifold and $\pi^{-1}\mathcal{A}$ is its sheaf of affine functions.
To prove the existence of the natural isomorphism, we use Cartan-Leray spectral sequence, of which $E_{2}$-page terms are
$E_{2}^{p,q}=\mathrm{H}^{p}(G,\mathrm{H}^{q}(\tilde{M};\pi^{-1}\mathcal{A}))\implies\mathrm{H}^{p+q}(M;\mathcal{A}).$
Recall from homological algebra that the following sequence
$0\longrightarrow E^{1,0}_{2}\longrightarrow\mathrm{H}^{1}(M;\mathcal{A})\longrightarrow E^{0,1}_{2}\longrightarrow E^{2,0}_{2}\longrightarrow\cdots$
is exact. Since $\tilde{M}$ is simply connected, by Lemma 2.1 $\mathrm{H}^{1}(\tilde{M};\pi^{-1}\mathcal{A})=0$. Therefore we have that
$E^{0,1}_{2}=\mathrm{H}^{0}(G;0)=0.$
This proves that
$\mathrm{H}^{1}(M;\mathcal{A})=E^{1,0}_{2}=\mathrm{H}^{1}(G,\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))$
as required. ∎
The following vanishing theorem of sheaf cohomology group turns out to be useful for our purpose:
Proposition 2.3.
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. If the fundamental group of $M$ is finite, then the first sheaf cohomology group $\mathrm{H}^{1}(M;\mathcal{A})$ vanishes.
Proof.
Take any universal covering $\pi\colon\tilde{M}\rightarrow M$. Then by Lemma 2.2
$\mathrm{H}^{1}(M;\mathcal{A})\cong\mathrm{H}^{1}(G;\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))$
as $\mathbb{R}$-vector spaces, where $G:=\pi_{1}(M)$ is the fundamental group of $M$. But, recall from the theory of group cohomology that, the first cohomology of a finite group is a torsion group. We therefore conclude that $\mathrm{H}^{1}(G;\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))=0$. ∎
We are now ready to state and prove the main theorem in this section.
Theorem 2.4.
Let $((M,\nabla),h)$ be a Hessian manifold. If the fundamental group of $M$ is finite, then there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$.
Proof.
By the definition of a Hessian metric, there exists an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a smooth function $f_{\alpha}\in C^{\infty}(U_{\alpha})$, such that for each $\alpha\in I$ it holds that $h_{\alpha}=\nabla(df_{\alpha})$, where $h_{\alpha}$ is the restriction of $h$ on $U_{\alpha}$.
For every $\alpha,\beta\in I$, define $g_{\alpha\beta}:=f_{\alpha}-f_{\beta}\in C^{\infty}(U_{\alpha}\cap U_{\beta})$. Then $g=\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ is a Čech 1-cocycle with coefficients in the sheaf $\mathcal{A}$ of affine functions on $M$. It remains to prove that the Čech cohomology class $g]\in\check{\mathrm{H}}^{1}(M;\mathcal{A})$ vanishes. But this immediately follows from Proposition 2.3. ∎
Example 2.5.
There does exist Hessian manifold with finite but non-trivial fundamental group. The simplest example is the punctured affine space $\mathbb{R}^{3}-\{0\}$ quotient by the antipodal map $\textup{{x}}\mapsto-\textup{{x}}$, of which fundamental group is precisely $\mathbb{Z}/2\mathbb{Z}$.
Before we proceed to the next section, also notice that we have the following obvious example for the non-existence of global potential for a pseudo-Riemannian metric.
Proposition 2.6.
Let $(M,g)$ be a compact pseudo-Riemannian manifold, and let $\nabla$ be a connection on the cotangent bundle $T^{*}M\rightarrow M$. Then $g\neq\nabla(df)$ for all $f\in C^{\infty}(M)$.
Proof.
Take any $f\in C^{\infty}(M)$ such that the Hessian quadratic form $\nabla(df)$ of $f$ is everywhere non-degenerate. In particular $f$ is not a constant function. Since $M$ is compact, by the extreme value theorem, there exists distinct points $p,q\in M$ such that $f$ attains a global maximum at $p$ and attains a global minimum at $q$. But then by the second derivative test, the signatures of the Hessian quadratic form of $f$ must be different at points $p$ and $q$. ∎
Proposition 2.6 leads immediately to an interesting result:
Corollary 2.7.
The fundamental group of a compact Hessian manifold is infinite.
Proof.
Let $h$ be a Hessian metric on affine manifold $(M,\nabla)$. Assume the contrary that $\pi_{1}(M)$ is a finite group. Then by Proposition 2.4, there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. But this contradicts Proposition 2.6 as $M$ is assumed to be compact. ∎
Remark 2.8.
In [22], it is proved that there exists a unique closed orientable flat Riemannian 3-manifold with vanishing first Betti number, called the Hantzsche–Wendt manifold. Therefore, in particular, although by Corollary 2.7 a compact Hessian manifold admits an infinite fundamental group, its first Betti number can still be zero.
However, in contrast, we have the following proposition:
Proposition 2.9.
Let $(M,\nabla)$ be an affine manifold, and let $h$ be a Riemannian metric on $M$. Suppose that there exists $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$. Then $\eta$ is a closed 1-form. Moreover, if it is assumed further that $M$ is compact, then $\eta$ is not exact, and in particular the first Betti number of $M$ is positive.
Proof.
Since $\nabla$ is torsion-free, the exterior differential $d\eta$ is the anti-symmetrization of $\nabla\eta$. But the anti-symmetrization of $\nabla\eta$ is identically zero as $\nabla\eta=h$ is symmetric. Therefore, $\eta$ is a closed 1-form.
Now, assume that $M$ is compact. By Hodge decomposition theorem, there exists a smooth function $f\in C^{\infty}(M)$ and a harmonic 1-form $\theta$ such that $\eta=df+\theta$. Since
$\nabla\theta=\nabla\eta-\nabla(df)=h-\nabla(df)\neq 0$
by Proposition 2.6, we conclude that $\theta\neq 0$. Now it suffices to recall from Hodge theory that $\mathrm{H}^{1}(M;\mathbb{R)}$ is isomorphic to the space of the harmonic 1-forms on $M$. ∎
Proposition 2.9 motivates the following definition.
Definition 2.10.
Let $(M,\nabla)$ be an affine manifold, and let $h$ be a Riemannian metric on $M$. The pair $((M,\nabla),h)$ is termed a Hessian manifold of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$, and in this case, the 1-form $\eta$ is termed a $1$-form potential for $((M,\nabla),h)$.
By Poincaré’s lemma, Hessian manifolds of Koszul type are indeed Hessian. As explained in the introduction section, this is, actually, the ultimate origin of the notion of Hessian structures.
2.2 Fundamental Group and Euler Number
For each $n\in\mathbb{N}$, we denote by $\operatorname{Aff}_{n}(\mathbb{R})$ the group of affine transformations on $\mathbb{R}^{n}$.
Definition 2.11.
A Hessian manifold $((M,\nabla),h)$ is said to be complete, if $h$ is a complete Riemannian metric on $M$.
First, we recall the definition of affine mappings: For affine manifolds $(M_{1},\nabla_{1})$ and $(M_{2},\nabla_{2})$, a smooth mapping $\Phi:M_{1}\rightarrow M_{2}$ is said to be affine if $\hat{\nabla}(d\Phi)=0$, where
$\hat{\nabla}=\nabla_{1}\otimes\operatorname{Id}+\operatorname{Id}\otimes\Phi^{*}\nabla_{2}$
is the induced connection on the tensor bundle $T^{*}M_{1}\otimes\Phi^{*}TM_{2}$.
Let $(M,\nabla)$ be an affine manifold. It is readily seen that, a vector-valued smooth function
$F\colon M\longrightarrow\mathbb{R}^{m},\quad p\longmapsto(F_{1}(p),\dots,F_{m}(p))$
is an affine mapping between $(M,\nabla)$ and the flat Euclidean space, if and only if $F_{1},\dots,F_{m}$ are global affine functions on $M$.
For an affine manifold $(M,\nabla)$, since $\nabla$ is torsion-free, the exterior differential $d\omega$ of a $1$-form $\omega$ is the anti-symmetrization of the covariant derivative $\nabla\omega$. Therefore, on an affine manifold, all parallel 1-forms are closed.
Moreover, for an affine manifold $(M,\nabla)$ with finite fundamental group, parallel 1-forms admit no periods as the first de-Rham cohomology group of $M$ vanishes. Therefore, the following notion of affine development is well-defined:
Definition 2.12.
Let $(M,\nabla)$ be a simply connected $n$-dimensional affine manifold, and fix a reference point $p_{0}\in M$. The affine development of $(M,\nabla)$ is the smooth mapping $\Phi\colon M\rightarrow\mathcal{P}(M)^{\vee}=\mathbb{R}^{n}$ defined by
$\langle\omega,\Phi(p)\rangle=\int_{\gamma}\omega$
for $\omega\in\mathcal{P}(M)$ and $p\in M$, where the bracket denotes the canonical pairing between $\mathcal{P}(M)$ and its dual space $\mathcal{P}(M)^{\vee}$, and where $\gamma\colon[0,1]\rightarrow M$ is a smooth curve such that $\gamma(0)=p_{0}$ and $\gamma(1)=p$.
See for example [15] for a proof of the fact that an affine development is indeed an affine mapping from an affine manifold to the flat Euclidean space.
In [12], using the concept of affine development, H.Shima and K.Yagi proved the following important theorem on convexity of complete Hessian metrics.
Theorem 2.13 (Shima-Yagi, 1997[12]).
Let $((M,\nabla),h)$ be a simply connected complete Hessian manifold. Then the affine development $\Phi\colon M\rightarrow\mathbb{R}^{n}$ of $(M,\nabla)$ is a diffeomorphism onto its image such that $\Phi(M)$ is a convex domain in $\mathbb{R}^{n}$.
Theorem 2.13 provides a strong restriction on the topology of complete Hessian manifolds.
Corollary 2.14.
Let $((M,\nabla),h)$ be a complete Hessian manifold. Then $M$ is aspherical, i.e. the homotopy group $\pi_{k}(M)$ vanishes for all $k\geq 2$.
Proof.
Recall that being an aspherical space is equivalent to possessing a contractible universal cover. The desired result then follows immediately from Theorem 2.13. ∎
In [14] and [13], Kostant et al. and Fried verified Conjecture 1.6 in some special cases.
Theorem 2.15 (Kostant-Sullivan,1975[13]).
Let $(M,\nabla)$ be a compact affine manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. If $\nabla$ is geodesically complete, or equivalently the affine development $\Phi\colon\tilde{M}\rightarrow\mathbb{R}^{n}$ of $(\tilde{M},\pi^{*}\nabla)$ is a diffeomorphism, then the Euler characteristic $\chi(M)=0$.
Theorem 2.16 (Fried,1982[14]).
[…]
Combine Theorem 2.13 and Theorem 2.152.16, we immediately get the following result, confirming Conjecture 1.6 for Hessian manifolds.
Proposition 2.17.
The Euler characteristic of a compact Hessian manifold is zero.
Proof.
[…]
Corollary 2.7 shows that compact Hessian manifolds has infinite fundamental group. We shall now improve the result in Corollary 2.7.
Lemma 2.18.
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$, and let $G$ be a group acting freely on $\Omega$. Then $G$ is torsion-free, that is, it contains no non-trivial element of finite order.
Proof.
Suppose that $g\in G$ enjoys finite order $k\in\mathbb{N}$. Take any arbitrary point $\textbf{x}\in\Omega$. Since $\Omega$ is convex and $g^{1}(\textbf{x}),\dots,g^{k}(\textbf{x})\in\Omega$, the barycenter
$\textbf{b}:=\frac{1}{k}\sum_{i=0}^{k-1}g^{i}(\textbf{x})$
is also a point in $\Omega$. But by the very construction we have that $g(\textbf{b})=\textbf{b}$. Therefore $g$ is the identity element of $G$, as the $G$-action is by assumption free. ∎
The following results are direct consequences of Lemma 2.18:
Proposition 2.19.
The fundamental group a complete Hessian manifold is torsion-free.
Proof.
By Theorem 2.13, the fundamental group of a complete Hessian manifold acts freely on a convex domain. The desired result then follows from Lemma 2.18. ∎
Corollary 2.20.
The fundamental group a complete Hessian manifold is either trivial or infinite.
Proof.
This follows from Proposition 2.19. ∎
Remark 2.21.
Let $((M,\nabla),h)$ be a compact Hessian manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. Then $((\tilde{M},\pi^{*}\nabla),\pi^{*}h)$ is a complete Hessian manifold. Provided that the fundamental group of $M$ is finite, then $\tilde{M}$ is compact, which contradicts Theorem 2.13. This gives another proof of Corollary 2.7.
Before we close this section, we note that Proposition 2.19 also straightforwardly implies the following interesting theorem:
Theorem 2.22.
Let $((M,\nabla),h)$ be a compact Hessian manifold of dimension $n$. If the fundamental group of $M$ is Abelian, then $M$ is homeomorphic to the $n$-torus $\mathbb{T}^{n}$.
Proof.
By Proposition 2.19, the fundamental group $\pi_{1}(M)$ of $M$ is isomorphic to $\mathbb{Z}^{k}$ for some $k\in\mathbb{N}$. By Corollary 2.14 and the uniqueness of Eilenberg-MacLane space, $M$ is homotopy equivalent to the $k$-torus $\mathbb{T}^{k}$. Comparing the top homology groups of $M$ and $\mathbb{T}^{k}$ yields that $k=n$. The desired result then follows from Theorem A.3 in the appendix. ∎
2.3 Flat Line Bundles
2.3.1 Locally Constant Picard Group
In this section, we study Hessian manifolds using the concept of flat line bundles. We also define and study the locally constant Picard group of a differentiable manifold, and use it to prove that a curved compact orientable Hessian manifold has positive first Betti number.
We shall now recall the notion of a flat $\mathbb{R}$-line bundle:
Definition 2.23.
For a $\mathbb{R}$-vector bundle $\mathscr{E}\rightarrow M$ over a differentiable manifold $M$, a pair $(\mathscr{E},D)$ is termed a flat $\mathbb{R}$-vector bundle over $M$ if $D$ is a flat connection on $\mathscr{E}\rightarrow M$. A flat $\mathbb{R}$-vector bundle of rank 1 is then called a flat $\mathbb{R}$-line bundle. In particular, we call $(M\times\mathbb{R},d)$ the trivial flat $\mathbb{R}$-line bundle over $M$.
To fix the terminology, we recall the definition of a local frame field, and that of its gluing 1-cocycles.
Let $\mathscr{L}\rightarrow M$ be a $\mathbb{R}$-line bundle over a differentiable manifold $M$. By the local triviality of $\mathscr{L}\rightarrow M$, there exists an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a local section $s_{\alpha}$ of $\mathscr{L}\rightarrow M$ defined on $U_{\alpha}$, such that $s_{\alpha}$ is non-vanishing for all $\alpha\in I$. The collection $\{s_{\alpha}\}_{\alpha\in I}$ is traditionally termed a local frame field of $\mathscr{L}\rightarrow M$. For each $\alpha,\beta\in I$, there exists a unique nowhere vanishing smooth function $g_{\alpha\beta}\in C^{\infty}(U_{\alpha}\cap U_{\beta})$ such that
$s_{\beta}(p)=s_{\alpha}(p)\cdot g_{\alpha\beta}(p)$
for all $p\in U_{\alpha}\cap U_{\beta}$, where $U_{\alpha}$ here is the domain of definition of $s_{\alpha}$ and $U_{\beta}$ is the domain of definition of $s_{\beta}$. The collection $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ is termed the gluing 1-cocycle of the local frame field $\{s_{\alpha}\}_{\alpha\in I}$ of $\mathscr{L}$.
We then investigate the most basic properties of flat $\mathbb{R}$-line bundles.
Lemma 2.24.
Let $M$ be a differentiable manifold and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Then $\mathscr{L}\rightarrow M$ admits a local frame field $\{s_{\alpha}\}_{\alpha\in I}$ such that $Ds_{\alpha}=0$ for all $\alpha\in I$, and that the gluing 1-cocycle of the local frame field $\{s_{\alpha}\}_{\alpha\in I}$ consists of locally constant functions.
Proof.
Since the connection $D$ is flat, its connection 1-forms are locally exact. without loss of generality, we may assume that there exists $f\in C^{\infty}(U_{\alpha})$ such that the restriction of $D$ on $U_{\alpha}$ is $d+df$.
Now, it suffices to show that, up to a real constant multiple, the first order partial differential equation $ds+sdf=0$ for $s\in C^{\infty}(U_{\alpha})$ has a unique positive solution. But by the method of integration factor, we have
$e^{f}(ds+sdf)=d(e^{f}s)$
and hence the general solution of differential equation $ds+sdf=0$ is precisely $s=Ce^{-f}$, where $C\in\mathbb{R}$ is an integration constant. This proves the desired result. ∎
Lemma 2.24 motivates the following natural definition.
Definition 2.25.
Let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over a differentiable manifold $M$. A local frame field $\{s_{\alpha}\}_{\alpha\in I}$ of $\mathscr{L}\rightarrow M$ is termed an affine local frame field of $(\mathscr{L},D)$, if $Ds_{\alpha}=0$ for all $\alpha\in I$. The gluing 1-cocycle of some affine local frame field of $(\mathscr{L},D)$ is called a 1-cocycle of $(\mathscr{L},D)$.
Now we introduce the isomorphism classes in the category of flat $\mathbb{R}$-line bundle over a fixed differentiable manifold $M$.
Definition 2.26.
Two flat $\mathbb{R}$-line bundles $(\mathscr{L}_{1},D_{1})$ and $(\mathscr{L}_{2},D_{2})$ over a differentiable manifold $M$ are said to be gauge equivalent, if there exists a $\mathbb{R}$-vector bundle isomorphism $\phi\colon\mathscr{L}_{1}\rightarrow\mathscr{L}_{2}$ such that $D_{1}-\phi^{*}D_{2}\in\Omega^{1}(M)$ is an exact 1-form on $M$.
Let $M$ be a differentiable manifold. The collection of gauge equivalence classes of flat $\mathbb{R}$-line bundle over $M$ equipped with the tensor product $\otimes$ as binary operation forms a multiplicative group.
Definition 2.27.
Let $M$ be a differentiable manifold. The group of gauge equivalence classes of flat $\mathbb{R}$-line bundles over $M$ is termed the locally constant Picard group of $M$.
The following theorem, which relates the locally constant Picard group with cohomologies, is of fundamental importance.
Lemma 2.28.
Let $M$ be a differentiable manifold. The mapping that sends every flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$ to the Čech cohomology class of a 1-cocycle of $(\mathscr{L},D)$ in $\check{\mathrm{H}}(M;\mathbb{R}^{\times})$ is surjective, and two flat $\mathbb{R}$-line bundles over $M$ share the same image under this map if and only if they are gauge equivalent.
Proof.
This is a very special case of much more general statements in gauge theory. For a comprehensive proof, see [23]. ∎
We shall now have a closer look at the structure of the locally constant Picard group.
Proposition 2.29.
Let $M$ be a differentiable manifold. The induced homomorphism $w_{1}\colon\mathrm{H}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}(M;\mathbb{Z}/2\mathbb{Z})$ in the long exact sequence
$\cdots\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R}^{\times})\longrightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\longrightarrow\cdots$
of cohomologies induced from the short exact sequence
$0\longrightarrow\mathbb{R}\overset{\operatorname{exp}}{\longrightarrow}\mathbb{R}^{\times}\overset{\operatorname{sgn}}{\longrightarrow}\mathbb{Z}/2\mathbb{Z}\longrightarrow 0$
of Abelian groups is precisely the first Stiefel-Whitney class.
Proof.
This statement follows from standard algebraic topology. ∎
A crucial difference between the exponential sequence in complex geometry and our short exact sequence
$0\longrightarrow\mathbb{R}\longrightarrow\mathbb{R}^{\times}\longrightarrow\mathbb{Z}/2\mathbb{Z}\longrightarrow 0$
is that the latter one splits. This allow us to define a new invariant for flat $\mathbb{R}$-line bundles.
Definition 2.30.
Let $M$ be a differentiable manifold, and let $\varphi\colon\mathbb{R}^{\times}\rightarrow\mathbb{R}$ be the even extension of the logarithm function, that is, $\varphi(x)=\operatorname{log}|x|$ for all $x\in\mathbb{R}^{\times}$. The induced homomorphism $\operatorname{sh}:=\varphi_{*}\colon\mathrm{H}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}(M;\mathbb{R})$ in the long exact sequence
$\cdots\longrightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R}^{\times})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})\longrightarrow\cdots$
of cohomologies induced from the short exact sequence
$0\longrightarrow\mathbb{Z}/2\mathbb{Z}\longrightarrow\mathbb{R}^{\times}\overset{\operatorname{\varphi}}{\longrightarrow}\mathbb{R}\longrightarrow 0$
of Abelian groups is called the first Koszul class.
As an immediate consequence of Proposition 2.29 and Definition 2.30, the first Koszul class and the first Stiefel-Whitney class together classify the locally constant Picard group.
Corollary 2.31.
Let $M$ be a differentiable manifold. The mapping
$w_{1}\oplus\operatorname{sh}\colon\mathrm{H}^{1}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\oplus\mathrm{H}^{1}(M;\mathbb{R}),\quad(\mathscr{L},D)\mapsto(w_{1}(\mathscr{L}),\operatorname{sh}(D))$
is an isomorphism of Abelian groups.
Proof.
Apply the splitting lemma and use the exactness of the cohomology functor. ∎
We finally arrived at the structure theorem of locally constant Picard groups.
Corollary 2.32.
Let $M$ be a differentiable manifold. The locally constant Picard group of $M$ is isomorphic to the additive group $\mathrm{H}^{1}(M;\mathbb{R})\oplus\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$.
Proof.
The desired result follows from Lemma 2.28 and Corollary 2.31. ∎
In the case when the underlying manifold is aspherical, we may give an explicit description of the action of first Koszul class as an induced homomorphism.
Proposition 2.33.
Let $M$ be an aspherical differentiable manifold and $\pi_{1}(M)$ its fundamental group, so that we identify $\operatorname{Hom}(\pi_{1}(M),\mathbb{R})\equiv\mathrm{H}^{1}(M;\mathbb{R})$ via holonomy representation. Let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$, and let $\rho\colon\pi_{1}(M)\rightarrow\mathbb{R}^{\times}$ be its holonomy representation. Then, the homomorphism
$\pi_{1}(M)\rightarrow\mathbb{R},\quad g\mapsto\operatorname{log}|\rho(g)|$
as a group cohomology class in $\mathrm{H}^{1}(\pi_{1}(M);\mathbb{R})$ is precisely equal to the first Koszul class $\operatorname{sh}(D)$ of $(\mathscr{L},D)$.
Proof.
Denote by $G:=\pi_{1}(M)$ the fundamental group of $M$. Since $M$ is aspherical, the classifying map $M\to BG$ induces an isomorphism between the sheaf cohomology with locally constant coefficients and the corresponding group cohomology:
$\mathrm{H}^{1}(M;A)\cong\mathrm{H}^{1}(\pi_{1}(M);A)$
for any abelian group $A$.
Under this canonical isomorphism, the gauge equivalent class of $(\mathscr{L},D)$ in $\mathrm{H}^{1}(M;\mathbb{R}^{\times})$ corresponds to its holonomy representation $\rho\in\mathrm{H}^{1}(G;\mathbb{R}^{\times})\cong\operatorname{Hom}(G,\mathbb{R}^{\times})$.
Define group homomorphism $\varphi\colon\mathbb{R}^{\times}\rightarrow\mathbb{R}$ by $\varphi(x)=\operatorname{log}|x|$ for all $x\in\mathbb{R}^{\times}$. Then it readily seen that the following diagram commutes
$\begin{array}[]{ccc}\mathrm{H}^{1}(M;\mathbb{R}^{\times})&\overset{\operatorname{sh}}{\longrightarrow}&\mathrm{H}^{1}(M;\mathbb{R})\\ \downarrow&&\downarrow\\ \mathrm{H}^{1}(G;\mathbb{R}^{\times})&\overset{\varphi_{*}}{\longrightarrow}&\mathrm{H}^{1}(G;\mathbb{R})\end{array}$
and the desired property follows. ∎
We shall also be interested in a trivially criterion of elements of the locally constant Picard group. Recall that a line bundle is trivial if and only if it admits a nowhere vanishing global section. The following proposition gives a refinement in the case of flat $\mathbb{R}$-line bundles.
Proposition 2.34.
Let $M$ be an differentiable manifold, and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Then $(\mathscr{L},D)$ is gauge equivalent to the trivial flat $\mathbb{R}$-line bundle over $M$ if and only if it admits a nowhere vanishing parallel global section.
Proof.
Assume that $(\mathscr{L},D)$ is gauge equivalent to the trivial flat $\mathbb{R}$-line bundle over $M$. Then we identify $\mathscr{L}\equiv M\times\mathbb{R}$, and there exists an exact form $\theta\in\Omega^{1}(M)$ such that $D=d+\theta$. Take $f\in C^{\infty}(M)$ such that $\theta=df$. Then the global section
$s=\operatorname{exp}(-f)\in\mathrm{H}^{0}(M;M\times\mathbb{R})=C^{\infty}(M)$
is nowhere vanishing and satisfies that
$Ds=d\operatorname{exp}(-f)+e^{-f}\theta=e^{-f}(\theta-df)=0.$
To prove the converse statement, we assume that $(\mathscr{L},D)$ admits a nowhere vanishing parallel global section $s$. Then in particular we have $\mathscr{L}\cong M\times\mathbb{R}$. Moreover, since
$D(fs)=df\otimes s+f\cdot Ds=df\otimes s$
for all $f\in C^{\infty}(M)$, we conclude that $D=d$. ∎
To illustrate the power of the structure theorem of locally constant Picard group, we compare two proofs of the following torsion-free property.
Proposition 2.35.
Let $M$ be a differentiable manifold, and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Suppose that $\mathscr{L}$ is a trivial $\mathbb{R}$-line bundle, and there exists a positive integer $k$ such that $(\mathscr{L}^{\otimes k},D)$ is a trivial flat $\mathbb{R}$-line bundle. Then $(\mathscr{L},D)$ itself is a trivial flat $\mathbb{R}$-line bundle.
Proof.
Let $s\in\mathrm{H}^{0}(M;\mathscr{L})$ be a nowhere-vanishing smooth section. Then by assumption and Proposition 2.34, there exists $f\in C^{\infty}(M)$ such that $\operatorname{exp}(kf)s^{\otimes k}$ is parallel. Then by Leibniz rule of differentiation,
$k(\operatorname{exp}(f)s)^{\otimes k-1}\otimes D(\operatorname{exp}(f)s)=D(\operatorname{exp}(kf)s^{\otimes k})=0$
and hence $D(\operatorname{exp}(f)s)=0$, as $k(\operatorname{exp}(f)s)^{\otimes k-1}$ is nowhere vanishing. Therefore $\operatorname{exp}(f)s$ is a nowhere vanishing parallel section of $(\mathscr{L},D)$. Again by Proposition 2.34, we conclude that $(\mathscr{L},D)$ is trivial as a flat $\mathbb{R}$-line bundle. ∎
Proof.
Since $\mathscr{L}$ is a trivial $\mathbb{R}$-line bundle, its Stiefel-Whitney class $w_{1}(\mathscr{L})=0$. Therefore the gauge equivalent class of $(\mathscr{L},D)$ is contained in the subgroup $\mathrm{H}^{1}(M;\mathbb{R})$ of the locally constant Picard group of $M$. Since $(\mathscr{L}^{\otimes k},D)$ is trivial and $\mathrm{H}^{1}(M;\mathbb{R})$ is torsion-free, we conclude that $(\mathscr{L},D)$ is also trivial. ∎
2.3.2 Applications of Koszul Forms
Amongst all flat $\mathbb{R}$-line bundle over an affine manifold, the following one, which will be called the canonical line bundle, is of great importance.
Definition 2.36.
Let $(M,\nabla),$ be an affine manifold, and denote also by $\nabla$ the connection on tensor bundles of $M$ induced by $\nabla$. The flat $\mathbb{R}$-line bundle
$K_{\nabla}=(\bigwedge^{n}(T^{*}M),\nabla)$
over $M$ is termed the canonical line bundle of $(M,\nabla)$.
Note that, although the top exterior power of the cotangent bundle of an orientable manifold is always trivial, the canonical line bundle of an orientable affine manifold in general is non-trivial as a flat $\mathbb{R}$-line bundle.
Recall that a Calabi-Yau manifold is a Kähler manifold with trivial canonical line bundle, or equivalently, admitting a holomorphic volume form. The following corollary exhibits an close analogy.
Corollary 2.37.
Let $(M,\nabla)$ be an affine manifold of dimension $n$. Then the canonical line bundle $K_{\nabla}$ of $(M,\nabla)$ is trivial if and only if $M$ admits a nowhere vanishing $n$-form $\Omega$ such that $\nabla\Omega=0$.
Proof.
This directly follows from Proposition 2.34. ∎
In this section, we shall explore various applications of flat $\mathbb{R}$-line bundles and their first Koszul forms, including a powerful structure theorem of compact Hessian manifolds.
We now introduce the notion of the first Koszul form of a Hessian line bundle.
Definition 2.38.
Let $M$ be a differentiable manifold. A Hessian line bundle over $M$ is a pair $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, where $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$ and $\langle\cdot,\cdot\rangle$ is a bundle metric of $\mathscr{L}\rightarrow M$.
Definition 2.39.
Let $M$ be a differentiable manifold, and let $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ be a Hessian line bundle over $M$. The first Koszul form $\varkappa\in\Omega^{1}(M)$ of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is defined by
$\varkappa|_{U_{\alpha}}:=\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
where $\{s_{\alpha}\}_{\alpha\in I}$ is an affine local frame of $(\mathscr{L},D)$ and $U_{\alpha}$ is the domain of definition of $s_{\alpha}$ for each $\alpha\in I$.
By the properties of affine local frame, the first Koszul form of a Hessian line bundle is well-defined, and is closed as it is by definition locally exact.
A natural question is to compare the first Koszul class of a flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ and the cohomology class of the first Koszul form of Hessian line bundle $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, where $\langle\cdot,\cdot\rangle$ is some bundle metric on $\mathscr{L}$. Surprisingly, the cohomology class of the first Koszul form is independent of the auxiliary bundle metric and always coincides with the first Koszul class.
Proposition 2.40.
Let $M$ be a differentiable manifold, and let $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ be a Hessian line bundle over $M$. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, then we have $\operatorname{sh}(D)=[\varkappa]\in\mathrm{H}^{1}(M;\mathbb{R})$ is the first Koszul class of $(\mathscr{L},D)$.
Proof.
Let $\{s_{\alpha}\}_{\alpha\in I}$ be an affine local frame field for $(\mathscr{L},D)$ and let $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ be its gluing 1-cocycle.
By the very definition, the first Koszul class $\operatorname{sh}(D)$ sends the class of the cocycle $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ in $\mathrm{H}^{1}(M;\mathbb{R}^{\times})$ to the class of the cocycle $\{\operatorname{log}|g_{\alpha\beta}|\}_{\alpha,\beta\in I}$ in $\mathrm{H}^{1}(M;\mathbb{R})$.
By Definition 2.39, the first Koszul form $\varkappa$ is a global 1-form defined locally by
$\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
for $\alpha\in I$. The de Rham cohomology class $\varkappa]$ corresponds to a Čech cohomology class in $\mathrm{H}^{1}(M;\underline{\mathbb{R}})$ via the de Rham isomorphism. For any $\alpha\in I$, define
$f_{\alpha}:=\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle).$
Then for any $\alpha\in I$, we have
$d(f_{\alpha})=\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle).$
Straightforward computation yields that
$f_{\beta}-f_{\alpha}$ $=\frac{1}{2}\operatorname{log}(\langle s_{\beta},s_{\beta}\rangle)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\frac{1}{2}\operatorname{log}(\langle s_{\alpha}g_{\alpha\beta},s_{\alpha}g_{\alpha\beta}\rangle)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\frac{1}{2}\operatorname{log}(g_{\alpha\beta}^{2}\langle s_{\alpha},s_{\alpha}\rangle)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\frac{1}{2}\left(\operatorname{log}(g_{\alpha\beta}^{2})+\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)\right)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\frac{1}{2}\operatorname{log}(g_{\alpha\beta}^{2})$
$=\operatorname{log}|g_{\alpha\beta}|.$
Therefore, a Čech cocycle representing $\varkappa]$ is $\{\operatorname{log}|g_{\alpha\beta}|\}_{\alpha,\beta\in I}$, and hence $\operatorname{sh}(D)=[\varkappa]$. ∎
On a Hessian manifold, we have a canonical Koszul form, which we will now define.
Definition 2.41.
Let $((M,\nabla),h)$ be a Hessian manifold and denote by $\langle\cdot,\cdot\rangle_{h}$ the bundle metric on tensor bundles of $M$ induced by $h$. The first Koszul form of the Hessian line bundle $(K_{\nabla},\langle\cdot,\cdot\rangle_{h})$ is also said to be the first Koszul form of $((M,\nabla),h)$.
For the convenience of the reader, we review the standard notion of first Koszul forms of a Hessian manifold.
For an orientable Riemannian manifold $(M,h)$, we denote by $\operatorname{vol}_{h}$ its volume form.
Proposition 2.42.
Let $((M,\nabla),h)$ be an oriented Hessian manifold, and let $\varkappa$ be its first Koszul form. Then it holds that $2\varkappa=d\operatorname{log}(|h|)$, and
$\nabla_{X}\operatorname{vol}_{h}=\varkappa(X)\operatorname{vol}_{h}$
for every vector field $X$ on $M$, where $|h|$ is the determinant of the matrix of which $(i,j)$-entry is $h(\partial/\partial x^{i},\partial/\partial x^{j})$ for some oriented affine coordinates $x^{1},\dots,x^{n}$ of the oriented affine manifold $(M,\nabla)$.
Proof.
The proof is straightforward and hence omitted. ∎
In [12], H.Shima and K.Yagi established the following criterion for flatness of Hessian manifolds in terms of the first Koszul form.
Proposition 2.43 (Shima-Yagi, 1997[12]).
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold. Let $\bar{\nabla}$ be the Levi-Civita connection of $(M,h)$, and let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((M,\nabla),h)$. Then it holds that $\bar{\nabla}\varkappa=0$.
Theorem 2.44 (Shima-Yagi, 1997[12]).
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold of dimension $n$. Then the following statements are equivalent:
  1. 1.
    $(M,h)$ is Riemannian flat;
  2. 2.
    The first Koszul form of $((M,\nabla),h)$ vanishes identically;
  3. 3.
    $\nabla$ is the Levi-Civita connection of $(M,h)$;
  4. 4.
    $M$ admits a nowhere vanishing $n$-form $\Omega$ such that $\nabla\Omega=0$.
Recall that, in Remark 2.8, we found an example of an orientable compact Hessian manifold $((M,\nabla),h)$ with vanishing first Betti number, but in that case $(M,h)$ is in fact a flat Riemannian manifold. This is not just a coincidence.
Theorem 2.45.
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold. If the first Betti number of $M$ is zero, then $(M,h)$ is a flat Riemannian manifold and $\nabla$ is its Levi-Civita connection.
Proof.
Since $\mathrm{H}^{1}(M;\mathbb{R})=0$, by Proposition 2.32 we have that the Picard group of $M$ is isomorphic to $\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$, and hence is 2-torsion. In particular, the canonical line bundle $K$ of $(M,\nabla)$ satisfies $K^{\otimes 2}=0$. Since $M$ is orientable, the $\mathbb{R}$-line bundle $\bigwedge^{n}(T^{*}M)$ is trivial. Therefore, by virtue of Lemma 2.35, the canonical line bundle of $M$ is trivial as a flat $\mathbb{R}$-line bundle. Corollary 2.37 then implies that $M$ admits a volume form $\Omega$ such that $\nabla\Omega=0$. The desired result follows immediately from Theorem 2.44. ∎
As an application of the first Koszul forms, we suggest several important splitting theorems of Hessian manifolds, which all relies on the following fundamental result of Tischler on fibration over circles.
Theorem 2.46 (Tischler, 1970[19]).
Let $M$ be a compact differentiable manifold. If $M$ admits a nowhere vanishing closed 1-form, then $M$ is a mapping torus.
Theorem 2.47.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then there exists a fiber bundle $\varphi\colon M\rightarrow\mathbb{S}^{1}$.
Proof.
Let $\eta\in\Omega^{1}(M)$ be a $1$-form potential of $((M,\nabla),h)$. View $\eta\colon M\rightarrow T^{*}M$ as an embedding. Then, $\eta$ vanishes at $p\in M$ if and only if $p$ is an intersection point of $\eta\colon M\rightarrow T^{*}M$ and the zero section $\sigma$ of the cotangent bundle of $M$. Since $\nabla\eta=h>0$, each intersection point of $\eta$ and $\sigma$ is a nondegenerate transverse intersection, and the intersection numbers at those intersection points are all positive. But, on the other hand, the sum of the intersection numbers at the intersection points of $\eta$ and $\sigma$ is precisely
$\langle e(T^{*}M),[M]\rangle=\chi(M)=0$
by Proposition 2.17. This proves that the sections $\eta$ and $\sigma$ must have no intersection, i.e. $\eta$ is a nowhere vanishing closed 1-form on $M$. The desired result then follows from Theorem 2.46. ∎
An stronger result can be obtained if the manifold under consideration is orientable.
Theorem 2.48.
Let $((M,\nabla),h)$ be an oriented compact Hessian manifold. Then either $(M,h)$ is a flat Riemannian manifold, or there exists a fiber bundle $\varphi\colon M\rightarrow\mathbb{S}^{1}$ with connected and oriented fiber.
Proof.
Let $\varkappa$ be the first Koszul form of $((M,\nabla),h)$. If $\varkappa=0$, then $(M,h)$ is a flat Riemannian manifold by Theorem 2.44. We therefore assume $\varkappa\neq 0$. But then since $\bar{\nabla}\varkappa=0$ by Proposition 2.43, we have that $\varkappa$ is nowhere vanishing by parallel transport.
Since $\mathrm{H}^{1}(M;\mathbb{Q})$ is dense in $\mathrm{H}^{1}(M;\mathbb{R})$, there exists a closed $1$-form $\hat{\varkappa}\in\Omega^{1}(M)$ such that $\hat{\varkappa}]\in\mathrm{H}^{1}(M;\mathbb{Q})$ and $\hat{\varkappa}$ is also nowhere vanishing. Define
$m:=\operatorname{min}\{k\in\mathbb{R}:k>0,k[\hat{\varkappa}]\in\mathrm{H}^{1}(M;\mathbb{Z})\}.$
Fix a base point $q$ on $M$, and define the smooth mapping $\varphi\colon M\rightarrow\mathbb{S}^{1}$ via
[…]
where $\gamma\colon[0,1]\rightarrow M$ is a smooth curve with $\gamma(0)=q$ and $\gamma(1)=p$. Since the periods of $m[\hat{\varkappa}$] are integers, the map $\varphi$ is well-defined. Moreover, since $\hat{\varkappa}$ is nowhere vanishing, we have that $\varphi$ is a submersion. By Ehresmann’s fibration theorem, we have that $\varphi\colon M\rightarrow\mathbb{S}^{1}$ is a fiber bundle. Denote by $F$ the fiber of $\varphi\colon M\rightarrow\mathbb{S}^{1}$.
Identifying each homotopy class in $\pi_{1}(\mathbb{S}^{1})$ with its winding number, it is readily seen that the induced homomorphism $\varphi_{*}\colon\pi_{1}(M)\rightarrow\pi_{1}(\mathbb{S}^{1})=\mathbb{Z}$ satisfies
$\varphi_{*}([\gamma])=\int_{\gamma}m\hat{\varkappa}$
for all $\gamma]\in\pi_{1}(M)$. Since $\hat{\varkappa}\neq 0$, we have that $\operatorname{Im}(\varphi_{*})\neq 0$. Therefore, there exists a positive integer $d\in\mathbb{Z}$ such that $\operatorname{Im}(\varphi_{*})=d\mathbb{Z}$. But by the minimality of the factor $m$, we must have that $d=1$ and hence $\operatorname{Im}(\varphi_{*})=\mathbb{Z}$. Now, since $M$ is connected and $\varphi_{*}$ is surjective, the long exact sequence
$\cdots\longrightarrow\pi_{1}(M)\longrightarrow\pi_{1}(\mathbb{S}^{1})\longrightarrow\pi_{0}(F)\longrightarrow\pi_{0}(M)\longrightarrow\cdots$
of homotopy groups yields that $F$ is also a connected topological space.
Moreover, since $\varphi\colon M\rightarrow\mathbb{S}^{1}$ is a submersion between compact differentiable manifolds, we have that $F$ is also a compact differentiable manifold.
By the very definition of $\varphi$, it is readily seen that the vector field $\hat{\varkappa}^{\sharp}$ is transverse to the fibers of $\varphi\colon M\rightarrow\mathbb{S}^{1}$. Consider the (n1)(n-1)-form $\Omega$ on $M$ defined by
$\Omega(X_{1},\dots,X_{n-1}):=\operatorname{vol}_{h}(X_{1},\dots,X_{n-1},\hat{\varkappa}^{\sharp})$
where $X_{1},\dots,X_{n-1}$ are arbitrary vector fields on $M$. Then by construction $\Omega|_{F}$ is nowhere vanishing. Therefore we conclude that $F$ is oriented. ∎
As yet another application of the first Koszul forms, resembling the notion of positive holomorphic line bundles in complex geometry, we now define the positivity of a flat $\mathbb{R}$-line bundle, and use it to characterize Hessian metrics of Koszul type.
Definition 2.49.
Let $(M,\nabla)$ be an affine manifold. A flat line bundle $(\mathscr{L},D)$ over $M$ is said to be a positive flat line bundle over $(M,\nabla)$ if there exists a bundle metric $\langle\cdot,\cdot\rangle$ of $\mathscr{L}\rightarrow M$ such that $\nabla\varkappa>0$, where $\varkappa$ is the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$.
Proposition 2.50.
An affine manifold admits a Hessian metric of Koszul type if and only of it admits a positive flat line bundle.
Proof.
Let $(M,\nabla)$ be an arbitrary but fixed affine manifold.
Assume that there exists a Hessian metric $h$ on $(M,\nabla)$ and a closed $1$-form $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$. Consider an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a smooth function $f_{\alpha}\in C^{\infty}(U_{\alpha})$, such that for each $\alpha\in I$ it holds that $\eta_{\alpha}=df_{\alpha}$, where $\eta_{\alpha}$ is the restriction of $\eta$ on $U_{\alpha}$. The flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$,
Let $\mathscr{L}\rightarrow M$ be the $\mathbb{R}$-line bundle over $M$ defined by the 1-cocycle $\{\operatorname{exp}(f_{\alpha}-f_{\beta})\}_{\alpha,\beta\in I}$ and let $D$ be the flat connection on $\mathscr{L}\rightarrow M$ defined by the connection 1-forms $\{df_{\alpha}\}_{\alpha\in I}$.
Let $\{s_{\alpha}\}_{\alpha\in I}$ be an affine local frame of $(\mathscr{L},D)$. Without loss of generality, for each $\alpha\in I$, let $U_{\alpha}$ be the domain of definition of $s_{\alpha}$. Define a bundle metric $\langle\cdot,\cdot\rangle$ on $\mathscr{L}$ by
$\langle s_{\alpha},s_{\alpha}\rangle:=\operatorname{exp}(-2f_{\alpha})$
for every $\alpha\in I$. Then, $\langle\cdot,\cdot\rangle$ is well-defined, and by construction the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is precisely $\eta$. Therefore $\nabla\eta=h>0$, and hence $(\mathscr{L},D)$ is positive.
Conversely, if $(M,\nabla)$ admits a positive flat line bundle $(\mathscr{L},D)$, then by definition there exists a bundle metric $\langle\cdot,\cdot\rangle$ such that $h:=\nabla\varkappa>0$ is a Hessian metric of Koszul type, where $\varkappa$ is the first Koszul form of the Hessian line bundle $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$. ∎
3 Interaction with Classical Geometric Structures
3.1 Digression on Complex Geometry
This digression demonstrates the utility of our framework beyond affine geometry. By lifting the Hessian structure to the tangent bundle, we can leverage powerful theorems from complex analysis.
We prove that every positive flat line bundles over an affine manifold $(M,\nabla)$ can be lifted to a positive holomorphic line bundles over the Sasaki lifting of $(M,\nabla)$, and that the Sasaki lifting of a complete Hessian manifold is Stein, which finally results in an analytic embedding theorem for complete Hessian manifolds.
Let $(M,\nabla)$ be an affine manifold, and let $\pi\colon TM\rightarrow M$ be the tangent bundle of $M$. Denote by $\mathcal{A}$ the sheaf of affine functions on $(M,\nabla)$.
Consider any coordinate map
$U\longrightarrow\mathbb{R}^{n},\quad p\longmapsto(x^{1}(p),\dots,x^{n}(p))$
of $M$ such that $x^{1},\dots,x^{n}\in\mathcal{A}(U)$ are affine functions, and that $TM\rightarrow M$ is trivialized over $U$.
Then, it is readily seen that
$\pi^{-1}(U)\longrightarrow\mathbb{R}^{2n},\quad(p,v^{1}\partial/\partial x^{1}+\cdots+v^{n}\partial/\partial x^{n})\longmapsto(x^{1}(p),\dots,x^{n}(p);v^{1},\dots,v^{n})$
is a coordinate map of $TM$.
Now, the canonical almost complex structure $J_{\nabla}$ on the tangent bundle $T(TM)\rightarrow TM$ is defined by
$J_{\nabla}(\partial/\partial x^{i}):=-\partial/\partial v^{i},\quad J_{\nabla}(\partial/\partial v^{i}):=\partial/\partial x^{i}$
for all $i=1,\dots,n$. By the classical construction in [24], the linear operator $J_{\nabla}$ is a well-defined integrable complex structure, i.e. $(TM,J_{\nabla})$ is a complex manifold. The complex manifold $(TM,J_{\nabla})$ is said to be the Sasaki lifting of $(M,\nabla)$, and the holomorphic coordinates
$z^{1}:=x^{1}+\sqrt{-1}v^{1},\dots,z^{n}:=x^{n}+\sqrt{-1}v^{n}$
of $(TM,J_{\nabla})$ are termed the canonical complexifications of the affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$.
Moreover, for each flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$, it is readily seen that the anti-holomorphic part $\bar{\partial}_{D}$ of the $\mathbb{C}$-linear extension of the connection $D$ is a Dolbeault operator on the pullback bundle $\mathscr{L}_{\mathbb{C}}:=\pi^{*}(\mathscr{L}\otimes\mathbb{C})$.
We thus arrived at the following definition:
Definition 3.1.
For an affine manifold $(M,\nabla)$, the complex manifold $(TM,J_{\nabla})$ constructed above is termed the Sasaki lifting of the affine manifold $(M,\nabla)$, and moreover, for a flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$, the holomorphic line bundle $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ over $(TM,J_{\nabla})$ constructed above is termed the holomorphic pullback of $(\mathscr{L},D)$ along $\nabla$, where a holomorphic line bundle here is equivalently defined as a pair consists of a smooth $\mathbb{C}$-line bundle and a Dolbeault operator on it.
For the complex geometric aspect of affine manifolds, the following result is of fundamental importance.
Lemma 3.2.
Let $(M,\nabla)$ be an affine manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Suppose that $(\mathscr{L},D)$ is a positive flat line bundle over $(M,\nabla)$. Then the holomorphic pullback $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ of $(\mathscr{L},D)$ along $\nabla$ is a positive holomorphic line bundle over the complex manifold $(TM,J_{\nabla})$.
Proof.
By the positivity of $(\mathscr{L},D)$, there exists a bundle metric $\langle\cdot,\cdot\rangle$ of $\mathscr{L}\rightarrow M$ such that the first Koszul form $\varkappa$ of the Hessian line bundle $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ satisfies $\nabla\varkappa>0$.
Take any arbitrary point $p\in M$. Since the first Koszul form $\varkappa$ of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is closed, by Poincaré’s lemma there exists an open neighborhood $U$ of $p$ in $M$ and a smooth function $f\in C^{\infty}(U)$ such that $\varkappa|_{U}=df$. Then, by the very definition of the first Koszul form of a Hessian line bundle, there exists a non-vanishing parallel local section $s$ of $\mathscr{L}\rightarrow M$ defined on $U$, such that $\langle s,s\rangle=\operatorname{exp}(-2f)$.
Denote by $\pi$ the bundle map of the tangent bundle $TM\rightarrow M$. We equip $(\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D})$ with the pullback bundle metric $\langle\cdot,\cdot\rangle_{\mathbb{C}}:=\pi^{*}\langle\cdot,\cdot\rangle$. Then the curvature 2-form $\Theta$ of Chern connection of the Hermitian line bundle $((\mathscr{L}_{\mathbb{C}},\bar{\partial}_{D}),\langle\cdot,\cdot\rangle_{\mathbb{C}})$ is given by
$2\pi\sqrt{-1}\Theta=\partial\bar{\partial}\operatorname{log}(\langle\pi^{*}s,\pi^{*}s\rangle_{\mathbb{C}})=2\partial\bar{\partial}(\pi^{*}f).$
Now it remains to prove that $\pi^{*}f$ is a plurisubharmonic function. But this follows immediately from the assumption that $\nabla(df)=\nabla\varkappa|_{U}>0$. ∎
After the notion of pullback of line bundles being successfully defined, we also want to discover a proper way to lift smooth functions from the base. Instead of naively pullback smooth functions on an affine manifold to its tangent bundle via the bundle projection, there exists a much more purposeful method of lifting smooth functions.
Proposition 3.3.
Let $(M,\nabla)$ be an affine manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Let $f\colon M\rightarrow\mathbb{R}$ and $F\colon TM\rightarrow\mathbb{C}$ be smooth functions satisfying
$F(x^{1},\dots,x^{n};v^{1},\dots,v^{n})=f(x^{1},\dots,x^{n})+\sqrt{-1}\sum_{i=1}^{n}\frac{\partial f}{\partial x^{i}}(x^{1},\dots,x^{n})v^{i}$
where here
$z^{1}=x^{1}+\sqrt{-1}v^{1},\dots,z^{n}=x^{n}+\sqrt{-1}v^{n}$
are the canonical complexifications of some affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. Then $f\in C^{\infty}(M)$ is affine if and only of $F\in C^{\infty}(TM)\otimes\mathbb{C}$ is holomorphic.
Proof.
Apply Cauchy-Riemann equation to the specific complex structure $J_{\nabla}$. ∎
Proposition 3.3 then motivates the following definition for general smooth functions.
Definition 3.4.
Let $(M,\nabla)$ be an affine manifold, and let $(TM,J_{\nabla})$ be its Sasaki lifting. Let
$z^{1}=x^{1}+\sqrt{-1}v^{1},\dots,z^{n}=x^{n}+\sqrt{-1}v^{n}$
be the canonical complexifications of some affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. For any smooth function $f\in C^{\infty}(M)$ on $M$, The complex-valued smooth function $F\in C^{\infty}(TM)\otimes\mathbb{C}$ defined by
$F(x^{1},\dots,x^{n};v^{1},\dots,v^{n}):=f(x^{1},\dots,x^{n})+\sqrt{-1}\sum_{i=1}^{n}\frac{\partial f}{\partial x^{i}}(x^{1},\dots,x^{n})v^{i}$
is termed the canonical complexification of $f$ along $\nabla$.
Note that, the notion of canonical complexifications of affine coordinates is consistent with the definition of canonical complexifications of smooth functions, which also justifies the usage of terminology.
Although a tangent bundle of positive dimension is never compact, we may hope that the Sasaki liftings are Stein. Fortunately, for the the Sasaki liftings of complete Hessian manifolds, this is indeed true.
Proposition 3.5.
Let $((M,\nabla),h)$ be a complete Hessian manifold, and $(TM,J_{\nabla})$ its Sasaki lifting. Then $(TM,J_{\nabla})$ is Stein.
Proof.
Let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. By considering the Sasaki lifting of $(\tilde{M},\pi^{*}\nabla)$, it is readily seen that there exists a holomorphic covering map $T\tilde{M}\rightarrow TM$. Therefore it suffices to show that $T\tilde{M}$ is Stein.
But, by Theorem 2.13, $T\tilde{M}$ is a convex tube domain, and is therefore pesudoconvex by complex analysis of several variables. ∎
The following totally real embedding theorem for complete Hessian manifolds is a direct application of Proposition 3.5.
Theorem 3.6.
Let $((M,\nabla),h)$ be an $n$-dimensional complete Hessian manifold. Then there exists a proper analytic embedding of $M$ into $\mathbb{R}^{d}$ for any integer $d>(3n+1)/2$.
Proof.
By Proposition 3.5, the Sasaki lifting $(TM,J_{\nabla})$ of $(M,\nabla)$ is Stein. By Stein’s embedding theorem [25], there exists a proper holomorphic embedding $\Phi$ of $(TM,J_{\nabla})$ into $\mathbb{C}^{d}$ for any integer $d>(3n+1)/2$. Now, it suffices to restrict the real part $\operatorname{Re}(\Phi)$ of $\Phi$ to the totally real submanifold $M$ of $(TM,J_{\nabla})$. ∎
Remark 3.7.
Recall that Whitney’s embedding theorem states that all real analytic $n$-manifolds can be analytically embedded into $\mathbb{R}^{2n}$. Theorem 3.6 suggests a more efficient bound on analytical embedding dimension for complete Hessian manifolds.
3.2 Legendre Transformation
It turns out that there exists a natural duality between radiant Hessian manifolds and Hessian manifolds of Koszul type. Also, for a radiant Hessian manifold $((M,\nabla),h)$ of Koszul type, we prove that its dual $((M,\nabla^{*}),h)$ is also a radiant Hessian manifold of Koszul type, and provide an explicit expression of the $1$-form potential of $((M,\nabla^{*}),h)$ in terms of that of $((M,\nabla),h)$, resembling the classical Legendre transformation.
In this section, we introduce radiant manifolds and then study Hessian manifold in terms of the generalized Legendre duality mentioned above.
Definition 3.8.
Let $(M,\nabla)$ be an affine manifold. A vector field $H$ on $M$ is termed an Euler vector field of $(M,\nabla)$, if $\nabla_{X}H=X$ for every vector field $X$ on $M$. An affine manifold is said to be radiant, if it admits an Euler vector field.
Remark 3.9.
It is readily seen that, if $X$ and $Y$ are vector fields on an affine manifold such that $X$ is Eulerian, then $Y$ is also Eulerian if and only if XYX-Y is parallel.
Definition 3.10.
Let $M$ be a differentiable manifold. An atlas $\mathscr{A}=\left\{\left(U_{\alpha},\varphi_{\alpha}\right)\mid\alpha\in I\right\}$ of $M$ is said to be a radiant affine atlas, if for any $\alpha,\beta\in I$ there exist $A\in\operatorname{GL}_{n}(\mathbb{R})$ such that
$\varphi_{\beta}\left(\varphi_{\alpha}^{-1}(\bf{x})\right)=A{\bf{x}}$
for all ${\bf{x}}\in\varphi_{\alpha}\left(U_{\alpha}\cap U_{\beta}\right)$. An element of a radiant affine atlas of $M$ is called a radiant affine chart of $M$, and the local coordinates associated to a radiant affine chart of $M$ are termed radiant affine coordinates.
The following theorem shows that the above two definitions of radiant affine manifolds are equivalent to each other.
Theorem 3.11.
An affine manifold $(M,\nabla)$ is radiant if and only if the adapted affine atlas of $\nabla$ is a radiant affine atlas of $M$.
Proof.
See for example [15] and [26]. ∎
The following propositions characterize the universal coverings of compact radiant affine manifolds as open cones in the Euclidean spaces.
Proposition 3.12.
[…]
Proof.
By Theorem 2.13, we may identify $\Omega:=\Phi(\tilde{M})=\tilde{M}$.
Since, up to a translation, $\Omega$ is a cone with apex at the origin, all affine automorphisms of $(\Omega,\pi^{*}\nabla)$ are in fact linear. Now, it suffices to descend the standard Euler vector field
$x^{1}\frac{\partial}{\partial x^{1}}+\cdots+x^{n}\frac{\partial}{\partial x^{n}}$
defined on $(\Omega,\pi^{*}\nabla)$ to an Euler vector field of $(M,\nabla)$. ∎
The converse statement also holds if we assume compactness.
Proposition 3.13.
[…]
Proof.
By Theorem 2.13, we may identify $\Omega:=\Phi(\tilde{M})=\tilde{M}$. An Euler vector field $H$ of $(M,\nabla)$ lifts to a standard Euler vector field
$\pi^{*}H=\sum_{i=1}^{n}(x^{i}-a^{i})\frac{\partial}{\partial x^{i}}$
on $(\Omega,\pi^{*}\nabla)$, where $a^{1},\dots,a^{n}\in\mathbb{R}$. Moreover, by Poincaré-Hopf index theorem, we have $(a^{1},\dots,a^{n})\notin\Omega$. Up to an affine change of coordinates, we may assume $a^{1},\dots,a^{n}=0$. Since $M$ is compact, it follows that $H$ is a complete vector field. Since $\pi\colon\Omega\rightarrow M$ is a covering space, we obtain that $\pi^{*}H$ is also a complete vector field as $H$ is. Now it remains to recall that the integral curves of the standard Euler vector field
$x^{1}\frac{\partial}{\partial x^{1}}+\cdots+x^{n}\frac{\partial}{\partial x^{n}}$
are rays emitting from the origin of $\mathbb{R}^{n}$. ∎
In [26], D.Fried et al. studied the existence of parallel tensors on compact radiant affine manifolds.
Theorem 3.14 (Fried-Goldman-Hirsch,1981[26]).
A compact radiant affine manifold $(M,\nabla)$ does not admit a nonzero parallel $1$-form.
Theorem 3.15 (Fried-Goldman-Hirsch,1981[26]).
A compact radiant affine manifold $(M,\nabla)$ of dimension $n$ does not admit a nonzero parallel $n$-form.
Adapting the theorems above to our framework immediately yields the following corollaries.
Corollary 3.16.
Let $(M,\nabla)$ be a compact radiant manifold. Then the Euler vector field of $(M,\nabla)$ is unique.
Proof.
Combine Theorem 3.14 with Remark 3.9. ∎
Corollary 3.17.
Let $(M,\nabla)$ be a compact radiant manifold. Then the canonical line bundle $K_{\nabla}$ of $(M,\nabla)$ is non-trivial.
Proof.
Combine Theorem 3.15 with Corollary 2.37. ∎
Proposition 3.18.
Let $((M,\nabla),h)$ be a compact radiant Hessian manifold. Then the Riemannian manifold $(M,h)$ is not flat.
Proof.
By considering the orientation double cover of $M$, without loss of generality we may assume $M$ is orientable. Now the desired result follows immediately from Theorem 3.15 and Theorem 2.44. ∎
We shall also recall the core concept of information geometry, that is, the notion of dual connections.
Definition 3.19.
Let $(M,g)$ be a pseudo-Riemannian manifold. For a connection $D$ on the tangent bundle $TM\rightarrow M$, its conjugate connection $D^{*}$ with respect to $g$ is defined to be the unique affine connection on the tangent bundle of $M$ such that the equation
$Zg(X,Y)=g\left(D_{Z}X,Y\right)+g\left(X,D_{Z}^{*}Y\right)$
holds for all smooth vector fields $X,Y,Z$ on $M$.
See [11] for a proof of the fact that the conjugate connection $D^{*}$ is well-defined and that $D^{**}=D$ for every connection $D$ on the tangent bundle $TM\rightarrow M$.
Theorem 3.20.
Let $(M,\nabla)$ be an affine manifold, and let $g$ be a pseudo-Riemannian metric on $(M,\nabla)$. Then $g$ is a Hessian metric on $(M,\nabla)$ if and only if $\nabla^{*}$ is also flat and torsion-free, or equivalently, the $(0,3)$-tensor $\nabla g$ is totally symmetric. Moreover, in this case
$\bar{\nabla}:=\frac{1}{2}\left(\nabla^{*}+\nabla\right)$
is the Levi-Civita connection of $(M,g)$.
Proof.
See [15] for a proof of this statement. ∎
Lemma 3.21.
Let $(M,g)$ be a pseudo-Riemannian manifold, and let $\nabla$ be a connection on the tangent bundle $TM\rightarrow M$. Then for any 1-form $\theta\in\Omega^{1}(M)$ and any vector field $X$ on $M$, it holds that $\nabla_{X}(\theta^{\sharp})=(\nabla^{*}_{X}\theta)^{\sharp}$, and in particular $\nabla^{*}\theta=0$ if and only if $\nabla(\theta^{\sharp})=0$, where $\sharp$ is the musical isomorphism.
Proof.
Take arbitrary vector fields $X,Y$ on $M$. Then
$(\nabla^{*}_{X}\theta)(Y)=X(\theta(Y))-\theta(\nabla^{*}_{X}Y).$
Now, since $\theta(Y)=g(\theta^{\sharp},Y)$, we have
$(\nabla^{*}_{X}\theta)(Y)=X(g(\theta^{\sharp},Y))-g(\theta^{\sharp},\nabla^{*}_{X}Y).$
Moreover, since
$X(g(\theta^{\sharp},Y))=g(\nabla_{X}\theta^{\sharp},Y)+g(\theta^{\sharp},\nabla^{*}_{X}Y),$
substituting this into the previous expression, we obtain
$(\nabla^{*}_{X}\theta)(Y)=(g(\nabla_{X}\theta^{\sharp},Y)+g(\theta^{\sharp},\nabla^{*}_{X}Y))-g(\theta^{\sharp},\nabla^{*}_{X}Y)=g(\nabla_{X}\theta^{\sharp},Y).$
This proves that $\nabla_{X}\theta^{\sharp}=(\nabla^{*}_{X}\theta)^{\sharp}$ holds for all vector fields $X$ on $M$. ∎
The following two fundamental theorems illustrate the duality between radiant Hessian manifolds and Hessian manifolds of Koszul type.
Proposition 3.22.
Let $((M,\nabla),h)$ be a Hessian manifold of Koszul type. Then $((M,\nabla^{*}),h)$ is a radiant Hessian manifold. In fact, if $\eta$ is a $1$-form potential of $((M,\nabla),h)$, then $\eta^{\sharp}$ is an Euler vector field of $(M,\nabla^{*})$, where $\sharp$ is the musical isomorphism.
Proof.
Consider affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. For each $i=1,\dots,n$, define $p_{i}:=\eta(\partial/\partial x^{i})$, so that
$\eta=\sum_{i=1}^{n}p_{i}dx^{i}.$
Since $\partial/\partial x^{i}$ is parallel, we obtain that
$h=\nabla\eta=\sum_{i=1}^{n}dp_{i}\otimes dx^{i}.$
Then for every $i,j\in\{1,\dots,n\}$ we have that $h(\partial/\partial x^{i},\partial/\partial p_{j})=\delta^{i}_{j}$ and hence
$h(\nabla^{*}_{X}(\partial/\partial p_{j}),\partial/\partial x^{i})=$ $h(\nabla_{X}(\partial/\partial x^{i}),\partial/\partial p_{j})+h(\partial/\partial x^{i},\nabla^{*}_{X}(\partial/\partial p_{j}))$
$=$ $Xh(\partial/\partial x^{i},\partial/\partial p_{j})$
$=$ $X\delta^{i}_{j}$
$=$ $0$
for every vector field $X$ on $M$. Therefore $(dx^{i})^{\sharp}=\partial/\partial p_{i}$ and $\nabla^{*}(\partial/\partial p_{i})=0$ for all $i=1,\dots,n$. This proves that
$\eta^{\sharp}=\sum_{i=1}^{n}p_{i}(dx^{i})^{\sharp}=\sum_{i=1}^{n}p_{i}\frac{\partial}{\partial p_{i}}$
and it readily seen that the latter is an Euler vector field of $(M,\nabla^{*})$. ∎
Proposition 3.23.
Let $((M,\nabla),h)$ be a radiant Hessian manifold. Then $((M,\nabla^{*}),h)$ is a Hessian manifold of Koszul type. In fact, if $H$ is an Euler vector field of $(M,\nabla)$, then $H^{\flat}$ is a $1$-form potential of $((M,\nabla^{*}),h)$, where $\flat$ is the musical isomorphism.
Proof.
Consider radiant affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. For each $i=1,\dots,n$, since $\nabla(dx^{i})=0$, by Lemma 3.21 we have $\nabla^{*}(\partial/\partial x^{i})^{\flat}=0$. Recall that parallel 1-forms on an affine manifold is closed. By Poincaré’s lemma, there exist local coordinates $p_{1},\dots,p_{n}$ of $M$, such that $dp_{i}=(\partial/\partial x^{i})^{\flat}$ for all $i=1,\dots,n$. Therefore we arrive at the expression
$h=\sum_{i=1}^{n}dp_{i}\otimes dx^{i}$
for the Hessian metric $h$.
Since $x^{1},\dots,x^{n}$ are radiant affine, it is readily seen that the vector field
$X:=\sum_{i=1}^{n}x^{i}\frac{\partial}{\partial x^{i}}$
is a well-defined Euler vector field of $(M,\nabla)$. By the uniqueness of Euler vector field, there exists a parallel vector field $Y$ on $(M,\nabla)$, such that $H=X+Y$. Therefore, again by Lemma 3.21, we obtain that
$\nabla^{*}(H^{\flat})=\nabla^{*}(Y^{\flat})+\nabla^{*}\sum_{i=1}^{n}x^{i}(\partial/\partial x^{i})^{\flat}=0+\sum_{i=1}^{n}dx^{i}\otimes dp_{i}=\sum_{i=1}^{n}dp_{i}\otimes dx^{i}=h$
as required. ∎
Corollary 3.24.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type. Then the Hessian manifold $((M,\nabla^{*}),h)$ is also radiant of Koszul type.
Proof.
This follows immediately from Proposition 3.22 and Proposition 3.23. ∎
Now, in view of Corollary 3.24, for a radiant Hessian manifold $((M,\nabla),h)$ of Koszul type, it is natural to ask for an explicit expression for the $1$-form potential for its dual $((M,\nabla^{*}),h)$. We will now give an satisfactory answer to this question.
For a vector field $X$, as per usual, we denote by $\mathcal{L}_{X}$ the Lie derivative along $X$.
Definition 3.25.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type, with Euler vector field $H$ and $1$-form potential $\eta$. The differential form $\eta^{*}:=-\eta+\mathcal{L}_{H}\eta$ is said to be the Legendre transform of $\eta$ along direction $H$.
Theorem 3.26.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type, with Euler vector field $H$ and $1$-form potential $\eta$. Then the Legendre transform $\eta^{*}$ of $\eta$ along direction $H$ satisfies $h=\nabla^{*}\eta^{*}$.
Proof.
Consider radiant affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$ so that
$H=\sum_{i=1}^{n}x^{i}\frac{\partial}{\partial x^{i}}$
is Eulerian. For each $i\in\{1,\dots,n\}$ define $p_{i}:=\eta(\partial/\partial x^{i})$. Then
$\eta=\sum_{i=1}^{n}p_{i}dx^{i}$
and $\nabla^{*}(dp_{1})=\cdots=\nabla^{*}(dp_{n})=0$. By Cartan’s magic formula, we obtain
$\eta^{*}=-\eta+\mathcal{L}_{H}\eta=d(\eta(H))-\eta=d(x^{1}p_{1}+\cdots+x^{n}p_{n})-\sum_{i=1}^{n}p_{i}dx^{i}=\sum_{i=1}^{n}x^{i}dp_{i}$
and hence
$\nabla^{*}\eta^{*}=\sum_{i=1}^{n}dx^{i}\otimes dp_{i}=h$
as required. ∎
Remark 3.27.
If the Hessian manifold $((M,\nabla),h)$ under consideration satisfies
  1. 1.
    $M=\Omega$ is a convex domain in $\mathbb{R}^{d}$;
  2. 2.
    $\nabla$ is the Levi-Civita connection of the $d$-dimensional flat Euclidean space;
  3. 3.
    $h$ is the Hessian quadratic form of a strictly convex potential,
then the construction in Definition 3.25 recovers the Legendre transformation in classical mechanics.
3.3 Space-like Lagrangian Embeddings
In this section, we characterize Hessian manifolds of Koszul type as space-like Lagrangians, and then study their minimality as submanifolds of their own cotangent bundles.
Throughout this section we use Einstein’s summation convention.
Definition 3.28 (Patterson-Walker,1952[16]).
For a differentiable manifold $M$ and a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$ with Christoffel symbol $\Gamma^{k}_{ij}$, the pseudo-Riemannian metric
$g:=\frac{1}{2}(dx^{i}\otimes dp_{i}+dp_{i}\otimes dx^{i})-\Gamma^{k}_{ij}p_{k}dx^{i}dx^{j}$
on $T^{*}M$ is called the Riemann extension of $\nabla$, where $x^{1},\dots,x^{n},p_{1},\dots,p_{n}$ here are canonical coordinates on $T^{*}M$.
Note that, the Riemann extension of a flat torsion-free connection is flat of neutral signature.
Definition 3.29.
For a differentiable manifold $M$ and a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$, a $1$-form $\eta$ on $M$, i.e. a smooth section $\eta:M\rightarrow T^{*}M$ of the cotangent bundle $T^{*}M\rightarrow M$, is said to be a space-like Lagrangian embedding of $(M,\nabla)$, if $\eta:M\rightarrow T^{*}M$ is an embedded Lagrangian submanifold of $(T^{*}M,\omega)$ and the symmetric $(0,2)$-tensor $\eta^{*}g$ is positive definite, where $\omega$ is the standard symplectic form on $T^{*}M$ and $g$ is the Riemann extension of $\nabla$.
Remark 3.30.
Let $M$ be a differentiable manifold, and $\nabla$ a torsion-free connection on the tangent bundle $TM\rightarrow M$. Then any space-like Lagrangian embedding of $(M,\nabla)$ is by definition an exact Lagrangian, and hence admits a vanishing Maslov class $\mu\in\mathrm{H}^{2}(M;\mathbb{Z})$.
The following theorem shows that, in the category of affine manifolds, space-like embedded Lagrangians are precisely those Hessian manifolds of Koszul type.
Theorem 3.31.
Let $M$ be a differentiable manifold, and let $\nabla$ be a torsion-free connection $\nabla$ on the tangent bundle $TM\rightarrow M$. Then, for a $1$-form $\eta$ on $M$, the following statements are equivalent:
  1. 1.
    $\eta$ is a space-like Lagrangian embedding of $(M,\nabla)$;
  2. 2.
    $\eta$ is closed, and $\nabla\eta>0$.
Moreover, for a closed $1$-form $\eta$, the $(0,3)$-tensor $\nabla(\nabla\eta)$ is totally symmetric if and only if $R^{m}_{ijk}\eta_{m}=0$, where $R$ here is the curvature tensor of $\nabla$.
Proof.
Recall from differential topology that a smooth cross-section of a fiber bundle is always an embedding.
Denote by $\tau$ the tautological 1-form of $T^{*}M$, and denote by $\omega$ the standard symplectic form on $T^{*}M$. Then we have $\omega=-d\tau$ and $\eta^{*}\tau=\eta$. Since $d\eta=d(\eta^{*}\tau)=\eta^{*}d\tau=-\eta^{*}\omega$, we obtain that $\eta\in\Omega^{1}(M)$ is closed if and only if $\eta:M\rightarrow T^{*}M$ is a Lagrangian embedding.
Consider canonical coordinates $x^{1},\dots,x^{n},p_{1},\dots,p_{n}$ on $T^{*}M$. Denote by $g$ the Riemann extension of $\nabla$. Now, it suffices to notice that, by the very definition, we have
$\nabla\eta=(\partial_{i}\eta_{j}-\Gamma^{k}_{ij}\eta_{k})dx^{i}\otimes dx^{j}=d\eta_{i}\otimes dx^{i}-\Gamma^{k}_{ij}\eta_{k}dx^{i}dx^{j}=\eta^{*}g$
where $\Gamma^{k}_{ij}$ is the Christoffel symbol of $\nabla$. Denote $h:=\nabla\eta$ and denote $e_{i}:=\partial/\partial x^{i}$. Then
$\nabla_{e_{k}}h=(\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik})dx^{i}\otimes dx^{j}.$
Therefore, $\nabla h$ is totally symmetric if and only if $\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik}$ is invariant under permutations of indices $i,j,k$. Define $S^{m}_{ijk}:=\partial_{k}\Gamma^{m}_{ij}-\Gamma^{m}_{il}\Gamma^{l}_{jk}-\Gamma^{m}_{jl}\Gamma^{l}_{ik}$. Then a straightforward computation yields that
$\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik}=\partial_{i}\partial_{j}\eta_{k}-S^{m}_{ijk}\eta_{m}-(\Gamma^{m}_{ij}\partial_{k}\eta_{m}+\Gamma^{m}_{jk}\partial_{i}\eta_{m}+\Gamma^{m}_{ik}\partial_{j}\eta_{m}).$
Notice that $\partial_{i}\partial_{j}\eta_{k}$ and $\Gamma^{m}_{ij}\partial_{k}\eta_{m}+\Gamma^{m}_{jk}\partial_{i}\eta_{m}+\Gamma^{m}_{ik}\partial_{j}\eta_{m}$ are already invariant under permutations of indices $i,j,k$. Therefore, the expression $\partial_{k}h_{ij}-h_{il}\Gamma^{l}_{jk}-h_{jl}\Gamma^{l}_{ik}$ is invariant under permutations of indices $i,j,k$ if and only if $S^{m}_{ijk}\eta_{m}$ is. But, by construction, we have $S^{m}_{ijk}-S^{m}_{ikj}=-R^{m}_{ijk}$. This concludes the proof. ∎
Corollary 3.32.
Let $(M,\nabla)$ be an affine manifold, and $\eta$ a $1$-form on $M$. Then, the following statements are equivalent:
  1. 1.
    $\eta$ is a space-like Lagrangian embedding of $(M,\nabla)$;
  2. 2.
    $\eta$ is closed, and $\nabla\eta>0$, i.e. $((M,\nabla),\nabla\eta)$ is a Hessian manifold of Koszul type.
Proof.
Recall Theorem 3.20 and then apply Theorem 3.31. ∎
We remark that, in [27], J.Zhang and G.Khan derived a similar condition for totally symmetry of the $(0,3)$-tensor described in Theorem 3.31.
Now, one may naturally ask how a space-like Lagrangian is bended inside the ambient space, and under what condition a space-like Lagrangian is a minimal submanifold. To address this question, we first derive a formula for the Laplace-Beltrami operator on a Hessian manifold, and then study the submanifold geometry of space-like Lagrangian embeddings.
Lemma 3.33.
Let $((M,\nabla),h)$ be a Hessian manifold, and let $\Delta_{h}$ be the Laplace-Beltrami operator of $(M,h)$. Then, in affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$, we have
$\Delta_{h}\phi=h^{ij}(\partial_{i}\partial_{j}\phi-\varkappa_{i}\partial_{j}\phi)$
for all $\phi\in C^{\infty}(M)$, where $\partial_{i}:=\partial/\partial x^{i}$ for $i\in\{1,\dots,n\}$ and $\varkappa\in\Omega^{1}(M)$ is the first Koszul form of $((M,\nabla),h)$.
Proof.
By the very definition of the Laplace-Beltrami operator, we have
$\Delta_{h}\phi=\frac{1}{\sqrt{|h|}}\frac{\partial}{\partial x^{i}}\left(\sqrt{|h|}h^{ij}\partial_{j}\phi\right)=h^{ij}\partial_{i}\partial_{j}\phi+\frac{\partial_{j}\phi}{\sqrt{|h|}}\frac{\partial}{\partial x^{i}}\left(\sqrt{|h|}h^{ij}\right)$
for all $\phi\in C^{\infty}(M)$.
Recall form Riemannian geometry that
$h^{ij}\Gamma^{k}_{ij}=-\frac{1}{\sqrt{|h|}}\frac{\partial(\sqrt{|h|}h^{kl})}{\partial x^{l}}$
where $\Gamma^{k}_{ij}=\frac{1}{2}h^{lk}\left(\partial_{i}h_{jl}+\partial_{j}h_{il}-\partial_{l}h_{ij}\right)$ for $i,j,k\in\{1,\dots,n\}$. But, since $h_{ij}=h_{ji}=\partial_{i}\eta_{j}$, we have that
$\partial_{i}h_{jk}=\partial_{j}h_{ik}=\partial_{k}h_{ij}$
and hence $\Gamma^{k}_{ij}=\frac{1}{2}h^{kl}\partial_{i}h_{jl}$ for all $i,j,k\in\{1,\dots,n\}$. Therefore, we arrive at
$2h^{ij}\Gamma^{k}_{ij}=h^{ij}h^{kl}\partial_{i}h_{jl}=h^{kl}h^{ij}\partial_{l}h_{ij}=2h^{kj}\Gamma^{i}_{ij}$
for all $k=1,\dots,n$. Since $\varkappa=d\operatorname{log}(\sqrt{|h|})=\Gamma^{i}_{ij}dx^{j}$ is the first Koszul form of $((M,\nabla),h)$, we finally obtain
$\Delta_{h}\phi=h^{ij}(\partial_{i}\partial_{j}\phi-\varkappa_{i}\partial_{j}\phi)$
for all $\phi\in C^{\infty}(M)$. ∎
By establishing a Witten type Laplacian equation on the Hessian manifold under consideration, the following theorem characterize all space-like Lagrangians that are minimal submanifolds.
Theorem 3.34.
Let $(M,\nabla)$ be an $n$-dimensional affine manifold, and $\eta$ a space-like Lagrangian embedding of $(M,\nabla)$, so that $h:=\nabla\eta>0$. Let $\operatorname{Ric}$ be the Ricci curvature of $(M,h)$ and let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((M,\nabla),h)$. Then $\eta(M)$ is a minimal submanifold of $(T^{*}M,g)$ if and only if
$2\Delta\eta=\varkappa+\operatorname{Ric}(\eta^{\sharp},-)$
where $\Delta$ is the Hodge Laplacian and $\sharp$ is the musical isomorphism induced by the Riemann extension $g$ of $\nabla$.
Proof.
Consider affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$, and extend to canonical coordinates $x^{1},\dots,x^{n},p_{1},\dots,p_{n}$ of $(T^{*}M,\omega)$, where $\omega$ is the standard symplectic form of $T^{*}M$. Also, write $\eta=\eta_{i}dx^{i}$. Then it is readily seen that there exists a local frame of tangent vectors on $\eta(M)$ consists of
$e_{i}:=\frac{\partial}{\partial x^{i}}+\frac{\partial\eta_{j}}{\partial x^{i}}\frac{\partial}{\partial p_{j}}$
for $i=1,\dots,n$. Denote $\partial_{i}:=\partial/\partial x^{i}$ for $i\in\{1,\dots,n\}$ and denote by $D$ the Levi-Civita connection of $(T^{*}M,g)$. Straightforward computation then yields
$D_{e_{i}}e_{j}=(\partial_{i}\partial_{j}\eta_{k})\partial/\partial p_{k}$
and hence the second fundamental form $\operatorname{II}$ of $\eta\colon M\rightarrow T^{*}M$ satisfies
$\operatorname{II}(e_{i},e_{j})=(\partial_{i}\partial_{j}\eta_{k})\partial/\partial p_{k}-\frac{1}{2}h^{kl}(\partial_{i}\partial_{j}\eta_{k})e_{l}$
for all $i,j\in\{1,\dots,n\}$. Therefore the mean curvature vector of submanifold $\eta(M)$ in $(T^{*}M,g)$ is precisely
$\frac{1}{n}h^{ij}\operatorname{II}(e_{i},e_{j})=\frac{1}{2n}(2h^{ij}(\partial_{i}\partial_{j}\eta_{k})\partial/\partial p_{k}-h^{ij}h^{kl}(\partial_{i}\partial_{j}\eta_{k})e_{l}).$
Since $e_{1},\dots,e_{n},\partial/\partial p_{1},\dots,\partial/\partial p_{n}$ are $\mathbb{R}$-linearly independent, the submanifold $\eta(M)$ is minimal in $(T^{*}M,g)$ if and only if $h^{ij}\partial_{i}\partial_{j}\eta_{k}=0$ for all $k\in\{1,\dots,n\}$, that is, the $1$-form $\theta:=h^{ij}(\partial_{i}\partial_{j}\eta_{k})dx^{k}$ vanishes.
Denote by $\Delta_{h}$ the Laplace-Beltrami operator of $(M,h)$. Then by Lemma 3.33, we have that
$\Delta_{h}(\eta_{k})=h^{ij}(\partial_{i}\partial_{j}\eta_{k}-\varkappa_{i}\partial_{j}\eta_{k})=h^{ij}\partial_{i}\partial_{j}\eta_{k}-h^{ij}\varkappa_{i}h_{jk}=-\varkappa_{k}+h^{ij}\partial_{i}\partial_{j}\eta_{k}$
for all $k\in\{1,\dots,n\}$. Now, recall that Weitzenböck formula reads
$(\Delta\eta)_{k}=-\Delta_{h}(\eta_{k})+\nabla_{k}(\operatorname{div}(\eta^{\sharp}))+h^{ij}\eta_{j}\operatorname{Ric}_{ik}.$
By carefully rearranging terms, we arrive at
$\theta=-2\Delta\eta+\operatorname{Ric}(\eta^{\sharp},-)+\varkappa.$
Therefore, $\eta(M)$ is a minimal submanifold of $(T^{*}M,g)$ if and only if
$2\Delta\eta=\varkappa+\operatorname{Ric}(\eta^{\sharp},-)$
as required. ∎
The beautiful Theorem 3.34 shows that the extrinsic minimality condition is equivalent to an intrinsic Witten type Laplacian equation on the Hessian manifold itself, which reveals that the minimality of the space-like Lagrangian embedding is governed by a delicate balance between the curvature property and the second-order behavior of the $1$-form potential.
3.4 Para-Kähler Host Theorem
In this section, we shall finally prove the para-Kähler host theorem, which generalizes the concept of space-like Lagrangian embeddings into cotangent bundles, and at the same time gives another powerful construction of general Hessian manifolds.
We first formally introduce flat para-Kähler manifolds and the space-like Lagrangian transversals as their very special submanifolds.
Definition 3.35.
A para-Kähler manifold $((M,\omega),g)$ is said to be flat if the underlying pseudo-Riemannian manifold $(M,g)$ is flat.
Definition 3.36.
A Lagrangian immersion $\varphi\colon L\rightarrow M$ into a para-Kähler manifold $((M,\omega),g)$ is said to be space-like if $\varphi^{*}g>0$, and is said to be a transversal if $\varphi$ is transverse to at least one of the bi-Lagrangian foliations of $((M,\omega),g)$.
On a flat para-Kähler manifold, we shall be interested in a specific kind of Darboux coordinate system.
Lemma 3.37.
Let $((M,\omega),g)$ be a flat para-Kähler manifold, and let $D$ be the Levi-Civita connection of $(M,g)$. Then, for any $p\in M$, there exists Darboux coordinates $x^{1},\dots,x^{n},y_{1},\dots,y_{n}$ of $(M,\omega)$ at the vicinity of $p\in M$ such that
$g=\frac{1}{2}\sum_{i=1}^{n}(dx^{i}\otimes dy_{i}+dy_{i}\otimes dx^{i})$
and $x^{1},\dots,x^{n},y_{1},\dots,y_{n}$ are locally defined affine functions on $(M,D)$.
Proof.
Since $(M,g)$ is flat, we have that $(M,D)$ is affine. Consider any affine coordinates $u^{1},\dots,u^{2n}$ of $(M,D)$.
Since $((M,\omega),g)$ is para-Kähler, we have $D\omega=0$, and hence the tensor components $\omega_{ij}$ are all constant in coordinates $u^{1},\dots,u^{2n}$. Now, by linear algebra, up to an linear change of coordinates, we may reduce the symplectic matrix of which $(i,j)$-entry equals $\omega_{ij}$ to its canonical form. In particular, there exists Darboux coordinates $x^{1},\dots,x^{n},y_{1},\dots,y_{n}$ of $(M,\omega)$ that are also affine coordinates of $(M,D)$.
Let $J$ be the para-complex structure of $((M,\omega),g)$. Then, by the properties of Darboux coordinates, we have that $\partial/\partial x^{1},\dots,\partial/\partial x^{n}$ are the eigenvectors of $J$ with eigenvalue $1$, and that $\partial/\partial y_{1},\dots,\partial/\partial y_{n}$ are the eigenvectors of $J$ with eigenvalue 1-1. Properties of the compatible triple then yield that
$g=\frac{1}{2}\sum_{i=1}^{n}(dx^{i}\otimes dy_{i}+dy_{i}\otimes dx^{i})$
as required. ∎
Definition 3.38.
For a flat para-Kähler manifold, Darboux coordinates described in Lemma 3.37 are termed affine Darboux coordinates.
We then recall from last section that Hessian manifolds of Koszul type are naturally space-like Lagrangian transversals in flat para-Kähler manifolds.
Corollary 3.39.
Let $((L,\nabla),h)$ be a Hessian manifold of Koszul type. Then there exists an isometric transverse Lagrangian embedding of $(L,h)$ into some flat para-Kähler manifold $((M,\omega),g)$, such that the affine coordinates on $(L,\nabla)$ extends to an affine Darboux coordinates on $((M,\omega),g)$
Proof.
Denote by $\omega$ the standard symplectic form of $T^{*}L$ and denote by $g$ the Riemann extension of $\nabla$. Take any $1$-form potential $\eta$ of $((L,\nabla),h)$. Then, by Corollary 3.32, the smooth mapping $\eta\colon L\rightarrow T^{*}L$ satisfies all the required properties. ∎
As a final preparation, we investigate the transition functions on flat para-Kähler manifolds.
Lemma 3.40.
Let $((M,\omega),g)$ be a flat para-Kähler manifold, and let $D$ be the flat Levi-Civita connection of $(M,g)$. Suppose that both $x^{1},\dots,x^{n},y_{1},\dots,y_{n}$ and $\tilde{x}^{1},\dots,\tilde{x}^{n},\tilde{y}_{1},\dots,\tilde{y}_{n}$ are affine Darboux coordinates of $((M,\omega),g)$ at the vicinity of $p\in M$. Then there exist locally defined affine functions $u_{1},\dots u_{n}$ on $(M,D)$ such that $\tilde{x}^{i}=u_{i}(x^{1},\dots,x^{n})$ for $i=1,\dots,n$.
Proof.
Recall that, by the very definition of affine Darboux coordinates, the transition function from $x^{1},\dots,x^{n},y_{1},\dots,y_{n}$ to $\tilde{x}^{1},\dots,\tilde{x}^{n},\tilde{y}_{1},\dots,\tilde{y}_{n}$ is simultaneously an affine isometry and a symplectomorphism. A computation using block matrices then yields the desired formula. ∎
We can now prove the para-Kähler host theorem, which states that space-like Lagrangian transversals in flat para-Kähler manifolds are Hessian manifolds.
Theorem 3.41.
Let $((M,\omega),g)$ be a flat para-Kähler manifold. Then for any space-like transversal Lagrangian embedding $\varphi\colon L\rightarrow M$ there exists an affine structure $\nabla$ on $L$ such that $((L,\nabla),\varphi^{*}g)$ is a Hessian manifold.
Proof.
Let $x^{1},\dots,x^{n},y_{1},\dots,y_{n}$ be affine Darboux coordinates of $((M,\omega),g)$ at vicinity of $p\in M$. Then, by Weinstein’s Lagrangian neighbourhood theorem and the transversality assumption, there exists a smooth function germ $f$ at $p\in M$ such that $L$ is locally defined by equations
$y_{1}=f_{1}(x^{1},\dots,x^{n}),\cdots,y_{n}=f_{n}(x^{1},\dots,x^{n})$
at vicinity of $p\in M$, where $f_{i}:=\partial f/\partial x^{i}$ for all $i\in\{1,\cdots,n\}$. Now, by Lemma 3.40, we have that the connection on the tangent bundle $TM\rightarrow M$ defined by
$\nabla(dx^{1})=\cdots=\nabla(dx^{n})$
is an affine structure on $L$. Moreover, the identity
$\varphi^{*}g=\sum_{i=1}^{n}df_{i}\otimes dx^{i}=\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{\partial^{2}f}{\partial x^{i}\partial x^{j}}dx^{i}\otimes dx^{j}$
shows that $\varphi^{*}g$ is a Hessian metric on $(L,\nabla)$. ∎
Remark 3.42.
Together with the examples provided by Corollary 3.39, the para-Kähler host theorem, i.e. Theorem 3.41, are of significant interest, as they greatly generalize the ambient space of space-like Lagrangians, allowing them to be compact, and therefore initiate the study of global symplectic theory of Hessian manifolds.
3.5 Hyperbolic Affine Manifolds
For completeness, we recall yet another necessary and sufficient condition for a compact affine manifold to admit a Hessian metric of Koszul type.
Definition 3.43.
A convex domain $\Omega$ in $\mathbb{R}^{d}$ is said to be regular, if it contains no straight line, i.e. a $1$-dimensional affine subspace of $\mathbb{R}^{d}$.
Definition 3.44.
Let $(M,\nabla)$ be an affine manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. The affine manifold $(M,\nabla)$ is said to be hyperbolic, if the affine development $\Phi\colon\tilde{M}\rightarrow\mathbb{R}^{n}$ of $(\tilde{M},\pi^{*}\nabla)$ is a diffeomorphism onto its image, and its image is a regular convex domain in $\mathbb{R}^{n}$.
Theorem 3.45.
A compact affine manifold is hyperbolic if and only if it admits a Hessian metric of Koszul type.
Proof.
This theorem is proved in [15]. ∎
To summarize, we suggest the following theorem:
Theorem 3.46.
For a compact affine manifold, the following are equivalent:
  1. 1.
    it is hyperbolic;
  2. 2.
    it admits a Hessian metric of Koszul type;
  3. 3.
    it admits a positive flat line bundle;
  4. 4.
    it admits a space-like Lagrangian embedding.
Proof.
Combine Proposition 2.503.32 with Theorem 3.45. ∎
We shall now introduce an important property of compact hyperbolic affine manifolds.
[…]
Definition 3.47.
The first Koszul class of the canonical line bundle of an affine manifold $(M,\nabla)$ is said to be the canonical class of $(M,\nabla)$.
We recall that, by Proposition 2.40, when the affine manifold under consideration is equipped with a Hessian metric, its first Koszul form represents precisely the canonical class.
The following corollary is natural.
Corollary 3.48.
Let $((M,\nabla),h)$ be a complete oriented Hessian manifold. Suppose that the canonical class of $(M,\nabla)$ is $K$. Then the canonical class of $(M,\nabla^{*})$ is K-K.
Proof.
The desired result follows immediately from Proposition 2.40 & 2.33 and the fact that the determinant of $h$ in affine coordinates of $(M,\nabla^{*})$ is the reciprocal of the determinant of $h$ in affine coordinates of $(M,\nabla)$. ∎
We shall show that, on an oriented compact manifold, an hyperbolic affine structure can always be perturbed into another hyperbolic affine structure of which canonical class is rational.
To proceed, we adapt a theorem of J.-L.Koszul to our setting.
Theorem 3.49 (Koszul, 1968[17]).
[…]
Applying Theorem 3.49 together with group cohomologies with Lie algebra coefficients, we can prove the following density theorem:
Theorem 3.50.
Let $(M,\nabla)$ be a compact oriented hyperbolic affine manifold. Then there exists a flat and torsion-free connection $\nabla^{\prime}$ on the tangent bundle $TM\rightarrow M$ such that $(M,\nabla^{\prime})$ is a hyperbolic affine manifold of which canonical class is rational.
Proof.
[…]
$\operatorname{Hom}(G,\operatorname{GL}^{+}_{n}(\mathbb{R}))\rightarrow\operatorname{Hom}(G,\mathbb{R}^{+}),\quad\rho\mapsto(g\mapsto\operatorname{det}(\rho(g)))$
is a submersion, or equivalently, the induced map
$\operatorname{tr}_{*}\colon\mathrm{H}^{1}(G;\mathfrak{gl}_{n}(\mathbb{R}))\rightarrow\mathrm{H}^{1}(G;\mathbb{R})$
of the trace homomorphism
$\operatorname{tr}\colon\mathfrak{gl}_{n}(\mathbb{R})\rightarrow\mathbb{R}$
is surjective, which is a direct consequence of the fact that the short exact sequence
$0\rightarrow\mathfrak{sl}_{n}(\mathbb{R})\rightarrow\mathfrak{gl}_{n}(\mathbb{R})\rightarrow\mathbb{R}\rightarrow 0$
splits. This concludes the proof. ∎
4 Geometric Analysis of Hessian Metrics
4.1 Effects on Curvature Tensors
While previous sections focus on the topology of Hessian manifolds, in this section we delve into geometric aspect.
Throughout this section we use Einstein’s summation convention.
The following notion is classical in information geometry.
Definition 4.1.
The Amari–Chentsov tensor $A$ of a Hessian manifold $((M,\nabla),h)$ is defined to be the $(0,3)$-tensor satisfying
$A(X,Y,Z)=(\nabla_{Z}h)(X,Y)=h(\nabla^{*}_{X}Y-\nabla_{X}Y,Z)$
for all vector fields $X,Y,Z$ on $M$.
We first note that the Riemann curvature tensor of a Hessian manifold $((M,\nabla),h)$ enjoys a simplified form, which also provides a trivial bound for Ricci curvature in terms of the Amari–Chentsov tensor.
Lemma 4.2.
Let $((M,\nabla),h)$ be a Hessian manifold, and $A$ its Amari–Chentsov tensor. Then $A$ is totally symmetric and
$4R_{ijkl}=A_{il}^{m}A_{jkm}-A_{ik}^{m}A_{jlm},$
where $A_{ij}^{k}:=A_{ijl}h^{kl}$, and $R$ is the Riemann curvature tensor of $(M,h)$.
Proof.
See for example [28] for a proof. ∎
The followings are immediate consequences of Lemma 4.2.
Corollary 4.3.
Let $((M,\nabla),h)$ be an $n$-dimensional Hessian manifold, and let $A$ be its Amari–Chentsov tensor. Then, in an orthonormal frame, the Riemann curvature tensor $R$ and Ricci tensor $\operatorname{Ric}$ of $(M,h)$ satisfies $4R_{ijkl}=[A_{i},A_{j}]_{kl}$ and
$4\operatorname{Ric}_{ij}=\operatorname{tr}(A_{i}A_{j})-\sum_{k=1}^{n}\operatorname{tr}(A_{k})A_{ijk}$
where $A_{k}$ is the $n\times n$ matrix of which $(i,j)$-th entry is $A_{ijk}$.
Corollary 4.4.
Let $f$ be a smooth function defined on a domain in $\mathbb{R}^{2}$. Suppose that
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a pseudo-Riemannian metric. Then Gaußian curvature $K$ of the metric $g$ satisfies
$K=-\frac{\left|\begin{array}[]{ccc}f_{xx}&f_{xxx}&f_{xxy}\\ f_{xy}&f_{xxy}&f_{xyy}\\ f_{yy}&f_{xyy}&f_{yyy}\end{array}\right|}{4\left|\begin{array}[]{cc}f_{xx}&f_{xy}\\ f_{xy}&f_{yy}\end{array}\right|^{2}}=\frac{f_{xx}\{f_{yy},f_{xy}\}+f_{yy}\{f_{xy},f_{xx}\}+f_{xy}\{f_{xx},f_{yy}\}}{4(f_{xx}f_{yy}-f_{xy}f_{xy})^{2}}$
where $\{\cdot,\cdot\}$ is the standard Poisson bracket normalized by $\{x,y\}=1$.
Corollary 4.5.
Let $((M,\nabla),h)$ be an $n$-dimensional Hessian manifold, and let $A$ be its Amari–Chentsov tensor. Then the Ricci tensor $\operatorname{Ric}$ of $M$ satisfies
$\frac{n-1}{2}|A|^{2}h(X,X)\leq\operatorname{Ric}(X,X)\leq\frac{n-1}{4}|A|^{2}h(X,X)$
for every vector field $X$ on $M$.
Proof.
Apply the Cauchy–Schwarz inequality to Lemma 4.3. ∎
The following simple observation on sectional curvature shows that there exists restrictions on other Riemannian metrics that a Hessian manifold can possess.
Proposition 4.6.
Let $((M,\nabla),h)$ be a complete Hessian manifold. Then for any Riemannian metric $g$ on $M$, then there exists no positive real number $k>0$ such that the Ricci tensor $\operatorname{Ric}^{g}$ of $(M,g)$ satisfies $\operatorname{Ric}^{g}\geq k$.
Proof.
If the Ricci tensor satisfies $\operatorname{Ric}^{g}\geq k$ for some positive real number $k>0$, then by Bonnet–Myers theorem, $M$ is compact with finite fundamental group, which contradicts Corollary 2.7. ∎
Corollary 4.7.
Let $((M,\nabla),g)$ be a compact Hessian manifold. Then for any Riemannian metric $h$ on $M$, the Riemannian manifold $(M,h)$ is not of positive sectional curvature.
Proof.
This is a direct consequence of Proposition 4.6. ∎
We note a very important application of de Rham splitting theorem to Hessian manifolds, which leads to structure theorems for universal coverings of Hessian manifolds.
Proposition 4.8.
Let $((M,\nabla),h)$ be a $(n+1)$-dimensional compact Hessian manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. Then there exists a complete Riemannian metric $g$ on $\mathbb{R}^{n}$ such that $(\tilde{M},\pi^{*}h)$ is isometric to the product of Riemannian manifolds $(\mathbb{R}^{n},g)$ and $(\mathbb{R},dt^{2})$.
Proof.
By considering the orientation double cover of $M$, without loss of generality we may assume $M$ is orientable.
If $(M,h)$ is a flat Riemannian manifold, then $(\tilde{M},\pi^{*}h)$ is isometric to the $(n+1)$-dimensional flat Euclidean space, so the required property is trivially satisfied.
From now on, we assume that the Riemannian manifold $(M,h)$ is not flat. By Theorem 2.13, up to an diffeomorphism, we may assume further that $\tilde{M}=\mathbb{R}^{n+1}$. Denote by $\bar{\nabla}$ the Levi-Civita connection of $(M,h)$. Then $\pi^{*}\bar{\nabla}$ is the Levi-Civita connection of $(\mathbb{R}^{n+1},\pi^{*}h)$. By Proposition 2.43 and Theorem 2.44, the first Koszul form $\varkappa$ of $((M,\nabla),h)$ satisfies $\varkappa\neq 0$ and $\bar{\nabla}\varkappa=0$. Now, apply de Rham splitting theorem to the non-trivial parallel vector field $\pi^{*}\varkappa$ on the complete Riemannian manifold $(\mathbb{R}^{n+1},\pi^{*}h)$. ∎
In particular, we obtain the following immediate corollary:
Corollary 4.9.
Let $((M,\nabla),h)$ be a compact Hessian manifold of dimension $n\geq 2$. Then $(M,h)$ is not of negative sectional curvature.
Proof.
Let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. Assume the contrary that $(M,h)$ is of negative sectional curvature. Then $(\tilde{M},\pi^{*}h)$ is also of negative sectional curvature, but this contradicts Proposition 4.8. ∎
Corollary 4.9 successfully shows that a Hessian metric on a compact manifold cannot be of negative sectional curvature. But, actually, using technique of harmonic mappings and Bochner formula, we can further prove that a compact Hessian manifold cannot support any metric of negative sectional curvature.
Theorem 4.10.
Let $((M,\nabla),g)$ be a compact Hessian manifold of dimension $n\geq 2$. Then for any Riemannian metric $h$ on $M$, the Riemannian manifold $(M,h)$ is not of negative sectional curvature.
Proof.
By considering the orientation double cover of $M$, without loss of generality we may assume $M$ is orientable.
Assume the contrary that $h$ is a Riemannian metric on $M$ such that $(M,h)$ is of negative sectional curvature.
Suppose that $(M,g)$ is a flat Riemannian manifold. By Bieberbach theorem, there exists a finite covering map $\pi\colon\mathbb{T}^{n}\rightarrow M$. Therefore $(\mathbb{T}^{n},\pi^{*}h)$ is of negative sectional curvature. But then by Preissman’s theorem every non-trivial abelian subgroup of $\pi_{1}(\mathbb{T}^{n})=\mathbb{Z}^{n}$ is isomorphic to the inifinte cyclic group, which is absurd as $n\geq 2$.
From now on, we assume that $(M,g)$ is not a flat Riemannian manifold. Since $(M,h)$ is of non-positive sectional curvature, by Eells-Sampson theorem, there exists a harmonic map $f$ from $(M,g)$ to $(M,h)$ of mapping degree 1 homotopic to the identity map, and hence is in particular surjective.
Denote by $\nabla^{g}$ the Levi-Civita connection of $(M,g)$, and denote by $\varkappa\in\Omega^{1}(M)$ the first Koszul form of $((M,\nabla),g)$. Then by Proposition 2.43 we have $\nabla^{g}\varkappa=0$. Since $(M,h)$ is a compact Riemannian manifold of negative sectional curvature, we obtain that $f_{*}(\varkappa^{\sharp})\equiv 0$ by Theorem B.2.
By Proposition 2.44, we have that $\varkappa^{\sharp}$ is nowhere vanishing as $(M,g)$ is not flat. Therefore $\operatorname{dim}\operatorname{Ker}(df)\geq 1$, i.e. the differential $df$ is everywhere degenerated, and hence $f$ cannot be surjective by Sard’s Lemma, which contradicts the fact that $\operatorname{deg}(f)=1$.
This concludes the proof. ∎
The readers are invited to compare carefully Corollary 4.7 with Theorem 4.10.
Finally, as a highlight of this article, we prove that all compact Einstein-Hessian manifolds are flat.
Proposition 4.11.
Let $((M,\nabla),h)$ be a compact Hessian manifold of nonnegative Ricci curvature. Then $(M,h)$ is Riemannian flat.
Proof.
Recall from Theorem 2.13 that the universal covering of $M$ is contractible. The desired property then follows from Cheeger-Gromoll splitting theorem [29]. ∎
Definition 4.12.
A Hessian manifold $((M,\nabla),h)$ is said to be a Hessian-Einstein manifold, if $h$ is an Einstein metric on $M$.
Theorem 4.13.
All compact Hessian-Einstein manifolds are Riemannian flat.
Proof.
Consider any compact Hessian-Einstein manifold $((M,\nabla),h)$ of dimension $n+1$, and let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$.
By Proposition 4.8, there exist a complete Riemannian metric $g$ on $\mathbb{R}^{n}$ such that $(\tilde{M},\pi^{*}h)$ is isometric to the product of Riemannian manifolds $(\mathbb{R}^{n},g)$ and $(\mathbb{R},dt^{2})$. Since $(\tilde{M},\pi^{*}h)$ is Einstein as $(M,h)$ is, and one of its factor $(\mathbb{R},dt^{2})$ is Einstein with cosmological constant $\Lambda=0$, by Riemannian geometry we obtain that $(\tilde{M},\pi^{*}h)$ should also be Einstein with cosmological constant $\Lambda=0$, that is, $(\tilde{M},\pi^{*}h)$ is Ricci flat. Therefore, the Ricci curvature of $(M,h)$ is nonnegative.
By Proposition 4.11, we conclude that $(M,h)$ is Riemannian flat. ∎
4.2 The Cheng-Yau Metric
Amongst all compact orientable affine manifolds, those being simultaneously radiant and hyperbolic are special, as they admit canonical Hessian metrics. We shall explain this observation in the present section.
For a smooth function $u$ defined on a domain in $\mathbb{R}^{n}$, we write
$\nabla u:=\left[\begin{array}[]{ccc}\partial_{1}u\\ \vdots\\ \partial_{n}u\end{array}\right],\quad\operatorname{Hess}(u):=\left[\begin{array}[]{ccc}\partial_{1}\partial_{1}u&\cdots&\partial_{1}\partial_{n}u\\ \vdots&\ddots&\vdots\\ \partial_{n}\partial_{1}u&\cdots&\partial_{n}\partial_{n}u\end{array}\right],$
where $\partial_{i}=\partial/\partial x^{i}$ for $i=1,\dots,n$, and $x^{1},\dots,x^{n}$ are the standard coordinates of $\mathbb{R}^{n}$.
Recall the following existence and uniqueness result on real Monge–Ampère equation due to Cheng and Yau.
Theorem 4.14 (Cheng-Yau, 1982[18]).
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$. Then there exists a unique convex function $u\in C^{\infty}(\Omega)$ satisfying the following properties:
  1. 1.
    $\operatorname{det}(\operatorname{Hess}(u))=\operatorname{exp}(2u)^{n}$, and $u(\textup{{x}})\rightarrow\infty$ as $\textup{{x}}\rightarrow\partial\Omega$;
  2. 2.
    $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a complete Riemannian metric on $\Omega$;
  3. 3.
    $L^{*}g=g$ for every linear automorphism $L$ of $\Omega$.
Definition 4.15.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$. The convex function $u\in C^{\infty}(\Omega)$ satisfying the properties listed in Theorem 4.14 is termed the Cheng-Yau potential of $\Omega$.
The differential of a Cheng-Yau potential enjoys one more important property, namely, it is always a unit covector field.
Lemma 4.16.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$, and let $u\in C^{\infty}(\Omega)$ be the Cheng-Yau potential of $\Omega$. Then the holds that $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv 1$.
Proof.
For any $t>0$, define function $u_{t}\in C^{\infty}(\Omega)$ by $u_{t}(\textup{{x}}):=u(t\cdot\textup{{x}})+\operatorname{log}(t)$. Then straightforward computation yields that $u_{t}$ is also a Cheng-Yau potential of $\Omega$ for every $k>0$. Therefore, by the uniqueness part of Theorem 4.14, we obtain the functional equation
$u(\textup{{x}})-u(t\cdot\textup{{x}})=\operatorname{log}(t)$
where $\textup{{x}}\in\Omega$ and $k>0$. Now, for any $\textup{{x}}\in\Omega$, since
$\frac{u(\textup{{x}}+t\cdot\textup{{x}})-u(\textup{{x}})}{t}=-\operatorname{log}(1+t)/t\rightarrow-1$
as $t\rightarrow 0$, we have that
$\langle\nabla u(\textup{{x}}),\textup{{x}}\rangle\equiv-1.$ (1)
Differentiate both sides of equation( 1) again, we obtain that
$\operatorname{Hess}(u)^{-1}\nabla u(\textup{{x}})=-\textup{{x}}$ (2)
holds for all $\textup{{x}}\in\Omega$. Substitute equation( 2) back into equation( 1), we finally arrive at $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv 1$ as required. ∎
Using Lemma 4.16, we can prove that Cheng-Yau potentials are harmonic functions.
Lemma 4.17.
Let $\Omega$ be a domain in $\mathbb{R}^{d}$, and let $u\in C^{\infty}(\Omega)$. Suppose that $\operatorname{Hess}(u)$ is nowhere degenerated, so that $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a pseudo-Riemannian metric on $\Omega$. Then, the Laplace–Beltrami operator $\Delta^{g}$ of $(\Omega,g)$ satisfies
$\Delta_{g}\phi=\operatorname{tr}(\operatorname{Hess}(u)^{-1}\operatorname{Hess}(\phi))-\frac{1}{2}\langle\nabla\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u))),\operatorname{Hess}(u)^{-1}\nabla\phi\rangle$
for all $\phi\in C^{\infty}(\Omega)$.
Proof.
This is nothing else but a coordinate-dependent and local version of the formula in Lemma 3.33. ∎
Proposition 4.18.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{d}$, and let $u\in C^{\infty}(\Omega)$ be the Cheng-Yau potential of $\Omega$, so that $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a Riemannian metric on $\Omega$. Then $u$ is a harmonic function on $(\Omega,g)$.
Proof.
Since $u$ is a Cheng-Yau potential, we have $2d\cdot u=\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u)))$. Denote by $\Delta^{g}$ the Laplace–Beltrami operator of $(\Omega,g)$. Then by Lemma 4.16 and Lemma 4.17,
$\Delta_{g}u=$ $\operatorname{tr}(\operatorname{Hess}(u)^{-1}\operatorname{Hess}(u))-\frac{1}{2}\langle\nabla\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u))),\operatorname{Hess}(u)^{-1}\nabla u\rangle$
$=$ $\operatorname{tr}(\operatorname{Id})-d\cdot\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle$
$=$ $d-d\cdot 1$
$=$ $0$
as required. ∎
Adapt Cheng and Yau’s result to our setting, we immediately obtain the following crucial observation.
Proposition 4.19.
Let $(M,\nabla)$ be an $n$-dimensional compact orientable affine manifold. Suppose that $(M,\nabla)$ is both radiant and hyperbolic. Then, there exists a unique space-like Lagrangian embedding $\eta$ of $(M,\nabla)$ such that $\varkappa=n\eta$, where $\varkappa$ is the first Koszul form of the Hessian manifold $((M,\nabla),\nabla\eta)$.
Proof.
[…]
$n\cdot df=\frac{1}{2}d\operatorname{log}(\operatorname{det}(\operatorname{Hess}(f)))$
descends to $n\cdot\eta=\varkappa$ on $(M,\nabla)$ so that $\eta$ is a space-like Lagrangian embedding of $(M,\nabla)$ and $\varkappa$ is the first Koszul form of $((M,\nabla),\nabla\eta)$. The uniqueness of $\eta$ also follows immediately from the uniqueness of the Cheng-Yau potential. ∎
Proposition 4.19 motivates the following more general definition.
Definition 4.20.
Let $((M,\nabla),h)$ be a Hessian manifold, and let $\varkappa\in\Omega^{1}(M)$ be its first Koszul form. The metric $h$ is said to be a Cheng-Yau metric if there exists $\eta\in\Omega^{1}(M)$ and a real constant $k>0$ such that $h=\nabla\eta$ and $\varkappa=k\eta$.
4.3 Koszul’s Uncertainty Principle
In this very short section, we introduce a geometric uncertainty principle on compact Hessian manifolds of Koszul type.
Throughout this section, we use Einstein’s summation convention.
We shall first suggest several lemmata, which will turn out to be useful later.
Lemma 4.21.
Let $((M,\nabla),h)$ be a Hessian manifold and let $\varkappa\in\Omega^{1}(M)$ be its first Koszul form. Let $A$ be the Amari–Chentsov tensor of $((M,\nabla),h)$. Then, for any vector field $Y$ on $M$, the identity
$\varkappa(Y)=\frac{1}{2}\sum_{i=1}^{n}A(e_{i},e_{i},Y)$
holds in orthonormal frame fields $e_{1},\dots,e_{n}$ of $(M,h)$.
Proof.
Fix affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. Denote $\partial_{i}:=\partial/\partial x^{i}$ for $i=1,\dots,n$. By Jacobi’s formula for the derivative of a determinant, we have that
$2\varkappa_{k}=\partial_{k}(\operatorname{log}(|h|))=\operatorname{tr}(h^{-1}\partial_{k}h)=h^{ij}\partial_{k}h_{ij}$
where $|h|$ is the determinant of the matrix of which $(i,j)$-entry is $h_{ij}$. Moreover, by definition, we have $A_{ijk}=\partial_{k}h_{ij}$ for all $i,j,k\in\{1,\dots,n\}$. By computing the trace of $Y^{k}A_{ijk}$, we obtain
$2\varkappa_{k}Y^{k}=h^{ij}\partial_{k}h_{ij}Y^{k}=h^{ij}A_{ijk}Y^{k}$
as required. ∎
Lemma 4.22.
Let $((M,\nabla),h)$ be an $n$-dimensional Hessian manifold of Koszul type, and let $\eta\in\Omega^{1}(M)$ be its $1$-form potential. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((M,\nabla),h)$. Then the identity
$\operatorname{div}_{h}(\eta^{\sharp})=-\varkappa(\eta^{\sharp})+n$
holds, where $\sharp$ is the musical isomorphism induce by $h$.
Proof.
Denote by $A$ the Amari–Chentsov tensor of $((M,\nabla),h)$ and denote by
$\bar{\nabla}:=\frac{1}{2}(\nabla+\nabla^{*})$
the Levi-Civita connection of $(M,h)$. Then, by definition, we have
$(\bar{\nabla}_{X}\eta)(Y)=h(X,Y)-\frac{1}{2}A(X,Y,\eta^{\sharp}).$
Therefore, by Lemma 4.21, we conclude that
$\operatorname{div}_{h}(\eta^{\sharp})=-\star d\star\eta=-\varkappa(\eta^{\sharp})+n$
where $\star$ is the Hodge star operator on $(M,h)$. ∎
As usual, for a compact Riemannian manifold $(M,g)$, we denote by $\operatorname{Vol}_{g}(M)$ is volume.
Corollary 4.23.
Let $((M,\nabla),h)$ be an $n$-dimensional compact Hessian manifold of Koszul type, and let $\eta\in\Omega^{1}(M)$ be its $1$-form potential. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((M,\nabla),h)$. Then the integral identity
$\operatorname{Vol}_{h}(M)=\frac{1}{n}\int_{M}\varkappa\wedge\star\eta$
holds, where $\star$ is the Hodge star operator on $(M,h)$.
Proof.
Apply divergence theorem to the formula in Lemma 4.22. ∎
With the preparations done, the geometric uncertainty principle can now be derived easily as a corollary:
Corollary 4.24.
Let $((M,\nabla),h)$ be an $n$-dimensional compact Hessian manifold of Koszul type, and let $\eta\in\Omega^{1}(M)$ be its $1$-form potential. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((M,\nabla),h)$. Then the integral inequality
$\operatorname{Vol}_{h}(M)^{2}\leq\frac{1}{n^{2}}\left(\int_{M}|\varkappa|^{2}\operatorname{vol}_{h}\right)\left(\int_{M}|\eta|^{2}\operatorname{vol}_{h}\right)$
holds on $M$, with equality if and only if $\varkappa=k\cdot\eta$ for some positive real number $k\in\mathbb{R}$, that is, $h$ is a Cheng-Yau metric.
Proof.
Applying Cauchy-Schwarz inequality to the integral formula in Corollary 4.23, we immediately obatin the desired integral inequality. The equality holds exactly when there exists a real constant $k\in\mathbb{R}$ such that $\varkappa=k\cdot\eta$. By Corollary 4.23, we have
$\operatorname{Vol}_{h}(M)=\frac{1}{n}\int_{M}\varkappa\wedge\star\eta=\frac{k}{n}\int_{M}\eta\wedge\star\eta=\frac{k}{n}\int_{M}|\eta|^{2}\operatorname{vol}_{h}.$
Therefore the constant $k\in\mathbb{R}$ must be positive. ∎
A facile application of Corollary 4.24 is the following non-vanishing theorem for Riemann curvature tensors of compact Hessian manifolds of Koszul type.
Corollary 4.25.
Compact Hessian manifolds of Koszul type are not Riemannian flat.
Proof.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type, and let $\varkappa\in\Omega^{1}(M)$ be its first Koszul form.
Assume the contrary that $(M,h)$ is Riemannian flat. Then, by Theorem 2.44, we have that $\varkappa=0$. But Corollary 4.24 then yields $\operatorname{Vol}_{h}(M)\leq 0$, which is absurd. ∎
Remark 4.26.
The integral inequality stated in Corollary 4.24 is an analog of Heisenberg uncertainty principle in quantum mechanics: For a given dimension and volume, a $1$-form potential small in $L^{2}$-norm must be compensated by a first Koszul form large in $L^{2}$-norm, and vice-versa. This phenomenon indicates a highly non-trivial relation between a $1$-form potential and the first Koszul form of a Hessian manifold. We name this integral inequality as Koszul’s uncertainty principle, in honor of the great French mathematician Jean-Louis Koszul.
Remark 4.27.
On a compact Hessian manifold $((M,\nabla),h)$ of Koszul type, the integral inequality of Koszul’s uncertainty principle is sharp, and the equality is achieved exactly when $h$ is a Cheng-Yau metric. This reveals a deep connection with Theorem 4.14 and Proposition 4.19.
5 Topics in Low Dimensional Hessian Geometry
5.1 Exotic Potentials for the Euclidean Plane
Hessian potential functions for a metric, when exist, are in general not unique. For instance, besides the half norm square, the flat metric tensor of Euclidean or Minkowski spaces may well be the Hessian quadratic form of other potential functions.
Here, we shall provide a convenient equivalent condition for a smooth function defined on a planar domain to be a Hessian potential of a flat metric.
Lemma 5.1.
Let $\Omega$ be a domain in $\mathbb{R}^{2}$, and let
$\textup{{x}}\colon\Omega\rightarrow\mathbb{R}^{3},\quad(u,v)\mapsto\textup{{x}}(u,v)$
be a parametrization of a regular smooth surface $S$ in $\mathbb{R}^{3}$. Suppose that $S$ does not pass through the origin, and the vectors $\textup{{x}},\textup{{x}}_{u},\textup{{x}}_{v}$ are everywhere $\mathbb{R}$-linearly dependent. Then, there exists a homogeneous function $F$ in three variables such that $F(\textup{{x}})\equiv 0$. Consequently, $S$ lies on a cone of which apex is the origin.
Proof.
This is a well-known fact in classical differential geometry in $\mathbb{R}^{3}$. We nevertheless give a sketch of proof.
Consider the smooth mapping $\phi:=\textbf{x}/\|\textbf{x}\|$. By the assumption on vectors $\textbf{x},\textbf{x}_{u},\textbf{x}_{v}$, a straightforward computation shows that $\phi_{u}\times\phi_{v}\equiv 0$. Therefore, the image $\phi(\Omega)$ is a smooth curve on the $2$-sphere $\mathbb{S}^{2}:=\{(x,y,z)\in\mathbb{R}^{3}|x^{2}+y^{2}+z^{2}=1\}$. By the implicit function theorem, there exists a smooth function $f\colon\mathbb{S}^{2}\rightarrow\mathbb{R}$ such that $\phi(\Omega)$ is locally the zero locus of $f$. Now, define homogeneous function $F$ of degree zero via
$F(x,y,z):=f(x/r,y/r,z/r)$
where $r=\sqrt{x^{2}+y^{2}+z^{2}}$. Then, by the very construction, we have $F(\textbf{x})\equiv 0$. ∎
Proposition 5.2.
Let $f$ be a smooth function defined on a domain in $\mathbb{R}^{2}$ such that $\operatorname{Hess}(f)$ is everywhere non-degenerate, so in particular
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a pseudo-Riemannian metric. Then, $g$ is flat if and only if there exists a homogeneous function $F$ in three variables such that $F(f_{xx},f_{xy},f_{yy})\equiv 0$.
Proof.
Combine Corollary 4.4 with Lemma 5.1. ∎
Example 5.3.
Let $f$ be a smooth homogeneous function of degree $d$ defined on a domain in $\mathbb{R}^{2}$ such that its Hessian quadratic form $H:=\operatorname{Hess}(f)$ is everywhere non-degenerate. Then, by Euler’s theorem on homogeneous functions, we have
$(2-d)H+x\frac{\partial H}{\partial x}+y\frac{\partial H}{\partial y}=0$
that is, the matrices $H,H_{x},H_{y}$ are everywhere $\mathbb{R}$-linearly dependent. Therefore, by Corollary 4.4, we have that
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a flat pseudo-Riemannian metric.
The following interesting example shows that nonlinear harmonic functions are Hessian potentials for the flat $1+1$ Minkowski plane $\mathbb{R}^{1,1}$.
Example 5.4.
Let $\Omega$ be a domain in $\mathbb{C}^{1}$, and let $f\in C^{\infty}(\Omega)$ be a nonlinear harmonic function. Then, $\operatorname{tr}(\operatorname{Hess}(f))=f_{xx}+f_{yy}=0$ and
$\operatorname{det}(\operatorname{Hess}(f))=f_{xx}f_{yy}-f_{xy}^{2}=-f_{yy}^{2}-f_{xy}^{2}<0.$
By Cayley–Hamilton theorem, we obtain that $\operatorname{Hess}(f)$ is non-degenerate and indefinite, and hence
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}$
is a Lorentzian metric. Moreover, by Proposition 5.2, $g$ is flat, i.e. locally the metric of the $1+1$ Minkowski space.
We also suggest the following very concrete example of a Hessian potential for the flat Euclidean plane, which is constructed using the method introduced in Theorem 1.8.
Example 5.5.
Let $\mathbb{H}^{2}:=\{(x,y)\in\mathbb{R}^{2}:y>0\}$ be the upper half plane and define $f\in C^{\infty}(\mathbb{H}^{2})$ by
$f(x,y):=\frac{x^{2}}{2y}+\frac{1}{4}\operatorname{log}(y)y.$
Then the coordinate change $(x(r,\theta),y(r,\theta))=(r^{2}\theta,r^{2})$ brings the Hessian quadratic form
$g:=f_{xx}dx^{2}+2f_{xy}dxdy+f_{yy}dy^{2}=\frac{1}{y}dx^{2}-\frac{2x}{y^{2}}dxdy+\frac{4x^{2}+y^{2}}{4y^{3}}dy^{2}$
of $f$ to the flat metric tensor $dr^{2}+r^{2}d\theta^{2}$ of the Euclidean plane in its polar coordinates. The flatness of $g$ can be proved also by noticing that
$f_{xx}^{2}-4f_{xx}f_{yy}+4f_{xy}^{2}\equiv 0$
and then apply Proposition 5.2.
Recall from Example 5.3 that, in dimension $2$, all smooth homogeneous functions with non-singular Hessian are potentials for flat metrics. However, a computation in [30] indicates that this is no longer true in higher dimension.
Before we close this section, we cite from [30] an example given by Dubrovin, to show that techniques from algebraic geometry, for example the Clebsch covariant of projective plane curves, might be useful for finding homogeneous potentials.
Example 5.6.
Let $f(x,y,z):=x^{6}+y^{6}+z^{6}-10(x^{3}y^{3}+y^{3}z^{3}+z^{3}x^{3})$ be the Maschke sextic. Then the Hessian quadratic form
$g=\sum_{i=1}^{3}\sum_{j=1}^{3}\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}}dx_{i}\otimes dx_{j}$
of $f$ is a flat metric on $\mathbb{R}^{3}$.
For more discussion on homogeneous polynomial potential of flat spaces, we refer to the work of B.Totaro [30].
5.2 Complete Hessian Surfaces
It is a well-known fact that an orientable closed Hessian surface is homeomorphic to the 2-torus $\mathbb{T}^{2}$. The ultimate objective of this section is to classify all complete Hessian surfaces topologically. To achieve this goal, the following preparations are needed:
Definition 5.7.
Let $(M,\nabla)$ be an affine manifold. We denote by $\operatorname{Aff}(M;\nabla)$ the group of affine automorphisms of $(M,\nabla)$. Whenever $M=\Omega$ is a convex domain in $\mathbb{R}^{d}$, and $\nabla$ is the Levi-Civita connection of the $d$-dimensional flat Euclidean space, by convention we omit $\nabla$ to write $\operatorname{Aff}(\Omega):=\operatorname{Aff}(\Omega;\nabla)$.
Lemma 5.8.
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$ and $\operatorname{recc}(\Omega)$ its recession cone. Let $f\in\operatorname{Aff}_{2}(\mathbb{R})$ and let $T$ be the linear part of $f$, that is, $f(\textup{{x}})=T\textup{{x}}+f(0)$ for all $\textup{{x}}\in\mathbb{R}^{d}$. Suppose that $\Omega=f(\Omega)$. Then $T\textup{{x}}\in\operatorname{recc}(\Omega)$ for all $\textup{{x}}\in\operatorname{recc}(\Omega)$.
Proof.
Take any $\textbf{x}\in\operatorname{recc}(\Omega)$ and $\textbf{y}\in\Omega$. By the definition of a recession cone, it suffices to prove that $T\textbf{x}+\textbf{y}\in\Omega$. But indeed, we have
$T\textbf{x}+\textbf{y}=T\textbf{x}+Tf^{-1}(\textbf{y})+f(0)=f(\textbf{x}+f^{-1}(\textbf{y}))\in\Omega$
as required. ∎
The following proposition concerning compactness of the affine automorphism group of a bounded convex domain is itself interesting and useful.
Proposition 5.9.
Let $\Omega$ be a bounded convex domain in $\mathbb{R}^{n}$. Then $\operatorname{Aff}(\Omega)$ is isomorphic to a compact Lie subgroup of $\operatorname{GL}_{n}(\mathbb{R})$.
Proof.
It is readily seen that $G:=\operatorname{Aff}(\Omega)$ is a closed subgroup of $\operatorname{Aff}_{n}(\mathbb{R})$. By Cartan’s theorem on closed subgroups, $G$ is a Lie subgroup of $\operatorname{Aff}_{n}(\mathbb{R})$.
Denote by $\overline{\Omega}$ the closure of $\Omega$. By Heine-Borel theorem and the assumption on $\Omega$, we have that $\overline{\Omega}$ is a convex body. Denote by $E_{\Omega}$ the inner Löwner–John ellipsoid of $\overline{\Omega}$. By the uniqueness of the inner Löwner–John ellipsoid, $E_{\Omega}=g(E_{\Omega})$ for every $g\in G$. By linear algebra, if a non-degenerate affine transform $g$ maps an ellipsoid $E$ onto $E$ itself, then $g$ fix the center of $E$. It follows that all elements of $G$ share a common fixed point, namely the center of the ellipsoid $E_{\Omega}$. without loss of generality, we may assume that the ellipsoid $E_{\Omega}$ is centered at the origin $0\in\mathbb{R}^{n}$. Therefore $G$ is a subgroup of $\operatorname{GL}_{n}(\mathbb{R})$.
Denote by $\|\cdot\|_{\operatorname{op}}$ the operator norm. Denote by $\ell_{1}$ and $\ell_{2}$ the diameter and the width of $E_{\Omega}$, respectively. Then it is readily seen that $\|g\|_{\operatorname{op}}\leq\ell_{1}/\ell_{2}$ for all $g\in G$. In particular, $G$ is a bounded subset of $\mathbb{R}^{n\times n}$. Now, again by Heine-Borel theorem, $G$ is compact. ∎
We will also use the following classical theorem on crystallographic groups of D. Fried and W. Goldman.
Theorem 5.10 (Fried-Goldman,1983[31]).
Let $G$ be a subgroup of $\operatorname{Aff}(\mathbb{R}^{2})$. Suppose that $G$ acts proper discontinuously on $\mathbb{R}^{2}$ by affine transform. Then $G$ admits a solvable subgroup of finite index.
Proof.
This is a main theorem proved in [31]. ∎
Now we are fully prepared to prove the main theorem of this section.
Theorem 5.11.
Let $((M,\nabla),h)$ be a complete Hessian surface. Then $M$ is diffeomorphic to one of the following smooth manifolds:
  1. 1.
    the plane $\mathbb{R}^{2}$;
  2. 2.
    the cylinder $\mathbb{S}^{1}\times\mathbb{R}$;
  3. 3.
    the Möbius strip;
  4. 4.
    the 2-torus $\mathbb{T}^{2}$;
  5. 5.
    the Klein bottle.
Proof.
Since $((M,\nabla),h)$ is a complete Hessian manifold of dimension 2, by Theorem 2.13, there exists a planar convex domain $\Omega$ and a discrete subgroup $G$ of the affine group $\operatorname{Aff}_{2}(\mathbb{R})$, such that $G$ acts freely and proper discontinuously on $\Omega$ by affine transform, and that $M\cong\Omega/G$.
If $G$ is trivial, then $M=\Omega$ is homeomorphic to $\mathbb{R}^{2}$. If $M$ is closed, then by Theorem 2.17, we have that $\chi(M)=0$ and hence $M$ is homeomorphic to the 2-torus or the Klein bottle. Recall from low-dimensional topology that the fundamental group of a non-compact surface is always free. From now on, we assume that $M$ is non-compact and $G$ is a non-trivial free group.
Denote by $\operatorname{recc}(\Omega)$ the recession cone of $\Omega$. By elementary Euclidean geometry, any convex cone in $\mathbb{R}^{2}$, and hence in particular $\operatorname{recc}(\Omega)$, is congruent to one of the following sets:
  1. 1.
    a point in $\mathbb{R}^{2}$;
  2. 2.
    a straight line, i.e. an 1-dimensional affine subspace of $\mathbb{R}^{2}$;
  3. 3.
    a wedge, i.e. the set $W_{\theta}:=\{z\in\mathbb{C}^{1}:0\leq\operatorname{Arg}(z)\leq\theta\}$ for $0\leq\theta\leq\pi$;
  4. 4.
    the whole plane $\mathbb{R}^{2}$.
Assume that $\operatorname{recc}(\Omega)$ is a point. Then, by convex geometry, we have that $\Omega$ is bounded. Therefore, by Proposition 5.9, $\operatorname{Aff}(\Omega)$ is a compact Lie group. But, by assumption, $G$ is a infinite discrete subgroup of $\operatorname{Aff}(\Omega)$, which leads to a contradiction. Therefore $\operatorname{recc}(\Omega)$ cannot be a point.
Assume that $\operatorname{recc}(\Omega)$ is the whole plane $\mathbb{R}^{2}$. Then we have $\Omega=\operatorname{recc}(\Omega)=\mathbb{R}^{2}$. Therefore $G$ admits a solvable subgroup $G_{0}$ of finite index by Theorem 5.10.
Assume that $\operatorname{recc}(\Omega)$ is a straight line in $\mathbb{R}^{2}$ or a wedge $W_{\theta}$ with $0\leq\theta\leq\pi$. Consider the group homomorphism
$\phi\colon\operatorname{Aff}_{2}(\mathbb{R})\rightarrow\operatorname{GL}_{2}(\mathbb{R}),\quad f\mapsto(\textbf{x}\mapsto f(\textbf{x})-f(0))$
that is, the map sending every affine transform to its linear part. Then by Lemma 5.8, we have that $\phi(G)$ acts on $\operatorname{recc}(\Omega)$ by linear automorphism. In particular, $\phi(G)$ preserves the apex and the boundary of $\operatorname{recc}(\Omega)$. Therefore, $\phi(G)$ admits a subgroup $H$ of index at most 2, such that all elements of $H$ share at least one eigenvector, and hence $G_{0}:=\phi^{-1}(H)\cap G$ is isomorphic to a subgroup of the group $B$ of invertible upper-triangular real $3\times 3$ matrices. Recall from linear algebra that the length of the derived sequence of Heisenberg group $H_{3}(\mathbb{R})$ is exactly $2$, and observe that $B/H_{3}(\mathbb{R})\cong(\mathbb{R}^{\times})^{3}$ is Abelian. By group theory, we obtain that $B$ is a solvable group, and hence $G_{0}$ ditto. Therefore, $G$ admits a solvable subgroup $G_{0}$ of index at most 2.
This proves that $G$ always admits a solvable subgroup $G_{0}$ of finite index.
Since $G$ is infinite, we have $G_{0}$ ditto. By Nielsen–Schreier theorem, $G_{0}$ is also free. Since any free group of rank $r>1$ is not solvable, the rank of $G_{0}$ is at most 1. Therefore, we conclude that $G_{0}\cong\mathbb{Z}$. Finally, by Schreier’s index formula, we arrive at that in fact the rank of $G$ is also 1, i.e. $G$ is isomorphic to $\mathbb{Z}$.
Therefore, the non-compact surface $M$ has fundamental group $\pi_{1}(M)=\mathbb{Z}$.
By the classification theorem of non-compact surfaces, we have that $M$ is a closed surface $N$ with $k\geq 1$ punctures.
Assume that $N$ is orientable of genus $g$. Then by Seifert-van Kampen theorem, $\pi_{1}(M)$ is a free group of rank r=2g+k1r=2g+k-1. Since $r=1$ and $k>0$, we have that $g=0$ and $k=2$. Therefore $M$ is a sphere with $2$ punctures, i.e. a cylinder.
Assume that $N$ is non-orientable. Then, by the classification theorem of closed surfaces, $N$ is the connected sum of $l$ real projective planes for some integer $l\geq 1$. Again by Seifert-van Kampen theorem, $\pi_{1}(M)$ is a free group of rank r=l+k1r=l+k-1. Since $r=1$ and $k>0$, we obtain $l=k=1$. Therefore $M$ is a punctured $\mathbb{R}P^{2}$, i.e. a Möbius strip.
Quod erat demonstrandum. ∎
Example 5.12.
A complete Hessian surface of constant Gaußian curvature 1-1 can be isometric to a hyperbolic funnel, that is, the quotient of the Lobachevsky plane by an expanding homothety $z\mapsto\lambda\cdot z$ where $\lambda>1$.
In the polar coordinates $(r,\theta)$ of the complex plane, the $1$-form potential of a hyperbolic funnel is precisely
$\eta=\frac{dr}{2r}+\left(\frac{1}{2}\frac{\operatorname{sin}(\theta)}{\operatorname{cos}(\theta)}-\frac{\operatorname{sin}(\theta)}{\operatorname{cos}(\theta)^{3}}\right)d\theta$
defined on an open annulus centered at the origin. A straightforward computation shows that $\eta$ is closed, but not exact.
5.3 Compact Orientable Hessian 3-folds
The flat tori, or more generally, Bieberbach manifolds are examples of compact Hessian manifolds, but they are trivial in the sense that their affine structures are given by their Levi-Civita connections. In is section, we construct concretely a family of compact radiant Hessian manifolds of Koszul type with negative scalar curvature.
Denote by $\mathbb{H}^{2}:=\{z\in\mathbb{C}:\operatorname{Im}(z)>0\}$ the upper half plane in $\mathbb{C}^{1}$. Consider the product Riemannian metric
$ds^{2}:=d\rho^{2}+\frac{|d\tau|^{2}}{\operatorname{Im}(\tau)^{2}}$
on $\mathbb{H}^{2}\times\mathbb{R}$, where $\tau$ is the coordinate on $\mathbb{H}^{2}$ and $\rho$ is the coordinate on $\mathbb{R}$.
[…]
Consider the quadratic function
$Q\colon\mathbb{R}^{3}\rightarrow\mathbb{R},\quad(x,y,t)\mapsto t^{2}-x^{2}-y^{2}$
and denote $\Omega:=\{(x,y,t)\in\mathbb{R}^{3}:Q(x,y,t)>0\}$. Also, define the smooth function $f\in C^{\infty}(\Omega)$ by
$f(x,y,t):=-\frac{1}{2}\operatorname{log}(t^{2}-x^{2}-y^{2}).$
Then computation yields that the Hessian quadratic form
$2\frac{(tdt-xdx-ydy)^{2}}{(t^{2}-x^{2}-y^{2})^{2}}-\frac{dt^{2}-dx^{2}-dy^{2}}{t^{2}-x^{2}-y^{2}}$
of $f$ is a Riemannian metric on $\Omega$.
Denote by $\nabla$ the Levi-Civita connection of the Euclidean 3-space. Then, by construction, $\nabla(df)$ is a Hessian metric on $(\Omega,\nabla)$.
A tedious but straightforward computation shows that
$\Phi\colon\mathbb{H}^{2}\times\mathbb{R}\rightarrow\Omega,\quad(\tau,\rho)\mapsto\frac{\operatorname{exp}(\rho)}{2\operatorname{Im}(\tau)}(\tau+\bar{\tau},|\tau|^{2}-1,|\tau|^{2}+1)$
is an isometry between $(\mathbb{H}^{2}\times\mathbb{R},ds^{2})$ and $(\Omega,\nabla(df))$.
By the additivity of scalar curvature, we obtain that the scalar curvature of $(\Omega,\nabla(df))$ is constantly 1+0=1-1+0=-1.
Denote by $\tilde{\nabla}:=\Phi^{*}\nabla$ the pullback connection. Then $((\mathbb{H}^{2}\times\mathbb{R},\tilde{\nabla}),ds^{2})$ is a Hessian manifold as $((\Omega,\nabla),\nabla(df))$ is.
From now on, for simplicity, we identify $((\Omega,\nabla),\nabla(df))$ with $((\mathbb{H}^{2}\times\mathbb{R},\tilde{\nabla}),ds^{2})$.
Consider any arbitrary cocompact Fuchsian group action
$(\Gamma,\mathbb{H}^{2})\rightarrow\mathbb{H}^{2},\quad(\gamma,\tau)\mapsto\gamma\cdot\tau$
on $\mathbb{H}^{2}$ so that $\mathbb{H}^{2}/\Gamma=\Sigma_{g}$ is a closed orientable surface of genus $g>1$. Let $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ be a periodic mapping class. By Nielsen’s realization theorem, without loss of generality, we may assume that $\varphi$ lifts to an $\Gamma$-equivariant isometry $\tilde{\varphi}$ of the Lobachevsky plane. Then $G:=\Gamma\rtimes_{\varphi}\mathbb{Z}$ acts on $\mathbb{H}^{2}\times\mathbb{R}$ by isometries via
$G\times(\mathbb{H}^{2}\times\mathbb{R})\rightarrow\mathbb{H}^{2}\times\mathbb{R},\quad((\gamma,n),(\tau,\rho))\mapsto(\gamma\cdot\varphi^{n}(\tau),\rho+n).$
The $G$-action is by construction free and properly discontinuous.
Recall from the theory of Lie groups that the isometry group $\operatorname{PSL}_{2}(\mathbb{R})$ of $\mathbb{H}^{2}$ is isomorphic to the Lorentz group $\operatorname{SO}^{+}(2,1)$, whilst the latter is a Lie subgroup of $\operatorname{GL}_{3}(\mathbb{R})$. It follows that $G$ acts on the convex cone $\Omega$ by linear automorphisms. Therefore, $\tilde{\nabla}$ descends to a flat and torsion-free connection $D$ on the tangent bundle of the mapping torus $M_{\varphi}=(\mathbb{H}^{2}\times\mathbb{R})/G$ of $\varphi$, so that $(M_{\varphi},D)$ is an affine manifold. Moreover, $(M_{\varphi},D)$ is radiant by Proposition 3.12.
Since it always holds that
$(|z|^{2}+1)^{2}-(|z|^{2}-1)^{2}-(z+\bar{z})^{2}=4\operatorname{Im}(z)^{2}$
for every $z\in\mathbb{C}$, we obtain that $Q(\Phi(\gamma\cdot\varphi^{n}(\tau),\rho+n))=Q(\Phi(\tau,\rho))e^{2n}$, and hence
$f(\Phi(\gamma\cdot\varphi^{n}(\tau),\rho+n))=-\frac{1}{2}\operatorname{log}(Q(\Phi(\tau,\rho))\operatorname{exp}(2n))=f(\Phi(\tau,\rho))-n$
for every $(\gamma,n)\in G$ and every $(\tau,\rho)\in\mathbb{H}^{2}\times\mathbb{R}$. Therefore the 1-form potential $df$ of $((\Omega,\nabla),\nabla(df))$ is invariant under the $G$-action, and thus descends to the quotient $M_{\varphi}$. In particular, the product Riemannian metric $ds^{2}$ on $\mathbb{H}^{2}\times\mathbb{R}$ descends to a Riemannian metric $h$ on $M_{\varphi}$.
This proves that $((M_{\varphi},D),h)$ is a compact radiant Hessian manifold of Koszul type with constant negative scalar curvature 1-1, as claimed.
In particular, we arrived at the following theorem:
Proposition 5.13.
Let $\Sigma_{g}$ be a closed orientable surface of genus $g>1$, and let $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ be a periodic mapping class. Then there exist a Riemannian metric $h$ on the mapping torus $M_{\varphi}$ of $\varphi$ and a flat torsion-free connection $D$ on the tangent bundle $T(M_{\varphi})\rightarrow M_{\varphi}$, such that $((M_{\varphi},D),h)$ is a closed and orientable radiant Hessian $3$-manifold of Koszul type with negative scalar curvature.
Under the scope of Thurston’s geometrization, the high genus examples in the proof of Proposition 5.13 are those locally modeled by $\mathbb{H}^{2}\times\mathbb{R}$. A natural next step is to construct similar examples for the genus one case. We will now prove that these examples indeed exist, and they all admit flat models.
Proposition 5.14.
Let $\varphi]\in\operatorname{Mod}(\mathbb{T}^{2})$ be a periodic mapping class. Then there exist a Riemannian metric $g$ on the mapping torus $M_{\varphi}$ of $\varphi$, such that $(M_{\varphi},g)$ is a closed orientable Bieberbach $3$-manifold.
Proof.
Recall from the arithmetic of $\operatorname{SL}_{2}(\mathbb{Z})\cong\operatorname{Mod}(\mathbb{T}^{2})$ that every finite order element of $\operatorname{SL}_{2}(\mathbb{Z})$ is conjugate to either $\pm\operatorname{Id}$ or one of the following matrices
$A_{1}=\left[\begin{array}[]{cc}0&-1\\ 1&-1\end{array}\right],\quad A_{2}=\left[\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right],\quad A_{3}=\left[\begin{array}[]{cc}0&-1\\ 1&1\end{array}\right]$
or their inverses. Now, it suffices to notice that $\pm\operatorname{Id},A_{2}$ are isometries of the rhombic elliptic curve $\mathbb{C}^{1}/(\mathbb{Z}\oplus i\mathbb{Z})$, and $A_{1},A_{3}$ are isometries of the hexagonal elliptic curve $\mathbb{C}^{1}/(\mathbb{Z}\oplus\omega\mathbb{Z})$, where $\omega\in\mathbb{C}$ is the $3$-rd primitive root of the unity. ∎
The examples constructed above are consistent with the results established earlier. As predicted by Theorem 2.46, for a closed orientable surface $\Sigma_{g}$ of genus $g\geq 1$, the mapping torus $M_{\varphi}$ of a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ is indeed a fiber bundle over $\mathbb{S}^{1}$. The fundamental group of $M_{\varphi}$ is a semidirect product of $\mathbb{Z}$ and $\pi_{1}(\Sigma_{g})$, which is indeed infinite and torsion free, as stated in Proposition 2.7 and Proposition 2.19.
However, given the nature of all examples that we constructed, one can be curious about whether there exists an example of closed orientable Hessian $3$-manifold that is a mapping torus of a mapping class of infinite order. We now prove that this is impossible, which leads to a classification theorem of closed Hessian $3$-manifolds.
Theorem 5.15.
Let $((M,\nabla),h)$ be a closed orientable Hessian $3$-manifold. Then there exists a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ such that $M$ is diffeomorphic to the mapping torus of $\varphi$, where $\Sigma_{g}$ is the closed orientable surface of genus $g\geq 1$.
Proof.
In the special case that $(M,h)$ is a flat Riemannian manifold, the desired theorem is proved by the classification of space groups.
From now on, we assume that $(M,h)$ is not flat. Then by Theorem 2.48, there exists a $\Sigma_{g}$-bundle $f\colon M\rightarrow\mathbb{S}^{1}$ for some closed orientable surface $\Sigma_{g}$ of genus $g$.
Since $M$ is aspherical by Corollary 2.14, we obtain that $\Sigma_{g}$ ditto. Therefore, in particular, $\Sigma_{g}$ is not homeomorphic to $\mathbb{S}^{2}$, and hence $g\geq 1$.
In particular, we have $\pi_{1}(M)=\pi_{1}(\Sigma_{g})\rtimes_{\varphi}\mathbb{Z}$ where $\varphi$ is the monodromy of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$. Let $\pi\colon\tilde{M}\rightarrow M$ be a universal covering of $M$. Then there exists a complete Riemannian metric $g$ on $\mathbb{R}^{n}$ such that $(\tilde{M},\pi^{*}h)$ is isometric to the product of Riemannian manifolds $(\mathbb{R}^{n},g)$ and $(\mathbb{R},dt^{2})$. Denote by $\hat{\varphi}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ the lifting of $\varphi\colon\Sigma_{g}\rightarrow\Sigma_{g}$ to the universal covering spaces. Since $\pi_{1}(M)$ acts on $(\mathbb{R}^{n}\times\mathbb{R},g\oplus dt^{2})$ by isometry, in particular the smooth mapping
$\Phi\colon\mathbb{R}^{n}\times\mathbb{R}\rightarrow\mathbb{R}^{n}\times\mathbb{R},\quad(\textbf{x},t)\mapsto(\hat{\varphi}(\textbf{x}),t+1)$
is an isometry. Therefore we conclude that $\hat{\varphi}$ and hence $\varphi$ are isometries.
Denote $i\colon\Sigma_{g}\rightarrow M$ the inclusion map of a fiber of the fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$. Consider the subgroup $A$ of the isometry group $G$ of $(\Sigma_{g},i^{*}h)$ generated by $\varphi$, and denote by $\bar{A}$ its closure in $G$. Then it is readily seen that $\bar{A}$ is a compact abelian subgroup of $G$, and hence the identity component of $\bar{A}$ is a torus, Therefore, $\varphi$ is isotopic in $G$ to a group element of finite order.
Quod erat demonstrandum. ∎
Remark 5.16.
Combine Proposition 5.13 and Proposition 5.14, we obtain that the converse statement of Theorem 5.15 is also true: For every closed orientable surface $\Sigma_{g}$ of genus $g\geq 1$ and every periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$, there exist a Riemannian metric $h$ on the mapping torus $M_{\varphi}$ of $\varphi$ and a flat torsion-free connection $D$ on the tangent bundle $T(M_{\varphi})\rightarrow M_{\varphi}$, such that $((M_{\varphi},D),h)$ is a Hessian manifold. These thus gives a complete topological classification of closed orientable Hessian $3$-manifolds.
Theorem 5.15 is also consistent with earlier discoveries: Consider a closed orientable surface $\Sigma_{g}$ of genus $g$. Recall that pseudo-Anosov maps and Dehn twists are of infinite order. By Nielsen–Thurston classification, the mapping torus $M_{\varphi}$ of a pseudo-Anosov mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ admits a hyperbolic metric, which is forbidden by our previous Theorem 4.10. For the very same reason, we obtain that the nilpotent Heisenberg manifold cannot support any Hessian metric.
We close this section with a final remark: It is obvious that the product of two Hessian manifolds is naturally a Hessian manifold. However, the converse statement is far from being correct, as in the example above, the product of the circle $\mathbb{S}^{1}$ and a closed orientable surface $\Sigma_{g}$ of genus $g>1$ admits an affine structure, but $\Sigma_{g}$ itself does not admit any affine structure if $g\neq 1$ by the proof of Chern’s conjecture in dimension two. Therefore, in particular, the examples presented in this section are different to those constructed using Example 1.7 and Theorem 1.8.
5.4 Topology of Hessian 4-Manifolds
In this section, we list several properties of Hessian manifolds in dimension four.
For an oriented closed topological 4-manifold $M$, as per usual, we denote by $Q_{M}$ its intersection form.
Lemma 5.17.
Let $((M,\nabla),h)$ be an oriented closed Hessian 4-manifold. Then $\chi(M)=\operatorname{sign}(Q_{M})=0$. In particular, the first Betti number of $M$ is positive and the second Betti number of $M$ is even.
Proof.
The first claim follows from Proposition 1.52.17 and Hirzebruch’s signature theorem. Moreover, since $\operatorname{dim}M=4$ is even and $\chi(M)=0$, by Poincaré duality the first Betti number of $M$ is positive and the second Betti number of $M$ is even. ∎
Corollary 5.18.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. Then $M$ is null-cobordant, i.e. there exist an orientable smooth 5-manifold $N$ with bundary $\partial N=M$.
Proof.
See [32] for a proof of the fact that an oriented closed smooth 4-manifold is null-cobordant if and only if it has signature zero. ∎
Recall that all closed orientable 4-manifolds have $\operatorname{spin}^{c}$ structures, and therefore their third integral Stiefel-Whitney classes vanish. The following property is specific for Hessian 4-manifolds.
Proposition 5.19.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. If $M$ is spin then it admits an almost complex structure.
Proof.
Recall that $M$ is spin if its second Stiefel–Whitney class $w_{2}(TM)\in\mathrm{H}^{2}(M;\mathbb{Z}/2\mathbb{Z})$ vanishes.
Moreover, by Ehresmann-Wu theorem, $M$ admits an almost complex structure if and only if there exists $c\in\mathrm{H}^{2}(M;\mathbb{Z})$ such that
$3\operatorname{sign}\left(Q_{M}\right)+2\chi(M)=c^{2}$ (3)
and $w_{2}(TM)$ is the $\bmod 2$ reduction of $c$. For our case, by Lemma 5.17, $\operatorname{sign}\left(Q_{M}\right)+2\chi(M)=0$. If $w_{2}(TM)=0$, then $c=0$ satisfies equation( 3) and its $\bmod 2$ reduction is trivially equal to $w_{2}(TM)$. ∎
Furthermore, the following proposition shows that there are only two possibilities for the intersection form of an oriented closed Hessian 4-manifold, depending on the existence of $\operatorname{spin}$ structure.
Proposition 5.20.
Let $((M,\nabla),h)$ be an oriented closed Hessian 4-manifold with the second Betti number $b_{2}(M)=2m$. If $M$ is spin, then it holds that $Q_{M}=H^{\oplus m}$ where
$H:=Q_{\mathbb{S}^{2}\times\mathbb{S}^{2}}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right];$
if $M$ is not spin, then $Q_{M}=\operatorname{diag}(\underbrace{1,\dots,1}_{m},\underbrace{-1,\dots,-1}_{m})$.
Proof.
Recall that $M$ is spin if its second Stiefel–Whitney class $w_{2}:=w_{2}(TM)$ vanishes. By Wu’s formula, we have $Sq^{2}(a)=a\smile a=a\smile w_{2}$ for all $a\in\mathrm{H}^{2}(M;\mathbb{Z}/2\mathbb{Z})$. It is then readily seen that $Q_{M}$ is even if and only if $w_{2}=0$. The desired result then follows from Hasse-Minkowski classification theorem of indefinite unimodular symmetric bilinear form over $\mathbb{Z}$, and the fact that $\operatorname{sign}(Q_{M})=0$. ∎
To sum up, we suggest several Riemannian geometric properties of Hessian 4-manifolds, showing that a Hessian metric puts severe restrictions on other Riemannian metrics that a $4$-manifold can possess.
Proposition 5.21.
Let $((M,\nabla),h)$ be a compact Hessian 4-manifold. Then for any Riemannian metric $g$ on $M$, the Riemannian 4-manifold $(M,g)$ is not of negative sectional curvature.
Proof.
If $(M,g)$ is of negative sectional curvature, then, by Hopf’s conjecture in dimension four which is proved using Gauß–Bonnet formula, we have that $\chi(M)>0$, a contradiction to Proposition 2.17. ∎
Proposition 5.22.
Let $((M,\nabla),h)$ be a closed Hessian $4$-manifold. Suppose that $g$ is an Einstein metric on $M$. Then $(M,g)$ is a flat Riemannian manifold.
Proof.
By considering the orientation double cover, we may assume without loss of generality that $M$ is orientable.
Now, assume the contrary that the Einstein manifold $(M,g)$ is not flat. Then, since by Lemma 5.17
$\chi(M)=0=\frac{3}{2}\left|\operatorname{sign}\left(Q_{M}\right)\right|,$
by Hitchin–Thorpe inequality we have that the universal covering of $M$ is $\pi\colon X\rightarrow M$ where $X$ is a K3 surface. Since K3 surfaces are compact and without loss of generality assume $\pi$ is orientation-preserving, we have that $d:=\operatorname{deg}(\pi)<\infty$ is a positive integer. But as $\chi(X)=d\cdot\chi(M)$ by Riemann-Hurwitz formula, and it is well-known in algebraic geometry that $\chi(X)=24$, we arrive at $\chi(M)>0$, which contradicts Proposition 2.17. ∎
Appendix A On Homotopy Tori
Theorem 2.22 crucially uses the result in geometric topology that a homotopy torus is in fact homeomorphic to the flat torus, which we will now produce a proof for completeness. We shall first recall the following two propositions from low-dimensional topology.
Proposition A.1.
Let $M$ be a closed $3$-manifold. Suppose that there exists a closed surface $B$ with infinite fundamental group, and a finitely generated normal subgroup $G$ of $\pi_{1}(M)$, such that the following sequence
$1\longrightarrow G\longrightarrow\pi_{1}(M)\longrightarrow\pi_{1}(B)\longrightarrow 1$
is exact. Then $G=\mathbb{Z}$ and $M$ is an $\mathbb{S}^{1}$-bundle over $B$.
Proof.
See [33] for a combinatorial proof of this statement assuming Poincaré’s conjecture, while the latter is confirmed by G.Perelman in [34],[35], and [36]. See also [37]. ∎
Proposition A.2.
Let $f\colon M\rightarrow N$ be a homotopy equivalence between two closed aspherical 4-manifolds with polycyclic fundamental group. Then $f$ is homotopic to a homeomorphism.
Proof.
This is proved in [38]. ∎
Now we are ready to prove the main theorem regarding homotopy tori.
Theorem A.3.
Let $M$ be a closed topological manifold. If $M$ is homotopy equivalent to the flat $n$-torus $\mathbb{T}^{n}$, then $M$ is actually homeomorphic to $\mathbb{T}^{n}$.
Proof.
Comparing the top homology groups yields that $M$ is orientable and $\operatorname{dim}M=n$. For $n\leq 2$, it is well-known that closed $n$-manifolds are diffeomorphic if they are of the same homotopy type. For $n\geq 5$, the desired result is proved in [39]. It remains to consider the case when $n=3$ or $4$.
Assume that $n=4$. Since $\mathbb{Z}^{4}$ is polycyclic and the closed 4-manifold $\mathbb{T}^{4}$ is aspherical, by Proposition A.2 we conclude that $M$ is homeomorphic to $\mathbb{T}^{4}$.
Assume that $n=3$. By Proposition A.1, it is readily seen that $M$ is a $\mathbb{S}^{1}$-bundle over $\mathbb{T}^{2}$. But by the theory of 3-manifolds, two $\mathbb{S}^{1}$-bundles over $\mathbb{T}^{2}$ are homeomorphic if and only if their Euler class coincide. Assume that $M$ is the $\mathbb{S}^{1}$-bundles over $\mathbb{T}^{2}$ with Euler class $k\in\mathbb{Z}$. Then $\mathrm{H}_{1}(M;\mathbb{Z})=\mathbb{Z}^{2}\oplus\mathbb{Z}/k\mathbb{Z}$. But by Hurewicz’s theorem, $\mathrm{H}_{1}(M;\mathbb{Z})=\mathbb{Z}^{3}$ as the Abelianization of $\pi_{1}(M)$. Therefore we obtain that $k=0$, and hence $M$ is the trivial $\mathbb{S}^{1}$-bundles over $\mathbb{T}^{2}$, i.e. $M$ is homeomorphic to the 3-torus $\mathbb{T}^{3}$. ∎
Appendix B Pushforward Parallel Vectors along Harmonic Map
In this section, we prove that the length square of the pushforward of a parallel vector field under a harmonic mapping into a negatively curved space is subharmonic, which is very crucially used in the proof of Theorem 4.10.
We first introduce a preparatory lemma.
Lemma B.1.
Let $(M,g)$ and $(N,h)$ be Riemannian manifolds. Let $f\colon M\rightarrow N$ be a smooth mapping. Let $\hat{\nabla}:=\operatorname{Id}\otimes f^{*}\nabla^{h}+\nabla^{g}\otimes\operatorname{Id}$ be the induced connection on the tensor bundle $\mathscr{E}:=T^{*}M\otimes f^{*}TN$, where $\nabla^{g}$ is the Levi-Civita connection of $(M,g)$ and $\nabla^{h}$ is the Levi-Civita connection of $(N,h)$. Then, for any vector fields $X,Y,Z$ on $M$ and smooth section $\alpha$ of $\mathscr{E}\rightarrow M$, it holds that
$(\hat{\nabla}_{X}\alpha)(Y)=(f^{*}\nabla^{h})_{X}(\alpha(Y))-\alpha(\nabla^{g}_{X}Y)$
and that
$(\hat{\nabla}_{X}\hat{\nabla}_{Y}\alpha)(Z)=$ $(f^{*}\nabla^{h})_{X}(f^{*}\nabla^{h})_{Y}(\alpha(Z))-(f^{*}\nabla^{h})_{X}(\alpha(\nabla^{g}_{Y}Z))$
$-(f^{*}\nabla^{h})_{Y}(\alpha(\nabla^{g}_{X}Z))+\alpha(\nabla^{g}_{Y}\nabla^{g}_{X}Z).$
Proof.
The first identity follows directly from Leibniz rule and the definition of $\hat{\nabla}$.
Repeatedly applying the first identity yields
$(\hat{\nabla}_{X}\hat{\nabla}_{Y}\alpha)(Z)=$ $(f^{*}\nabla^{h})_{X}((\hat{\nabla}_{Y}\alpha)(Z))-(\hat{\nabla}_{Y}\alpha)(\nabla^{g}_{X}Z)$
$=$ $(f^{*}\nabla^{h})_{X}((f^{*}\nabla^{h})_{Y}(\alpha(Z))-\alpha(\nabla^{g}_{Y}Z))$
$-((f^{*}\nabla^{h})_{Y}(\alpha(\nabla^{g}_{X}Z))-\alpha(\nabla^{g}_{Y}\nabla^{g}_{X}Z))$
$=$ $(f^{*}\nabla^{h})_{X}(f^{*}\nabla^{h})_{Y}(\alpha(Z))-(f^{*}\nabla^{h})_{X}(\alpha(\nabla^{g}_{Y}Z))$
$-(f^{*}\nabla^{h})_{Y}(\alpha(\nabla^{g}_{X}Z))+\alpha(\nabla^{g}_{Y}\nabla^{g}_{X}Z)$
as desired. ∎
The following theorem as it stated is not only a tool to prove Theorem 4.10, but also reveals an important property of harmonic mappings.
Theorem B.2.
Let $(M,g)$ and $(N,h)$ be Riemannian manifolds with $\operatorname{dim}(N)\geq 2$. Let $f\colon M\rightarrow N$ be a harmonic mapping. Suppose that $(N,h)$ is of non-positive sectional curvature, and that $X$ is a parallel vector field on $(M,g)$. Then the energy density
$\frac{1}{2}(f^{*}h)(f_{*}X,f_{*}X)\in C^{\infty}(M)$
is a subharmonic function on $(M,g)$. Moreover, if we assume further that $M$ is compact and $(N,h)$ is of negative sectional curvature, then $f_{*}X\equiv 0$.
Proof.
Consider orthonormal frame fields $e_{1},\dots,e_{n}$ on $(M,g)$.
Let $\hat{\nabla}:=\operatorname{Id}\otimes f^{*}\nabla^{h}+\nabla^{g}\otimes\operatorname{Id}$ be the induced connection on the tensor bundle $\mathscr{E}:=T^{*}M\otimes f^{*}TN$, where $\nabla^{g}$ is the Levi-Civita connection of $(M,g)$ and $\nabla^{h}$ is the Levi-Civita connection of $(N,h)$.
Since $X$ is a parallel vector field on $(M,g)$ and $\hat{\nabla}(df)$ is symmetric, we have
$(\hat{\nabla}_{e_{i}}\hat{\nabla}_{X}(df))(e_{j})=$ $(f^{*}\nabla^{h})_{e_{i}}((\hat{\nabla}_{X}(df))(e_{j}))-(\hat{\nabla}_{X}(df))(\nabla^{g}_{e_{i}}e_{j})$
$=$ $(f^{*}\nabla^{h})_{e_{i}}((\hat{\nabla}_{e_{j}}(df))(X))-(\hat{\nabla}_{e_{j}}(df))(\nabla^{g}_{e_{i}}X)-(\hat{\nabla}_{X}(df))(\nabla^{g}_{e_{i}}e_{j})$
$=$ $(\hat{\nabla}_{e_{i}}\hat{\nabla}_{e_{j}}(df))(X)-(\hat{\nabla}_{X}(df))(\nabla^{g}_{e_{i}}e_{j})$
by Lemma B.1. Denote by $\hat{\Delta}$ the connection Laplacian of $\hat{\nabla}$, that is,
$(\hat{\Delta}(df))(X)=\sum_{i=1}^{n}(\hat{\nabla}_{e_{i}}\hat{\nabla}_{e_{i}}(df))(X)-\sum_{i=1}^{n}(\hat{\nabla}_{X}(df))(\nabla^{g}_{e_{i}}e_{j}).$
Again, since $X$ is a parallel vector field on $(M,g)$, we obtain $\operatorname{Ric}^{g}(-,X)=0$ and hence
$\sum_{i=1}^{n}R^{g}(X,e_{i})e_{i}=0$
where $R^{g}$ is the Riemann curvature tensor and $\operatorname{Ric}^{g}$ is the Ricci tensor of $(M,g)$.
Therefore, by the product rule for curvature tensors, we have
$\sum_{i=1}^{n}(\hat{R}(X,e_{i})(df))(e_{i})=$ $\sum_{i=1}^{n}R^{h}(f_{*}X,f_{*}e_{i})(f_{*}e_{i})-f_{*}\sum_{i=1}^{n}R^{g}(X,e_{i})e_{i}$
$=$ $\sum_{i=1}^{n}R^{h}(f_{*}X,f_{*}e_{i})(f_{*}e_{i})$
where $R^{h}$ is the curvature tensor of $\nabla^{h}$, and $\hat{R}$ is the curvature tensor of $\hat{\nabla}$.
Now, since $f$ is harmonic, a computation yields
$0=$ $\sum_{i=1}^{n}(f^{*}\nabla^{h})_{X}((\hat{\nabla}_{e_{i}}(df))(e_{i}))$
$=$ $\sum_{i=1}^{n}(\hat{\nabla}_{e_{i}}\hat{\nabla}_{X}(df))(e_{i})+\sum_{i=1}^{n}(\hat{R}(X,e_{i})(df))(e_{i})$
$=$ $(\hat{\Delta}(df))(X)+\sum_{i=1}^{n}R^{h}(f_{*}X,f_{*}e_{i})(f_{*}e_{i}).$
For simplicity, for any sections $s,s_{1},s_{2}$ of $f^{*}TN$, we write $|s|:=(f^{*}h)(s,s)$ and
$\langle s_{1},s_{2}\rangle:=(f^{*}h)(s_{1},s_{2}).$
Since $(N,h)$ is of non-positive sectional curvature, we have
$\langle(\hat{\Delta}(df))(X),f_{*}X\rangle=\sum_{i=1}^{n}\langle R^{h}(f_{*}e_{i},f_{*}X)(f_{*}e_{i}),f_{*}X\rangle\geq 0.$
Since $X$ is a parallel vector field on $(M,g)$, Weitzenböck identity reads
$\frac{1}{2}\Delta_{g}|f_{*}X|^{2}=\langle(\hat{\Delta}(df))(X),f_{*}X\rangle+\sum_{i=1}^{n}|(f^{*}\nabla^{h})_{e_{i}}(f_{*}X)|^{2}\geq 0$
where $\Delta_{g}$ here is the Laplace-Beltrami operator on $(M,g)$. The proof of the first claim is completed.
Now, we assume further that $M$ is compact. Then, by Liouville’s theorem, the energy density
$u:=\frac{1}{2}(f^{*}h)(f_{*}X,f_{*}X)\in C^{\infty}(M)$
is a constant function. Since $\Delta_{g}u=0$ and $\langle(\hat{\Delta}(df))(X),f_{*}X\rangle\geq 0$, we obtain
$\frac{1}{2}\Delta_{g}|f_{*}X|^{2}=\langle(\hat{\Delta}(df))(X),f_{*}X\rangle=\sum_{i=1}^{n}|(f^{*}\nabla^{h})_{e_{i}}(f_{*}X)|^{2}=0$
and hence
$\sum_{i=1}^{n}\langle R^{h}(f_{*}e_{i},f_{*}X)(f_{*}e_{i}),f_{*}X\rangle=0.$
Since $\langle R^{h}(f_{*}e_{i},f_{*}X)(f_{*}e_{i}),f_{*}X\rangle\geq 0$, we also have that
$\langle R^{h}(f_{*}e_{i},f_{*}X)(f_{*}e_{i}),f_{*}X\rangle=0$
for all $i=1,\dots,n$. Therefore, in particular, for any smooth vector field $Y$ on $N$, we have $\langle R^{h}(Y,f_{*}X)(Y),f_{*}X\rangle=0$. If $f_{*}X$ is non-vanishing at some point $p\in M$, then by the additional assumption on sectional curvatures, for any vector $\xi\in T_{f(p)}N$ that is not proportional to $f_{*}X$ at $f(p)\in N$, we have
$\langle R^{h}(\xi,f_{*}X)(\xi),f_{*}X\rangle>0$
which is a contradiction. Therefore we conclude that $f_{*}X\equiv 0$. ∎
Acknowledgement
The author wants to express the author’s most sincere gratitude to Weiyi Zhang for very useful discussions and general advices along the whole preparation process of this manuscript, and to Maxwell Stolarski for a careful proof reading with various valuable suggestions. The author also wants to thank Jun Zhang for pointing out a potential connection with Riemann extension, which results in an improvement of section3.2.
Special thanks should also go to Saul Schleimer who inspires the author’s work on Hessian 3-folds, and to Ruadhaí Dervan for a discussion about Kähler geometry of Hessian manifolds.
This research is completed while the author is studying at the mathematics institute of University of Warwick. The author therefore would like to thank the hospitality of University of Warwick.
The author is in debt to the Russian mathematical society, especially, to all the professors teaching the math in Moscow program, and most importantly, to the author’s supervisor Alexander Petrovich Veselov at Loughborough University, as they cultivated the author’s mathematical literacy and maturity.
Finally, the author would love to thank the author’s family and friends, for their kindness and steady help.
Statements and Declarations
No funding was received to assist with the preparation of this manuscript.
The author certifies that the author has no affiliations with or involvement in any other organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author and the source, provide a link to the Creative Commons licence, and indicate if changes were made. The images or other third party material in this article are included in the article’s Creative Commons licence, unless indicated otherwise in a credit line to the material. If material is not included in the article’s Creative Commons licence and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder.
The author hereby provide consent for the publication of the manuscript detailed above.