### v6

# On Topology of Compact Hessian Manifolds

Abstract
For a differentiable manifold $M$, a pair $(M,\nabla)$ is called an affine manifold if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. A Riemannian metric $g$ on $M$ is said to be a Hessian metric on $(M,\nabla)$ if the $(0,3)$-tensor $\nabla g$ is totally symmetric.
We investigate general topological properties of compact Hessian manifolds, in particular confirm a long-standing conjecture of S.S.Chern in this special case, and then apply these constraints to low dimensions, providing a thorough topological classification of complete Hessian surfaces and closed orientable Hessian manifolds of dimension less than $5$.
Keywords: Hessian Metric, Affine Manifold, Low-Dimensional Topology
1 Basics on Hessian Manifolds
1.1 Introduction and Background
The study of Hessian manifolds originates with the foundational work of J.-L. Koszul in 1961 [18]. While their roots lie in affine differential geometry, in recent decades, these geometric structures have attracted significant attention and emerged as indispensable tools across a diverse spectrum of mathematical and scientific disciplines. Their influence spans from the theoretical underpinnings of information geometry and complex analysis to practical applications in optimization and mathematical statistics, and even touches upon the abstract realms of homological mirror symmetry. This resurgence of interest highlights the deep connections these manifolds forge between seemingly disparate areas, offering a unifying geometric language to explore sophisticated phenomena.
In information geometry, the parameter space of a multi-dimensional smooth family of probability density functions is oftentimes a Hessian manifold, with its metric defined by the relative entropy, and the Amari–Chentsov tensor codifies its divergence function. Hessian metrics also arise naturally in convex optimization, where barrier functions induce complete Hessian metrics on convex domains. For more details, we refer to [3] and [20].
Also, for applications to mirror symmetry, the seminal work of Hitchin [14], Gross–Siebert [11], and Kontsevich–Soibelman [17] has shown that compact Hessian manifolds model the base of torus fibrations in the Strominger-Yau-Zaslow conjecture.
On the other hand, Cheng and Yau’s solution [6] to the real Monge–Ampère equation constructs complete affine hyperspheres of which Blaschke metrics are precisely global Hessian metrics with prescribed volume forms. Remarkably, these examples are linked to convex projective manifolds via the work of Loftin [21], Benoist–Hulin [5], and that of Danciger–Guéritaud–Kassel [7].
However, despite this rich tapestry, fundamental questions about Hessian manifolds remain open, particularly regarding the interplay between global topology and algebraic obstructions. In this manuscript, we study Hessian manifolds from topological perspectives:
After briefly summarizing the history and background of the theory of Hessian manifolds, our investigation begins by answering the natural question that in which case a Hessian manifold $((M,\nabla),h)$ may admit a global potential $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. Using techniques from sheaf cohomology and spectral sequences, we proved that A Hessian manifold admits a global potential if its fundamental group is finite (Theorem 2.4). However, foundational results by Shima and Yagi [24] in 1997 established that complete Hessian manifolds are aspherical, as their universal coverings are convex domains, which then immediately implies that the fundamental group of a compact Hessian manifold is infinite and torsion-free (Corollary 2.6 and Proposition 2.16). This contrast brings a historically significant class of Hessian manifolds into sharp focus: those of Koszul type, which admit global 1-form potentials.
Results from the literature [23] manifest that a compact affine manifold admits a Hessian metric of Koszul type if and only if its universal covering is developed into a salient convex cone. Compact affine manifolds of this kind are traditionally said to be hyperbolic. Utilizing tools from convex analysis, we prove a central structural result: any compact Hessian manifold with dimension no greater than $6$ is finitely covered by the product of a flat torus and a hyperbolic affine manifold (Theorem 4.15):
Theorem 1.
[…]
$((E,f^{*}\nabla),f^{*}h)=((B\times\mathbb{T}^{n-k},\hat{\nabla}\oplus D),g+ds^{2})$
[…]
This result is complemented by powerful fibration theorems, which state that any compact orientable Hessian manifold must be a mapping torus or a Bieberbach manifold (Proposition 4.2) and any compact hyperbolic affine manifold must admit a fibration over $\mathbb{S}^{1}$ with periodic monodromy (Corollary 4.12):
Theorem 2.
Let $(M,\nabla)$ be a compact oriented hyperbolic affine manifold. Then there exists a fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$ with connected and oriented fiber, such that the monodromy of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$ is periodic.
Finally, as one of the highlights of this article, we apply these general constraints to low dimensions. We provide the first topological classification of all complete Hessian surfaces (Theorem 5.6) and topological classification of closed orientable Hessian 3-manifolds and 4-manifolds (Theorem 5.9 and Theorem 5.15):
Theorem 3.
Let $((M,\nabla),h)$ be a complete Hessian surface. Then $M$ is diffeomorphic to a complete flat Riemannian manifold of dimension two.
Theorem 4.
Let $((M,\nabla),h)$ be a closed orientable Hessian $3$-manifold. Then there exists a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ such that $M$ is diffeomorphic to the mapping torus of $\varphi$, where $\Sigma_{g}$ is a closed orientable surface of genus $g\geq 1$.
Theorem 5.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. Then $M$ is diffeomorphic to one of the following smooth manifolds:
  1. 1.
    a Bieberbach 4-manifold;
  2. 2.
    the mapping torus of an isometry of a hyperbolic 3-manifold;
  3. 3.
    a $\Sigma_{g}$-bundle over $\mathbb{T}^{2}$ with periodic monodromy,
where $\Sigma_{g}$ here is the closed orientable surface of genus $g>1$.
Conversely, we also construct explicit examples, to show that indeed all $3$-manifolds mentioned above do admit Hessian structures (Proposition 5.7 and Proposition 5.8). Similar constructions in dimension four is left to the reader.
As an easy corollary of our fibration theorems, our work confirms Chern’s longstanding 1955 conjecture in this particular setting by proving that the Euler characteristic of any compact Hessian manifold must vanish (Proposition 4.3).
While our primary lens is topological, we also explore necessary analytical and geometric tools, including the Cheng-Yau metric, and a novel duality between Koszul type and radiant Hessian manifolds, resembling the classical Legendre transform in theoretical mechanics.
1.2 Affine and Hessian Manifolds
Throughout this article, all smooth manifolds are tacitly assumed to be connected and without boundary.
For each $i\in\{1,\ldots,n\}$, denote by $\pi^{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ the projection onto the $i$-th coordinate.
Let $M$ be an $n$-dimensional differentiable manifold. As usual, for any chart $(U,\varphi)$ of $M$, the smooth functions $x^{1},\ldots,x^{n}\in C^{\infty}(U)$ defined by
$x^{i}:U\longrightarrow\mathbb{R},\quad p\longmapsto\pi^{i}(\varphi(p))$
are said to be the local coordinates associated to the chart $(U,\varphi)$. An atlas
$\mathscr{A}=\left\{\left(U_{\alpha},\varphi_{\alpha}\right)\mid\alpha\in I\right\}$
of $M$ is said to be an affine atlas of $M$, if for any $\alpha,\beta\in I$ there exist $A\in\operatorname{GL}_{n}(\mathbb{R})$ and $\bf{a}\in$ $\mathbb{R}^{n}$ such that
$\varphi_{\beta}\left(\varphi_{\alpha}^{-1}(\bf{x})\right)=A{\bf{x}}+{\bf{a}}$
for all ${\bf{x}}\in\varphi_{\alpha}\left(U_{\alpha}\cap U_{\beta}\right)$. An element of an affine atlas of $M$ is then called an affine chart of $M$, and the local coordinates associated to an affine chart of $M$ are termed affine coordinates.
For any two affine atlases $\mathscr{A}_{1}$ and $\mathscr{A}_{2}$ of $M$, we write $\mathscr{A}_{1}\sim\mathscr{A}_{2}$ if and only if $\mathscr{A}_{1}\cup\mathscr{A}_{2}$ is also an affine atlas of $M$. An affine atlas of $M$ is said to be maximal, if for any other affine atlas $\mathscr{A}^{\prime}$ of $M$ we have that $\mathscr{A}^{\prime}\sim\mathscr{A}$ implies $\mathscr{A}^{\prime}\subseteq\mathscr{A}$. Traditionally, a maximal affine atlas on $M$ is also termed an affine structure of the smooth manifold $M$.
Recall that a connection on a vector bundle over $M$ is said to be flat, if its curvature 2-form vanishes identically.
Consider an arbitrary connection $D$ on the tangent bundle $TM\rightarrow M$. By Cartan-Ambrose-Hicks theorem, if $D$ is flat and torsion-free then there exists a maximal affine atlas $\mathscr{A}_{D}:=\left\{\left(U_{\alpha},\varphi_{\alpha}\right)\mid\alpha\in I\right\}$ of $M$, such that
$D(dx_{\alpha}^{1})=\cdots=D(dx_{\alpha}^{n})=0$
for all $\alpha\in I$, where $x_{\alpha}^{1},\ldots,x_{\alpha}^{n}$ are the local coordinates associated to the chart $\left(U_{\alpha},\varphi_{\alpha}\right)$. Moreover, the map $D]\longmapsto\mathscr{A}_{D}$ is a well-defined bijection from the set of all gauge equivalent classes of flat and torsion-free connections on the tangent bundle $TM\rightarrow M$ to the collection of all maximal affine atlases of $M$, and the atlas $\mathscr{A}_{D}$ is termed the adapted affine atlas of $D$. Therefore, fixing an affine structure on $M$ is equivalent to specifying a gauge equivalent class of flat and torsion-free connection on the tangent bundle $TM\rightarrow M$. In view of this fact, by a slightly abuse of notation we also refer to a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$ as an affine structure of $M$, and refer to a pair $(M,\nabla)$ as an affine manifold, if $\nabla$ is a flat and torsion-free connection on the tangent bundle $TM\rightarrow M$.
Remark 1.1.
All affine manifolds are in particular analytic, as their transition functions are by definition real analytic mappings.
On an affine manifold, there exist two important sheaves of Abelian groups:
Definition 1.2.
Let $(M,\nabla)$ be an affine manifold. The sheaf $\mathcal{A}:=\mathcal{A}_{M}$ of affine functions on $(M,\nabla)$ is defined by
$\mathcal{A}(U):=\{u\in C^{\infty}(U)\mid\nabla(du)=0\}$
where $U$ is any open subset of $M$. Sections of $\mathcal{A}$ are termed affine functions.
Definition 1.3.
Let $(M,\nabla)$ be an affine manifold. We denote by $\mathcal{P}:=\mathcal{P}_{M}$ the sheaf of parallel 1-forms on $(M,\nabla)$.
Recall from differential geometry that, on an $n$-dimensional affine manifold, the sheaf of parallel 1-forms is locally constant of rank $n$.
Moreover, since $\nabla$ is always assumed to be torsion-free, the exterior differential $d\omega$ of a $1$-form $\omega$ is the anti-symmetrization of the covariant derivative $\nabla\omega$. Therefore, in particular, on an affine manifold, all parallel 1-forms are closed.
We shall also recall the following standard definition of Hessian manifolds:
Definition 1.4.
Let $(M,\nabla)$ be an affine manifold. A Riemannian metric $h$ on $M$ is called a Hessian metric on $(M,\nabla)$, if for any point $p\in M$, there exists an open neighborhood $U$ of $p$ in $M$, such that $h|_{U}=\nabla(df)$ for some $f\in C^{\infty}(U)$; or equivalently, the $(0,3)$-tensor $\nabla h$ is totally symmetric. A pair $((M,\nabla),h)$ is termed a Hessian manifold, if $h$ is a Hessian metric on the affine manifold $(M,\nabla)$.
As Hessian manifolds are affine, we first present a theorem on topological invariants of affine manifolds, which of course applicable to all Hessian manifolds.
Proposition 1.5.
All Pontryagin classes of an affine manifold must vanish.
Proof.
Note that the tangent bundle of an affine manifold admits a connection of which curvature vanishes identically. The proof of the desired result is then a direct application of the Chern–Weil construction. ∎
It is also worth noting the following famous conjecture of Chern.
Conjecture 1.6 (Chern,1955).
The Euler characteristic of a compact affine manifold is zero.
Chern’s conjecture is known to hold in various cases. For example, the 2-dimensional case is settled by John Milnor [22] in 1957. Moreover, in [16], Bruno Klingler proved that if a compact affine manifold admits a parallel nowhere vanishing top differential form then its Euler characteristic is zero. In later sections we will also confirm this conjecture for all compact Hessian manifolds.
1.3 Concepts from Information Geometry
In this short section, we list several preliminary results that will become useful later.
We first recall the core concept of information geometry, that is, the notion of dual connections.
Definition 1.7.
Let $(M,g)$ be a pseudo-Riemannian manifold. For a connection $D$ on the tangent bundle $TM\rightarrow M$, its conjugate connection $D^{*}$ with respect to $g$ is defined to be the unique affine connection on the tangent bundle of $M$ such that the equation
$Zg(X,Y)=g\left(D_{Z}X,Y\right)+g\left(X,D_{Z}^{*}Y\right)$
holds for all smooth vector fields $X,Y,Z$ on $M$.
See [4] for a proof of the fact that the conjugate connection $D^{*}$ is well-defined and that $D^{**}=D$ for every connection $D$ on the tangent bundle $TM\rightarrow M$.
Theorem 1.8.
Let $(M,\nabla)$ be an affine manifold, and let $g$ be a pseudo-Riemannian metric on $(M,\nabla)$. Then $g$ is a Hessian metric on $(M,\nabla)$ if and only if $\nabla^{*}$ is also flat and torsion-free, or equivalently, the $(0,3)$-tensor $\nabla g$ is totally symmetric. Moreover, in this case
$\bar{\nabla}:=\frac{1}{2}\left(\nabla^{*}+\nabla\right)$
is the Levi-Civita connection of $(M,g)$.
Proof.
See [23] for a proof of this statement. ∎
Lemma 1.9.
Let $(M,g)$ be a pseudo-Riemannian manifold, and let $\nabla$ be a connection on the tangent bundle $TM\rightarrow M$. Then for any 1-form $\theta\in\Omega^{1}(M)$ and any vector field $X$ on $M$, it holds that $\nabla_{X}(\theta^{\sharp})=(\nabla^{*}_{X}\theta)^{\sharp}$, and in particular $\nabla^{*}\theta=0$ if and only if $\nabla(\theta^{\sharp})=0$, where $\sharp$ is the musical isomorphism.
Proof.
Take arbitrary vector fields $X,Y$ on $M$. Then
$(\nabla^{*}_{X}\theta)(Y)=X(\theta(Y))-\theta(\nabla^{*}_{X}Y).$
Now, since $\theta(Y)=g(\theta^{\sharp},Y)$, we have
$(\nabla^{*}_{X}\theta)(Y)=X(g(\theta^{\sharp},Y))-g(\theta^{\sharp},\nabla^{*}_{X}Y).$
Moreover, since
$X(g(\theta^{\sharp},Y))=g(\nabla_{X}\theta^{\sharp},Y)+g(\theta^{\sharp},\nabla^{*}_{X}Y),$
substituting this into the previous expression, we obtain
$(\nabla^{*}_{X}\theta)(Y)=(g(\nabla_{X}\theta^{\sharp},Y)+g(\theta^{\sharp},\nabla^{*}_{X}Y))-g(\theta^{\sharp},\nabla^{*}_{X}Y)=g(\nabla_{X}\theta^{\sharp},Y).$
This proves that $\nabla_{X}\theta^{\sharp}=(\nabla^{*}_{X}\theta)^{\sharp}$ holds for all vector fields $X$ on $M$. ∎
The following notion is also classical in information geometry.
Definition 1.10.
The Amari–Chentsov tensor $A$ of a Hessian manifold $((M,\nabla),h)$ is defined to be the $(0,3)$-tensor satisfying
$A(X,Y,Z)=(\nabla_{Z}h)(X,Y)=h(\nabla^{*}_{X}Y-\nabla_{X}Y,Z)$
for all vector fields $X,Y,Z$ on $M$.
Corollary 1.11.
Let $((M,\nabla),h)$ be a compact orientable Hessian manifold, and let $A$ be its Amari–Chentsov tensor. Suppose that $(M,h)$ is a flat Riemannian manifold. Then it holds that $A\equiv 0$.
Proof.
Since $(M,h)$ is a flat Riemannian manifold, we have that $\nabla$ is identical to the Levi-Civita connection of $(M,h)$, and hence $A=\nabla h=0$ by metric compatibility. ∎
2 Topological Generalities
2.1 Methods of Sheaf Cohomology
In this section, we study the very natural question that in which case a Hessian manifold $((M,\nabla),h)$ may admit a global potential $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. Before we can give an answer to this question, we need to state and prove the following two lemmata from algebraic topology and sheaf theory.
Lemma 2.1.
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. If $M$ is simply connected, then the first sheaf cohomology group $\mathrm{H}^{1}(M;\mathcal{A})$ vanishes.
Proof.
Since the holonomy of $(M,\nabla)$ is trivial and the sheaf $\mathcal{P}$ of parallel 1-forms on $M$ is a locally constant sheaf of rank $n$, by abuse of notation we may identify $\mathcal{P}$ with $\underline{\mathbb{R}}^{n}$. It is then readily seen that there exists a short exact sequence
$0\longrightarrow\mathbb{R}\longrightarrow\mathcal{A}\overset{d}{\longrightarrow}\mathbb{R}^{\oplus n}=\mathcal{P}\longrightarrow 0$
of sheaves, which induces a long exact sequence
$\cdots\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})\longrightarrow\mathrm{H}^{1}(M;\mathcal{A})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})^{\oplus n}\longrightarrow\cdots$
of cohomology groups. But since $M$ is simply connected, we have that the first de Rham cohomology group of $M$ vanishes. Therefore, we obtain $\mathrm{H}^{1}(M;\mathcal{A})=0$. ∎
Lemma 2.2.
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. Suppose that $\pi\colon\tilde{M}\rightarrow M$ is the universal covering. Then $\pi^{-1}\mathcal{A}$ is the sheaf of affine functions on the affine manifold $(\tilde{M},\pi^{*}\nabla)$, and there exists a natural isomorphism
$\mathrm{H}^{1}(M;\mathcal{A})\cong\mathrm{H}^{1}(G;\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))$
of $\mathbb{R}$-vector spaces, where $G:=\pi_{1}(M)$ is the fundamental group of $M$.
Proof.
It is routine to check that $(\tilde{M},\pi^{*}\nabla)$ is an affine manifold and $\pi^{-1}\mathcal{A}$ is its sheaf of affine functions.
To prove the existence of the natural isomorphism, we use Cartan-Leray spectral sequence, of which $E_{2}$-page terms are
$E_{2}^{p,q}=\mathrm{H}^{p}(G,\mathrm{H}^{q}(\tilde{M};\pi^{-1}\mathcal{A}))\implies\mathrm{H}^{p+q}(M;\mathcal{A}).$
Recall from homological algebra that the following sequence
$0\longrightarrow E^{1,0}_{2}\longrightarrow\mathrm{H}^{1}(M;\mathcal{A})\longrightarrow E^{0,1}_{2}\longrightarrow E^{2,0}_{2}\longrightarrow\cdots$
is exact. Since $\tilde{M}$ is simply connected, by Lemma 2.1 $\mathrm{H}^{1}(\tilde{M};\pi^{-1}\mathcal{A})=0$. Therefore we have that
$E^{0,1}_{2}=\mathrm{H}^{0}(G;0)=0.$
This proves that
$\mathrm{H}^{1}(M;\mathcal{A})=E^{1,0}_{2}=\mathrm{H}^{1}(G,\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))$
as required. ∎
The following vanishing theorem of sheaf cohomology group turns out to be useful for our purpose:
Proposition 2.3.
Let $(M,\nabla)$ be an affine manifold, and let $\mathcal{A}$ be the sheaf of affine functions on $(M,\nabla)$. If the fundamental group of $M$ is finite, then the first sheaf cohomology group $\mathrm{H}^{1}(M;\mathcal{A})$ vanishes.
Proof.
Take any universal covering $\pi\colon\tilde{M}\rightarrow M$. Then by Lemma 2.2
$\mathrm{H}^{1}(M;\mathcal{A})\cong\mathrm{H}^{1}(G;\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))$
as $\mathbb{R}$-vector spaces, where $G:=\pi_{1}(M)$ is the fundamental group of $M$. But, recall from the theory of group cohomology that, the first cohomology of a finite group is a torsion group. We therefore conclude that $\mathrm{H}^{1}(G;\mathrm{H}^{0}(\tilde{M};\pi^{-1}\mathcal{A}))=0$. ∎
We are now ready to state and prove the main theorem in this section.
Theorem 2.4.
Let $((M,\nabla),h)$ be a Hessian manifold. If the fundamental group of $M$ is finite, then there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$.
Proof.
By the definition of a Hessian metric, there exists an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a smooth function $f_{\alpha}\in C^{\infty}(U_{\alpha})$, such that for each $\alpha\in I$ it holds that $h_{\alpha}=\nabla(df_{\alpha})$, where $h_{\alpha}$ is the restriction of $h$ on $U_{\alpha}$.
For every $\alpha,\beta\in I$, define $g_{\alpha\beta}:=f_{\alpha}-f_{\beta}\in C^{\infty}(U_{\alpha}\cap U_{\beta})$. Then $g=\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ is a Čech 1-cocycle with coefficients in the sheaf $\mathcal{A}$ of affine functions on $M$. It remains to prove that the Čech cohomology class $g]\in\check{\mathrm{H}}^{1}(M;\mathcal{A})$ vanishes. But this immediately follows from Proposition 2.3. ∎
Before we proceed to the next section, also notice that we have the following obvious example for the non-existence of global potential for a pseudo-Riemannian metric.
Proposition 2.5.
Let $(M,g)$ be a compact pseudo-Riemannian manifold, and let $\nabla$ be a connection on the tangent bundle $TM\rightarrow M$. Then $g\neq\nabla(df)$ for all $f\in C^{\infty}(M)$.
Proof.
Take any $f\in C^{\infty}(M)$ such that the Hessian quadratic form $\nabla(df)$ of $f$ is everywhere non-degenerate. In particular $f$ is not a constant function. Since $M$ is compact, by the extreme value theorem, there exists distinct points $p,q\in M$ such that $f$ attains a global maximum at $p$ and attains a global minimum at $q$. But then by the second derivative test, the signatures of the Hessian quadratic form of $f$ must be different at points $p$ and $q$. ∎
Proposition 2.5 leads immediately to an interesting result:
Corollary 2.6.
The fundamental group of a compact Hessian manifold is infinite.
Proof.
Let $h$ be a Hessian metric on affine manifold $(M,\nabla)$. Assume the contrary that $\pi_{1}(M)$ is a finite group. Then by Proposition 2.4, there exists $f\in C^{\infty}(M)$ such that $h=\nabla(df)$. But this contradicts Proposition 2.5 as $M$ is assumed to be compact. ∎
Remark 2.7.
In [13], it is proved that there exists a unique closed orientable flat Riemannian 3-manifold with vanishing first Betti number, called the Hantzsche–Wendt manifold. Therefore, in particular, although by Corollary 2.6 a compact Hessian manifold admits an infinite fundamental group, its first Betti number can still be zero.
However, in contrast, we have the following proposition:
Proposition 2.8.
Let $(M,\nabla)$ be an affine manifold, and let $h$ be a Riemannian metric on $M$. Suppose that there exists $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$. Then $\eta$ is a closed 1-form. Moreover, if it is assumed further that $M$ is compact, then $\eta$ is not exact, and in particular the first Betti number of $M$ is positive.
Proof.
Since $\nabla$ is torsion-free, the exterior differential $d\eta$ is the anti-symmetrization of $\nabla\eta$. But the anti-symmetrization of $\nabla\eta$ is identically zero as $\nabla\eta=h$ is symmetric. Therefore, $\eta$ is a closed 1-form.
Now, assume that $M$ is compact. By Hodge decomposition theorem, there exists a smooth function $f\in C^{\infty}(M)$ and a harmonic 1-form $\theta$ such that $\eta=df+\theta$. Since
$\nabla\theta=\nabla\eta-\nabla(df)=h-\nabla(df)\neq 0$
by Proposition 2.5, we conclude that $\theta\neq 0$. Now it suffices to recall from Hodge theory that $\mathrm{H}^{1}(M;\mathbb{R)}$ is isomorphic to the space of the harmonic 1-forms on $M$. ∎
Proposition 2.8 motivates the following definition.
Definition 2.9.
Let $(M,\nabla)$ be an affine manifold, and $h$ a Riemannian metric on $M$. The pair $((M,\nabla),h)$ is termed a Hessian manifold of Koszul type, if there exists a $1$-form $\eta\in\Omega^{1}(M)$ such that $h=\nabla\eta$, and in this case, the 1-form $\eta$ is termed a $1$-form potential for $((M,\nabla),h)$.
By Poincaré lemma, Hessian manifolds of Koszul type are indeed Hessian. As explained in the introduction section, this is, actually, the ultimate origin of the notion of Hessian structures.
2.2 More on Fundamental Group
For each $n\in\mathbb{N}$, we denote by $\operatorname{Aff}_{n}(\mathbb{R})$ the group of affine transformations on $\mathbb{R}^{n}$.
Definition 2.10.
A Hessian manifold $((M,\nabla),h)$ is said to be complete, if $h$ is a complete Riemannian metric on $M$.
First, we recall the definition of affine mappings: For affine manifolds $(M_{1},\nabla_{1})$ and $(M_{2},\nabla_{2})$, a smooth mapping $\Phi:M_{1}\rightarrow M_{2}$ is said to be affine if $\hat{\nabla}(d\Phi)=0$, where
$\hat{\nabla}=\nabla_{1}\otimes\operatorname{Id}+\operatorname{Id}\otimes\Phi^{*}\nabla_{2}$
is the induced connection on the tensor bundle $T^{*}M_{1}\otimes\Phi^{*}TM_{2}$.
Let $(M,\nabla)$ be an affine manifold. It is readily seen that, a vector-valued smooth function
$F\colon M\longrightarrow\mathbb{R}^{m},\quad p\longmapsto(F_{1}(p),\dots,F_{m}(p))$
is an affine mapping between $(M,\nabla)$ and the flat Euclidean space, if and only if $F_{1},\dots,F_{m}$ are global affine functions on $M$.
For an affine manifold $(M,\nabla)$, since $\nabla$ is torsion-free, the exterior differential $d\omega$ of a $1$-form $\omega$ is the anti-symmetrization of the covariant derivative $\nabla\omega$. Therefore, on an affine manifold, all parallel 1-forms are closed.
Moreover, for an affine manifold $(M,\nabla)$ with finite fundamental group, parallel 1-forms admit no periods as the first de Rham cohomology group of $M$ vanishes. Therefore, the following notion of affine development is well-defined:
Definition 2.11.
Let $(M,\nabla)$ be a simply connected $n$-dimensional affine manifold, and fix a reference point $p_{0}\in M$. The affine development of $(M,\nabla)$ is the smooth mapping $\Phi\colon M\rightarrow\mathcal{P}(M)^{\vee}=\mathbb{R}^{n}$ defined by
$\langle\omega,\Phi(p)\rangle=\int_{\gamma}\omega$
for $\omega\in\mathcal{P}(M)$ and $p\in M$, where the bracket denotes the canonical pairing between $\mathcal{P}(M)$ and its dual space $\mathcal{P}(M)^{\vee}$, and where $\gamma\colon[0,1]\rightarrow M$ is a smooth curve such that $\gamma(0)=p_{0}$ and $\gamma(1)=p$.
See for example [23] for a proof of the fact that an affine development is indeed an affine mapping from an affine manifold to the flat Euclidean space.
In [24], using the concept of affine development, H.Shima and K.Yagi proved the following important theorem on convexity of complete Hessian metrics.
Theorem 2.12 (Shima-Yagi, 1997[24]).
Let $((M,\nabla),h)$ be a simply connected complete Hessian manifold. Then the affine development $\Phi\colon M\rightarrow\mathbb{R}^{n}$ of $(M,\nabla)$ is a diffeomorphism onto its image such that $\Phi(M)$ is a convex domain in $\mathbb{R}^{n}$.
Corollary 2.13.
[…]
Theorem 2.12 provides a strong restriction on the topology of complete Hessian manifolds.
Corollary 2.14.
Let $((M,\nabla),h)$ be a complete Hessian manifold. Then $M$ is aspherical, i.e. the homotopy group $\pi_{k}(M)$ vanishes for all $k\geq 2$.
Proof.
Recall that being an aspherical space is equivalent to possessing a contractible universal cover. The desired result then follows immediately from Theorem 2.12. ∎
Corollary 2.6 shows that compact Hessian manifolds have infinite fundamental group. We shall now improve the result in Corollary 2.6.
Lemma 2.15.
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$, and $G$ a group acting freely on $\Omega$. Then $G$ is torsion-free, that is, it contains no non-trivial element of finite order.
Proof.
Suppose that $g\in G$ enjoys finite order $k\in\mathbb{N}$. Take any arbitrary point $\textbf{x}\in\Omega$. Since $\Omega$ is convex and $g^{1}(\textbf{x}),\dots,g^{k}(\textbf{x})\in\Omega$, the barycenter
$\textbf{b}:=\frac{1}{k}(g(\textbf{x})+g^{2}(\textbf{x})+\cdots+g^{k}(\textbf{x}))$
is also a point in $\Omega$. But by the very construction we have that $g(\textbf{b})=\textbf{b}$. Therefore $g$ is the identity element of $G$, as the $G$-action is by assumption free. ∎
The following results are direct consequences of Lemma 2.15:
Proposition 2.16.
The fundamental group of a complete Hessian manifold is torsion-free.
Proof.
By Theorem 2.12, the fundamental group of a complete Hessian manifold acts freely on a convex domain. The desired result then follows from Lemma 2.15. ∎
Corollary 2.17.
The fundamental group of a complete Hessian manifold is either trivial or infinite.
Proof.
This follows from Proposition 2.16. ∎
Remark 2.18.
Let $((M,\nabla),h)$ be a compact Hessian manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be the universal covering of $M$. Then $((\tilde{M},\pi^{*}\nabla),\pi^{*}h)$ is a complete Hessian manifold. Provided that the fundamental group of $M$ is finite, then $\tilde{M}$ is compact, which contradicts Theorem 2.12. This gives another proof of Corollary 2.6.
Before we close this section, we note that Proposition 2.16 also straightforwardly implies the following interesting theorem:
Theorem 2.19.
Let $((M,\nabla),h)$ be a compact Hessian manifold of dimension $n$. If the fundamental group of $M$ is Abelian, then $M$ is homeomorphic to the $n$-torus $\mathbb{T}^{n}$.
Proof.
By Proposition 2.16, the fundamental group $\pi_{1}(M)$ of $M$ is isomorphic to $\mathbb{Z}^{k}$ for some $k\in\mathbb{N}$. By Corollary 2.14 and the uniqueness of Eilenberg-MacLane space, $M$ is homotopy equivalent to the $k$-torus $\mathbb{T}^{k}$. Comparing the top homology groups of $M$ and $\mathbb{T}^{k}$ yields that $k=n$. The desired result then follows from the Borel conjecture, which is proved for homotopy tori. See for example [15] and [8]. ∎
2.3 Flat Line Bundles
In this section, we study Hessian manifolds using the concept of flat line bundles. We also define and study the locally constant Picard group of a differentiable manifold, and use it to prove that a curved compact orientable Hessian manifold has positive first Betti number.
We shall now recall the notion of a flat $\mathbb{R}$-line bundle:
Definition 2.20.
For a $\mathbb{R}$-vector bundle $\mathscr{E}\rightarrow M$ over a differentiable manifold $M$, a pair $(\mathscr{E},D)$ is termed a flat $\mathbb{R}$-vector bundle over $M$ if $D$ is a flat connection on $\mathscr{E}\rightarrow M$. A flat $\mathbb{R}$-vector bundle of rank 1 is then called a flat $\mathbb{R}$-line bundle. In particular, we call $(M\times\mathbb{R},d)$ the trivial flat $\mathbb{R}$-line bundle over $M$.
To fix the terminology, we recall the definition of a local frame field, and that of its gluing 1-cocycles.
Let $\mathscr{L}\rightarrow M$ be a $\mathbb{R}$-line bundle over a differentiable manifold $M$. By the local triviality of $\mathscr{L}\rightarrow M$, there exists an open covering $\{U_{\alpha}\}_{\alpha\in I}$ of $M$ and for each $\alpha\in I$ a local section $s_{\alpha}$ of $\mathscr{L}\rightarrow M$ defined on $U_{\alpha}$, such that $s_{\alpha}$ is non-vanishing for all $\alpha\in I$. The collection $\{s_{\alpha}\}_{\alpha\in I}$ is traditionally termed a local frame field of $\mathscr{L}\rightarrow M$. For each $\alpha,\beta\in I$, there exists a unique nowhere vanishing smooth function $g_{\alpha\beta}\in C^{\infty}(U_{\alpha}\cap U_{\beta})$ such that
$s_{\beta}(p)=s_{\alpha}(p)\cdot g_{\alpha\beta}(p)$
for all $p\in U_{\alpha}\cap U_{\beta}$, where $U_{\alpha}$ here is the domain of definition of $s_{\alpha}$ and $U_{\beta}$ is the domain of definition of $s_{\beta}$. The collection $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ is termed the gluing 1-cocycle of the local frame field $\{s_{\alpha}\}_{\alpha\in I}$ of $\mathscr{L}$.
We then investigate the most basic properties of flat $\mathbb{R}$-line bundles.
Lemma 2.21.
Let $M$ be a differentiable manifold and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Then $\mathscr{L}\rightarrow M$ admits a local frame field $\{s_{\alpha}\}_{\alpha\in I}$ such that $Ds_{\alpha}=0$ for all $\alpha\in I$, and that the gluing 1-cocycle of the local frame field $\{s_{\alpha}\}_{\alpha\in I}$ consists of locally constant functions.
Proof.
Since the connection $D$ is flat, its connection 1-forms are locally exact. without loss of generality, we may assume that there exists $f\in C^{\infty}(U_{\alpha})$ such that the restriction of $D$ on $U_{\alpha}$ is $d+df$.
Now, it suffices to show that, up to a real constant multiple, the first order partial differential equation $ds+sdf=0$ for $s\in C^{\infty}(U_{\alpha})$ has a unique positive solution. But by the method of integration factor, we have
$e^{f}(ds+sdf)=d(e^{f}s)$
and hence the general solution of differential equation $ds+sdf=0$ is precisely $s=Ce^{-f}$, where $C\in\mathbb{R}$ is an integration constant. This proves the desired result. ∎
Lemma 2.21 motivates the following natural definition.
Definition 2.22.
Let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over a differentiable manifold $M$. A local frame field $\{s_{\alpha}\}_{\alpha\in I}$ of $\mathscr{L}\rightarrow M$ is termed an affine local frame field of $(\mathscr{L},D)$, if $Ds_{\alpha}=0$ for all $\alpha\in I$. The gluing 1-cocycle of some affine local frame field of $(\mathscr{L},D)$ is called a 1-cocycle of $(\mathscr{L},D)$.
Now we introduce the isomorphism classes in the category of flat $\mathbb{R}$-line bundle over a fixed differentiable manifold $M$.
Definition 2.23.
Two flat $\mathbb{R}$-line bundles $(\mathscr{L}_{1},D_{1})$ and $(\mathscr{L}_{2},D_{2})$ over a differentiable manifold $M$ are said to be gauge equivalent, if there exists a $\mathbb{R}$-vector bundle isomorphism $\phi\colon\mathscr{L}_{1}\rightarrow\mathscr{L}_{2}$ such that $D_{1}-\phi^{*}D_{2}\in\Omega^{1}(M)$ is an exact 1-form on $M$.
Let $M$ be a differentiable manifold. The collection of gauge equivalence classes of flat $\mathbb{R}$-line bundle over $M$ equipped with the tensor product $\otimes$ as binary operation forms a multiplicative group.
Definition 2.24.
Let $M$ be a differentiable manifold. The group of gauge equivalence classes of flat $\mathbb{R}$-line bundles over $M$ is termed the locally constant Picard group of $M$.
The following theorem, which relates the locally constant Picard group with cohomologies, is of fundamental importance.
Lemma 2.25.
Let $M$ be a differentiable manifold. The mapping that sends every flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ over $M$ to the Čech cohomology class of a 1-cocycle of $(\mathscr{L},D)$ in $\check{\mathrm{H}}(M;\mathbb{R}^{\times})$ is surjective, and two flat $\mathbb{R}$-line bundles over $M$ share the same image under this map if and only if they are gauge equivalent.
Proof.
This is a very special case of much more general statements in gauge theory. For a comprehensive proof, see [12]. ∎
We shall now have a closer look at the structure of the locally constant Picard group.
Proposition 2.26.
Let $M$ be a differentiable manifold. The induced homomorphism
$w_{1}\colon\mathrm{H}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}(M;\mathbb{Z}/2\mathbb{Z})$
in the long exact sequence
$\cdots\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R}^{\times})\longrightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\longrightarrow\cdots$
of cohomologies induced from the short exact sequence
$0\longrightarrow\mathbb{R}\overset{\operatorname{exp}}{\longrightarrow}\mathbb{R}^{\times}\overset{\operatorname{sgn}}{\longrightarrow}\mathbb{Z}/2\mathbb{Z}\longrightarrow 0$
of Abelian groups is precisely the first Stiefel-Whitney class.
Proof.
This statement follows from standard algebraic topology. ∎
A crucial difference between the exponential sequence in complex geometry and our short exact sequence
$0\longrightarrow\mathbb{R}\longrightarrow\mathbb{R}^{\times}\longrightarrow\mathbb{Z}/2\mathbb{Z}\longrightarrow 0$
is that the latter one splits. This allow us to define a new invariant for flat $\mathbb{R}$-line bundles.
Definition 2.27.
Let $M$ be a differentiable manifold, and let $\varphi\colon\mathbb{R}^{\times}\rightarrow\mathbb{R}$ be the even extension of the logarithm function, that is, $\varphi(x)=\operatorname{log}|x|$ for all $x\in\mathbb{R}^{\times}$. The induced homomorphism $\kappa_{1}:=\varphi_{*}\colon\mathrm{H}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}(M;\mathbb{R})$ in the long exact sequence
$\cdots\longrightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R}^{\times})\longrightarrow\mathrm{H}^{1}(M;\mathbb{R})\longrightarrow\cdots$
of cohomologies induced from the short exact sequence
$0\longrightarrow\mathbb{Z}/2\mathbb{Z}\longrightarrow\mathbb{R}^{\times}\overset{\operatorname{\varphi}}{\longrightarrow}\mathbb{R}\longrightarrow 0$
of Abelian groups is called the first Koszul class.
As an immediate consequence of Proposition 2.26 and Definition 2.27, the first Koszul class and the first Stiefel-Whitney class together classify the locally constant Picard group.
Corollary 2.28.
Let $M$ be a differentiable manifold. The mapping
$w_{1}\oplus\kappa_{1}\colon\mathrm{H}^{1}(M;\mathbb{R}^{\times})\rightarrow\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})\oplus\mathrm{H}^{1}(M;\mathbb{R}),\quad(\mathscr{L},D)\mapsto(w_{1}(\mathscr{L}),\kappa_{1}(D))$
is an isomorphism of Abelian groups.
Proof.
Apply the splitting lemma and use the exactness of the cohomology functor. ∎
We finally arrived at the structure theorem of locally constant Picard groups.
Corollary 2.29.
Let $M$ be a differentiable manifold. The locally constant Picard group of $M$ is isomorphic to the additive group $\mathrm{H}^{1}(M;\mathbb{R})\oplus\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$.
Proof.
The desired result follows from Lemma 2.25 and Corollary 2.28. ∎
We shall also be interested in a triviality criterion of elements of the locally constant Picard group. Recall that a line bundle is trivial if and only if it admits a nowhere vanishing global section. The following proposition gives a refinement in the case of flat $\mathbb{R}$-line bundles.
Proposition 2.30.
Let $M$ be an differentiable manifold, and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Then $(\mathscr{L},D)$ is gauge equivalent to the trivial flat $\mathbb{R}$-line bundle over $M$ if and only if it admits a nowhere vanishing parallel global section.
Proof.
Assume that $(\mathscr{L},D)$ is gauge equivalent to the trivial flat $\mathbb{R}$-line bundle over $M$. Then we identify $\mathscr{L}\equiv M\times\mathbb{R}$, and there exists an exact form $\theta\in\Omega^{1}(M)$ such that $D=d+\theta$. Take $f\in C^{\infty}(M)$ such that $\theta=df$. Then the global section
$s=\operatorname{exp}(-f)\in\mathrm{H}^{0}(M;M\times\mathbb{R})=C^{\infty}(M)$
is nowhere vanishing and satisfies that
$Ds=d\operatorname{exp}(-f)+e^{-f}\theta=e^{-f}(\theta-df)=0.$
To prove the converse statement, we assume that $(\mathscr{L},D)$ admits a nowhere vanishing parallel global section $s$. Then in particular we have $\mathscr{L}\cong M\times\mathbb{R}$. Moreover, since
$D(fs)=df\otimes s+f\cdot Ds=df\otimes s$
for all $f\in C^{\infty}(M)$, we conclude that $D=d$. ∎
To illustrate the power of the structure theorem of locally constant Picard group, we present a proof of the following torsion-free property.
Proposition 2.31.
Let $M$ be a differentiable manifold, and let $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$. Suppose that $\mathscr{L}$ is a trivial $\mathbb{R}$-line bundle, and there exists a positive integer $k$ such that $(\mathscr{L}^{\otimes k},D)$ is a trivial flat $\mathbb{R}$-line bundle. Then $(\mathscr{L},D)$ itself is a trivial flat $\mathbb{R}$-line bundle.
Proof.
Since $\mathscr{L}$ is a trivial $\mathbb{R}$-line bundle, its Stiefel-Whitney class $w_{1}(\mathscr{L})=0$. Therefore the gauge equivalent class of $(\mathscr{L},D)$ is contained in the subgroup $\mathrm{H}^{1}(M;\mathbb{R})$ of the locally constant Picard group of $M$. Since $(\mathscr{L}^{\otimes k},D)$ is trivial and $\mathrm{H}^{1}(M;\mathbb{R})$ is torsion-free, we conclude that $(\mathscr{L},D)$ is also trivial. ∎
2.4 Applications of Koszul Forms
Amongst all flat $\mathbb{R}$-line bundle over an affine manifold, the following one, which will be called the canonical line bundle, is of great importance.
Definition 2.32.
Let $(M,\nabla),$ be an affine manifold, and denote also by $\nabla$ the connection on tensor bundles of $M$ induced by $\nabla$. The flat $\mathbb{R}$-line bundle
$K_{\nabla}=(\bigwedge^{n}(T^{*}M),\nabla)$
over $M$ is termed the canonical line bundle of $(M,\nabla)$.
Note that, although the top exterior power of the cotangent bundle of an orientable manifold is always trivial, the canonical line bundle of an orientable affine manifold in general is non-trivial as a flat $\mathbb{R}$-line bundle.
Recall that a Calabi-Yau manifold is a Kähler manifold with trivial canonical line bundle, or equivalently, admitting a holomorphic volume form. The following corollary exhibits a close analogy.
Corollary 2.33.
Let $(M,\nabla)$ be an affine manifold of dimension $n$. Then the canonical line bundle $K_{\nabla}$ of $(M,\nabla)$ is trivial if and only if $M$ admits a nowhere vanishing $n$-form $\Omega$ such that $\nabla\Omega=0$.
Proof.
This directly follows from Proposition 2.30. ∎
In this section, we shall explore various applications of flat $\mathbb{R}$-line bundles and their first Koszul forms, including a powerful structure theorem of compact Hessian manifolds.
We now introduce the notion of the first Koszul form of a Hessian line bundle.
Definition 2.34.
Let $M$ be a differentiable manifold. A Hessian line bundle over $M$ is a pair $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, where $(\mathscr{L},D)$ be a flat $\mathbb{R}$-line bundle over $M$ and $\langle\cdot,\cdot\rangle$ is a bundle metric of $\mathscr{L}\rightarrow M$.
Definition 2.35.
Let $M$ be a differentiable manifold, and let $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ be a Hessian line bundle over $M$. The first Koszul form $\varkappa\in\Omega^{1}(M)$ of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ is defined by
$\varkappa|_{U_{\alpha}}:=\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
where $\{s_{\alpha}\}_{\alpha\in I}$ is an affine local frame of $(\mathscr{L},D)$ and $U_{\alpha}$ is the domain of definition of $s_{\alpha}$ for each $\alpha\in I$.
By the properties of affine local frame, the first Koszul form of a Hessian line bundle is well-defined, and is closed as it is by definition locally exact.
A natural question is to compare the first Koszul class of a flat $\mathbb{R}$-line bundle $(\mathscr{L},D)$ and the cohomology class of the first Koszul form of Hessian line bundle $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, where $\langle\cdot,\cdot\rangle$ is some bundle metric on $\mathscr{L}$. Surprisingly, the cohomology class of the first Koszul form is independent of the auxiliary bundle metric and always coincides with the first Koszul class.
Proposition 2.36.
Let $M$ be a differentiable manifold, and let $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$ be a Hessian line bundle over $M$. Let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((\mathscr{L},D),\langle\cdot,\cdot\rangle)$, then we have $\kappa_{1}(D)=[\varkappa]\in\mathrm{H}^{1}(M;\mathbb{R})$ is the first Koszul class of $(\mathscr{L},D)$.
Proof.
Let $\{s_{\alpha}\}_{\alpha\in I}$ be an affine local frame field for $(\mathscr{L},D)$ and let $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ be its gluing 1-cocycle.
By the very definition, the first Koszul class $\kappa_{1}(D)$ sends the class of the cocycle $\{g_{\alpha\beta}\}_{\alpha,\beta\in I}$ in $\mathrm{H}^{1}(M;\mathbb{R}^{\times})$ to the class of the cocycle $\{\operatorname{log}|g_{\alpha\beta}|\}_{\alpha,\beta\in I}$ in $\mathrm{H}^{1}(M;\mathbb{R})$.
By Definition 2.35, the first Koszul form $\varkappa$ is a global 1-form defined locally by
$\varkappa_{\alpha}:=\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
for $\alpha\in I$. The de Rham cohomology class $\varkappa]$ corresponds to a Čech cohomology class in $\mathrm{H}^{1}(M;\underline{\mathbb{R}})$ via the de Rham isomorphism. For any $\alpha\in I$, define
$f_{\alpha}:=\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle).$
Then for any $\alpha\in I$, we have
$d(f_{\alpha})=\frac{1}{2}d\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle).$
Straightforward computation yields that
$f_{\beta}-f_{\alpha}$ $=\frac{1}{2}\operatorname{log}(\langle s_{\beta},s_{\beta}\rangle)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\frac{1}{2}\operatorname{log}(g_{\alpha\beta}^{2}\langle s_{\alpha},s_{\alpha}\rangle)-\frac{1}{2}\operatorname{log}(\langle s_{\alpha},s_{\alpha}\rangle)$
$=\operatorname{log}|g_{\alpha\beta}|.$
Therefore, a Čech cocycle representing $\varkappa]$ is $\{\operatorname{log}|g_{\alpha\beta}|\}_{\alpha,\beta\in I}$, and hence $\kappa_{1}(D)=[\varkappa]$. ∎
On a Hessian manifold, we have a canonical Koszul form, which we will now define.
Definition 2.37.
Let $((M,\nabla),h)$ be a Hessian manifold and denote by $\langle\cdot,\cdot\rangle_{h}$ the bundle metric on tensor bundles of $M$ induced by $h$. The first Koszul form of the Hessian line bundle $(K_{\nabla},\langle\cdot,\cdot\rangle_{h})$ is also said to be the first Koszul form of $((M,\nabla),h)$.
For the convenience of the reader, we review the standard notion of first Koszul forms of a Hessian manifold.
For an orientable Riemannian manifold $(M,h)$, we denote by $\operatorname{vol}_{h}$ its volume form.
Proposition 2.38.
Let $((M,\nabla),h)$ be an oriented Hessian manifold, and let $\varkappa$ be its first Koszul form. Then it holds that $2\varkappa=d\operatorname{log}(|h|)$, and
$\nabla_{X}\operatorname{vol}_{h}=\varkappa(X)\operatorname{vol}_{h}$
for every vector field $X$ on $M$, where $|h|$ is the determinant of the matrix of which $(i,j)$-entry is $h(\partial/\partial x^{i},\partial/\partial x^{j})$ for some affine coordinates $x^{1},\dots,x^{n}$ of the oriented affine manifold $(M,\nabla)$.
Proof.
The proof is straightforward and hence omitted. ∎
In [24], H.Shima and K.Yagi established the following criterion for flatness of Hessian manifolds in terms of the first Koszul form.
Proposition 2.39 (Shima-Yagi, 1997[24]).
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold. Let $\bar{\nabla}$ be the Levi-Civita connection of $(M,h)$, and let $\varkappa\in\Omega^{1}(M)$ be the first Koszul form of $((M,\nabla),h)$. Then it holds that $\bar{\nabla}\varkappa=0$.
Theorem 2.40 (Shima-Yagi, 1997[24]).
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold of dimension $n$. Then the following statements are equivalent:
  1. 1.
    $(M,h)$ is Riemannian flat;
  2. 2.
    The first Koszul form of $((M,\nabla),h)$ vanishes identically;
  3. 3.
    $\nabla$ is the Levi-Civita connection of $(M,h)$;
  4. 4.
    $M$ admits a nowhere vanishing $n$-form $\Omega$ such that $\nabla\Omega=0$.
Recall that, in Remark 2.7, we found an example of an orientable compact Hessian manifold $((M,\nabla),h)$ with vanishing first Betti number, but in that case $(M,h)$ is in fact a flat Riemannian manifold. This is not just a coincidence.
Corollary 2.41.
Let $((M,\nabla),h)$ be an orientable compact Hessian manifold. If the first Betti number of $M$ is zero, then $(M,h)$ is a flat Riemannian manifold and $\nabla$ is its Levi-Civita connection.
Proof.
Since $\mathrm{H}^{1}(M;\mathbb{R})=0$, by Proposition 2.29 we have that the Picard group of $M$ is isomorphic to $\mathrm{H}^{1}(M;\mathbb{Z}/2\mathbb{Z})$, and hence is 2-torsion. In particular, the canonical line bundle $K$ of $(M,\nabla)$ satisfies $K^{\otimes 2}=0$. Since $M$ is orientable, the $\mathbb{R}$-line bundle $\bigwedge^{n}(T^{*}M)$ is trivial. Therefore, by virtue of Proposition 2.31, the canonical line bundle of $M$ is trivial as a flat $\mathbb{R}$-line bundle. Corollary 2.33 then implies that $M$ admits a volume form $\Omega$ such that $\nabla\Omega=0$. The desired result follows immediately from Theorem 2.40. ∎
3 Essential Tools from Analysis
3.1 Radiant Hessian manifolds
In this section, we introduce radiant manifolds and then study Hessian manifold in terms of the generalized Legendre duality mentioned above.
Definition 3.1.
Let $(M,\nabla)$ be an affine manifold. A vector field $H$ on $M$ is termed an Euler vector field of $(M,\nabla)$, if $\nabla_{X}H=X$ for every vector field $X$ on $M$. An affine manifold is said to be radiant, if it admits an Euler vector field.
Remark 3.2.
It is readily seen that, if $X$ and $Y$ are vector fields on an affine manifold such that $X$ is Eulerian, then $Y$ is also Eulerian if and only if XYX-Y is parallel.
Definition 3.3.
Let $M$ be a differentiable manifold. An atlas
$\mathscr{A}=\left\{\left(U_{\alpha},\varphi_{\alpha}\right)\mid\alpha\in I\right\}$
of $M$ is said to be a radiant affine atlas, if for any $\alpha,\beta\in I$ there exist $A\in\operatorname{GL}_{n}(\mathbb{R})$ such that
$\varphi_{\beta}\left(\varphi_{\alpha}^{-1}(\bf{x})\right)=A{\bf{x}}$
for all ${\bf{x}}\in\varphi_{\alpha}\left(U_{\alpha}\cap U_{\beta}\right)$. An element of a radiant affine atlas of $M$ is called a radiant affine chart of $M$, and the local coordinates associated to a radiant affine chart of $M$ are termed radiant affine coordinates.
The following theorem shows that the above two definitions of radiant affine manifolds are equivalent to each other.
Theorem 3.4.
An affine manifold $(M,\nabla)$ is radiant if and only if the adapted affine atlas of $\nabla$ is a radiant affine atlas of $M$.
Proof.
See for example [23] and [10]. ∎
The following propositions characterize the universal coverings of compact radiant affine manifolds as open cones in the Euclidean spaces.
Proposition 3.5.
[…]
Proof.
By Theorem 2.12, we may identify $\Omega:=\Phi(\tilde{M})=\tilde{M}$.
Since, up to a translation, $\Omega$ is a cone with apex at the origin, all affine automorphisms of $(\Omega,\pi^{*}\nabla)$ are in fact linear. Now, it suffices to descend the standard Euler vector field
$x^{1}\frac{\partial}{\partial x^{1}}+\cdots+x^{n}\frac{\partial}{\partial x^{n}}$
defined on $(\Omega,\pi^{*}\nabla)$ to an Euler vector field of $(M,\nabla)$. ∎
The converse statement also holds if we assume compactness.
Proposition 3.6.
[…]
Proof.
By Theorem 2.12, there exists a convex domain $\Omega$ in $\mathbb{R}^{n}$ such that $\Omega:=\Phi(\tilde{M})=\tilde{M}$. An Euler vector field $H$ of $(M,\nabla)$ lifts to a standard Euler vector field
$\pi^{*}H=\sum_{i=1}^{n}(x^{i}-a^{i})\frac{\partial}{\partial x^{i}}$
on $(\Omega,\pi^{*}\nabla)$, where $a^{1},\dots,a^{n}\in\mathbb{R}$. Moreover, by Poincaré-Hopf index theorem, we have $(a^{1},\dots,a^{n})\notin\Omega$. Up to an affine change of coordinates, we may assume $a^{1},\dots,a^{n}=0$. Since $M$ is compact, it follows that $H$ is a complete vector field. Since $\pi\colon\Omega\rightarrow M$ is a covering space, we obtain that $\pi^{*}H$ is also a complete vector field as $H$ is. Now it remains to recall that the integral curves of the standard Euler vector field
$x^{1}\frac{\partial}{\partial x^{1}}+\cdots+x^{n}\frac{\partial}{\partial x^{n}}$
are rays emitting from the origin of $\mathbb{R}^{n}$. ∎
In [10], D.Fried et al. studied the existence of parallel tensors on compact radiant affine manifolds.
Theorem 3.7 (Fried-Goldman-Hirsch,1981[10]).
A compact radiant affine manifold $(M,\nabla)$ does not admit a nonzero parallel $1$-form.
Theorem 3.8 (Fried-Goldman-Hirsch,1981[10]).
A compact radiant affine manifold $(M,\nabla)$ of dimension $n$ does not admit a nonzero parallel $n$-form.
Adapting the theorems above to our framework immediately yields the following corollaries.
Corollary 3.9.
Let $(M,\nabla)$ be a compact radiant manifold. Then the Euler vector field of $(M,\nabla)$ is unique.
Proof.
Combine Theorem 3.7 with Remark 3.2. ∎
Proposition 3.10.
Let $((M,\nabla),h)$ be a compact radiant Hessian manifold. Then the Riemannian manifold $(M,h)$ is not flat.
Proof.
By considering the orientation double cover of $M$, without loss of generality we may assume $M$ is orientable. Now the desired result follows immediately from Theorem 3.8 and Theorem 2.40. ∎
3.2 Legendre Transformation
It turns out that there exists a natural duality between radiant Hessian manifolds and Hessian manifolds of Koszul type. Also, for a radiant Hessian manifold $((M,\nabla),h)$ of Koszul type, we prove that its dual $((M,\nabla^{*}),h)$ is also a radiant Hessian manifold of Koszul type, and provide an explicit expression of the $1$-form potential of $((M,\nabla^{*}),h)$ in terms of that of $((M,\nabla),h)$, resembling the classical Legendre transformation.
The following two fundamental theorems illustrate the duality between radiant Hessian manifolds and Hessian manifolds of Koszul type.
Proposition 3.11.
Let $((M,\nabla),h)$ be a Hessian manifold of Koszul type. Then $((M,\nabla^{*}),h)$ is a radiant Hessian manifold. In fact, if $\eta$ is a $1$-form potential of $((M,\nabla),h)$, then $\eta^{\sharp}$ is an Euler vector field of $(M,\nabla^{*})$, where $\sharp$ is the musical isomorphism.
Proof.
Consider affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. For each $i=1,\dots,n$, define $p_{i}:=\eta(\partial/\partial x^{i})$, so that
$\eta=\sum_{i=1}^{n}p_{i}dx^{i}.$
Since $\partial/\partial x^{i}$ is parallel, we obtain that
$h=\nabla\eta=\sum_{i=1}^{n}dp_{i}\otimes dx^{i}.$
Then for every $i,j\in\{1,\dots,n\}$ we have that $h(\partial/\partial x^{i},\partial/\partial p_{j})=\delta^{i}_{j}$ and hence
$h(\nabla^{*}_{X}(\partial/\partial p_{j}),\partial/\partial x^{i})=$ $h(\nabla_{X}(\partial/\partial x^{i}),\partial/\partial p_{j})+h(\partial/\partial x^{i},\nabla^{*}_{X}(\partial/\partial p_{j}))$
$=$ $Xh(\partial/\partial x^{i},\partial/\partial p_{j})$
$=$ $X\delta^{i}_{j}=0$
for every vector field $X$ on $M$. Therefore $(dx^{i})^{\sharp}=\partial/\partial p_{i}$ and $\nabla^{*}(\partial/\partial p_{i})=0$ for all $i=1,\dots,n$. This proves that
$\eta^{\sharp}=\sum_{i=1}^{n}p_{i}(dx^{i})^{\sharp}=\sum_{i=1}^{n}p_{i}\frac{\partial}{\partial p_{i}}$
and it readily seen that the latter is an Euler vector field of $(M,\nabla^{*})$. ∎
Proposition 3.12.
Let $((M,\nabla),h)$ be a radiant Hessian manifold. Then $((M,\nabla^{*}),h)$ is a Hessian manifold of Koszul type. In fact, if $H$ is an Euler vector field of $(M,\nabla)$, then $H^{\flat}$ is a $1$-form potential of $((M,\nabla^{*}),h)$, where $\flat$ is the musical isomorphism.
Proof.
Consider radiant affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. For each $i=1,\dots,n$, since $\nabla(dx^{i})=0$, by Lemma 1.9 we have $\nabla^{*}(\partial/\partial x^{i})^{\flat}=0$. Recall that parallel 1-forms on an affine manifold is closed. By Poincaré’s lemma, there exist local coordinates $p_{1},\dots,p_{n}$ of $M$, such that $dp_{i}=(\partial/\partial x^{i})^{\flat}$ for all $i=1,\dots,n$. Therefore we arrive at the expression
$h=\sum_{i=1}^{n}dp_{i}\otimes dx^{i}$
for the Hessian metric $h$.
Since $x^{1},\dots,x^{n}$ are radiant affine, it is readily seen that the vector field
$X:=\sum_{i=1}^{n}x^{i}\frac{\partial}{\partial x^{i}}$
is a well-defined Euler vector field of $(M,\nabla)$. By the uniqueness of Euler vector field, there exists a parallel vector field $Y$ on $(M,\nabla)$, such that $H=X+Y$. Therefore, again by Lemma 1.9, we obtain that
$\nabla^{*}(H^{\flat})=\nabla^{*}(Y^{\flat})+\nabla^{*}\sum_{i=1}^{n}x^{i}(\partial/\partial x^{i})^{\flat}=0+\sum_{i=1}^{n}dx^{i}\otimes dp_{i}=\sum_{i=1}^{n}dp_{i}\otimes dx^{i}=h$
as required. ∎
Corollary 3.13.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type. Then the Hessian manifold $((M,\nabla^{*}),h)$ is also radiant of Koszul type.
Proof.
This follows immediately from Proposition 3.11 and Proposition 3.12. ∎
Now, in view of Corollary 3.13, for a radiant Hessian manifold $((M,\nabla),h)$ of Koszul type, it is natural to ask for an explicit expression for the $1$-form potential for its dual $((M,\nabla^{*}),h)$.
For a vector field $X$, as per usual, we denote by $\mathcal{L}_{X}$ the Lie derivative along $X$.
Definition 3.14.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type, with Euler vector field $H$ and $1$-form potential $\eta$. The differential form
$\eta^{*}:=-\eta+\mathcal{L}_{H}\eta$
is said to be the Legendre transform of $\eta$ along direction $H$.
Theorem 3.15.
Let $((M,\nabla),h)$ be a radiant Hessian manifold of Koszul type, with Euler vector field $H$ and $1$-form potential $\eta$. Then the Legendre transform $\eta^{*}$ of $\eta$ along direction $H$ satisfies $h=\nabla^{*}\eta^{*}$.
Proof.
Consider radiant affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$ so that
$H=\sum_{i=1}^{n}x^{i}\frac{\partial}{\partial x^{i}}$
is Eulerian. For each $i\in\{1,\dots,n\}$ define $p_{i}:=\eta(\partial/\partial x^{i})$. Then
$\eta=\sum_{i=1}^{n}p_{i}dx^{i}$
and $\nabla^{*}(dp_{1})=\cdots=\nabla^{*}(dp_{n})=0$. By Cartan’s magic formula, we obtain
$\eta^{*}=-\eta+\mathcal{L}_{H}\eta=d(\eta(H))-\eta=d(x^{1}p_{1}+\cdots+x^{n}p_{n})-\sum_{i=1}^{n}p_{i}dx^{i}=\sum_{i=1}^{n}x^{i}dp_{i}$
and hence
$\nabla^{*}\eta^{*}=\sum_{i=1}^{n}dx^{i}\otimes dp_{i}=h$
as required. ∎
Being both radiant and of Koszul type may seem to be a strong assumption on Hessian manifolds. However, when the underlying manifold is compact, a subtle observation of J.Vey shows that the radiant condition is redundant.
Definition 3.16.
A convex domain $\Omega$ in $\mathbb{R}^{d}$ is said to be regular, if it contains no straight line, i.e. a $1$-dimensional affine subspace of $\mathbb{R}^{d}$.
Proposition 3.17 (Vey,1970[26]).
Let $\Omega$ be a regular convex domain in $\mathbb{R}^{n}$. Suppose there exists a discrete subgroup $G$ of $\operatorname{Aff}(\Omega)$ acting freely and proper discontinuously on $\Omega$ so that $\Omega/G$ is compact. Then $\Omega$ is a cone.
We also recall a hyperbolicity condition in the context of affine differential geometry.
Definition 3.18.
Let $(M,\nabla)$ be an affine manifold, and let $\pi\colon\tilde{M}\rightarrow M$ be the universal covering of $M$. The affine manifold $(M,\nabla)$ is said to be hyperbolic, if the affine development $\Phi\colon\tilde{M}\rightarrow\mathbb{R}^{n}$ of $(\tilde{M},\pi^{*}\nabla)$ is a diffeomorphism onto its image, and its image is a regular convex domain in $\mathbb{R}^{n}$.
We observe that the notion of affine hyperbolicity is closed related to that of a Hessian metric of Koszul type.
Theorem 3.19.
A compact affine manifold is hyperbolic if and only if it admits a Hessian metric of Koszul type.
Proof.
This theorem is proved in [23]. ∎
The following Corollaries are then natural and facile.
Corollary 3.20.
A compact Hessian manifold is radiant if and only if it is of Koszul type.
Proof.
By Theorem 3.19 and Proposition 3.17, a compact Hessian manifold of Koszul type is automatically radiant.
Let $((M,\nabla),h)$ be a compact radiant Hessian manifold. Then by Proposition 3.12, we have that $((M,\nabla^{*}),h)$ is a compact Hessian manifold of Koszul type, and hence is radiant. Now, since $\nabla^{**}=\nabla$, the Hessian manifold $((M,\nabla),h)$ is also of Koszul type by Corollary 3.13. ∎
Corollary 3.21.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then the $1$-form potential of $((M,\nabla),h)$ is unique.
Proof.
Recall Corollary 3.20, and then apply Corollary 3.9 and Propositon 3.11. ∎
Corollary 3.22.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then the Riemannian manifold $(M,h)$ is not flat.
Proof.
This assertion follows from Corollary 3.20 and Propositon 3.10. ∎
Remark 3.23.
If the Hessian manifold $((M,\nabla),h)$ under consideration satisfies
  1. 1.
    $M=\Omega$ is a convex domain in $\mathbb{R}^{d}$;
  2. 2.
    $\nabla$ is the Levi-Civita connection of the $d$-dimensional flat Euclidean space;
  3. 3.
    $h$ is the Hessian quadratic form of a strictly convex potential,
then the construction in Definition 3.14 recovers the Legendre transformation in classical mechanics.
3.3 The Cheng-Yau Metric
Amongst all compact affine manifolds, those being hyperbolic are special, as they admit canonical Hessian metrics. We shall explain this observation in the present section.
For a smooth function $u$ defined on a domain in $\mathbb{R}^{n}$, we write
$\nabla u:=\left[\begin{array}[]{ccc}\partial_{1}u\\ \vdots\\ \partial_{n}u\end{array}\right],\quad\operatorname{Hess}(u):=\left[\begin{array}[]{ccc}\partial_{1}\partial_{1}u&\cdots&\partial_{1}\partial_{n}u\\ \vdots&\ddots&\vdots\\ \partial_{n}\partial_{1}u&\cdots&\partial_{n}\partial_{n}u\end{array}\right],$
where $\partial_{i}=\partial/\partial x^{i}$ for $i=1,\dots,n$, and $x^{1},\dots,x^{n}$ are the standard coordinates of $\mathbb{R}^{n}$.
Recall the following existence and uniqueness result on real Monge–Ampère equation due to Cheng and Yau.
Theorem 3.24 (Cheng-Yau, 1982[6]).
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$. Then, for any positive real number $\Lambda\in\mathbb{R}$, there exists a unique convex function $u\in C^{\infty}(\Omega)$ satisfying the following properties:
  1. 1.
    $\operatorname{det}(\operatorname{Hess}(u))=\operatorname{exp}(2\Lambda u)$;
  2. 2.
    $u(\textup{{x}})\rightarrow\infty$ as $\textup{{x}}\rightarrow\partial\Omega$;
so that in particular $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a complete Riemannian metric on $\Omega$.
Definition 3.25.
For a regular convex cone $\Omega$ in $\mathbb{R}^{n}$ and a positive real number $\Lambda\in\mathbb{R}$, the convex function $u\in C^{\infty}(\Omega)$ satisfying the properties listed in Theorem 3.24 is termed the Cheng-Yau potential of $\Omega$ with cosmological constant $\Lambda>0$.
Example 3.26.
Denote by $Q(x_{1},\dots,x_{n};y):=y^{2}-(x_{1}^{2}+\cdots+x_{n}^{2})$ the Lorentzian quadratic form on $\mathbb{R}^{1,n}$. The Cheng-Yau potential for the right circular cone $\Omega:=\{(x_{1},\dots,x_{n};y)\in\mathbb{R}^{n}\times\mathbb{R}:Q(x_{1},\dots,x_{n};y)>0\}$ is
$u(x_{1},\dots,x_{n};y)=-\frac{n}{2\Lambda}\operatorname{log}(\Lambda\cdot Q(x_{1},\dots,x_{n};y)/n)$
where $\Lambda>0$ is the cosmological constant.
Proposition 3.27.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$, and let $u\in C^{\infty}(\Omega)$ be a Cheng-Yau potential of $\Omega$. Then $u$ satisfies $du=L^{*}du$ for every linear automorphism $L$ of $\Omega$.
Proof.
See [2], or see [6] for the original proof. ∎
The differential of a Cheng-Yau potential enjoys one more important property, namely, it is always a covector field of constant length.
Lemma 3.28.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$, and let $u\in C^{\infty}(\Omega)$ be the Cheng-Yau potential of $\Omega$ with cosmological constant $\Lambda>0$. Then $u\in C^{\infty}(\Omega)$ satisfies $\langle\nabla u(\textup{{x}}),\textup{{x}}\rangle\equiv-n/\Lambda$, and
$\operatorname{Hess}(u)^{-1}\nabla u(\textup{{x}})=-\textup{{x}}$
for all $\textup{{x}}\in\Omega$, that is, $\operatorname{Hess}(u)^{-1}\nabla u$ is the standard Euler vector field of $\Omega$. In particular, it holds that $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv n/\Lambda$.
Proof.
For any $t>0$, define function $u_{t}\in C^{\infty}(\Omega)$ by $u_{t}(\textup{{x}}):=u(t\cdot\textup{{x}})+n\cdot\operatorname{log}(t)/\Lambda$. Then straightforward computation yields that $u_{t}$ is also a Cheng-Yau potential of $\Omega$ for every $t>0$. Therefore, by the uniqueness part of Theorem 3.24, we obtain the functional equation
$u(\textup{{x}})-u(t\cdot\textup{{x}})=n\cdot\operatorname{log}(t)/\Lambda$
where $\textup{{x}}\in\Omega$ and $k>0$. Now, for any $\textup{{x}}\in\Omega$, since
$\frac{u(\textup{{x}}+t\cdot\textup{{x}})-u(\textup{{x}})}{(n/\Lambda)t}=-\operatorname{log}(1+t)/t\rightarrow-1$
as $t\rightarrow 0$, we have that
$\langle\nabla u(\textup{{x}}),\textup{{x}}\rangle\equiv-n/\Lambda.$ (1)
Differentiate both sides of equation( 1) again, we obtain that
$\operatorname{Hess}(u)^{-1}\nabla u(\textup{{x}})=-\textup{{x}}$ (2)
holds for all $\textup{{x}}\in\Omega$. Substitute equation( 2) back into equation( 1), we finally arrive at $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv n/\Lambda$ as required. ∎
Adapt Cheng and Yau’s result to our setting, we immediately obtain the following crucial observation.
Proposition 3.29.
Let $(M,\nabla)$ be a compact hyperbolic affine manifold. Then, for any positive real number $\Lambda\in\mathbb{R}$, there exists a unique space-like Lagrangian embedding $\eta\in\Omega^{1}(M)$ of $(M,\nabla)$ such that $\varkappa=\Lambda\eta$, where $\varkappa$ is the first Koszul form of the Hessian manifold $((M,\nabla),h)$ and here $h=\nabla\eta$.
Proof.
[…]
$\Lambda\cdot df=\frac{1}{2}d\operatorname{log}(\operatorname{det}(\operatorname{Hess}(f)))$
descends to $\Lambda\cdot\eta=\varkappa$ on $(M,\nabla)$ so that $\eta$ is a space-like Lagrangian embedding of $(M,\nabla)$ and $\varkappa$ is the first Koszul form of $((M,\nabla),\nabla\eta)$. The uniqueness of $\eta$ also follows immediately from the uniqueness of the Cheng-Yau potential. ∎
Proposition 3.29 motivates the following more general definition.
Definition 3.30.
Let $((M,\nabla),h)$ be a Hessian manifold, and let $\varkappa\in\Omega^{1}(M)$ be its first Koszul form. The metric $h$ is said to be a Cheng-Yau metric with cosmological constant $\Lambda>0$ if $\nabla\varkappa=\Lambda\cdot h$. A Hessian manifold $((M,\nabla),h)$ is termed a Cheng-Yau manifold if $h$ is a Cheng-Yau metric.
We observe a property of Cheng-Yau metrics that their $1$-form potentials are self-dual under Legendre transformation.
Proposition 3.31.
Let $((M,\nabla),h)$ be a Cheng-Yau manifold with cosmological constant $\Lambda>0$. Let $\eta$ be the $1$-form potential of $((M,\nabla),h)$, and $\eta^{*}$ its Legendre transform. Then $\eta^{\sharp}$ is the unique Euler field of $(M,\nabla)$, and it holds that $\eta^{*}=-\eta$, where $\sharp$ is the musical isomorphism.
Proof.
For the first assertion, it suffices to lift $\eta\in\Omega^{1}(M)$ to the universal covering and apply Lemma 3.28. The uniqueness follows from Corollary 3.9.
Again by Lemma 3.28, we have
$\eta(\eta^{\sharp})=|\eta|^{2}=\operatorname{dim}(M)/\Lambda$
is a constant function. By Cartan’s magic formula and the definition of Legendre transformation, we have
$\eta^{*}=d(\eta(-\eta^{\sharp}))-\eta=-\eta$
as required. ∎
4 Fibration and Splitting Theorems
4.1 A Naive Attempt via Tischler’s Construction
Before we delve deeply into the structure theorems of Hessian manifolds, we first suggest here several naive observations on splitting of Hessian manifolds, which all relies on the following fundamental result of Tischler on fibration over circles.
Theorem 4.1 (Tischler, 1970[25]).
Let $M$ be a compact differentiable manifold. If $M$ admits a nowhere vanishing closed 1-form, then $M$ is a mapping torus.
Proposition 4.2.
Let $((M,\nabla),h)$ be an oriented compact Hessian manifold. Then either $(M,h)$ is a flat Riemannian manifold, or there exists a fiber bundle $\varphi\colon M\rightarrow\mathbb{S}^{1}$ with connected and oriented fiber.
Proof.
Let $\varkappa$ be the first Koszul form of $((M,\nabla),h)$, which is closed by definition. If $\varkappa=0$, then $(M,h)$ is a flat Riemannian manifold by Theorem 2.40. We therefore assume $\varkappa\neq 0$. But then since $\bar{\nabla}\varkappa=0$ by Proposition 2.39, we have that $\varkappa$ is nowhere vanishing by parallel transport.
The desired result now follows from Theorem 4.1. ∎
We now can confirm Chern’s conjecture in our special case:
Corollary 4.3.
The Euler characteristic of a compact Hessian manifold is zero.
Proof.
By considering the orientation double cover of $M$, without loss of generality we may assume $M$ is orientable.
By Proposition 4.2, $M$ is homeomorphic to a mapping torus or a Bieberbach manifold, and hence $\chi(M)=0$ by the multiplicativity of Euler characteristic. ∎
By showing that the $1$-form potential is also nowhere vanishing, we give an alternative proof of the fibration theorem.
Proposition 4.4.
The $1$-form potential of a compact Hessian manifold of Koszul type is a nowhere vanishing closed 1-form.
Proof.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type, and let $\eta\in\Omega^{1}(M)$ be its $1$-form potential.
View $\eta\colon M\rightarrow T^{*}M$ as an embedding. Then, $\eta$ vanishes at $p\in M$ if and only if $p$ is an intersection point of $\eta\colon M\rightarrow T^{*}M$ and the zero section $\sigma$ of the cotangent bundle of $M$. Since $\nabla\eta=h>0$, each intersection point of $\eta$ and $\sigma$ is a nondegenerate transverse intersection, and the intersection numbers at those intersection points are all positive. But, on the other hand, the sum of the intersection numbers at the intersection points of $\eta$ and $\sigma$ is precisely
$\langle e(T^{*}M),[M]\rangle=\chi(M)=0$
by Proposition 4.3. This proves that the sections $\eta$ and $\sigma$ must have no intersection, i.e. $\eta$ is a nowhere vanishing closed 1-form on $M$. ∎
Corollary 4.5.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then there exists a fiber bundle $\varphi\colon M\rightarrow\mathbb{S}^{1}$ with connected fiber. Moreover, if $M$ is oriented, then the fiber of $\varphi\colon M\rightarrow\mathbb{S}^{1}$ is also oriented.
Proof.
The desired result follows immediately from Theorem 4.1 and Proposition 4.4. ∎
4.2 Rational Approximation of Affine Structures
Making use of the existence of Cheng-Yau metric, we now introduce an important property of compact hyperbolic affine manifolds.
Definition 4.6.
The first Koszul class of the canonical line bundle of an affine manifold $(M,\nabla)$ is said to be the canonical class of $(M,\nabla)$.
We recall that, by Proposition 2.36, when the affine manifold under consideration is equipped with a Hessian metric, its first Koszul form represents precisely its canonical class.
Compact orientable Hessian manifolds with rational canonical classes are particularly well-behaved, as their mapping tori structures are compatible with their Hessian metrics.
Theorem 4.7.
Let $((M,\nabla),h)$ be a compact oriented Hessian manifold with non-zero rational canonical class. Then there exists a Riemannian submersion $f\colon M\rightarrow\mathbb{S}^{1}$ with connected and oriented fiber $F$, such that the monodromy of $f\colon M\rightarrow\mathbb{S}^{1}$ is an isometry of $(F,h|_{F})$ and in particular is of finite order.
Proof.
Let $\varkappa$ be the first Koszul form and $\bar{\nabla}$ the Levi-Civita connection of $((M,\nabla),h)$. Since $\bar{\nabla}\varkappa=0$ by Proposition 2.39 and $\varkappa\neq 0$ by assumption, we have that $\varkappa$ is nowhere vanishing via parallel transport.
Since $\varkappa]\in\mathrm{H}^{1}(M;\mathbb{Q})$, the rational number
$m:=\operatorname{min}\{k\in\mathbb{R}:k>0,k[\varkappa]\in\mathrm{H}^{1}(M;\mathbb{Z})\}$
is well-defined. Fix a base point $q$ on $M$, and define the smooth mapping $f\colon M\rightarrow\mathbb{S}^{1}$ via
$f(p):=\operatorname{exp}\left(2\pi\sqrt{-1}\int_{\gamma}m\varkappa\right)$
where $\gamma\colon[0,1]\rightarrow M$ is a smooth curve with $\gamma(0)=q$ and $\gamma(1)=p$. Since the periods of $m[\varkappa]$ are integers, the map $f$ is well-defined. Moreover, since $\varkappa$ is nowhere vanishing, we have that $f$ is a submersion. By Ehresmann’s fibration theorem, we have that $f\colon M\rightarrow\mathbb{S}^{1}$ is a fiber bundle. Denote by $F$ the fiber of $f\colon M\rightarrow\mathbb{S}^{1}$.
Identifying each homotopy class in $\pi_{1}(\mathbb{S}^{1})$ with its winding number, it is readily seen that the induced homomorphism $f_{*}\colon\pi_{1}(M)\rightarrow\pi_{1}(\mathbb{S}^{1})=\mathbb{Z}$ satisfies
$f_{*}([\gamma])=\int_{\gamma}m\varkappa$
for all $\gamma]\in\pi_{1}(M)$. Since $\varkappa\neq 0$, we have that $\operatorname{Im}(f_{*})\neq 0$. Therefore, there exists a positive integer $d\in\mathbb{Z}$ such that $\operatorname{Im}(f_{*})=d\mathbb{Z}$. But by the minimality of the factor $m$, we must have that $d=1$ and hence $\operatorname{Im}(f_{*})=\mathbb{Z}$. Now, since $M$ is connected and $f_{*}$ is surjective, the long exact sequence
$\cdots\longrightarrow\pi_{1}(M)\longrightarrow\pi_{1}(\mathbb{S}^{1})\longrightarrow\pi_{0}(F)\longrightarrow\pi_{0}(M)\longrightarrow\cdots$
of homotopy groups yields that $F$ is also a connected topological space.
Moreover, since $f\colon M\rightarrow\mathbb{S}^{1}$ is a submersion between compact differentiable manifolds, we have that $F$ is also a compact differentiable manifold. Also, the fiber $F$ is oriented as $M$ is.
Denote by $\varphi$ the monodromy of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$. Let $\pi\colon\tilde{M}\rightarrow M$ be the universal covering of $M$. Then there exists a complete Riemannian metric $g$ on $\mathbb{R}^{n}$ such that $(\tilde{M},\pi^{*}h)$ is isometric to the product of Riemannian manifolds $(\mathbb{R}^{n},g)$ and $(\mathbb{R},dt^{2})$. Denote by $\hat{\varphi}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ the lifting of $\varphi\colon F\rightarrow F$ to the universal covering spaces. Since $\pi_{1}(M)=\pi_{1}(F)\rtimes_{\varphi}\mathbb{Z}$ acts on $(\mathbb{R}^{n}\times\mathbb{R},g\oplus dt^{2})$ by isometry, in particular the smooth mapping
$\Phi\colon\mathbb{R}^{n}\times\mathbb{R}\rightarrow\mathbb{R}^{n}\times\mathbb{R},\quad(\textbf{x},t)\mapsto(\hat{\varphi}(\textbf{x}),t+1)$
is an isometry. Therefore we conclude that $\hat{\varphi}$ and hence $\varphi$ are isometries. Furthermore, since $f\colon M\rightarrow\mathbb{S}^{1}$ lifts to the coordinate projection $\mathbb{R}^{n}\times\mathbb{R}\rightarrow\mathbb{R}$ on universal covering spaces, we conclude that $f$ is a Riemannian submersion.
Denote by $i\colon F\rightarrow M$ the inclusion map of the fiber of the fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$. Consider the subgroup $A$ of the isometry group $G$ of $(F,i^{*}h)$ generated by $\varphi$, and denote by $\bar{A}$ its closure in $G$. Then it is readily seen that $\bar{A}$ is a compact abelian subgroup of $G$, and hence the identity component of $\bar{A}$ is a torus, Therefore, $\varphi$ is isotopic in $G$ to a group element of finite order. ∎
We shall show that, on an oriented compact manifold, an hyperbolic affine structure can always be perturbed into another hyperbolic affine structure of which canonical class is rational.
We first introduce the moduli space of radiant affine structures on a differentiable manifold.
Lemma 4.8.
Let $M$ be an $n$-dimensional differentiable manifold and $G:=\pi_{1}(M)$ its fundamental group. Then the moduli space of radiant affine structures on $M$ is a real analytic subvariety of the character variety $\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))/\operatorname{GL}_{n}(\mathbb{R})$.
Proof.
The moduli space of flat connections on the frame bundle of $M$ is the character variety $\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))/\operatorname{GL}_{n}(\mathbb{R})$, whilst the torsion-free condition imposes a system of transcendental equations on $\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))/\operatorname{GL}_{n}(\mathbb{R})$. ∎
To proceed, we adapt a theorem of J.-L.Koszul to our setting.
Theorem 4.9 (Koszul, 1968[19]).
Let $(M,\nabla)$ be a compact hyperbolic affine manifold. Then the collection of hyperbolic affine structures on $M$ is an non-empty open subset of the moduli space of radiant affine structures on $M$.
Applying Theorem 4.9 together with group cohomologies with Lie algebra coefficients, we can prove the following density or approximation theorem:
Definition 4.10.
For an affine manifold $(M,\nabla)$, a sequence $\{\nabla^{i}\}_{i=1}^{\infty}$ of flat torsion-free connections on the tangent bundle $TM\rightarrow M$ is said to be a convergent sequence of differential operators with limit $\nabla$, if for any vector fields $X,Y$ on $M$, the sequence $\{\nabla^{i}_{X}Y\}_{i=1}^{\infty}$ of vector fields converges uniformly to $\nabla_{X}Y$.
Theorem 4.11.
Let $(M,\nabla)$ be a compact hyperbolic affine manifold. Then there exists a sequence $\{\nabla^{i}\}_{i=1}^{\infty}$ of hyperbolic affine structures converging to $\nabla$, such that for every integer $i\geq 1$ the canonical class of $(M,\nabla^{i})$ is rational.
Proof.
For simplicity, denote by $G:=\pi_{1}(M)$ the fundamental group of $M$. Let $\mathfrak{X}$ be the moduli space of radiant affine structures on $M$. By Lemma 4.8, $\mathfrak{X}$ is a real analytic subvariety of the character variety $\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))/\operatorname{GL}_{n}(\mathbb{R})$. By algebraic geometry, the singular locus of $\mathfrak{X}$ is of positive codimension. Also, recall that by Theorem 4.9 the collection of hyperbolic affine structures on $M$ is an non-empty open subset of $\mathfrak{X}$. Therefore, without loss of generality, we may assume that the gauge equivalent class $A:=[\nabla]$ of $\nabla$ is a smooth point of $\mathfrak{X}$.
The $\operatorname{GL}_{n}(\mathbb{R})$-invariant epimorphism
$\varphi\colon\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))\rightarrow\operatorname{Hom}(G,\mathbb{R}^{\times}),\quad\rho\mapsto(g\mapsto\operatorname{det}(\rho(g)))$
induces a surjective regular morphism
$\hat{\varphi}\colon\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))/\operatorname{GL}_{n}(\mathbb{R})\rightarrow\operatorname{Hom}(G,\mathbb{R}^{\times}),\quad[\rho]\mapsto\varphi(\rho).$
Since the induced homomorphism
$\operatorname{tr}_{*}\colon\mathrm{H}^{1}(G;\mathfrak{gl}_{n}(\mathbb{R}))\rightarrow\mathrm{H}^{1}(G;\mathbb{R})$
of the trace map
$\operatorname{tr}\colon\mathfrak{gl}_{n}(\mathbb{R})\rightarrow\mathbb{R}$
is surjective as the short exact sequence
$0\rightarrow\mathfrak{sl}_{n}(\mathbb{R})\rightarrow\mathfrak{gl}_{n}(\mathbb{R})\rightarrow\mathbb{R}\rightarrow 0$
[…]
Let $h$ be the Cheng-Yau metric of $(M,\nabla)$ with cosmological constant $\Lambda=1$, and let $\eta$ be its $1$-form potential. Recall from Proposition 3.31 that the Legendre transform $\eta^{*}$ of $\eta$ is precisely $\eta$, and hence $\eta^{\sharp}$ is the Euler vector field of $(M,\nabla)$ where $\sharp$ is the musical isomorphism induced by $h$. Consider the $(1,2)$-tensor $B$ defined by
$B(X,Y):=\eta(X)Y+\eta(Y)X-h(X,Y)\eta^{\sharp}$
for all vector fields $X,Y$ on $M$. Since $B(X,Y)=B(Y,X)$ and $(\nabla_{X}B)(Y,Z)=(\nabla_{Y}B)(X,Z)$ for all vector fields $X,Y,Z$ on $M$ by the very construction of $B$, it is readily seen that $B\in T_{A}(\mathfrak{X})$ is an infinitesimal deformation of hyperbolic affine structures on $M$. Moreover, a computation yields that
$\operatorname{tr}_{h}(B)=\eta^{*}+(n+1)\eta=\operatorname{dim}(M)\eta$
is closed but not exact by Proposition 2.8.
[…]
By restricting $\hat{\varphi}$ to a sufficiently small neighborhood of $A$ in $\mathfrak{X}$, the desired result follows immediately from the open mapping theorem for submersions and the density of rational cohomology classes in de Rham cohomology groups. ∎
Theorem 4.11 has profound consequence, which significantly strengthen the previous Corollary 4.5.
Corollary 4.12.
Let $(M,\nabla)$ be a compact oriented hyperbolic affine manifold. Then there exists a fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$ with connected and oriented fiber, such that the monodromy of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$ is periodic.
Proof.
By Corollary 3.22 and Theorem 2.40, the canonical class $K$ of $(M,\nabla)$ is non-trivial. By Theorem 4.11, we may assume $K$ is a non-trivial integral cohomology class. Now it suffices to apply Theorem 4.7 to $((M,\nabla),h)$ for the Cheng-Yau metric $h$ of $(M,\nabla)$. ∎
4.3 Convex Decomposition Lemma
We shall now prove yet another splitting theorem for compact Hessian manifolds, showing that up to a finite covering every compact Hessian manifold is a product of a hyperbolic affine manifold with a flat torus.
We start by introducing an easy observation:
Lemma 4.13.
Let $((M,D),g)$ be a Hessian manifold and $(N,\nabla)$ an affine manifold. Suppose that $i\colon N\rightarrow M$ is an affine immersion. Then $((N,\nabla),i^{*}g)$ is a Hessian manifold.
Proof.
Since $i\colon N\rightarrow M$ is affine, we have that $\nabla=i^{*}D$, and hence $\nabla(i^{*}g)=i^{*}(Dg)$. Since $((M,D),g)$ is a Hessian manifold, we obtain that $\nabla(i^{*}g)$ is totally symmetric by Theorem 1.8. Therefore $\nabla(i^{*}g)$ is totally symmetric as $Dg$ is. Again by Theorem 1.8, we conclude that $((N,\nabla),i^{*}g)$ is a Hessian manifold. ∎
We will also need the following result on the famous Auslander conjecture.
Theorem 4.14 (Abels et al.,2002[1]).
Let $(M,\nabla)$ be a compact affine manifold of dimension $n\leq 6$, and $\pi\colon\tilde{M}\rightarrow M$ the universal covering of $M$. Let $\Phi\colon\tilde{M}\rightarrow\mathbb{R}^{n}$ be the affine development of $(\tilde{M},\pi^{*}\nabla)$. Suppose that $\Phi$ is a diffeomorphism. Then $(M,\nabla)$ admits a nowhere vanishing $n$-form $\Omega$ such that $\nabla\Omega=0$.
Recall that every convex domain in $\mathbb{R}^{d}$ can be written as the Cartesian product of a regular convex domain and a affine subspace of $\mathbb{R}^{d}$. This fact motivates the following splitting theorem for compact Hessian manifolds.
Theorem 4.15.
[…]
$((E,f^{*}\nabla),f^{*}h)=((B\times\mathbb{T}^{n-k},\hat{\nabla}\oplus D),g+ds^{2})$
[…]
Proof.
[…]
Since $\Omega$ contains no affine subspaces of positive dimension, for any $\gamma\in\pi_{1}(M)$, there exists $\phi_{\gamma}\in\operatorname{Aff}(\Omega)$ and $U_{\gamma}\in\operatorname{GL}(\mathbb{R}^{d})$ such that
$\gamma(\textbf{x},\textbf{y})=(\phi_{\gamma}(\textbf{x}),T_{\gamma}\textbf{x}+U_{\gamma}\textbf{y}+\textbf{v}_{\gamma})$
for some real $(n-d)\times d$-matrix $T_{\gamma}$ and some vector $\textbf{v}_{\gamma}\in\mathbb{R}^{d}$. Consider the homomorphism
$\phi\colon\pi_{1}(M)\rightarrow\operatorname{Aff}(\Omega),\quad\gamma\mapsto\phi_{\gamma}.$
For each $\textbf{x}\in\Omega$ denote
$N_{\textbf{x}}:=\frac{\{\textbf{x}\}\times\mathbb{R}^{d}}{\operatorname{Ker}(\phi)}$
where the normal subgroup $K:=\operatorname{Ker}(\phi)$ of $\pi_{1}(M)$ acts naturally on the vertical fiber $\{\textbf{x}\}\times\mathbb{R}^{d}$ by affine transformation.
[…]
Therefore, by Theorem 2.4, there exists a convex function $\varphi\in C^{\infty}(\Omega\times\mathbb{R}^{d})$ such that
$\pi^{*}h=\sum_{i=1}^{n-d}\sum_{j=1}^{n-d}\frac{\partial^{2}\varphi}{\partial x_{i}\partial x_{j}}dx_{i}dx_{j}+2\sum_{i=1}^{n-d}\sum_{j=1}^{d}\frac{\partial^{2}\varphi}{\partial x_{i}\partial y_{j}}dx_{i}dy_{j}+\sum_{i=1}^{d}\sum_{j=1}^{d}\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}dy_{i}dy_{j}$
and moreover,
$(\pi^{*}h)|_{\{\textbf{a}\}\times\mathbb{R}^{d}}=\sum_{i=1}^{d}\sum_{j=1}^{d}\frac{\partial^{2}\varphi}{\partial y_{i}\partial y_{j}}(a_{1},\dots,a_{n-d},y_{1},\dots,y_{d})dy_{i}\otimes dy_{j}$
is a flat Riemannian metric for any $\textbf{a}=(a_{1},\dots,a_{n-d})\in\Omega$. Therefore, by Corollary 1.11, we obtain that
$\frac{\partial^{3}\varphi}{\partial y_{i}\partial y_{j}\partial y_{k}}=0$
for all $i,j,k\in\{1,\dots,d\}$, and hence there exist a smooth function $u\in C^{\infty}(\Omega)$, a vector-valued function $\xi\in C^{\infty}(\Omega)^{\oplus d}$, and a matrix-valued function $F\in C^{\infty}(\Omega)^{d\times d}$ such that $\varphi(\textbf{x},\textbf{y})=u(\textbf{x})+\langle\xi(\textbf{x}),\textbf{y}\rangle+\langle F(\textbf{x})\textbf{y},\textbf{y}\rangle$ for all $(\textbf{x},\textbf{y})\in\Omega\times\mathbb{R}^{d}$.
Recall that an element $\gamma\in K$ acts on $\Omega\times\mathbb{R}^{d}$ by pure translation $\gamma(\textbf{x},\textbf{y})=(\textbf{x},T_{\gamma}\textbf{x}+\textbf{y}+\textbf{v}_{\gamma})$. The $K$-invariance of the Hessian quadratic form of $\varphi$ then yields conservation laws
$\frac{\partial}{\partial x_{i}}\frac{\partial}{\partial x_{j}}(F(\textbf{x})\xi(\textbf{x}))=\frac{\partial}{\partial x_{k}}\left(F(\textbf{x})(T_{\gamma}\textbf{x}+\textbf{v}_{\gamma})\right)=0$
for all $\gamma\in K$ and all $i,j,k\in\{1,\dots,n-d\}$, which implies that there exist real $d\times d$-matrices $A_{0},\dots,A_{n-d}$ and an affine mapping $\phi\colon\mathbb{R}^{n-d}\rightarrow\mathbb{R}^{d}$ such that $F(\textbf{x})=L(\textbf{x})^{-1}$ and $\xi(\textbf{x})=L(\textbf{x})^{-1}\phi(\textbf{x})$ where
$L(x_{1},\dots,x_{n-d}):=A_{0}+\sum_{i=1}^{n-d}x_{i}A_{i}$
and hence, up to an affine change of coordinates, the potential $\varphi$ is of the form
$\varphi(\textbf{x},\textbf{y}):=u(\textbf{x})+\langle L(\textbf{x})^{-1}\textbf{y},\textbf{y}\rangle\in C^{\infty}(\Omega)\otimes\mathbb{R}[y_{1},\cdots,y_{d}].$
In particular, for any $\textbf{x}\in\Omega$, the matrix $L(\textbf{x})$ is symmetric and positive definite.
Take any arbitrary $\gamma\in\pi_{1}(M)$. Since the Hessian quadratic form of $\varphi\circ\gamma-\varphi$ vanishes, we have that $\varphi\circ\gamma-\varphi$ is an affine function. Now, we regard $\varphi(\textbf{x},\textbf{y})\in(C^{\infty}(\Omega))[y_{1},\cdots,y_{d}]$ as a quadratic polynomial. The quadratic term of $\varphi$ and $\varphi\circ\gamma$ then coincide, that is,
$L(S_{\gamma}\textbf{x})=U_{\gamma}^{\dagger}L(\textbf{x})U_{\gamma}$
for all $\textbf{x}\in\Omega$, where $U_{\gamma}^{\dagger}$ here is the transpose of $U_{\gamma}$. By considering the limit $\textbf{x}\rightarrow 0$, we have
$A_{0}=L(0)=U_{\gamma}^{\dagger}L(0)U_{\gamma}=U_{\gamma}^{\dagger}A_{0}U_{\gamma}$
and $A_{0}=L(0)\in\operatorname{GL}(\mathbb{R}^{d})$ is a positive definite symmetric matrix. Therefore, we conclude that
$U_{\gamma}\in\operatorname{SL}_{d}(\mathbb{Z})\cap\operatorname{O}(d).$
Since $\operatorname{SL}_{d}(\mathbb{Z})$ is discrete and $\operatorname{O}(d)$ is compact, we have that $\operatorname{SL}_{d}(\mathbb{Z})\cap\operatorname{O}(d)$ is a finite group. By lifting to a finite-sheet covering space of $M$, we may assume that $U_{\gamma}$ is the identity matrix for all $\gamma\in\pi_{1}(M)$, and hence $L(\textbf{x})=L(S_{\gamma}\textbf{x})$ for all $\textbf{x}\in\Omega$.
Now, denote $\textbf{v}:=(\operatorname{tr}(A_{1}),\dots,\operatorname{tr}(A_{n-d}))\in\mathbb{R}^{n-d}$. Then
$\operatorname{tr}(L(\textbf{x}))=\langle\textbf{x},\textbf{v}\rangle+\operatorname{tr}(A_{0}),$
and $L(\textbf{x})=L(S_{\gamma}\textbf{x})$ implies that $\langle\textbf{x},\textbf{v}\rangle=\langle S_{\gamma}\textbf{x},\textbf{v}\rangle$ for all $\textbf{x}\in\Omega$. Therefore, $C^{\infty}(\Omega)^{H}$ is $H$-invariant. Since the $H$-action on $\Omega$ is cocompact, we have that the affine function $L(\textbf{x})$ is bounded, and hence must be constant. Therefore, we obtain $\langle\textbf{x},\textbf{v}\rangle=0$ for all $\textbf{x}\in\Omega$ and hence $\textbf{v}=0$, i.e. $\operatorname{tr}(A_{i})=0$ for all $i=1,\dots,n-d$.
Take any $\textbf{x}=(x_{1},\dots,x_{n-d})\in\Omega$. If
$A(x_{1},\dots,x_{n-d}):=\sum_{i=1}^{n-d}x_{i}A_{i}$
admits a negative eigenvalue, then for sufficient large $t>0$, we have
$L(\textbf{x}\cdot t)/t=A(\textbf{x})+\frac{1}{t}A_{0}$
is not positive definite, contradiction.
Therefore, for any $\textbf{x}=(x_{1},\dots,x_{n-d})\in\Omega$, we have
$A(\textbf{x})=\sum_{i=1}^{n-d}x_{i}A_{i}\geq 0.$
But, since $\operatorname{tr}(A_{1})=\cdots=\operatorname{tr}(A_{n-d})=0$, we have $\operatorname{tr}(A(\textbf{x}))=0$ and hence $A(\textbf{x})=0$ for all $\textbf{x}\in\Omega$. Therefore, we conclude that $L(\textbf{x})\equiv A_{0}$ is a constant matrix-valued function, and hence, up to an affine change of coordinates, we may assume that
$\varphi(\textbf{x},\textbf{y})=u(\textbf{x})+\frac{1}{2}\|\textbf{y}\|^{2}.$
The equation
$\frac{\partial\,T_{\gamma}\textbf{x}}{\partial x_{i}}=\frac{1}{2}\frac{\partial}{\partial x_{i}}\left(F(\textbf{x})(T_{\gamma}\textbf{x}+\textbf{v}_{\gamma})\right)=0,$
where $i\in\{1,\dots,n-d\}$ and $\gamma\in\pi_{1}(M)$, now gives that $T_{\gamma}=0$ for all $\gamma\in\pi_{1}(M)$. In particular, the central extension
$1\longrightarrow K\longrightarrow\pi_{1}(M)\longrightarrow H\longrightarrow 1$
splits as a direct product $\pi_{1}(M)=H\times K$. The theorem is therefore proved. ∎
5 Classification Results in Low Dimensions
5.1 Complete Hessian Surfaces
It is a well-known fact that an orientable closed Hessian surface is homeomorphic to the 2-torus $\mathbb{T}^{2}$. The ultimate objective of this section is to classify all complete Hessian surfaces topologically. To achieve this goal, the following preparations are needed:
Definition 5.1.
For a convex domain $\Omega$ in $\mathbb{R}^{d}$, we denote by
$\operatorname{Aff}(\Omega):=\{f\in\operatorname{Aff}_{d}(\mathbb{R}):f(\Omega)=\Omega\}$
the group of affine automorphisms of $\Omega$.
Lemma 5.2.
Let $\Omega$ be a convex domain in $\mathbb{R}^{d}$ and $\operatorname{recc}(\Omega)$ its recession cone. Let $f\in\operatorname{Aff}_{2}(\mathbb{R})$ and let $A$ be the linear part of $f$, that is, $f(\textup{{x}})=A\textup{{x}}+f(0)$ for all $\textup{{x}}\in\mathbb{R}^{d}$. Suppose that $\Omega=f(\Omega)$. Then $A\textup{{x}}\in\operatorname{recc}(\Omega)$ for all $\textup{{x}}\in\operatorname{recc}(\Omega)$.
Proof.
Take any $\textbf{x}\in\operatorname{recc}(\Omega)$ and $\textbf{y}\in\Omega$. By the definition of a recession cone, it suffices to prove that $A\textbf{x}+\textbf{y}\in\Omega$. But indeed, we have
$A\textbf{x}+\textbf{y}=A\textbf{x}+Af^{-1}(\textbf{y})+f(0)=f(\textbf{x}+f^{-1}(\textbf{y}))\in\Omega$
as required. ∎
The following proposition concerning compactness of the affine automorphism group of a convex body is itself interesting and useful.
Proposition 5.3.
Let $\Omega$ be a bounded convex domain in $\mathbb{R}^{n}$. Then $\operatorname{Aff}(\Omega)$ is isomorphic to a compact Lie subgroup of $\operatorname{GL}_{n}(\mathbb{R})$.
Proof.
It is readily seen that $G:=\operatorname{Aff}(\Omega)$ is a closed subgroup of $\operatorname{Aff}_{n}(\mathbb{R})$. By Cartan’s theorem on closed subgroups, $G$ is a Lie subgroup of $\operatorname{Aff}_{n}(\mathbb{R})$.
Denote by $\overline{\Omega}$ the closure of $\Omega$. By Heine-Borel theorem and the assumption on $\Omega$, we have that $\overline{\Omega}$ is a convex body. Denote by $E_{\Omega}$ the inner Löwner–John ellipsoid of $\overline{\Omega}$. By the uniqueness of the inner Löwner–John ellipsoid, $E_{\Omega}=g(E_{\Omega})$ for every $g\in G$. By linear algebra, if a non-degenerate affine transform $g$ maps an ellipsoid $E$ onto $E$ itself, then $g$ fix the center of $E$. It follows that all elements of $G$ share a common fixed point, namely the center of the ellipsoid $E_{\Omega}$. without loss of generality, we may assume that the ellipsoid $E_{\Omega}$ is centered at the origin $0\in\mathbb{R}^{n}$. Therefore $G$ is a subgroup of $\operatorname{GL}_{n}(\mathbb{R})$.
Denote by $\|\cdot\|_{\operatorname{op}}$ the operator norm. Denote by $\ell_{1}$ and $\ell_{2}$ the diameter and the width of $E_{\Omega}$, respectively. Then it is readily seen that $\|g\|_{\operatorname{op}}\leq\ell_{1}/\ell_{2}$ for all $g\in G$. In particular, $G$ is a bounded subset of $\mathbb{R}^{n\times n}$. Now, again by Heine-Borel theorem, $G$ is compact. ∎
Lemma 5.4.
Let $\Omega$ be a bounded convex domain in $\mathbb{R}^{d}$ and let a subgroup $G$ of $\operatorname{Aff}(\Omega)$ act on $\Omega$ by affine transformation. Then the quotient space $\Omega/G$ is never compact.
Proof.
By Proposition 5.3, up to an affine change of coordinates, we may assume $G$ is a subgroup of the orthogonal group $\operatorname{O}(d)$, and that $0\in\mathbb{R}^{d}$ is the center of mass of $\Omega$. Denote by $\partial\Omega$ the boundary of $\Omega$ in $\mathbb{R}^{d}$. Then the $G$-action on $\Omega$ extends uniquely to $\partial\Omega$. Notice that
$\frac{\partial\Omega\times[0,1)}{\partial\Omega\times\{0\}}\cong\{t\textbf{x}\in\mathbb{R}^{d}:\textbf{x}\in\partial\Omega,0\leq t<1\}=\Omega.$
Then we conclude that
$\Omega/G\cong\frac{(\partial\Omega/G)\times[0,1)}{(\partial\Omega/G)\times\{0\}}$
is non-compact, as it admits a continuous surjective mapping onto the interval $[0,1)$. ∎
We will also use the following classical theorem on crystallographic groups of D. Fried and W. Goldman.
Theorem 5.5 (Fried-Goldman,1983[9]).
Let $G$ be a subgroup of $\operatorname{Aff}(\mathbb{R}^{2})$. Suppose that $G$ acts proper discontinuously on $\mathbb{R}^{2}$ by affine transform. Then $G$ admits a solvable subgroup of finite index.
Proof.
This is a main theorem proved in [9]. ∎
Now we are fully prepared to prove the main theorem of this section.
Theorem 5.6.
Let $((M,\nabla),h)$ be a complete Hessian surface. Then $M$ is diffeomorphic to a complete flat Riemannian manifold of dimension two, that is, one of the following smooth manifolds:
  1. 1.
    the plane $\mathbb{R}^{2}$;
  2. 2.
    the cylinder $\mathbb{S}^{1}\times\mathbb{R}$;
  3. 3.
    the Möbius strip;
  4. 4.
    the 2-torus $\mathbb{T}^{2}$;
  5. 5.
    the Klein bottle.
Proof.
Since $((M,\nabla),h)$ is a complete Hessian manifold of dimension 2, by Theorem 2.12, there exists a planar convex domain $\Omega$ and a discrete subgroup $G$ of the affine group $\operatorname{Aff}_{2}(\mathbb{R})$, such that $G$ acts freely and proper discontinuously on $\Omega$ by affine transform so that $M\cong\Omega/G$.
If $G$ is trivial, then $M=\Omega$ is homeomorphic to $\mathbb{R}^{2}$. If $M$ is closed, then by Proposition 4.3, we have that $\chi(M)=0$ and hence $M$ is homeomorphic to the 2-torus or the Klein bottle. Recall from low-dimensional topology that the fundamental group of a non-compact surface is always free. From now on, we assume that $M$ is non-compact and $G$ is a non-trivial free group.
Denote by $\operatorname{recc}(\Omega)$ the recession cone of $\Omega$. By elementary Euclidean geometry, any convex cone in $\mathbb{R}^{2}$, and hence in particular $\operatorname{recc}(\Omega)$, is congruent to one of the following sets:
  1. 1.
    a point in $\mathbb{R}^{2}$;
  2. 2.
    a straight line, i.e. an 1-dimensional affine subspace of $\mathbb{R}^{2}$;
  3. 3.
    a wedge, i.e. the set $W_{\theta}:=\{z\in\mathbb{C}^{1}:0\leq\operatorname{Arg}(z)\leq\theta\}$ for $0\leq\theta\leq\pi$;
  4. 4.
    the whole plane $\mathbb{R}^{2}$.
Recall from convex geometry that a convex domain is bounded if and only if its recession cone is a point. If $\operatorname{recc}(\Omega)$ is indeed a point then the boundedness of $\Omega$ violates Lemma 5.4. Therefore $\operatorname{recc}(\Omega)$ cannot be a point.
Assume that $\operatorname{recc}(\Omega)$ is the whole plane $\mathbb{R}^{2}$. Then we have $\Omega=\operatorname{recc}(\Omega)=\mathbb{R}^{2}$. Therefore $G$ admits a solvable subgroup $G_{0}$ of finite index by Theorem 5.5.
Assume that $\operatorname{recc}(\Omega)$ is a straight line in $\mathbb{R}^{2}$ or a wedge $W_{\theta}$ with $0\leq\theta\leq\pi$. Consider the group homomorphism
$\phi\colon\operatorname{Aff}_{2}(\mathbb{R})\rightarrow\operatorname{GL}_{2}(\mathbb{R}),\quad f\mapsto(\textbf{x}\mapsto f(\textbf{x})-f(0))$
that is, the map sending every affine transform to its linear part. Then by Lemma 5.2, we have that $\phi(G)$ acts on $\operatorname{recc}(\Omega)$ by linear automorphism. In particular, $\phi(G)$ preserves the apex and the boundary of $\operatorname{recc}(\Omega)$. Therefore, $\phi(G)$ admits a subgroup $H$ of index at most 2, such that all elements of $H$ share at least one eigenvector, and hence $G_{0}:=\phi^{-1}(H)\cap G$ is isomorphic to a subgroup of the group $B$ of invertible upper-triangular real $3\times 3$ matrices. Recall from linear algebra that the length of the derived sequence of Heisenberg group $H_{3}(\mathbb{R})$ is exactly $2$, and observe that $B/H_{3}(\mathbb{R})\cong(\mathbb{R}^{\times})^{3}$ is Abelian. By group theory, we obtain that $B$ is a solvable group, and hence $G_{0}$ ditto. Therefore, $G$ admits a solvable subgroup $G_{0}$ of index at most 2.
This proves that $G$ always admits a solvable subgroup $G_{0}$ of finite index.
Since $G$ is infinite, we have $G_{0}$ ditto. By Nielsen–Schreier theorem, $G_{0}$ is also free. Since any free group of rank $r>1$ is not solvable, the rank of $G_{0}$ is at most 1. Therefore, we conclude that $G_{0}\cong\mathbb{Z}$. Finally, by Schreier’s index formula, we arrive at that in fact the rank of $G$ is also 1, i.e. $G$ is isomorphic to $\mathbb{Z}$.
Therefore, the non-compact surface $M$ has fundamental group $\pi_{1}(M)=\mathbb{Z}$.
By the classification theorem of non-compact surfaces, we have that $M$ is a closed surface $N$ with $k\geq 1$ punctures.
Assume that $N$ is orientable of genus $g$. Then by Seifert-van Kampen theorem, $\pi_{1}(M)$ is a free group of rank r=2g+k1r=2g+k-1. Since $r=1$ and $k>0$, we have that $g=0$ and $k=2$. Therefore $M$ is a sphere with $2$ punctures, i.e. a cylinder.
Assume that $N$ is non-orientable. Then, by the classification theorem of closed surfaces, $N$ is the connected sum of $l$ real projective planes for some integer $l\geq 1$. Again by Seifert-van Kampen theorem, $\pi_{1}(M)$ is a free group of rank r=l+k1r=l+k-1. Since $r=1$ and $k>0$, we obtain $l=k=1$. Therefore $M$ is a punctured $\mathbb{R}P^{2}$, i.e. a Möbius strip.
Quod erat demonstrandum. ∎
5.2 Closed Orientable Hessian 3-Manifolds
The flat tori, or more generally, Bieberbach manifolds are examples of compact Hessian manifolds, but they are trivial in the sense that their affine structures are given by their Levi-Civita connections. In this section, we construct concretely a family of compact radiant Hessian manifolds of Koszul type with negative scalar curvature.
Denote by $\mathbb{H}^{2}:=\{z\in\mathbb{C}:\operatorname{Im}(z)>0\}$ the upper half plane in $\mathbb{C}^{1}$. Consider the product Riemannian metric
$ds^{2}:=d\rho^{2}+\frac{|d\tau|^{2}}{\operatorname{Im}(\tau)^{2}}$
on $\mathbb{H}^{2}\times\mathbb{R}$, where $\tau$ is the coordinate on $\mathbb{H}^{2}$ and $\rho$ is the coordinate on $\mathbb{R}$.
[…]
Consider the quadratic function
$Q\colon\mathbb{R}^{3}\rightarrow\mathbb{R},\quad(x,y,t)\mapsto t^{2}-x^{2}-y^{2}$
and denote $\Omega:=\{(x,y,t)\in\mathbb{R}^{3}:Q(x,y,t)>0\}$. Also, define a smooth function $f\in C^{\infty}(\Omega)$ by
$f(x,y,t):=-\frac{1}{2}\operatorname{log}(t^{2}-x^{2}-y^{2}).$
Then computation yields that the Hessian quadratic form
$2\frac{(tdt-xdx-ydy)^{2}}{(t^{2}-x^{2}-y^{2})^{2}}-\frac{dt^{2}-dx^{2}-dy^{2}}{t^{2}-x^{2}-y^{2}}$
of $f$ is a Riemannian metric on $\Omega$.
Denote by $\nabla$ the Levi-Civita connection of the Euclidean 3-space. Then, by construction, $\nabla(df)$ is a Hessian metric on $(\Omega,\nabla)$.
A tedious but straightforward computation shows that
$\Phi\colon\mathbb{H}^{2}\times\mathbb{R}\rightarrow\Omega,\quad(\tau,\rho)\mapsto\frac{\operatorname{exp}(\rho)}{2\operatorname{Im}(\tau)}(\tau+\bar{\tau},|\tau|^{2}-1,|\tau|^{2}+1)$
is an isometry between $(\mathbb{H}^{2}\times\mathbb{R},ds^{2})$ and $(\Omega,\nabla(df))$.
By the additivity of scalar curvature, we obtain that the scalar curvature of $(\Omega,\nabla(df))$ is constantly 1+0=1-1+0=-1.
Denote by $\tilde{\nabla}:=\Phi^{*}\nabla$ the pullback connection. Then $((\mathbb{H}^{2}\times\mathbb{R},\tilde{\nabla}),ds^{2})$ is a Hessian manifold as $((\Omega,\nabla),\nabla(df))$ is.
From now on, for simplicity, we identify $((\Omega,\nabla),\nabla(df))$ with $((\mathbb{H}^{2}\times\mathbb{R},\tilde{\nabla}),ds^{2})$.
Consider any arbitrary cocompact Fuchsian group action
$(\Gamma,\mathbb{H}^{2})\rightarrow\mathbb{H}^{2},\quad(\gamma,\tau)\mapsto\gamma\cdot\tau$
on $\mathbb{H}^{2}$ so that $\mathbb{H}^{2}/\Gamma=\Sigma_{g}$ is a closed orientable surface of genus $g>1$. Let $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ be a periodic mapping class. By Nielsen’s realization theorem, without loss of generality, we may assume that $\varphi$ lifts to an $\Gamma$-equivariant isometry $\tilde{\varphi}$ of the Lobachevsky plane. Then $G:=\Gamma\rtimes_{\varphi}\mathbb{Z}$ acts on $\mathbb{H}^{2}\times\mathbb{R}$ by isometries via
$G\times(\mathbb{H}^{2}\times\mathbb{R})\rightarrow\mathbb{H}^{2}\times\mathbb{R},\quad((\gamma,n),(\tau,\rho))\mapsto(\gamma\cdot\varphi^{n}(\tau),\rho+n).$
The $G$-action is by construction free and properly discontinuous.
Recall from the theory of Lie groups that the isometry group $\operatorname{PSL}_{2}(\mathbb{R})$ of $\mathbb{H}^{2}$ is isomorphic to the Lorentz group $\operatorname{SO}^{+}(2,1)$, whilst the latter is a Lie subgroup of $\operatorname{GL}_{3}(\mathbb{R})$. It follows that $G$ acts on the convex cone $\Omega$ by linear automorphisms. Therefore, $\tilde{\nabla}$ descends to a flat and torsion-free connection $D$ on the tangent bundle of the mapping torus $M_{\varphi}=(\mathbb{H}^{2}\times\mathbb{R})/G$ of $\varphi$, so that $(M_{\varphi},D)$ is an affine manifold. Moreover, $(M_{\varphi},D)$ is radiant by Proposition 3.5.
Since it always holds that
$(|z|^{2}+1)^{2}-(|z|^{2}-1)^{2}-(z+\bar{z})^{2}=4\operatorname{Im}(z)^{2}$
for every $z\in\mathbb{C}$, we obtain that $Q(\Phi(\gamma\cdot\varphi^{n}(\tau),\rho+n))=Q(\Phi(\tau,\rho))e^{2n}$, and hence
$f(\Phi(\gamma\cdot\varphi^{n}(\tau),\rho+n))=-\frac{1}{2}\operatorname{log}(Q(\Phi(\tau,\rho))\operatorname{exp}(2n))=f(\Phi(\tau,\rho))-n$
for every $(\gamma,n)\in G$ and every $(\tau,\rho)\in\mathbb{H}^{2}\times\mathbb{R}$. Therefore the 1-form potential $df$ of $((\Omega,\nabla),\nabla df)$ is invariant under the $G$-action, and thus descends to the quotient $M_{\varphi}$. In particular, the product Riemannian metric $ds^{2}$ on $\mathbb{H}^{2}\times\mathbb{R}$ descends to a Riemannian metric $h$ on $M_{\varphi}$.
This proves that $((M_{\varphi},D),h)$ is a compact radiant Hessian manifold of Koszul type with constant negative scalar curvature 1-1, as claimed.
We thus arrived at the following theorem:
Proposition 5.7.
Let $\Sigma_{g}$ be a closed orientable surface of genus $g>1$, and let $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ be a periodic mapping class. Then there exist a Riemannian metric $h$ on the mapping torus $M_{\varphi}$ of $\varphi$ and a flat torsion-free connection $D$ on the tangent bundle $T(M_{\varphi})\rightarrow M_{\varphi}$, such that $((M_{\varphi},D),h)$ is a closed and orientable Hessian $3$-manifold of Koszul type.
Under the scope of Thurston’s geometrization, the high genus examples in the proof of Proposition 5.7 are those locally modeled by $\mathbb{H}^{2}\times\mathbb{R}$. A natural next step is to construct similar examples for the genus one case. We will now prove that these examples indeed exist, and they all admit flat models.
Proposition 5.8.
Let $\varphi]\in\operatorname{Mod}(\mathbb{T}^{2})$ be a periodic mapping class. Then there exist a Riemannian metric $g$ on the mapping torus $M_{\varphi}$ of $\varphi$, such that $(M_{\varphi},g)$ is a closed orientable Bieberbach $3$-manifold.
Proof.
Recall from the arithmetic of $\operatorname{SL}_{2}(\mathbb{Z})\cong\operatorname{Mod}(\mathbb{T}^{2})$ that every finite order element of $\operatorname{SL}_{2}(\mathbb{Z})$ is conjugate to either $\pm\operatorname{Id}$ or one of the following matrices
$A_{1}=\left[\begin{array}[]{cc}0&-1\\ 1&-1\end{array}\right],\quad A_{2}=\left[\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right],\quad A_{3}=\left[\begin{array}[]{cc}0&-1\\ 1&1\end{array}\right]$
or their inverses. Now, it suffices to notice that $\pm\operatorname{Id},A_{2}$ are isometries of the rhombic elliptic curve $\mathbb{C}^{1}/(\mathbb{Z}\oplus i\mathbb{Z})$, and $A_{1},A_{3}$ are isometries of the hexagonal elliptic curve $\mathbb{C}^{1}/(\mathbb{Z}\oplus\omega\mathbb{Z})$, where $\omega\in\mathbb{C}$ is the $3$-rd primitive root of the unity. ∎
The examples constructed above are consistent with the results established earlier. As predicted by Theorem 4.1, for a closed orientable surface $\Sigma_{g}$ of genus $g\geq 1$, the mapping torus $M_{\varphi}$ of a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ is indeed a fiber bundle over $\mathbb{S}^{1}$. The fundamental group of $M_{\varphi}$ is a semidirect product of $\mathbb{Z}$ and $\pi_{1}(\Sigma_{g})$, which is indeed infinite and torsion free, as stated in Corollary 2.6 and Proposition 2.16.
However, given the nature of all examples that we constructed, one can be curious about whether there exists an example of closed orientable Hessian $3$-manifold that is a mapping torus of a mapping class of infinite order. We now prove that this is impossible, which leads to a classification theorem of closed Hessian $3$-manifolds.
Theorem 5.9.
Let $((M,\nabla),h)$ be a closed orientable Hessian $3$-manifold. Then there exists a periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$ such that $M$ is diffeomorphic to the mapping torus of $\varphi$, where $\Sigma_{g}$ is a closed orientable surface of genus $g\geq 1$.
Proof.
By Theorem 4.15, without loss of generality, we may assume that $(M,\nabla)$ is the product of a compact hyperbolic affine manifold with a flat torus of dimension $d\in\{0,\dots,3\}$.
If $d\geq 2$, then the desired result holds trivially.
If $d=1$, then $M$ is homeomorphic to the $3$-torus $\mathbb{T}^{3}$ by Theorem 5.6.
Now, assume that $d=0$. Then $(M,\nabla)$ itself is a compact orientable hyperbolic affine manifold. The desired result, in this case, follows directly from Corollary 4.12. ∎
Recall that a Dehn twist on a closed orientable surface $\Sigma_{g}$ of genus $g>0$ is of infinite order. Therefore, in particular, Theorem 5.9 implies that the nilpotent Heisenberg manifold
$M:=H_{3}(\mathbb{R})/H_{3}(\mathbb{Z})$
cannot support any Hessian metric.
Remark 5.10.
Combine Proposition 5.7 and Proposition 5.8, we obtain that the converse statement of Theorem 5.9 is also true: For every closed orientable surface $\Sigma_{g}$ of genus $g\geq 1$ and every periodic mapping class $\varphi]\in\operatorname{Mod}(\Sigma_{g})$, there exist a Riemannian metric $h$ on the mapping torus $M_{\varphi}$ of $\varphi$ and a flat torsion-free connection $D$ on the tangent bundle $T(M_{\varphi})\rightarrow M_{\varphi}$, such that $((M_{\varphi},D),h)$ is a Hessian manifold. These thus gives a complete topological classification of closed orientable Hessian $3$-manifolds.
5.3 Closed Orientable Hessian 4-Manifolds
In this final section, we classify all closed orientable Hessian $4$-manifolds topologically.
For an oriented closed topological 4-manifold $M$, as per usual, we denote by $Q_{M}$ its intersection form.
Lemma 5.11.
Let $((M,\nabla),h)$ be an oriented closed Hessian 4-manifold. Then $\chi(M)=\operatorname{sign}(Q_{M})=0$. In particular, the first Betti number of $M$ is positive and the second Betti number of $M$ is even.
Proof.
The first claim follows from Proposition 1.54.3 and Hirzebruch’s signature theorem. Moreover, since $\operatorname{dim}M=4$ is even and $\chi(M)=0$, by Poincaré duality the first Betti number of $M$ is positive and the second Betti number of $M$ is even. ∎
Recall that all closed orientable 4-manifolds have $\operatorname{spin}^{c}$ structures. The following property is specific for Hessian 4-manifolds.
Proposition 5.12.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. If $(M,h)$ is spin then $M$ admits an almost complex structure with vanishing first Chern class. Moreover, the second Chern classes of all almost complex structures on $M$ are zero.
Proof.
Recall that $(M,h)$ is spin if the second Stiefel–Whitney class $w_{2}(TM)\in\mathrm{H}^{2}(M;\mathbb{Z}/2\mathbb{Z})$ of $M$ vanishes.
Moreover, by Ehresmann-Wu theorem, $M$ admits an almost complex structure $J$ with first Chern class $c\in\mathrm{H}^{2}(M;\mathbb{Z})$ if and only if the following equation
$3\operatorname{sign}\left(Q_{M}\right)+2\chi(M)=c^{2}$ (3)
is satisfied and $w_{2}(TM)$ is the $\bmod 2$ reduction of $c$. For our case, by Lemma 5.11, $3\operatorname{sign}\left(Q_{M}\right)+2\chi(M)=0+0=0$. If $w_{2}(TM)=0$, then $c=0$ satisfies equation (3) and its $\bmod 2$ reduction is trivially equal to $w_{2}(TM)$.
Now, equip the tangent bundle $TM\rightarrow M$ with an almost complex structure $J$. Then, by the theory of characteristic classes, the first Pontryagin class can be expressed as
$p_{1}(TM)=c_{1}(TM)^{2}-2c_{2}(TM)$
in terms of Chern classes. Since $p_{1}(TM)=0$ by Proposition 1.5, and $c_{1}(TM)^{2}=0+0=0$ again by Ehresmann-Wu theorem, we conclude that $c_{2}(TM)=0$. ∎
Furthermore, the following result shows that there are only two possibilities for the intersection form of an oriented closed Hessian 4-manifold, depending on the existence of $\operatorname{spin}$ structure.
Proposition 5.13.
Let $((M,\nabla),h)$ be an oriented closed Hessian 4-manifold with the second Betti number $b_{2}(M)=2m$. If $(M,h)$ is spin, then it holds that $Q_{M}=H^{\oplus m}$ where
$H:=Q_{\mathbb{S}^{2}\times\mathbb{S}^{2}}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right];$
if $(M,h)$ is not spin, then $Q_{M}=\operatorname{diag}(\underbrace{1,\dots,1}_{m},\underbrace{-1,\dots,-1}_{m})$.
Proof.
Recall that $(M,h)$ is spin if the second Stiefel–Whitney class $w_{2}:=w_{2}(TM)$ of $M$ vanishes. By Wu’s formula, we have $Sq^{2}(a)=a\smile a=a\smile w_{2}$ for all $a\in\mathrm{H}^{2}(M;\mathbb{Z}/2\mathbb{Z})$. It is then readily seen that $Q_{M}$ is even if and only if $w_{2}=0$. The desired result then follows from Hasse-Minkowski classification theorem of indefinite unimodular symmetric bilinear form over $\mathbb{Z}$, and the fact that $\operatorname{sign}(Q_{M})=0$. ∎
In fact, a closed orientable Hessian 4-manifold is either spin or Riemannian flat:
Proposition 5.14.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. Suppose that the Riemannian manifold $(M,h)$ is not flat. Then $(M,h)$ is spin.
Proof.
By Proposition 4.2, there exist a fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$ with connected orientable fiber $N$. Denote by $dt\in\Omega^{1}(\mathbb{S}^{1})$ the volume form of the unit circle, and denote by $\phi]\in\operatorname{Mod}(N)$ the monodromy of $f\colon M\rightarrow\mathbb{S}^{1}$. Then, by Wang sequence, the second cohomology group $\mathrm{H}^{2}(M;\mathbb{Z})$ decomposes as
$\mathrm{H}^{2}(M;\mathbb{Z})=\mathrm{H}^{2}(N;\mathbb{Z})^{\phi}\oplus A$
where
$A:=\{[\alpha]\smile f^{*}[dt]:[\alpha]\in\mathrm{H}^{1}(N;\mathbb{Z})\}\cong\mathrm{H}^{1}(N;\mathbb{Z})/\operatorname{Im}(\phi^{*}-\operatorname{Id})$
and $\mathrm{H}^{2}(N;\mathbb{Z})^{\phi}$ is the submodule of $\mathrm{H}^{2}(N;\mathbb{Z})$ consists of $\phi$-invariant classes. It is readily seen that for any $\beta]\in\mathrm{H}^{2}(N;\mathbb{Z})^{\phi}$ or $\beta]\in A$, we have $\beta]\smile[\beta]=0$. Also, by equivariant Poincaré duality, the intersection form $Q_{M}$ restricts to a non-degenerated pairing between $\mathrm{H}^{2}(N;\mathbb{Z})^{\phi}$ and $A$. Therefore, in particular, we have $Q_{M}=H^{\oplus m}$ where
$H:=Q_{\mathbb{S}^{2}\times\mathbb{S}^{2}}=\left[\begin{array}[]{cc}0&1\\ 1&0\end{array}\right]$
and hence $(M,h)$ is spin by Proposition 5.13. ∎
Finally, with all preparations done, we now state and prove our main theorem:
Theorem 5.15.
Let $((M,\nabla),h)$ be a closed orientable Hessian 4-manifold. Then $M$ is diffeomorphic to one of the following smooth manifolds:
  1. 1.
    a Bieberbach 4-manifold;
  2. 2.
    the mapping torus of an isometry of a hyperbolic 3-manifold;
  3. 3.
    a $\Sigma_{g}$-bundle over $\mathbb{T}^{2}$ with periodic monodromy,
where $\Sigma_{g}$ here is the closed orientable surface of genus $g>1$.
Proof.
By Theorem 4.15, there exists a finite-sheet covering map $\pi\colon E\rightarrow M$ such that the lifting $(E,\pi^{*}\nabla)$ is the product of a compact hyperbolic affine manifold with a flat torus of dimension $d\in\{0,\dots,4\}$.
If $d\geq 3$, then the desired result holds trivially.
If $1\leq d\leq 2$, then $M$ is diffeomorphic to a Bieberbach 4-manifold or a $\Sigma_{g}$-bundle over $\mathbb{T}^{2}$ with periodic monodromy, by Theorem 5.6 and Theorem 5.9.
Now, suppose that $d=0$. Then $(M,\nabla)$ itself is a compact orientable hyperbolic affine manifold. By Corollary 4.12, there exists a fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$ of which fiber is a closed orientable $3$-manifold NN, such that the monodromy $\phi]\in\operatorname{Mod}(N)$ of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$ is periodic. Since $N$ is aspherical as $M$ is, in particular the second homotopy group $\pi_{2}(N)=0$, and hence $N$ is irreducible. If $N$ is atoroidal, then by the hyperbolization theorem, there exists a Riemannian metric $g$ on $N$ such that $(N,g)$ is an hyperbolic $3$-manifold, and that $\phi$ is isotopic to an isometry of $(N,g)$. From now on, we shall assume that $N$ is not atoroidal. Then by Thurston’s geometrization, we have that $N$ is a graph manifold. Therefore, the fundamental group $\pi_{1}(M)$ admits a solvable subgroup of finite index. In view of Lemma 5.11 and Proposition 5.12, the Enriques–Kodaira classification yields that $M$ is diffeomorphic to one of the following complex surfaces:
  1. 1.
    a complex 2-torus;
  2. 2.
    a primary Kodaira surface;
  3. 3.
    a hyperelliptic surface;
  4. 4.
    an Inoue surface,
while complex 2-tori and hyperelliptic surfaces are all diffeomorphic to Bieberbach 4-manifolds, and the primary Kodaira surfaces and Inoue surfaces are not mapping tori of any periodic mapping classes.
The desired result is therefore proved. ∎
Remark 5.16.
Analogously to Example 5.7 and Example 5.8, examples for Theorem 5.15 can all be produced. Here we omit the detailed construction and leave it to the cautious readers.
Acknowledgements
The author wants to express the author’s most sincere gratitude to Weiyi Zhang for very useful discussions and general advice along the whole preparation process of this manuscript, and to Maxwell Stolarski for a careful proofreading with various valuable suggestions. Special thanks should also go to Saul Schleimer who inspires the author’s work on Hessian 3-manifolds.
This research is completed while the author is studying at the mathematics institute of University of Warwick. The author therefore would like to thank the hospitality of University of Warwick.